A Complete Bibliography of Publications in Mathematics and Computers in Simulation: 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

22 August 2025 Version 1.00

Title word cross-reference

(2 + 1) [1879]. (3 + 1) [1689]. (t, e, s) [1369]. (t, s) [1369]. 3 × 3 [1231]. [−1, 1] [1562]. ^{+} [542]. ^{2} [664]. _{2} [62, 732, 1525, 359, 1686]. _{3} [62, 1686]. _{7} [62]. _{∞} [732]. α [509, 1660, 315]. β [315]. C [664, 1581, 1169]. C^{1} [343]. C^{∞} [1758]. C^{1} [1096, 856, 859]. d = 2 [1559]. δ [315]. ϵ [643]. γ [315]. H_{∞} [1666, 1738]. ∞ [1775]. k [1376, 402]. L [732]. L^{1} [1562]. L_{2} [1468]. M [27, 936, 279, 678]. F_{2} [876]. R^{4} [448]. R^{d} [25]. C [1822]. RT_{k} [1103]. n [1340, 1093, 1691, 1152, 402]. p [901, 1776, 1012]. q [671]. RSV [11]. S [1752, 375, 27]. T [375, 225, 27, 583, 274]. W [230]. x [901].

-Copula [225]. -dimensional [1689, 1879].

-estimator [936]. -helix [509]. -identical [1152]. -insensitive [643]. -iteration [1752]. -linear [876]. -means [1581, 1169]. -mixture [274]. -net [27]. -out-of- [402]. -Phase [1660]. [?]PowellFortes:2011:HFM,Fortes:2014:HFM. -quasilinear [901]. -regimes [279]. -sequences [1369]. -sided [671]. -th [1376]. -th-order [1691]. -three [794]. -two [794]. -values [1776]. -variate [1775].

[?]PrBaGomez:2010:FPC.

1 [1251, 1682]. 1-D [1682]. 102C [937]. 1d [1214, 797, 348, 432, 516].

2 [1434]. 2-D [1434]. 2010 [135, 222, 4, 108, 93, 151, 807]. 2011 [293, 383, 446, 251, 364, 307, 414]. 2012 [524, 464, 580, 539, 630, 620, 611]. 2013 [762, 668, 657, 730, 673]. 2014 [841, 867, 879, 820, 926, 907]. 2015 [955, 1032, 1040, 988, 1022, 1074, 1065, 1057]. 2016 [1163, 1210, 1135, 1120, 1181, 1149, 1172, 1255, 1247]. 2017 [1352, 1413, 1454, 1305, 1374, 1430, 1770]. 2018 [1521, 1587, 1631, 1491, 1567, 1576, 1622, 1612, 1597]. 2019 [1784, 1857, 1671, 1750, 1695, 1748, 1834, 1803]. 23 [1569]. 25-point [1080]. 2D [1589, 494, 1843, 1531, 55, 1409, 835].

3-dimensional [1602]. 3d [1418, 485, 917, 1822, 1728, 1825, 569, 990, 409, 1113, 31, 1561, 560].

4D [1526].

5-point [1556].

6th [1386].

7 [1569]. 7/23 [1569]. 78 [1490]. 7th [1818].

80 [153]. 82 [608]. 8th [404].

93 [1003].

abdominal [1756]. Abelian [1547]. Absorbing [1024, 124]. absorption [1627, 1392, 1117]. abstract [1198]. Aby [650]. AC [695, 1290, 1715]. ACA’2013 [1202]. accelerate [1848]. Accelerated [1139, 921, 918, 1713]. accelerated-time [921, 918]. acceleration [1223, 30, 1175]. accident [1569]. Accord [775, 777]. account [1302]. accounting [185]. accuracy [1688, 1438, 85]. Accurate [1460, 815, 1796]. acetate [1107]. Acoustic [1223, 573, 1235, 1564]. acquired [1501]. across [505]. activation [1646, 1656, 229, 371]. active [1794, 1741]. activities [323]. activity [1443]. actuarial [529]. actuators [1300]. acute [814]. ADALINE [691]. Adams [1625]. adaptation [409, 1633, 603, 1594]. adaption [471]. Adaptive [543, 864, 752, 137, 722, 883, 561, 1605, 1791, 1705, 299, 329, 232, 1313, 1165, 997, 1217, 932, 1426, 61, 103, 25, 1034, 376, 72, 417, 1875]. Adding [32, 1500]. additive [1487, 1646, 546, 351, 1656]. adhesive [1506]. adhesive-contact [1506]. adjoint [823]. adjustment [742]. admitted [1547]. advanced [361, 216]. Advancement [320]. Advances [1402]. advancing [862]. advection [458, 1405, 349, 1237, 1563, 157]. Aedes [1529]. aegypti [1529]. aeronautical [1721]. aerosols [268]. affect [754]. affine [210, 1691, 792]. after [947]. aftereffect [19]. against [1059]. age [330, 1329, 1330, 1340, 455, 810, 273]. age-structured [330, 1329, 1330, 1340, 455]. ageing [1299]. agent [15, 1343]. Agent-based [1343]. agents [983, 970]. Agglomeration [974]. aggregation [436]. agonists [1336]. agriculture [661]. ahead [1720, 1286]. AIDS [1831]. aimed [692]. air [1605, 700]. aircraft [1734, 1657, 714, 178]. airfoil [83]. airport [921]. Al [1686]. alarm [6]. algal [1887]. Algebra [1202, 29, 663, 1206, 396, 916]. Algebraic [400, 1394, 1710, 399, 962, 664, 1262, 1699, 402, 391, 162]. algebraically [1259]. algebras [1204]. Algorithm [1868, 1606, 1199, 1283, 1647, 715, 21, 950, 824, 702, 1725, 1552, 1774, 1648, 1141, 567, 1880, 690, 589, 836, 1077, 1687, 1531, 1258, 41, 622, 1742, 606, 139, 1877, 1314, 1408, 1113, 1436, 1705, 1651, 1681, 1658, 311]. Algorithm-II [1606]. Algorithmic [1204, 401, 163, 1143]. Algorithms [1764, 1367, 199, 205, 1486, 461, 466, 566, 990, 711, 422, 402, 1450, 599, 20]. alignments [903]. Allee [928, 1624, 1381, 830, 1347, 1795, 766]. Allen [1124]. allocation [1828]. alloy [56].

Almost [960, 884, 946, 1590, 1419, 371]. alone [1537]. along [625, 626, 1322]. alpha [1336]. alternate [219, 1740]. alternative [1668, 13, 96]. alternatives [1059]. alternator [1727]. aluminum [1298]. American [205, 1174, 103]. among [15]. amplifier [495]. AMS [704]. Amsterdam [15]. analyses [212, 812]. Analysis [19, 1415, 1401, 929, 1286, 1463, 1661, 147, 275, 1689, 694, 1189, 1718, 1089, 1199, 654, 373, 1272, 441, 1662, 1505, 1155, 1060, 600, 254, 875, 1308, 1099, 1100, 1604, 1646, 1548, 281, 1260, 1514, 1333, 1214, 1122, 1583, 937, 900, 902, 342, 1871, 1854, 456, 797, 1275, 1715, 347, 1530, 438, 1013, 1624, 437, 827, 1812, 1079, 677, 563, 217, 1578, 1212, 71, 1657, 979, 1252, 144, 1793, 1381, 1131, 1790, 794, 1731, 1359, 700, 1082, 830, 1029, 8, 209, 678, 1151, 164, 1288, 131, 1157, 1114, 440, 1570, 816, 754, 1656, 1293, 1528, 1533, 1882, 7]. analysis [1215, 423, 696, 1346, 1563, 1285, 1847, 1643, 526, 672, 676, 1865, 1084, 649, 1582, 542, 825, 1890, 1789, 371, 1569, 766, 493]. analytic [298, 1869, 920, 424]. Analytical [634, 1874, 699, 1546, 701, 562, 1728, 1146, 244, 1269, 1239, 1301, 696]. analyzing [1316]. ancillary [1731]. and/or [1720]. aneurysms [1756]. ANFIS [1169]. anisotropic [329, 1870, 409, 1230, 1633, 132]. annealing [1367]. annuity [1312]. annular [586]. annulus [1570]. anomalies [1873]. anomalous [1298]. ANOVA [1299]. ant [1687]. anti [1232]. anti-continuum [1232]. antiretroviral [1501, 456]. any [822, 685]. aorta [1747]. aortic [1756]. Aperiodic [313]. Apex [1798]. Apnea [320, 943]. appearance [853, 302]. appearance/disappearance [302]. appended [1071]. applicable [38]. Application [20, 216, 1102, 677, 812, 574, 468, 1071, 651, 1683, 1475, 1391, 333, 1868, 1760, 1625, 137, 1725, 430, 547, 266, 155, 1636, 751, 375, 369, 607, 227, 697, 1556, 1358, 147, 1660, 397, 545, 1027, 1084, 633, 1716, 1747, 1469, 857, 335, 1141, 1735, 568, 117, 1732, 1665, 645]. application-oriented [1556]. Applications [1202, 29, 904, 1387, 1444, 1886, 1284, 1822, 1549, 260, 1500, 1703, 909, 916, 1337, 737, 721, 689, 1300, 474, 240, 396, 422, 1482, 27, 1242, 1215, 1427, 1250, 838, 1544, 1111]. Applied [207, 599, 469, 803, 511, 1299, 855, 1690, 1563, 374, 1606, 1544, 404]. applying [1704]. approach [1387, 1390, 290, 469, 1443, 1584, 1332, 1155, 1478, 1760, 917, 595, 775, 997, 319, 1046, 1864, 76, 1377, 125, 1227, 1737, 1400, 1581, 663, 1757, 1887, 1158, 970, 416, 1677, 1174, 1129, 1518, 273, 1069, 1649, 1262, 82, 322, 721, 160, 1463, 25, 795, 1645, 231, 77, 843, 1699, 398, 691, 798, 718, 173, 672, 998, 1599, 1250, 1813]. approaches [1733, 787, 882, 623, 963, 1349]. Approximate [675, 1051, 1349]. Approximating [894, 1341, 1415, 243, 1809]. Approximation [1386, 1818, 1313, 854, 569, 1291, 207, 404, 311, 1469, 333, 1094, 1478, 1097, 571, 1546, 1197, 340, 1226, 897, 1825, 1776, 1342, 1470, 1109, 838, 1042, 1112, 356, 855, 83, 1562, 1198, 585, 408]. approximations [1487, 1688, 1757, 228, 128, 1476, 10, 1483, 422]. April [293, 524, 841, 1163, 1352, 1521]. aquifer [1089]. aquifers [1633]. arbitrary [451, 132]. architecture [1725, 1742, 27]. architectures [1684]. area [1648]. area-wide [1648]. arising [864, 247, 897, 1757, 1674]. arithmetic [1445]. arm [1740]. ARMA [1097]. armlets [1027]. Arnoldi [1165]. Arrangement [1868]. array [100, 527, 1277, 1485]. Array-RQMC [1485]. arrays [709, 1712, 1418, 1533]. arterial [491, 38]. artery [1885, 1354]. arthropod [397]. Article [759]. Artificial [443, 232, 457, 1069, 828, 1658]. Asian [287, 276]. aspects [238, 1334, 899, 1015, 418, 1082, 793, 1250]. asphalt [564]. assembled [1607]. Assessing [1553, 259]. Assessment [943, 1715, 1292, 1673, 282, 406, 1143]. assignment [1668]. assimilation [348]. Assimulo [1068]. assisted [1880]. associated [1204, 302, 182]. association [331, 851]. associative [1582, 960]. asymmetric [267, 288, 472, 1013, 1432, 583, 323]. asymmetrical [1271]. asymmetries [974]. Asymptotic [1078, 1160, 162, 12, 494, 125, 345, 929, 1778]. Asymptotically [1329, 1330]. Asynchronous [1507]. Atanassov [1321]. atherosclerotic [1885]. atmosphere [1069]. atmospheric [419, 359]. attenuation [514]. Attracting [591]. attraction [995, 1376, 912, 309]. attractivity [1008]. attractors [796, 874, 1316, 982, 379, 870, 992]. Attribute [399]. atypical [1887]. Augmented [37, 1046, 1443, 624]. August [135, 383, 762, 1210, 1413, 1587, 1784]. Australia [273, 754, 755]. Australian [614, 277]. automata [52]. Automated [34, 948, 720]. Automatic [1536, 615, 1203, 1281, 721]. automaton [1637]. automotive [1146]. autonomous [1684, 1316, 591, 1382, 255, 1324]. autoregressive [116, 1309, 253, 274, 744]. AUV [1153]. auxiliary [466]. avalanches [1760]. avascular [41]. average [543, 1091, 1573, 1309, 1872, 1775, 1525, 1001]. AVR [1813]. aware [771]. axial [1270]. axially [511]. axis [211, 1673, 1113]. axisymmetric [165, 555].

B [328, 814, 1141, 1079, 1355, 809, 835]. B-form [328]. B-spline [1355]. B-splines [328, 1141, 835]. B.E. [206]. back [1060, 1510]. back-EMF [1060]. Backward [1889, 1590, 816, 733, 1571]. baggage [921]. Bagley [1314]. balance [1223, 1838, 675]. balanced [558, 1864, 559, 1027]. balancing [1387]. BAM [1067, 144, 229]. Bandcount [789]. Banded [1758]. Bangladesh [1322]. banks [453]. Barca [1887]. barley [1697]. Barrier [1866]. Barzilai [1510]. Based [1619, 232, 941, 332, 852, 1043, 573, 366, 335, 695, 1093, 1287, 1502, 1478, 1647, 1096, 858, 600, 1323, 1309, 1080, 1130, 604, 174, 439, 1146, 702, 1142, 1864, 317, 97, 433, 1552, 192, 860, 829, 742, 1402, 1290, 549, 461, 1644, 1786, 437, 962, 1735, 567, 1166, 377, 892, 1827, 757, 1282, 1343, 990, 826, 851, 1406, 155, 409, 957, 1378, 1318, 1799, 1636, 182, 662, 1572, 1291, 1607, 375, 589, 1807, 1649, 43, 679, 1531, 721, 1140, 632, 1605, 1883, 1738, 771, 725, 749, 1003, 1867, 700, 1741, 175, 1692, 25, 1645, 724, 713, 1601]. based [933, 1471, 781, 800, 931, 1116, 560, 1499, 170, 835, 1726, 710, 718, 1427, 1564, 1534, 576, 1594, 719, 526, 315, 321, 649, 1599, 102, 825, 1123, 651, 534, 1538, 1108, 1742]. Basel [775, 777]. bases [1205, 1772, 1474]. Bashforth [1625]. Basins [995, 912, 1376, 982]. basis [566, 859, 1355, 1176]. Bat [1880]. batch [101]. battery [1282, 183, 1537]. Bayes [1393, 472, 1852]. Bayesian [775, 224, 1141, 1378, 1084, 1599, 283]. BDDC [561, 45]. beacon [1608]. beam [1300]. beams [180]. bearing [1273, 372]. bearings [1673]. bed [338, 1410]. Beddington [1335, 830, 1456]. bee [828]. behavior [1446, 627, 1667, 829, 1160, 930, 512, 544, 1347, 1249]. behaviors [1310]. behaviour [1036, 791]. Bell [1207]. benchmark [1401]. benchmarking [1409]. Benjamin [1226, 505, 519]. bent [1472]. Bermudan [205, 873]. Bermudan/American [205]. Bernoulli [1549, 1448, 1850]. Bernstein [1250]. berry [397]. Bertrand [1311]. Bertrand-type [1311]. Bessel [1444]. best [460, 408]. Best-Approximation [408]. beta [369]. betas [614]. between [1764, 1530, 529, 1111, 1738, 562, 304]. Bi [697, 1233, 451]. Bi-criteria [697]. bi-frames [451]. bi-integrable [1233]. Bias [270, 1615, 1463, 747]. Bias-corrected [270]. Biased [20]. BIBO [110]. bicriteria [1880]. bicubic [1825, 1828, 898]. bidirectional [1582, 449]. bifurcating [1137]. Bifurcation [790, 373, 797, 830, 1184, 973, 385, 726, 1624, 827, 1251, 636, 1167, 1578, 1615, 1810, 1157, 816, 1417, 1582]. bifurcations [158, 1602, 659, 234, 1152, 794, 163, 449, 1795]. bilinear [895]. bimodal [1532, 386]. bimolecular [1471]. binary [758, 286, 56]. binding [980]. binomial [1140]. bio [1034]. bio-reactors [1034]. biochips [573]. bioeconomic [1381]. biological [812, 661, 1349]. biomass [1034, 1345]. Biomath [1328, 1328]. biomedicine [1307]. Biorthogonal [1027]. birds [1324]. Birnbaum [116, 1532]. birth [1335]. Birthmark [1867]. Bisheh [1752]. Bivariate [1448, 1676]. black [1486]. bladed [86]. Blind [1502, 147]. block [1844, 276, 1860, 162]. blocks [571]. blood [1756, 1885, 46, 1747, 1354, 1686]. Board [5, 18, 50, 68, 94, 109, 122, 136, 152, 168, 195, 223, 237, 252, 264, 294, 308, 326, 365, 384, 395, 415, 428, 447, 465, 480, 501, 525, 540, 553, 581, 598, 612, 621, 631, 641, 658, 669, 674, 683, 707, 739, 740, 763, 785, 806, 821, 842, 849, 868, 880, 890, 908, 915, 927, 940, 956, 968, 989, 1006, 1023, 1033, 1041, 1049, 1058, 1066, 1075, 1087, 1121, 1136, 1150, 1164, 1173, 1182, 1192, 1211, 1221, 1248, 1256, 1267, 1279, 1306, 1327, 1353, 1362, 1375, 1385, 1414]. Board [1422, 1431, 1439, 1440, 1441, 1455, 1465, 1466, 1492, 1503, 1522, 1543, 1566, 1568, 1577, 1588, 1598, 1613, 1620, 1621, 1623, 1632, 1640, 1669, 1670, 1672, 1696, 1745, 1749, 1751, 1769, 1785, 1804, 1817, 1832, 1833, 1835, 1858]. bodies [88]. body [942, 1407, 1499]. Boltzmann [1680, 388, 623, 59]. bond [750, 1558, 1616]. bone [1155, 42]. bones [857]. boost [722, 726, 1723, 725]. Bootstrap [660, 1144, 227]. Border [158, 659, 234]. Border-collision [659]. born [745]. borne [1862]. Borwein [1510]. Bose [483]. both [229]. bound [1634, 1486, 1002, 1481, 1484]. boundaries [531, 950, 52, 403]. Boundary [1555, 1424, 329, 1024, 485, 1018, 124, 1545, 488, 212, 823, 319, 494, 1214, 64, 1854, 1513, 388, 420, 555, 129, 623, 1884, 1512, 1527, 1844, 1539, 1865, 1860, 1747, 162, 798]. boundary-value [1844]. bounded [21, 869, 390, 1779, 1592, 1780, 1789]. boundedness [1034]. bounds [1369, 1807, 353]. Boussinesq [486, 491, 494, 1224, 1227]. box [1486, 613]. Braess [659]. brain [1443]. Branch [1647]. break [271]. breaker [607]. breaks [112]. Bricard [1206]. BRICS [780]. bridges [1558]. bright [504]. Brinkman [1592, 1875]. broiler [416]. broken [304]. broken-trend [304]. Brownian [1366]. brush [184]. Brusselator [1331]. BSDEs [1174]. BSE [1169]. bubbles [425]. buck [535, 449]. Buckley [1388]. buckling [1275]. Bucy [314]. budding [817]. building [147, 717]. bulks [701]. bundle [1272]. Burbea [851]. burden [441]. Burgers [1547, 1786, 546, 1877, 1879]. Burr [936]. bus [1663]. business [1200, 1797]. butter [828]. butter-oil [828]. BVPs [1447]. bypass [84]. bypasses [46].

C [102]. C-means [102]. cable [172, 1275]. Cahn [1124, 1216]. calcium [1339]. Calculating [132]. Calculation [999, 720, 392, 701, 86, 467]. calculations [1514, 62, 1564]. calculus [829, 1435]. Calendar [17, 49, 67, 92, 107, 121, 134, 150, 167, 194, 221, 236, 250, 263, 292, 306, 325, 363, 382, 394, 413, 427, 445, 463, 479, 500, 523, 538, 552, 579, 597, 610, 619, 629, 640, 656, 667, 682, 706, 729, 738, 761, 784, 805, 819, 840, 848, 866, 878, 889, 906, 914, 925, 939, 954, 967, 987, 1005, 1021, 1031, 1039, 1048, 1056, 1064, 1073, 1086, 1119, 1134, 1148, 1162, 1171, 1180, 1191, 1209, 1219, 1246, 1254, 1264, 1304, 1326, 1351, 1361, 1373, 1384]. Calendar [1412, 1421, 1429, 1453, 1489, 1520, 1542, 1575, 1586, 1596, 1611, 1639, 1694, 1707, 1744, 1768, 1783, 1802, 1816, 1856, 1892, 1278]. Calender [1630]. Calibration [767]. Calques3D [31]. Camassa [1249]. canals [1584]. canard [1070]. canard-cycles [1070]. cancer [1836, 1183]. Canticles [33]. capabilities [32, 1794]. capacitance [23]. capacitor [1298]. capacitors [1881]. capacity [282]. capillary [1388, 484]. capillary-gravity [484]. capital [976]. capture [409]. Capturing [1028, 1754]. car [918]. carbon [815, 568, 1025]. carcinogenesis [1346]. cardiac [95, 1625, 1601]. care [990]. Carlo [1600, 267, 19, 1059, 200, 758, 1371, 911, 199, 205, 204, 1368, 1356, 1757, 1106, 23, 312, 197, 119, 1476, 206, 1773, 1781, 1473, 60, 380, 203, 61, 25, 26, 1076, 1539, 1471, 1705, 28, 1778, 1681]. Carreau [1449]. carrying [882]. cars [565]. Cartesian [1747]. CAS [918, 31]. cascade [508, 1293, 449]. cascaded [1265, 175]. case [1024, 1188, 350, 1775, 721, 1554, 1555, 397, 650, 476]. case-study [1554, 1555]. cases [320]. casting [1761]. cat [1496]. cataract [749, 1003]. catastrophic [972]. catchment [282]. categorization [436]. Cauchy [1094, 881, 26, 920, 586, 1261]. causal [514]. causality [154, 304]. causation [1569]. cavitation [1099, 1100]. cavity [313, 1462]. CBC [1486]. CD4 [542]. cell [330, 543, 1091, 1573, 810, 722, 1403, 1843, 1282, 1299, 1721, 1293, 47, 1189, 1737, 206]. cell-average [543, 1091, 1573]. cell-centered [1843]. Cell/B.E. [206]. cells [1473, 170, 1339, 1285, 542]. cellular [52, 1637, 475]. censored [1393, 1318, 1799, 421, 1580, 526]. censoring [1143]. Center [385, 663, 1257, 1137]. centered [1843, 1797]. centers [1497]. central [1433, 356]. centroidal [826]. ceramic [1558]. certain [407, 1077]. CEV [473, 998]. CFD [1617, 1554, 1555, 1754]. Chain [1600, 1551, 1106, 1483, 1061]. chains [1130, 204, 1368]. challenges [431]. Challenging [1544, 1145]. chance [8, 633]. change [758, 159, 113]. changes [253]. Channel [1853, 339, 1686]. Chaos [1462, 1319, 468, 789, 1184, 595, 997, 1168, 1757, 1634, 797, 314, 1061, 1789]. chaotic [796, 137, 874, 313, 1044, 883, 1316, 791, 375, 680, 379, 452, 1200, 931, 1319, 1035, 470, 299]. Characterising [746]. Characteristic [585, 1406, 1053, 96, 1765]. characteristic-nonconforming [1053]. characteristics [172, 338, 1663, 496, 1495, 958]. characteristics-mixed [958]. characterization [857, 1709, 892]. characterize [1083]. charge [1827]. chart [751]. chartists [971]. chassis [141]. Chebyshev [124, 492, 1758, 1156, 1601]. Checking [1776]. chemical [1460, 468]. chemicals [544]. chemostat [854, 1320, 1678]. chemotherapy [1183]. Chen [995]. Cherkas [1257]. Chernivtsi [1178]. Chikungunya [1417]. children [1178]. chiller [1658]. chimpanzees [814]. China [756, 754]. Chinese [778, 453, 374]. chip [695, 1730]. choice [316, 758]. cholera [1338]. Christov [1228]. chronification [587]. CIR [10]. circles [400]. Circuit [1044, 685, 1296, 1733, 692, 698, 607]. circuit-breaker [607]. circuit-type [1733]. circuits [720, 431, 175]. circular [1348]. circulation [602, 376]. citations [759]. Clarke [1667]. class [1573, 448, 1101, 894, 347, 929, 1158, 1166, 1652, 884, 1677, 1109, 434, 664, 1126, 1642, 633, 959, 1524]. classes [1775]. classic [182]. Classical [1378, 486, 461]. classification [1647]. classifications [331]. Clean [756]. Clifford [1098]. climate [282]. clinical [42]. clique [1608]. close [1185]. closed [302, 141]. closed-loop [141]. cloud [1637]. clustered [752]. clustering [776, 1647, 102]. CML [330, 1013]. CMOS [431]. co [1008, 359]. co-dimension [1008]. coagulation [1781]. coal [1410]. coal-bed [1410]. coalescing [242]. coarse [1688, 559]. coarsening [22]. coast [1322]. coastal [1633]. code [721, 1554, 1555]. codewords [1606]. Codimension [1795, 794]. Codimension-one [1795, 794]. codimension-two [1795]. coding [648]. coefficient [269, 1603, 735]. coefficients [1515, 1682, 1780, 163, 98]. Coercive [759]. Cogging [1628, 699]. Cohen [1067, 144, 536]. cointegrating [114, 312]. cointegration [660, 13, 425]. Cole [155]. collapse [1244]. collecting [268]. collective [529]. Collector [184]. Collision [517, 158, 659, 1680, 234, 474]. collisionless [513]. collisions [495]. collocation [1394, 1821, 546, 1557]. Colombia [1805]. colony [1687, 828]. color [1098]. Coloured [435]. combat [1861]. combination [210, 771]. combinatorial [1204]. combinatorics [400]. Combined [338, 59, 862, 1725]. combining [1733, 1169, 1651]. combustion [498]. Comment [153]. Comments [833]. commercial [453]. commodity [747]. Communication [1665]. commutation [1740]. Comp [937]. compact [454]. Comparative [1680, 1530, 1883, 254, 1583, 754]. compare [1759]. Comparing [316, 30, 244, 227, 1736, 472, 1450]. Comparison [339, 266, 873, 711, 1178, 1124, 607, 1269, 1626, 178, 779, 941, 1059, 1356, 660, 1738, 1838, 514, 963, 747, 349]. comparisons [1369, 1485]. compatible [564]. compensation [1323, 1711, 723, 651]. compensator [1722]. competing [1861, 1852]. competition [443, 1357]. competitive [1662, 1078, 1419, 1890]. Complementarity [576, 853]. complementary [1635]. Complete [1851, 1534, 46]. Complex [627, 302, 15, 982, 1382, 1589, 457, 548, 951, 1849, 829, 969, 217, 1666, 140, 680, 322, 104, 1495, 1001, 1185, 1436, 1010, 1698, 1679, 1747, 1789, 1307]. complex-number [829]. complex-valued [1679]. complexities [934]. Complexity [1370, 1526, 764, 532]. component [1614, 1553, 297, 1714, 321]. component-based [321]. components [1274, 704, 1690]. composite [176, 367, 632]. compositions [34]. Compound [1240, 529]. comprehensive [892, 1673]. compressible [1214, 1591, 1437, 1800, 341, 74, 1403, 350, 354, 958, 520]. Comput [1490, 153, 1003, 608]. Computation [1220, 1355, 1435, 163, 548, 1203, 224, 238, 1402, 556, 1129, 813, 1607, 1012, 1705, 45, 646]. Computational [212, 472, 817, 1212, 1511, 259, 402, 532, 201, 1143, 997, 943, 1551, 1334, 32, 1015, 1786, 1307, 419, 793, 418, 1250]. computations [1175]. Computer [1202, 64, 230, 1111, 29, 457, 916, 71, 1206, 1082, 396, 397, 1569, 663]. Computers [652, 969, 1364]. Computing [328, 1388, 577, 1544, 207, 404, 1242, 1852, 403, 1540, 1886, 1502, 1416, 1757, 23, 563, 1322, 1839, 599, 1608, 216, 438]. Concentration [22, 1034]. concentrations [1525]. concept [1839, 399]. ‘Concept-Attribute-Value’ [399]. Concordia [182]. concrete [82]. concretes [1798]. condensate [508]. condensates [483]. condition [1424, 945, 388, 1555, 1516, 521]. conditional [116, 287, 1807]. conditioned [871]. conditioning [1848]. conditions [1024, 219, 124, 991, 823, 494, 1214, 944, 1008, 385, 663, 1104, 1257, 419, 1513, 297, 129, 623, 1741, 1721, 920]. conduction [1854, 432]. conductivities [95]. conductors [1272]. cones [165, 241]. Conference [1386, 1818, 1220]. configuration [1432]. confined [1089, 1798]. confinement [489]. conforming [965]. congestive [441]. Conjugate [139]. connected [1583, 833, 1731, 712, 387, 723, 2, 1716]. Connecting [31]. connectivity [1710]. consensus [1122]. conservation [1593, 1094, 83]. conservative [1871, 1124, 14, 1765]. conserved [1212, 297]. conserving [1099, 1100, 388, 509]. considering [1719, 1761, 1805]. consistent [572, 1737, 461]. constant [1682, 1766, 1860, 1878]. constants [1334]. constitutive [1798]. constrained [366, 744]. constraint [792, 1499]. constraint-based [1499]. constraints [177, 853, 1397, 801, 1112, 1510, 559, 164, 887]. Constructing [997, 1316, 1026]. Construction [1154, 310, 1115, 623, 1472]. constructions [1365]. consumer [754]. consumption [1184, 1717]. contact [1101, 569, 1509, 566, 1506]. Contents [481, 502]. context [1720, 1237]. contexts [993]. contingent [978]. Continual [88]. Continuation [565, 298, 1509, 424]. Continuous [928, 1394, 1592, 1836, 822, 1876, 997, 310, 459, 204, 1460, 1780, 300, 1626, 1382, 1010]. Continuous-time [928, 1876, 1460, 1382]. continuously [1275, 975]. continuum [1232]. contours [600]. contracting [339]. contracts [824]. contrapuntal [34]. Contribution [1721, 1336]. control [330, 1584, 1535, 685, 1296, 1060, 1284, 1719, 810, 595, 1684, 1297, 875, 137, 247, 459, 430, 722, 709, 716, 693, 910, 58, 732, 1062, 314, 1724, 225, 390, 1044, 727, 434, 1578, 1657, 182, 375, 138, 1666, 140, 721, 1798, 689, 1156, 1605, 1738, 1791, 725, 1729, 1722, 845, 1741, 175, 240, 1151, 724, 792, 208, 661, 931, 1319, 1295, 7, 460, 1594, 719, 1177, 1310, 1841, 141, 476, 111, 1678, 1659, 1674, 475, 110, 1145, 1524, 530, 961, 1789, 585, 216, 1649]. control-variate [460]. Controllability [1840, 1194, 164, 1643]. controlled [1876, 1309]. controlled-signalized [1876]. controller [695, 1297, 1323, 687, 186, 188, 726, 932, 143, 632, 1034, 1153, 1123, 1813]. controllers [874, 1738, 1319]. Controlling [1061]. controls [1808, 1740]. convection [1395, 1821, 1728, 408, 665, 228, 1796, 1845, 454, 1053, 417, 1765, 585, 378]. convection-dominated [1395, 1053, 1765]. Convergence [1018, 1060, 1403, 297, 1590, 1554, 329, 1487, 1848, 1051, 341, 1774, 1485, 930, 1131, 113, 1034, 1126, 198, 672, 161, 1125]. Conversion [1777, 1738, 1713]. converter [1535, 1734, 722, 726, 1794, 714, 190, 1276, 1740, 1289, 992]. converter-storage [714]. converters [535, 1731, 1293, 449]. Convertible [1616]. convex [1166, 1510]. convexity [1556]. convexity-preservation [1556]. Conway [671]. cooled [700]. cooling [605]. cooperation [1846, 1797]. cooperative [1425]. coordinated [716]. coordinates [944, 990, 130]. Copula [225]. core [206, 719]. corner [881, 623]. cornering [1495]. corners [560]. coronary [441, 46]. corporate [750]. corpses [1808]. corrected [270]. correction [411, 1555]. corrector [727, 1077]. correlated [1759, 115]. correlation [758]. correlations [287]. Corrigendum [1490, 1003]. cosine [1470]. Cost [828, 846, 1295, 111]. costs [978]. Coulomb [569, 1509, 577]. count [1676]. Counting [1205]. country [780, 285]. Coupled [1826, 1331, 692, 897, 567, 970, 71, 1232, 452, 1661, 1495, 77, 647, 1243, 1879, 541, 1809, 146, 517]. Coupling [197, 79, 1425, 555, 1537]. couplings [1233]. Cournot [1311, 869, 977, 99, 261, 1789]. covariance [1199]. cover [922]. covers [923]. Cox [613]. cracked [1589]. Craig [378]. Cramér [1109]. Crank [1531]. crashworthiness [645]. Creating [650, 1113]. Credit [972, 748]. criteria [1813, 1636, 697, 1266, 14, 1035, 229]. criterion [959]. critical [488, 385, 521, 1528, 959, 1244, 934]. CRONE [1738]. cross [331, 497, 105, 1712]. cross-classifications [331]. cross-diffusion [497]. Crossed [933]. Crossed-derivative [933]. crowd [1873]. Crowley [1790]. crystal [1883]. CS [893]. CS-RBF [893]. CSI [1823]. CT [1581]. CTL [142]. Cu [62, 1686]. cube [23, 25]. cube-based [25]. Cubic [1821, 1447, 504, 663, 1257, 1828, 1344, 859, 1883]. cultural [7, 961]. Currency [774]. current [100, 1272, 695, 685, 1, 1490, 1719, 186, 726, 1710, 1715, 187, 882, 1418, 724, 1699, 1594]. currents [172, 76, 1718]. curvature [1212]. curve [1037, 418, 1258]. curved [1869, 1686]. curves [1546, 302, 1824, 1212, 550]. curvilinear [1424]. CVT [1717]. cycle [846, 1735, 1200, 163]. cycle-based [1735]. cycles [790, 158, 448, 1070, 385, 1137, 301]. cyclic [872, 1014, 1293]. Cyclicity [796, 663]. cycling [810]. cylinder [1274]. cylindrical [1437, 990, 1697].

D [1079, 1682, 1434]. daily [1525]. damage [772, 1609, 1533]. Damped [1125, 1840, 1871]. damping [1165, 1083]. dangerous [1262]. Darcy [1592]. dark [483]. Data [1260, 1309, 348, 1683, 335, 257, 1822, 1546, 271, 964, 1122, 260, 1703, 549, 1825, 1644, 567, 757, 893, 1799, 1657, 1252, 369, 1653, 1787, 8, 545, 1534, 672]. data-driven [672]. database [27]. datasets [214]. Davydov [509]. day [1720]. DBD [190, 1841]. DC [1734, 1731, 1289, 178, 449, 992]. DC-link [1731]. DC/DC [1734, 992]. dead [1808]. dead-infectious [1808]. deal [319]. DeAngelis [1335, 830, 1456]. death [1340, 1335, 1189]. December [222, 446, 1454, 1631, 1857]. Decentralized [716]. Decentralized-coordinated [716]. Decision [261, 398]. declination [651]. decomposition [1836, 366, 858, 1552, 860, 127, 1405, 559, 549, 226, 1690]. deconvolution [1502]. dedicated [720, 1739]. defined [319, 403]. Definition [800]. deflated [593]. deformable [568]. deformation [1812, 557, 520]. Degenerate [350, 1627, 1820, 1137, 355]. degradation [1287, 948]. degree [1333, 1497]. delamination [1506]. Delaporte [1337]. Delaunay [1545, 946]. delay [1184, 1308, 1646, 1046, 869, 1132, 787, 1013, 1062, 368, 811, 450, 813, 1615, 9, 244, 1456, 1201, 1198, 831, 1679, 145, 638, 1186, 542, 1860, 111, 162, 1317, 530]. delay-driven [1308]. delay-integro-differential [831]. delayed [636, 1167, 1578, 1152, 138, 1790, 647, 676, 475, 825, 1317, 371]. delays [1840, 1662, 1646, 1377, 732, 1251, 1067, 144, 536, 1381, 591, 975, 1810, 1645, 14, 624, 1656, 229, 1417, 1582, 1459, 960]. delivery [42]. demagnetization [1715]. demand [1668, 281, 977, 277]. demographic [1347]. dengue [1650]. Denoising [1550, 1105, 836, 1027]. densities [1759]. density [100, 863, 350, 1773, 701, 1418, 361]. density-driven [863]. dependence [1592, 283]. dependent [1700, 1395, 1792, 1515, 937, 900, 902, 74, 787, 1530, 1509, 233, 1773, 270, 1112, 1560, 423, 303]. depends [350]. depth [1108]. Derivation [1227, 1437, 1800]. Derivative [662, 1376, 1664, 1844, 933, 1434, 378]. Derivative-based [662]. derivative-to-derivative [1376]. derivatives [370, 813]. Derive [34]. derived [1517]. describe [564]. described [562]. describing [638, 1529]. Description [1239, 1409]. descriptor [1873, 387]. Design [1323, 1870, 860, 757, 535, 714, 431, 1722, 1314, 410, 1293, 846, 177, 181, 337, 1546, 715, 698, 176, 687, 174, 874, 702, 192, 693, 1282, 1677, 727, 550, 471, 632, 1739, 697, 1269, 44, 1721, 190, 696, 718, 1606, 1599, 1716, 646, 645, 1619]. designed [81]. designs [752]. destructive [857]. detailed [779]. detecting [1262]. detection [1297, 1290, 590, 474, 795, 1873, 401, 1699, 1651, 102]. detectors [119]. determinant [1444]. Determination [177, 531, 903, 438, 312, 1599]. determine [96]. Determining [1261, 1764, 1654]. deterministic [1477, 104, 226, 1450]. detrended [1215]. Development [1823, 420, 756, 584, 702]. Developments [1427]. devices [864, 702, 714, 1730]. DFIG [177]. DFT [671]. DGFEM [665]. DGS [31]. diabetes [1882]. diagnosis [1296, 1581, 1715, 1554, 1295, 170]. diagnostic [1727]. diagonal [1139, 28]. Diagonalization [1682]. Diagonally [528]. diagram [56]. diameters [947]. diametric [400]. DIEBM [807]. diesel [1605]. difference [1837, 1154, 1080, 945, 1187, 1851, 1845, 454, 896, 1763, 575, 103, 1877, 798, 1563, 1349, 1889, 1674, 1317]. differences [1159]. differencing [1229]. different [1389, 692, 1214, 962, 584, 1741, 1533, 178]. Differential [1594, 1423, 1469, 172, 1092, 788, 1487, 1018, 1313, 803, 1760, 448, 1602, 1052, 310, 248, 239, 1667, 1849, 246, 802, 1015, 347, 929, 390, 1762, 1851, 813, 9, 1077, 1682, 1258, 1590, 1557, 164, 1560, 1345, 1643, 1198, 831, 148, 999, 1809, 145, 542, 1860, 1616, 162]. differential-algebraic [162]. differentialequations [1013]. differentiated [1311, 979]. differentiation [1043, 163]. diffuse [1399, 56]. diffusion [333, 1702, 1821, 245, 1846, 1298, 408, 665, 64, 125, 497, 126, 1228, 127, 228, 1000, 418, 1862, 1158, 1405, 990, 349, 587, 1796, 536, 1458, 591, 1654, 1884, 1159, 1845, 352, 1435, 105, 557, 1117, 417, 635, 1603, 1889, 733, 1571, 585, 378]. diffusion-deformation [557]. diffusion-reaction [417]. Diffusional [380]. diffusions [770, 300, 1483]. diffusive [1624, 811]. digital [1297, 353, 1461]. dilatational [1230]. dimension [824, 1224, 1008, 1559]. dimension-reduction [824]. dimensional [1469, 366, 796, 158, 1602, 258, 1217, 205, 1515, 1437, 605, 204, 1854, 234, 40, 885, 1853, 1316, 1252, 679, 354, 1382, 1493, 1863, 1873, 1877, 1689, 1483, 1450, 1879, 1746, 1244, 110, 378]. dimensions [1433, 73, 561, 356, 1481, 403]. diminishing [1077]. diode [1283]. dioxide [568]. Direct [1709, 1515, 228, 1, 1490, 1017, 81, 508, 1019, 725, 700, 1722, 845, 1741, 1285, 467, 1594]. direct-driven [700]. directional [1559]. directions [438]. director [1740]. Dirichlet [219, 1227]. Dirichlet-to-Neumann [1227]. disappearance [853, 302]. disc [634]. discontinuities [1423, 787, 971]. discontinuity [1007, 1008, 794]. Discontinuous [336, 786, 74, 795, 790, 796, 803, 1400, 965, 802, 1138, 1626, 800, 301]. discrepancy [1321, 1369, 873, 1468, 1365, 1481]. Discrete [1090, 1820, 1392, 73, 1425, 87, 332, 1676, 1433, 186, 1404, 228, 1880, 827, 1827, 1480, 1042, 243, 1435, 1200, 356, 1562, 519, 534, 934, 1795]. discrete-time [186]. discretely [459]. discretisation [533]. discretization [1424, 548, 1308, 1080, 1457, 1512, 1787, 637, 361]. discretizations [884, 1758, 509]. discs [1881]. disease [441, 1862, 1178, 1528, 1698, 7, 1496, 961]. diseases [1701]. disk [86]. dispersal [1618]. dispersion [495, 1344, 1484, 157]. dispersive [1238, 1529]. displacement [958]. Dissipative [1657, 490, 1227]. dissipativity [1656]. dissolution [333]. distance [1663, 1130]. distinct [584]. distortion [1711]. Distributed [61, 1840, 1323, 1132, 975, 1753, 1814, 530, 960]. Distribution [178, 156, 1683, 100, 331, 822, 1393, 1292, 1734, 1142, 317, 753, 260, 1703, 756, 834, 1109, 1829, 230, 1318, 1799, 421, 369, 737, 701, 203, 896, 1561, 1534, 1580, 1143, 844]. distributions [1500, 1652, 1636, 510, 1111, 227, 161, 1128]. disturbance [1691, 390, 471, 1605, 1741]. disturbances [627, 1677, 1791]. disturbed [632]. divergence [851, 1140]. division [1013]. divorce [289]. Dixon [30]. DNA [1607, 1606]. DNLS [1232]. Do [754]. does [124]. Domain [1552, 127, 1836, 366, 858, 1366, 1265, 1870, 965, 860, 1405, 1758, 1827, 1592, 1869, 559, 586, 309]. domains [1545, 950, 881, 319, 74, 1159, 146]. domestic [780, 277, 755]. dominated [1395, 454, 1053, 1606, 1765]. dose [1609]. doses [456]. Double [844, 1627, 334, 1197, 1777, 1724, 1344, 535, 801, 1882, 1145]. double-integrator [1145]. Double-looped [844]. double-precision [1777]. doubly [1265, 1318, 191, 700, 175]. doubly-fed [175]. Downside [1025]. DP [566]. DPC [1003, 749]. DPC-based [1003, 749]. drainage [1408]. drift [811]. drilling [295]. drive [695, 171, 186, 188, 845, 1295]. driven [573, 1487, 863, 1308, 313, 1462, 549, 546, 1212, 351, 700, 672, 746]. driver [141]. drives [177, 172, 1060, 686, 1297, 687, 1715, 690, 1699, 1726]. droop [1535]. drop [1236]. droplet [268, 296]. drops [1107]. drug [456, 42]. dry [791]. dry-friction [791]. drying [1874]. DSP [724]. DSP-based [724]. dual [1449, 910, 614, 557]. dual-betas [614]. dual-mixed [1449]. dual-porous [557]. duality [671]. duals [1472]. due [1322]. duopoly [977, 582]. duration [116, 781]. duration-based [781]. during [1697]. duty [1717]. Dynamic [1188, 189, 975, 688, 183, 978, 934, 1388, 1199, 1311, 625, 626, 1813, 917, 774, 627, 1583, 1737, 767, 549, 1644, 261, 1129, 1045, 42, 1517, 272, 832, 980, 1310, 1071, 32]. Dynamical [238, 869, 1381, 7, 1347, 177, 1840, 786, 1099, 1100, 137, 1007, 1036, 764, 243, 175, 795, 1029, 1493, 1863, 1194, 1814]. Dynamics [814, 799, 736, 1010, 1357, 1496, 530, 974, 330, 1340, 267, 1043, 1702, 1078, 1579, 943, 286, 872, 1014, 1501, 977, 302, 897, 969, 228, 780, 1861, 53, 1195, 1336, 957, 1081, 1578, 419, 1701, 984, 1178, 512, 1200, 474, 795, 1382, 1461, 1001, 397, 800, 1637, 1339, 517, 1243, 793, 541, 1797, 638, 1186, 1529, 825, 1789].

e-learning [443]. E-training [442]. each [350]. East [276]. Ebola [1463]. echo [1299]. ecological [1446, 627, 825, 530]. Economic [973, 1635, 778, 1043, 441, 1292, 1359, 275, 7, 1716, 1250]. economical [1720]. economics [969, 1200]. economies [276]. eddy [1272, 76]. edge [1545, 590, 606]. Editorial [196, 1363, 1504, 1467, 1442, 1770, 51, 169, 429, 1007, 5, 18, 50, 68, 94, 109, 122, 136, 152, 168, 195, 223, 237, 252, 264, 294, 308, 326, 365, 384, 395, 415, 428, 447, 465, 480, 501, 525, 540, 553, 581, 598, 612, 621, 631, 641, 658, 669, 674, 683, 707, 739, 740, 763, 785, 806, 821, 842, 849, 868, 880, 890, 908, 915, 927, 940, 956, 968, 989, 1006, 1023, 1033, 1041, 1049, 1058, 1066, 1075, 1087, 1121, 1136, 1150, 1164, 1173, 1182, 1192, 1211, 1221, 1248, 1256]. Editorial [1267, 1279, 1306, 1327, 1353, 1362, 1375, 1385, 1414, 1422, 1431, 1439, 1440, 1441, 1455, 1465, 1466, 1492, 1503, 1522, 1543, 1566, 1568, 1577, 1588, 1598, 1613, 1620, 1621, 1623, 1632, 1640, 1669, 1670, 1672, 1696, 1745, 1749, 1751, 1769, 1785, 1804, 1817, 1832, 1833, 1835, 1858, 238, 1193]. EEG [373]. Ef [1017]. Ef-Gaussian [1017]. Effect [1766, 1618, 200, 187, 1624, 764, 450, 323, 1381, 614, 830, 1243, 1144, 1347, 1795, 766, 1495]. effective [1746]. Effects [495, 114, 1686, 837, 928, 1662, 602, 1446, 943, 439, 899, 1862, 22, 377, 1511, 1028, 985, 423, 46, 1348, 1756, 1160]. Efficiency [1229, 275]. Efficient [1593, 951, 772, 1796, 71, 1513, 1572, 1477, 454, 1878, 1755, 78, 692, 702, 218, 411, 1869, 622, 1153]. EIA [824]. eigenvalue [1090, 360]. eigenvalues [132]. Einstein [483]. elastic [772, 1081]. elasto [1798, 574]. elasto-plastic [1798, 574]. elastodynamics [461]. Electric [100, 178, 487, 1, 1490, 1292, 1734, 605, 58, 1735, 1714, 1717, 697, 410, 1533]. electrical [715, 1734, 1298, 1710, 1714, 1281, 714, 190, 696, 1716]. Electro [1733]. Electro-thermal [1733]. electrochemical [1604, 1533]. electrode [100, 527, 1418, 1533]. electrodeposition [245]. electrodynamic [1273]. electrolytic [1298]. electromagnetic [1761, 702, 231]. electromagnetics [965]. electromechanical [1730, 1083]. Electron [200]. electronic [720, 1514, 1277, 1730, 54]. electronics [1739, 719]. electrophysiology [1625]. electrotherapy [100, 527]. element [329, 1332, 370, 1308, 1260, 604, 944, 1217, 73, 1449, 192, 897, 994, 128, 1081, 1458, 679, 1531, 1512, 160, 1495, 77, 72, 1053, 958, 83, 417, 1410, 534, 1765, 585, 1054]. elements [1593, 1398, 1103, 344, 555, 87]. elementwise [570]. elimination [1545]. ellipsoids [474]. elliptic [100, 1515, 1843, 1851, 1159, 355, 1494, 1243]. elliptically [517]. Embedded [1730, 715, 471, 714, 689]. embodied [846]. emergency [714]. emergent [1001]. EMF [1060]. emission [1841]. emissions [359]. emphasis [1364]. empirical [436, 1406, 1485, 65, 745]. emulation [1723]. enabling [861]. enclosed [1271]. encoder [1297]. encryption [1035]. end [441, 686, 1663, 690, 1276]. end-stage [441]. endemic [309]. Endogenous [301, 981]. endpoint [1758]. ends [131]. Energy [1736, 174, 1282, 1741, 713, 1740, 846, 181, 1593, 600, 1811, 1725, 722, 1036, 461, 509, 1281, 183, 1011, 1732, 1153, 1035, 83, 710, 1713, 1716, 1658, 1709]. Energy-Power [1709]. engine [720, 1605]. engineering [1460, 1512]. Enhanced [1644, 1712, 653].

Enhancing [1407]. ensemble [1196, 1581]. ensembles [1194]. enter [1890]. Entropic [1107]. entropy [1839]. entry [971]. enumerating [1205]. envelopes [1203]. envelopment [1122, 8]. Environment [1386, 1818, 736, 443, 1359, 1866, 1650]. environmental [335, 1703, 275]. environments [1203, 8]. enzyme [216]. EOQ [1201]. EP [315]. EP-based [315]. epicardial [209]. epidemic [1154, 875, 1792, 1808, 318, 827, 1167, 450, 1498, 15, 1463, 1459, 1379, 309]. epidemiological [1347, 638]. epilepsy [373]. equal [745]. equality [887]. equation [904, 1627, 1424, 1388, 1837, 790, 1392, 1024, 124, 1196, 489, 854, 504, 490, 1438, 1080, 1224, 945, 458, 1226, 1102, 215, 1849, 1871, 994, 1786, 128, 834, 1762, 592, 546, 506, 507, 1344, 1216, 129, 1654, 10, 1883, 1838, 514, 1314, 26, 1076, 164, 131, 521, 360, 963, 1053, 391, 1434, 417, 635, 1563, 98, 518, 519, 562, 1479, 586, 1603, 1865, 1889, 1616, 1878, 735, 1765, 1261, 157, 1571]. Equations [1220, 1423, 1469, 852, 1090, 1820, 1092, 334, 78, 1487, 1394, 1760, 488, 1821, 1017, 1547, 1052, 494, 964, 248, 1515, 1667, 1227, 246, 802, 1015, 1843, 787, 929, 1229, 1158, 1138, 155, 1771, 587, 1796, 1851, 1124, 813, 1232, 1077, 1531, 1159, 244, 1019, 1235, 1590, 352, 1877, 1539, 855, 1601, 422, 1560, 1126, 1345, 1643, 1879, 1198, 831, 1809, 145, 959, 1860, 1850, 162, 534, 1317, 1681, 146, 1054, 378, 517, 1243]. equilibria [1602, 1037, 1013, 870]. Equilibrium [582, 290, 1761, 1416, 753, 1648, 633, 309]. equity [266]. Equivalence [944]. equivalent [1199, 175, 624]. Era [431]. Erratum [937, 608]. Error [389, 1457, 347, 131, 1260, 1070, 937, 900, 902, 1871, 1486, 80, 1476, 1077, 1807, 243, 1830, 72, 576, 1875]. essential [1198]. estimate [602, 1097, 1260, 843]. estimates [779, 1457, 1779, 80, 1830, 1516, 72, 576, 1875]. Estimating [331, 279, 337, 936, 311].

Estimation [1093, 991, 823, 616, 179, 1827, 367, 1534, 1580, 653, 1332, 1297, 1052, 911, 1070, 935, 1634, 1290, 1644, 466, 909, 348, 1378, 421, 1460, 269, 1773, 1473, 1281, 1258, 352, 203, 614, 300, 105, 44, 1525, 1699, 358, 747, 1778, 1746, 844, 343, 567]. estimator [1691, 1367, 202, 845, 113, 936, 1852, 12]. estimators [1393, 886, 896, 745, 460]. estrogen [1336]. EU [1025]. Euclidean [1130]. Euler [496, 802, 550, 1235, 1590, 1780]. European [397]. eutrophication [825]. evaluate [1617]. evaluating [199]. Evaluation [749, 1003, 634, 1864, 1404, 407, 769, 614]. evaluations [1636]. event [1475, 1645]. event-triggered [1645]. Events [17, 49, 67, 92, 107, 121, 134, 150, 167, 194, 221, 236, 250, 263, 292, 306, 325, 363, 382, 394, 413, 427, 445, 463, 479, 500, 523, 538, 552, 579, 597, 610, 619, 629, 640, 656, 667, 682, 706, 729, 738, 761, 784, 805, 819, 840, 848, 866, 878, 889, 906, 914, 925, 939, 954, 967, 987, 1005, 1021, 1031, 1039, 1048, 1056, 1064, 1073, 1086, 1119, 1134, 1148, 1162, 1171, 1180, 1191, 1209, 1219, 1246, 1254, 1264, 1304, 1326, 1351, 1361, 1373, 1384]. Events [1412, 1421, 1429, 1453, 1489, 1520, 1542, 1575, 1586, 1596, 1611, 1630, 1639, 1694, 1707, 1744, 1768, 1783, 1802, 1816, 1856, 1892, 1278]. evidence [758, 312, 112, 416]. Evolution [1540, 1627, 1220, 1258, 322, 1239, 670, 57, 1594]. evolutionary [808, 979, 1436, 646]. exact [1508, 1468]. exactly [509]. example [278]. excavators [1811]. excess [266]. exchange [1299, 754, 170, 755]. excilamps [1841]. excitation [1284, 1724, 689, 1738]. Excited [483, 1300, 541]. exclusion [312]. exercise [1409]. exhibiting [1316]. Existence [504, 1667, 1112, 1419, 498, 1884, 355, 870, 959, 1662]. existing [903]. exit [1366, 1652, 406]. exogeneity [276]. expansion [1549, 468]. expansions [12]. expectation [232, 1141]. expectations [1311, 99]. expected [117, 1653]. experienced [1653]. experiment [493]. Experimental [715, 44, 1091, 567, 1841, 449, 1282]. experiments [319, 38, 1721, 1713]. expert [399, 962, 1262, 1259]. experts [962]. explaining [15]. explicit [872, 1014, 856, 1481, 1649]. Exploitation [979]. exploiting [1742]. exploration [435]. Explorer [919]. exponent [1340, 1102, 959]. Exponential [1625, 247, 1642, 470, 1679, 145, 475, 1662, 1142, 1415, 1667, 1500, 1229, 536, 528, 454, 1534, 229, 1143, 1317, 644, 161]. exponentiality [1059]. exponentially [768]. Exponentiated [737]. exponents [1392]. expressions [1886, 675]. Extended [654, 208, 1309, 1591, 688, 1605, 1463]. Extension [1398, 1556]. externalities [625, 626]. Extracting [214]. extraction [1108, 1794, 892]. extrapolating [1531]. extrapolation [1175]. extrapolative [582]. extraproximal [1416]. Extreme [782, 1475, 777, 273, 283]. extremes [161]. Extremum [430].

Face [560]. Face-based [560]. facility [368]. factor [759, 727, 719]. Factoring [920]. factorization [993, 1508, 451]. factors [472]. fails [1758]. failure [441, 368, 526, 1145, 311]. Falkner [1844]. Falkner-type [1844]. fall [1510]. fall-back [1510]. families [1316]. family [1091, 995, 874, 663, 1376, 930]. fan [751]. farm [275]. Fast [776, 1406, 1302, 1527, 1288, 115, 172, 1338, 1546, 693, 947, 1517, 1301, 432]. Faster [1779]. fatigue [39]. fault [177, 1296, 1684, 1297, 795, 1699, 1293, 1295]. fault-operation [1293]. fault-tolerant [1295]. Faults [1295, 1297, 692, 893]. Faure [201]. FDTD [1883]. FE [433, 557]. feasible [1205]. feature [1887]. featured [1432]. features [406, 1651]. February [251, 668, 1135, 1305, 1491, 1671]. fed [1265, 101, 191, 175, 1276]. fed-batch [101]. feedback [727, 1666, 632, 1319, 1345, 1123, 1348]. Feedforward [434]. Feir [505]. Feller [1434]. FEM [826]. femoral [1155, 46]. fermentation [101]. ferry [1735]. FETI [566]. FETI-DP [566]. fever [1650]. fiber [39]. field [200, 1823, 1277, 1400, 1474, 1432, 511, 59, 845, 422, 1570, 56, 1533, 835, 173, 57]. field-programmable [1277]. field-weakening [845]. fields [487, 458, 58, 1872, 838, 1564]. Filippov [1523]. Fill [81]. Fill-ins [81]. filled [1824]. Filling [1104, 1824, 343, 856, 1793]. filter [1196, 1463, 1741, 724, 315, 314]. filtered [1560]. filtering [1309, 208]. filters [1105, 268, 357]. filtration [1627]. fin [1728]. finance [1469, 969, 202]. financial [972, 1634, 377, 206, 1215, 971]. finders [1376]. finding [1752, 870]. Finite [333, 945, 192, 897, 130, 801, 1410, 1524, 1387, 329, 1837, 1819, 1332, 1154, 1593, 95, 338, 339, 863, 1095, 862, 370, 1308, 1398, 1260, 604, 1550, 944, 1691, 180, 1217, 1103, 73, 1449, 994, 547, 128, 1474, 1187, 1081, 114, 1458, 555, 679, 1531, 1159, 82, 1845, 160, 896, 1763, 575, 103, 1495, 1028, 72, 798, 1053, 958, 83, 417, 1563, 87, 719, 1349, 1674, 534, 1765, 585, 1054, 1843]. finite-difference [1563]. finite-length [180]. finite-set [719]. Finite-time [1524]. finite-volume [339, 862]. fire [1345]. fire-induced [1345]. firm [367]. firms [750]. First [62, 974, 124, 1807, 1560, 526]. first-failure-censored [526]. First-principles [62]. fish [1658]. Fisher [1129]. fisheries [1184]. fishery [1692]. fit [894, 81]. FIT-type [81]. fitted [248, 349, 1845, 1763, 768, 1747]. fitting [1822, 1546, 344]. FitzHugh [452]. five [211, 1719, 1715, 690]. five-axis [211]. five-phase [1719, 1715, 690]. fixed [1668, 186, 799]. Flexibility [1206]. Flicker [191]. flight [1657]. floating [1445, 789, 1777]. floating-point [1445, 1777]. flood [6]. Flow [97, 837, 1617, 1389, 1390, 1819, 485, 219, 405, 615, 1505, 1395, 37, 558, 991, 491, 1848, 1433, 1214, 1591, 1800, 341, 1756, 74, 1402, 1794, 165, 75, 350, 1853, 1212, 568, 1458, 1237, 160, 59, 622, 1626, 1527, 1689, 1288, 1885, 1409, 1315, 594, 83, 46, 562, 717, 361, 1071, 1747, 1354, 1686]. flow-based [1402]. flow-induced [83]. flowfield [212, 406]. flows [338, 339, 863, 1095, 862, 1397, 1399, 1225, 1680, 1449, 313, 1106, 1187, 89, 85, 420, 104, 130, 354, 356, 84, 90]. flu [1186]. fluctuating [588]. fluctuation [1471, 1215]. fluctuation-limited [1471]. Fluctuations [53, 1348]. fluid [1617, 943, 1214, 439, 1437, 1800, 1680, 547, 564, 507, 419, 1518, 420, 1592, 104, 1237, 160, 59, 1626]. fluidic [1225]. fluids [89, 1240]. flux [438, 1724, 1827, 1851, 1302, 701, 1270, 1261]. Fokker [854]. followed [1183]. following [1, 1490]. food [1551, 1061]. force [189, 1432]. forced [1728, 794, 700, 86, 507]. forced-air-cooled [700]. Forchheimer [1592]. forcing [484, 809]. forecast [1823, 743]. Forecasting [781, 232, 775, 266, 1169, 1286, 751, 1299, 1787, 1525, 272]. forecasts [742]. foresight [582]. forest [897, 957]. Foreword [1202, 123, 969, 916, 29, 482, 503, 1222]. form [328, 1444, 163]. formal [993]. formation [56]. formations [354]. forms [70, 1096, 1355, 1516, 1116]. formula [215, 671, 1468]. formulae [345, 1889]. formulation [366, 1668, 772, 1735, 679]. formulations [1217, 1380, 1405]. forward [1590, 207]. fossil [1025]. four [664, 919, 1126]. four-point [1126]. four-valued [919]. Fourier [1098, 1758, 425]. fourth [995, 129, 930, 912]. fourth-order [930]. FPGA [695, 687, 721, 725, 718, 719]. FPGA-based [695, 721, 725]. fractal [1230, 1111, 734]. fractional [1154, 1840, 1755, 1702, 1438, 1168, 1667, 660, 1871, 246, 802, 1015, 1786, 1158, 1044, 770, 1077, 1654, 1791, 1557, 795, 1877, 1314, 843, 1525, 1831, 1434, 635, 1879, 1603, 675, 1809, 1889, 542, 733, 157, 1571]. fractional-order [1791, 1809]. fracture [1389, 1589]. fractured [1402, 1830, 354]. fractures [1400]. fragile [1657, 1666]. fragmentation [1760, 1781]. Frame [1665, 1660]. frames [1550, 550, 451]. framework [1068, 1016, 1312, 1234, 1739, 792, 1797]. Fredholm [852, 215, 1820]. free [258, 1663, 79, 1380, 677, 1432, 1645, 1408]. free-weighting-matrix [1645]. freeway [1653]. Frequency [1644, 1869, 1536, 720, 186, 1291, 1185, 1035]. Frequency-domain [1869]. fresh [1551]. friction [569, 1509, 577, 791, 1495]. front [1663, 1130, 512]. front-end [1663]. frontal [45]. frontier [937, 900, 902]. fronts [497, 498]. Fuel [1737, 722, 1282, 1717, 1299, 1721, 1025, 170, 1285]. Full [1051, 1457, 1853]. full-dimensional [1853]. Fully [567, 534, 1819, 720, 1165, 1735, 104, 1071]. fully-appended [1071]. function [373, 1646, 1323, 601, 1376, 1158, 1342, 1406, 230, 982, 896]. functional [1700, 1132, 1335, 1079, 764, 1127, 1790, 830, 624, 1456, 1860]. functionally [654]. functions [1340, 858, 246, 1141, 436, 1166, 1758, 1342, 1470, 1775, 1472, 452, 1002, 1656, 1115, 1242, 1345, 229, 371]. Fundamental [239, 531, 256, 650]. fundamentalists [971]. Further [617, 112]. fusion [357, 208]. future [822, 317, 1307]. futures [775, 287, 269]. fuzzistics [1787]. Fuzzy [1755, 232, 715, 1323, 687, 174, 1864, 97, 1581, 1825, 732, 6, 1169, 375, 1266, 931, 102, 475, 110, 651, 1250]. fuzzy-logic [715, 1266]. FV [341].

G [678, 112, 678, 425]. G-7 [112, 425]. G/G/ [678]. gain [471]. gains [471]. Galerkin [1487, 336, 1400, 965, 74, 461, 1758, 1138, 420, 1881, 1626, 800, 1889]. Galerkin-based [461]. Gallager [1807]. game [869, 582, 671, 1789, 290]. Game-theoretic [290]. games [1043, 1876, 1416, 99, 261]. gamma [418, 844]. gamma-type [418]. GARCH [776, 279, 774, 472, 377, 283]. Gas [853, 1689, 1409]. gaseous [562]. gases [1223]. gate [1277]. Gaussian [1683, 776, 1017, 300]. gbent [1472]. GCD [1502]. gene [1195, 1348]. general [548, 370, 1646, 477, 1016, 1550, 753, 368, 884, 144, 130, 1844, 1656, 229, 676, 1189, 662]. general-purpose [370]. generalised [583]. Generalization [1523, 490]. generalizations [802]. Generalized [595, 226, 1528, 1690, 643, 156, 1445, 1393, 997, 1142, 1853, 369, 1645, 402, 1563, 1534, 98, 844, 549, 1062]. Generalizing [922]. generated [100, 1252, 1237, 1418, 1533, 1859]. Generating [379]. Generation [207, 404, 1475, 34, 484, 615, 950, 180, 722, 1794, 225, 320, 826, 230, 1477, 721, 1787, 1114]. generator [1284, 176, 797, 1715, 1271, 191, 689, 1738, 700, 175, 1123]. generators [1776, 876, 1704, 24, 451]. generic [1335, 217]. Genetic [1606, 1283, 715, 702, 1725, 1648, 1687]. genetics [911]. geodesic [1313]. Geoelectric [118]. geographical [1561]. Geography [973]. geological [951]. Geometric [209, 1024, 531, 224, 951, 535]. Geometrical [60]. geometries [547, 1747]. Geometry [32, 917, 286, 320, 1236]. geophysical [1187, 284]. GFC [777]. GFC-robust [777]. GHz [431]. Giant [515]. glacier [1407]. Global [1338, 636, 1067, 1167, 1252, 1884, 1790, 541, 676, 1177, 638, 1186, 1317, 766, 1165, 21, 1548, 1070, 1737, 1401, 1675, 662, 1714, 227, 355, 1002, 1450, 959, 1582, 371]. globalization [972, 753]. GMRES [593]. GNSS [1113]. Goal [80]. Goal-oriented [80]. Gompertz [1604, 834]. good [946]. goodness [894]. goodness-of-fit [894]. Gordon [489]. governed [1052]. Governmental [1861]. Gower [1324, 830]. GPS [653]. GPS/IMU [653]. GPU [1675, 990]. GPUs [767, 1364, 1572]. graded [654]. gradient [1309, 1130, 1139, 838, 1145, 139]. gradient-based [1130]. gradient-type [1139]. gradients [703]. grain [52, 55]. Granger [304]. granular [1839, 1315]. grape [397]. Graph [1333, 571]. graph-Laplacians [571]. Graph-theoretic [1333]. graphs [904, 1371]. grass [1345]. Grassmannian [1753]. gravitational [837]. gravity [484]. grazing [1810]. Greeks [1435]. Green [601, 1158]. Grid [1290, 1731, 207, 404, 615, 1313, 1292, 1583, 1458, 712, 723, 717, 1716, 1747]. grid-connected [1583, 712, 723]. grids [1600, 1535, 211, 1720, 1883, 1722, 1071]. grinding [1558]. Gröbner [1205, 922]. Gronwall [1480]. Grossberg [1067, 144, 536]. Grossberg-type [144]. groundwater [757, 1458]. group [1122, 126, 1252, 1753, 35]. groups [1806, 948]. growth [1341, 1, 1490, 389, 527, 418, 1251, 41, 982, 55, 575, 981, 301, 476, 1618, 644]. Guaranteed [1830, 1312, 1486, 111]. Guarantees [1034]. Guest [238, 1007, 1193]. Gumbel [1703].

Haar [1837, 1849, 1867, 1809]. half [1393]. half-logistic [1393]. Halton [1474]. Halton-type [1474]. Hamiltonian [1497, 247, 509, 1233, 1011, 768]. Hammersley [1468]. Hamming [671]. hand [901]. handling [1252, 141]. Hankel [1502]. haptic [1046]. hardening [910, 88]. hardware [695]. hardware/software [695]. harmonic [1711, 675]. harmonics [185, 723]. Harr [1651]. harvest [1692]. harvesting [1456]. Hassell [764]. Hastings [1367]. having [512]. hazard [1378, 526]. head [671]. head-or-tail [671]. health [1298, 990]. healthy [100, 1418]. heap [991]. heart [441, 209]. heat [1728, 945, 1854, 567, 882, 607, 1885, 432, 303, 1261]. heating [1869]. heavy [1823, 899, 1703, 1084, 1128]. heavy-tailed [1084]. hedging [774, 269]. helix [509]. Helmholtz [531, 1080, 601, 1076]. hepatitis [814, 809]. herbivore [1766]. herd [1446]. Hermite [1758, 1355]. Hessian [1432]. Hessian-free [1432]. Heston [1042]. heterogeneity [1045]. Heterogeneous [970, 837, 1311, 983, 341, 126, 582, 1111, 361, 1650, 1655]. Heuristic [1868, 1736]. hexapod [469]. Hidden [493, 1404, 1316, 1114, 1241, 870, 992]. Hiding [1486]. High [558, 965, 1560, 1499, 1427, 1424, 1469, 232, 1536, 720, 595, 176, 1438, 1526, 1046, 186, 498, 932, 1274, 1302, 1252, 471, 10, 1237, 1626, 1185, 1609, 1027, 1746, 649]. high-frequency [186]. high-gain [471]. high-impedance [1046]. High-order [965, 1560, 1427, 595, 1626, 649]. high-speed [176, 1274]. higher [1593, 73, 85, 454, 1240, 518]. highly [310, 1590]. Hilbert [1665, 1660]. Hilliard [1124, 1216]. hip [45]. HIV [1501, 456, 1251, 15, 1831, 1189, 1177, 542]. HIV-1 [1251]. HIV/AIDS [1831]. HLS [1742]. HLS-based [1742]. hodograph [550]. hog [288, 416, 1651]. Hölder [1780]. holdup [97]. hole [343, 856]. holes [1104, 1824]. Holling [1700, 1456, 1324]. Holling-type [1324]. Holm [1249]. homoclinic [799]. homogeneous [1397, 1655]. Homogenization [1339]. homogenized [557]. homosexual [15]. Hong [274]. Hopf [1602, 1624, 1251, 1167, 155, 1578, 1810, 1417, 1582, 449]. Hopfield [649, 644]. Hopfield-type [649]. horizontal [97, 1673]. host [1417]. hours [1286]. Howard [1552]. hp [1217, 665, 376]. hp}-adaptive [376]. hp}-DGFEM [665]. HTS [701]. human [455, 439, 829, 1336, 1754, 1496]. humoral [676]. humps [1160]. hunting [1846]. HVdc [1535]. Hybrid [1581, 1872, 1687, 689, 846, 1837, 1284, 1811, 248, 1146, 1725, 965, 726, 1290, 1282, 434, 1714, 1381, 322, 714, 1717, 1738, 697, 622, 1269, 1266, 1732, 139, 8, 678, 724, 1716, 1537, 646, 644, 174, 713]. hybridizable [1626]. hybridization [1285]. hydraulic [991, 1811, 348, 1123, 1537]. hydro [716, 71]. hydro-power [716]. hydro-thermo-mechanical [71]. hydrodynamic [864, 348]. hydrodynamics [588]. hydrogen [1720]. hydrology [282]. hydropower [633]. Hyperbolic [1228, 124, 234, 563, 155]. hyperboloids [1764]. hyperchaotic [1037, 1526, 140, 931, 870]. hypercircles [401]. hyperelastic [1518, 520]. hyperinterpolation [1451]. hyperplasticity [1234]. hypersingular [407]. hypersonic [212]. Hyperthermia [1836]. hypothesis [1348]. hysteresis [1792, 1511, 832].

I-Delaporte [1337]. ICMOSPS [652]. ICT [1293]. identical [973, 1152]. Identification [419, 38, 97, 1737, 101, 549, 688, 1867, 691, 470, 1644]. identify [282]. Identifying [1886, 276, 1603, 1283]. IDS [65]. IGBT [172]. II [1606, 1330, 1175, 1393, 626, 1100, 1799, 421, 1555, 515, 1580, 1324]. III [936]. Ill [521, 316]. ill-posed [316]. Ill-posedness [521]. illiquid [572]. ILUE [1508]. IMACS [1220, 17, 49, 67, 92, 107, 121, 134, 150, 167, 194, 221, 236, 16, 48, 66, 91, 106, 120, 133, 149, 166, 193, 220, 235, 250, 263, 292, 306, 325, 363, 382, 394, 413, 427, 445, 463, 249, 262, 291, 305, 324, 362, 381, 393, 412, 426, 444, 462, 479, 500, 523, 538, 552, 579, 597, 610, 619, 629, 640, 656, 478, 499, 522, 537, 551, 578, 596, 609, 618, 628, 639, 655, 667, 682, 706, 729, 738, 761, 784, 666, 681, 705, 728, 731, 760]. IMACS [783, 805, 819, 840, 848, 866, 878, 889, 906, 914, 925, 939, 954, 804, 818, 839, 847, 865, 877, 888, 905, 913, 924, 938, 953, 967, 987, 1005, 1021, 1031, 1039, 1048, 1056, 1064, 1073, 1086, 1119, 966, 986, 1004, 1020, 1030, 1038, 1047, 1055, 1063, 1072, 1085, 1118, 1134, 1148, 1162, 1171, 1180, 1191, 1209, 1219, 1246, 1254, 1264, 1133, 1147, 1161, 1170, 1179, 1190, 1208, 1218, 1245, 1253, 1263, 1304, 1326, 1351, 1361, 1373, 1384, 1412, 1421, 1429, 1453, 1303, 1325, 1350, 1360, 1372]. IMACS [1383, 1411, 1420, 1428, 1452, 1464, 1489, 1520, 1542, 1575, 1586, 1596, 1611, 1630, 1639, 1488, 1519, 1541, 1565, 1574, 1585, 1595, 1610, 1629, 1638, 1694, 1707, 1744, 1768, 1783, 1802, 1816, 1856, 1892, 1693, 1706, 1743, 1767, 1782, 1801, 1815, 1855, 1891, 1278]. Image [857, 904, 941, 1091, 1502, 1098, 1548, 604, 1107, 1141, 1407, 836, 139, 357, 1027, 386]. image-based [604]. imagery [590]. images [543, 1581]. imatinib [1151]. imitation [1685]. immature [1184]. immigration [1766]. immiscible [837, 341, 350]. immune [1501, 142, 622, 996]. immunity [1528, 676]. immunization [1528]. immunotherapy [996]. Impact [1583, 758, 288, 759, 323, 584, 809, 794, 1112, 694, 698]. impacting [79]. impacts [282, 755]. impairing [1553]. impedance [1046, 1290, 231]. imperfect [1792, 1498, 816]. imperfection [972]. Implementation [1772, 1092, 1593, 1733, 687, 494, 310, 860, 757, 71, 206, 721, 1082, 61, 1742, 1877, 724, 397, 1637, 718, 719]. implemented [702]. implementing [399, 962, 1259]. Implicit [862, 1819, 1545, 802, 1877, 1145]. importance [935, 119, 198]. important [238]. improve [1108, 1722]. Improved [1298, 1131, 1294, 637, 1658, 1606]. improvement [1741, 1035]. Improving [1688, 742, 1704, 141]. impulse [627]. impulses [1317]. Impulsive [1662, 456, 764, 144, 536, 1664, 375, 736, 1127, 1345, 423, 299, 1310, 1679, 649, 1659, 530, 961, 1618]. IMSP [872, 1014]. IMU [653]. in-depth [1108]. in-plane [634]. incidence [1157, 676]. including [1013, 1720, 795]. inclusion [1761, 9]. inclusions [390]. inclusive [749, 1003]. Income [753]. Incomplete [1508]. incompressibility [1397, 47]. incompressible [219, 405, 1433, 1680, 89, 85, 59, 1626, 1516]. incorporating [636]. increase [1146, 1887]. incremental [214]. incrementing [789]. indefinite [1664]. independence [529]. independent [1806, 1759, 998]. index [1169, 259, 117, 274, 1143]. India [1358, 754]. indicators [1287, 277]. indices [1806, 199, 662]. Indirect [770, 358, 1276]. individual [957]. individual-based [957]. Induced [1570, 602, 1701, 42, 147, 83, 1345]. Induction [698, 1060, 692, 1277, 1265, 88, 191, 845, 175, 1276, 1699, 691, 1295, 1301, 694, 173]. inductorless [797]. industrial [437, 721]. industries [416]. industry [1146]. inequalities [1051, 1674]. inequality [753, 887]. inexact [1415]. infected [1213]. infection [814, 1501, 1251, 142, 809, 676, 1189, 1417, 542]. infections [587]. infectious [1808, 1701, 1459]. Inference [156, 937, 900, 1532, 1318, 1142, 97, 1799, 114, 1599, 283, 902, 770]. infinite [1758, 513]. infinitesimal [228]. infinity [904, 1445, 999]. inflation [983]. Influence [181, 527, 508, 1717, 56, 1270, 172, 1663, 58, 1714, 1001, 759]. influencing [277]. influenza [1341]. information [1792, 1129, 1701, 1561]. information-dependent [1792]. information-induced [1701]. informative [438]. inhomogeneous [422]. initial [1445, 1370, 1539, 521, 1865]. inlet [90]. inner [531]. innovation [1309, 753]. innovative [442, 985]. Input [1664, 466, 1062, 138, 1523, 792, 1123, 1618]. input-delayed [138]. Input-to-state [1664]. input/output [1123]. inrush [1793]. ins [81]. insect [1332]. insensitive [643]. INSGA [1606]. INSGA-II [1606]. inspired [829, 811]. Instability [985, 1564, 487, 497, 1000, 505, 1344]. instantaneous [724]. insulated [187]. insulation [1294]. integer [1607, 795]. integers [1474]. integrable [1233]. Integral [1771, 852, 1090, 1820, 485, 1394, 1017, 1691, 215, 244, 1019, 422, 360, 1809, 1850]. integrals [1197, 407, 1705, 403]. Integrated [1533, 469, 457, 1864, 1122, 613, 1282, 1271, 276, 717, 998, 141]. integrating [1062]. Integration [1199, 1470, 1391, 1024, 1535, 1292, 389, 247, 772, 1673, 660, 1758, 471, 25]. integrator [801, 1145]. integrators [1593, 1625]. integrity [653]. integro [1018, 1026, 1667, 1849, 1643, 831, 1616]. integro-differential [1018, 1667, 1849, 1643, 1616]. Integrodifferential [76]. Intelligent [440, 710, 1581]. inter [1293]. inter-cell [1293]. interacting [1446, 544, 374]. interaction [200, 547, 1518, 933, 717, 996]. interactions [1766, 1345, 1746]. Interactive [646, 1203]. interchannel [495]. interest [824, 769, 770, 300, 1863]. Interface [1826, 1399, 64, 62, 56]. Interfaces [1117, 1627]. interfacial [268]. interfacing [1592]. interference [1310]. interior [488, 932, 1763, 697, 1314]. interleaved [722]. interlocking [398, 321]. intermittent [1729, 475]. Internal [1036, 1238, 823, 1302]. International [1386, 1818, 1220]. interplay [1111]. interpolant [332, 1822]. interpolants [328, 852, 1391, 1026, 347, 213, 1115, 1116]. Interpolated [1888]. interpolating [898, 1556]. interpolation [941, 543, 1821, 1096, 949, 1827, 1828, 893, 1355, 360, 734]. interpolatory [353, 357]. interruptions [1655]. intersection [1648]. intersections [1876]. interval [1359, 614, 532, 1143]. intervals [822]. intracranial [373]. intraguild [1579, 1842]. Intrinsic [1232]. introducing [1456]. Introduction [1770]. introductory [52]. intrusion [1089, 1663, 1633, 355]. intrusion-free [1663]. invariance [1291]. Invariant [834, 1009, 302, 241, 591, 528, 12]. invariants [1098]. inventory [1655]. Inverse [118, 303, 531, 493, 218, 411, 549, 885, 434, 508, 1482, 1176, 157]. inverses [571]. inverted [1142]. inverter [1723]. investigate [172]. Investigating [277]. investigation [165, 256, 1841, 449]. investigations [1546, 1859, 201]. investment [769, 978]. involving [1664, 104, 1690]. ion [183, 1282]. iPIC3D [63]. iron [1302]. irradiation [1609]. irreducible [1772]. island [1720, 712]. islanded [1711]. islanding [1290]. isoelastic [977]. Isogeometric [1505, 1514, 1082, 1380, 547]. isolation [1699]. isothermal [1702]. isotonic [1140]. isotropic [95]. Issue [1386, 1818, 1770, 652, 1202, 969, 916, 29]. issues [1203]. issuing [750]. Italian [984]. items [1635]. iterated [532]. Iteration [852, 405, 1752, 1888]. iterative [1755, 1848, 297, 1507, 930, 1494, 1113]. IV [1700]. IX [599].

jamming [1888]. January [4, 464, 657, 955, 1120]. Japan [1003, 749, 289]. Japanese [750]. joining [288]. joint [40, 1635, 45]. journal [759, 372, 759]. judgmental [742]. July [364, 580, 879, 1040, 1576]. jump [1774, 732, 1435, 105, 1645]. jump-diffusion [1435]. jumps [1667]. junctions [1225]. June [108, 730, 867, 1032, 1181, 1374, 1567, 1750].

Jupiter [670].

Kac [1681]. Kaczmarz [139]. Kaldor [1200]. Kalman [314, 1196, 1463, 208]. Kantorovich [654]. KdV [507, 1241, 98]. keeping [141]. Keoladeo [1358]. kernel [1820]. kernelized [1581]. Kharitonov [410]. Khyber [1847]. kind [852, 1394, 215]. kinetic [65, 61, 1471, 54]. kinetics [216, 1604, 527, 1034, 1348]. kink [1239]. Kinks [1223]. Kirchhoff [1380]. knot [1854]. knots [1828]. Knowledge [993, 962]. Koch [950]. Kong [274]. KP [506]. Kriging [335]. Kriging-based [335]. Kruskal [493]. Krylov [1415]. Kulkarni [332]. Kung [153]. Kutta [929, 145, 768]. Kutta-Nyström [768].

laboratory [443, 442]. lag [808, 475].

Lagrange [949, 1549, 1067]. Lambert [230]. lamp [190]. Lanczos [492]. landmark [941]. landmark-based [941]. lane [141]. Langmuir [602]. language [1259, 1723]. Laplace [586, 1261]. Laplacian [904, 519]. Laplacians [571]. Large [1812, 942, 242, 1139, 1794, 1523, 474]. large-scale [1139, 1794]. laser [295, 910, 1869]. LASSO [1367]. last [1652]. latent [638]. lattice [1680, 388, 623, 1839, 59]. lattices [970, 1244]. law [1094, 772, 1449, 1495]. laws [948]. Lawson [247]. layer [863, 212, 1138, 507, 1763]. layers [1473, 510]. LBB [1516]. LDG [1688]. leaching [991]. lead [750]. leading [614, 277]. leaf [296]. leak [594]. leakage [1646]. Learning [436, 904, 1581, 1343, 1771, 443]. least [994, 1887, 1644, 466, 352, 376]. least-squares [994]. Lee [153, 386]. Leffler [1683]. leg [831]. Lehmus [922]. Leibniz [1204]. lemma [1480]. length [180, 203, 1001]. Leslie [1700, 830, 1324]. less [14]. leukemia [1151]. level [1, 1490, 1776, 690, 1553, 1216, 1322, 1054]. levels [1528, 971]. leverage [1144]. Leverett [1388]. Lévy [1487, 873]. Lewis [498]. library [1512, 718]. lid [313, 1462]. lid-driven [313, 1462]. Lidstone [1447]. Lidstone-Spline [1447]. Lie [948, 1689]. Life [846, 751, 833, 2]. life-span [833, 2]. lifespan [1294]. lifetime [317, 259, 369, 1143]. lifting [451, 148]. Lighthill [1223]. like [874, 1207, 507, 182]. likelihood [871, 1309, 844]. limestones [857]. Limit [1137, 790, 448, 385, 1232, 746]. limit-order [746]. limitations [181, 1719]. limited [1471]. Lindley [260, 230, 421, 1580]. line [949]. Linear [352, 1469, 1837, 1330, 788, 1755, 154, 1736, 254, 448, 159, 964, 1691, 76, 458, 234, 1275, 929, 876, 567, 390, 1762, 99, 1081, 1129, 9, 1458, 528, 1711, 1029, 1493, 1560, 1194, 1010, 28, 1345, 1083, 1865]. linear-hyperbolic [234]. linearization [1199, 1123]. linearized [1235]. linearly [647, 1243]. lines [1196, 1281, 1845, 361]. link [1731]. linkage [1647]. linkages [1206]. linked [1201]. Linking [54]. links [662]. Lipschitz [371]. liquid [79, 97, 60, 380, 1236, 268]. liquid-holdup [97]. liquid-phase [60, 380]. lithium [1720, 183, 1282]. lithium-ion [183]. liver [587]. living [42, 1339]. LMI [1677, 632, 229, 649]. LMI-based [632, 649]. Load [545, 1536, 1341, 176, 1583, 1711, 190]. Load-sharing [545]. loading [42, 1658]. lobe [373]. Local [941, 1666, 1582, 1548, 1402, 1110, 1508, 473, 213, 606, 148, 178]. locality [1483]. Localized [1516, 1232]. Locating [242]. location [495, 419, 1035, 12]. lock [672]. locomotion [439]. logic [34, 715, 174, 1266, 398]. logical [217]. logical-temporal [217]. logics [919, 919]. logistic [1341, 1393, 1251, 736, 982]. Long [312, 284, 544, 994, 745, 1564, 1249]. Long-run [312]. long-term [1564]. Long-time [544, 1249]. longest [606]. longest-edge [606]. longitudinal [1684]. loop [727, 376, 141]. looped [844]. Lorenz [1037, 874, 140]. Lorenz-like [874]. Lorenz-type [1037]. loss [846, 1554]. losses [1272, 1302, 1270]. Love [1380]. Low [1321, 1617, 1, 1490, 873, 1711, 1365, 1721, 1493, 1863, 90, 1873, 422, 1294, 1053, 1295, 1746]. low-cost [1295]. low-dimensional [1493, 1863]. Low-discrepancy [1321, 1365]. low-level [1, 1490]. low-order [1746]. low-voltage [1711]. lower [1002, 1481, 1484, 934]. LTI [1436]. lubrication [1099, 1100, 899, 372]. Lumped [700]. Lumped-parameter-based [700]. lung [1581]. Lyapunov [1416, 1132, 1377, 385, 1664, 452, 624, 735].

M [1746]. M-PCM-OFFD [1746]. Macaulay [30]. Mach [90]. Machine [433, 904, 698, 1277, 1265, 1887, 185, 179, 1724, 1771, 1302, 1714, 701, 688, 697, 1729, 1276, 1294, 173, 1718]. machines [1272, 685, 1719, 1274, 1270, 184, 694, 696]. machining [211]. Macro [570, 1733, 344, 104]. Macro-elementwise [570]. macro-modelling [1733]. macroeconomic [1168, 278]. MacWilliams [671]. made [359]. magic [400]. magnet [1272, 699, 685, 1719, 1297, 176, 188, 1715, 932, 1274, 1628, 1271, 883, 1302, 700, 697, 1270, 696]. magnetic [19, 864, 698, 176, 1673, 1511, 189, 511, 1570, 173]. magnetizing [845]. magneto [77]. magneto-mechanical [77]. magnetostatics [555]. maintained [47]. maintenance [224]. majority [1001]. majority-rule [1001]. Making [1538, 261]. mal [1536]. mal-operation [1536]. malaria [450, 1615]. Malliavin [1435]. MAMERN [1386, 1818]. man [359]. man-made [359]. management [1736, 171, 775, 773, 1551, 1725, 757, 1282, 777, 1281, 1692, 1732, 1537]. Manakov [1160]. Manifold [703, 1753]. manifolds [1009]. manipulation [746]. manufacturer [584]. many [257]. map [373, 158, 1078, 797, 970]. Maple [31, 398, 919]. mapped [1758]. mapping [942, 703]. maps [796, 495, 234, 437, 163, 532]. March [673, 820, 988, 1149, 1695]. marginal [869]. margins [1676]. marker [1403]. marker-and-cell [1403]. market [972, 778, 1078, 780, 1720, 269, 323, 750, 1787, 782, 971, 746, 1084, 1789, 971]. markets [572, 780, 970, 99, 984, 1025, 985, 374, 1084, 755, 425]. Markov [1600, 1876, 1130, 204, 1368, 1106, 1404, 732, 1652, 1645, 1483, 1114]. Markovian [1655, 957, 1642]. marks [1561]. MARS [1188]. Marshall [1534]. Martin [1790]. martingale [473]. Maruyama [254, 1780]. Maruyama- [254]. masks [210]. Mass [1099, 1100, 346, 1614, 388, 513, 1765]. mass-conservative [1765]. Mass-conserving [1099, 1100]. matched [1605, 1791]. matching [750]. MATCOM [1386, 1818]. material [65, 1495, 1599]. materials [604, 52, 1238]. Math [1490, 153, 937, 1003, 608, 240]. Mathematical [1, 1490, 1099, 456, 1862, 1805, 1558, 816, 594, 1865, 1176, 172, 11, 527, 1178, 372, 1183, 812, 431, 1200, 947, 359, 1698, 1842]. Mathematics [652, 969]. matrices [748, 242, 1377, 1757, 671, 1762, 1559, 735]. matrix [993, 239, 1415, 81, 246, 1644, 1129, 617, 1012, 1645, 1276, 1850]. mature [1184]. Maxillomandibular [320]. maximisation [710]. maximization [1141, 633]. maximize [1724]. Maximizing [1710]. Maximum [1871, 1309, 73, 711, 87, 844, 1608]. Maximum-norm [1871]. Maxwell [366, 78]. May [93, 307, 539, 1022, 1172, 1748]. MC [202]. MC/QMC [202]. MCDM [1864]. MCM [1770]. McMC [909]. Mean [536, 112, 254, 929, 1779, 1473, 622, 1113, 1866]. mean-reverting [1866]. mean-square [254, 929]. meaningful [164]. means [1196, 1205, 1581, 1169, 957, 535, 1239, 284, 1562, 102]. measure [901, 997, 935]. measurement [1312]. measurements [1341]. measures [662, 933]. Measuring [203, 923, 747]. mechanical [181, 392, 71, 1495, 77, 691]. mechanics [439]. mechanism [693, 756]. mechanistic [1609]. mechatronic [469]. mechatronics [442]. media [837, 1390, 1819, 857, 485, 95, 1433, 341, 346, 1762, 350, 568, 160, 59, 1516, 557, 1117, 1315, 958, 361]. medical [1069]. Medicine [119]. medium [1392, 1102, 714, 1717, 1830, 835]. medium-duty [1717]. Meeting [404]. Mellin [1098]. melt [1761]. membrane [1299, 170]. memory [21, 284, 745, 1582, 960]. MEMS [646]. men [15]. Mersenne [876, 1777]. Merton [153, 3]. Mesh [826, 1424, 950, 1102, 1824, 873, 320, 409, 1633, 947, 1875]. meshed [1710]. meshes [329, 942, 1688, 965, 606, 148]. Meshfree [588]. meshing [1545, 1426]. meshless [1155, 1395, 885, 420]. mesoscopic [22, 515]. meta [1146]. meta-model [1146]. Metabolic [1334]. metabolism [815]. metabolite [438]. Metaecoepidemic [1213]. metaheuristic [1450]. metamodeling [1188]. metamodelling [214]. metapopulation [872, 1014]. methane [1410]. method [1469, 654, 1837, 1589, 290, 332, 1090, 1820, 1092, 485, 1445, 1755, 366, 615, 1018, 336, 124, 1502, 1394, 1095, 1395, 1196, 531, 1752, 686, 1165, 1821, 1026, 858, 1204, 1260, 997, 823, 1309, 911, 1130, 1416, 1125, 1888, 1051, 298, 408, 1591, 1102, 1122, 218, 341, 965, 1449, 1139, 192, 343, 856, 860, 1854, 994, 802, 1290, 1843, 1257, 910, 1786, 547, 1827, 1138, 885, 233, 546, 217, 883, 1081, 1252, 388, 420, 1664, 1432, 256, 1458, 138, 1687, 555, 679, 1654, 1131, 1759, 1845, 1869, 1299, 1523, 118, 1838, 1763, 1881, 588, 513, 575, 61, 103, 1557]. method [1510, 559, 624, 8, 1527, 678, 231, 1844, 1539, 1471, 360, 781, 1053, 432, 958, 574, 96, 417, 1778, 1852, 586, 148, 675, 637, 1809, 1746, 361, 1123, 1747, 534, 733, 1145, 1765, 1681, 157, 1571, 1054, 1826]. Methodologies [713, 777, 599]. Methodology [1292, 174, 192, 1083, 476, 1716]. Methods [1386, 1818, 1836, 1424, 1617, 852, 1175, 19, 864, 570, 316, 338, 803, 862, 1736, 1546, 786, 254, 1017, 492, 1371, 995, 1848, 310, 248, 944, 1007, 1551, 1415, 205, 204, 1400, 1321, 1675, 1015, 1106, 127, 202, 266, 1758, 23, 873, 1614, 349, 1796, 85, 1513, 1364, 1485, 297, 1216, 1507, 930, 1019, 1156, 1590, 771, 614, 912, 1626, 1012, 1601, 1482, 1294, 800, 72, 1126, 599, 635, 831, 1349, 424, 145, 1889, 1860, 768, 361, 1788, 162, 1668, 1062]. Metric [1426, 290, 391, 409]. Metric-based [409]. metro [1281]. Metropolis [20, 1367]. MFE [341]. MFE-FV [341]. MFS [1443]. mice [1, 1490]. micelles [1225]. micro [1225, 269, 104, 275]. micro-fluidic [1225]. micro-market [269]. micro-simulation [275]. microalgae [44]. microarches [1036]. microbial [101]. microfluidic [573]. Microforecasting [1551]. microgrid [1711, 1288]. microgrids [1290, 723]. micromagnetism [78]. microorganism [575]. micropolar [1214, 1437, 1800]. Microstructure [43, 65, 270, 56, 57, 1084]. microstructured [1238, 516]. microwave [1690]. migration [1213]. migratory [1324]. milling [637]. Milstein [254]. Milstein-type [254]. mimetic [1424]. MIMO [1165, 632, 1436]. Mindlin [516]. Mindlin-type [516]. Minimal [1169, 1083, 923]. Minimalistic [829, 1345]. Minimally [871]. minimization [1584, 185, 828]. minimize [622]. minimum [1130, 1731, 1890]. mining [1260, 1653]. miscible [1399, 958]. Mises [1109]. mismatched [1791]. mission [1553]. mistakes [33]. mistuned [86]. mitigate [171]. mitigation [191]. Mittag [1683]. Mittag-Leffler-Gaussian [1683]. Mixed [958, 1662, 1416, 1549, 1217, 772, 1103, 1449, 860, 318, 994, 1405, 1067, 563, 1237, 1506, 14, 1176, 1765, 585, 161, 378]. mixed-mode [1506]. mixture [776, 1853, 274, 744]. mKdV [98]. mobile [829, 208]. mode [172, 1323, 137, 1691, 722, 932, 314, 1378, 1302, 688, 632, 1605, 725, 845, 1506, 240, 1009, 1123]. Model [1719, 271, 1649, 1692, 1493, 1863, 170, 1389, 290, 1329, 1330, 455, 232, 333, 1331, 1332, 11, 1341, 815, 1874, 1685, 1676, 1283, 1287, 1338, 116, 625, 626, 1165, 1184, 692, 245, 875, 1792, 1308, 565, 1099, 1100, 1604, 1579, 871, 491, 279, 758, 823, 1399, 39, 38, 295, 983, 973, 572, 1298, 153, 627, 659, 1168, 1808, 1146, 1591, 1437, 1800, 605, 1401, 716, 1532, 829, 742, 456, 318, 1275, 817, 1000, 1648, 418, 1013, 466, 1624, 399, 962, 827, 1251, 377, 1335, 811, 348, 142, 1079, 1167, 6, 434, 957, 1853, 1378, 419, 450, 1152]. model [3, 1302, 182, 1252, 372, 509, 1042, 1045, 1615, 375, 736, 1183, 1498, 60, 380, 1127, 41, 1790, 529, 1798, 183, 982, 1359, 771, 749, 1003, 1463, 1200, 830, 976, 44, 190, 1029, 1151, 1322, 1314, 275, 1240, 855, 843, 1001, 544, 1157, 1456, 359, 816, 1561, 1517, 56, 661, 931, 1609, 1637, 1831, 184, 545, 1061, 1528, 1301, 747, 1345, 832, 1866, 1346, 1410, 971, 173, 1347, 1201, 47, 719, 1479, 526, 676, 998, 1417, 1842, 274, 744, 1084, 321, 386, 633, 638, 1186, 542, 1320, 1616, 1678, 1608, 1459, 1496, 825, 1379, 309, 1324, 530, 1538, 1650, 737]. Model-based [1692, 170, 742, 375]. model-order [1165]. Modeling [1386, 1818, 1272, 1501, 1887, 1723, 185, 727, 273, 453, 175, 1358, 8, 832, 1847, 274, 1128, 573, 1, 1490, 37, 558, 1846, 1146, 75, 1735, 320, 1343, 52, 1271, 1511, 1805, 119, 1069, 809, 353, 55, 209, 704, 1558, 1294, 1315, 594, 1797, 1083, 689, 1525, 599]. Modelling [748, 741, 288, 287, 693, 268, 1300, 296, 575, 255, 1153, 423, 280, 1285, 996, 1320, 361, 837, 1702, 1733, 1078, 1728, 773, 1312, 1862, 40, 89, 282, 1714, 43, 82, 42, 65, 1269, 1721, 717]. models [1387, 776, 1340, 1700, 1154, 808, 548, 339, 810, 1823, 1446, 1052, 477, 1260, 997, 159, 872, 1014, 472, 1583, 1680, 73, 433, 767, 1727, 1404, 909, 278, 563, 155, 1460, 1178, 1518, 886, 367, 812, 13, 607, 104, 1435, 1506, 435, 981, 226, 1495, 1114, 1436, 387, 1010, 1499, 7, 696, 301, 1144, 1213, 583, 1189, 1177, 1529, 283, 961, 1250, 1128]. moderately [654, 1812]. modes [1232, 712]. Modified [159, 143, 253, 1187, 680, 332, 531, 1604, 186, 1356, 994, 507, 1318, 1799, 830, 574, 378]. Modular [1794, 1293, 1289]. modularized [1499]. modulated [1296, 1872]. Modulation [1276, 690, 694]. module [1287]. moisture [1045]. Molecular [1336, 1608, 53]. mollification [298]. MoMaS [1401]. moment [12]. moment-type [12]. Moments [1652, 1098, 964, 1838, 467]. momentum [461, 509]. monetary [983]. monitored [459]. monitoring [159, 1298, 147]. monocyclic [1329, 1330]. monolithic [1518]. monopoly [975]. monotone [1425]. Monte [1600, 267, 19, 1059, 200, 758, 1371, 911, 199, 205, 204, 1368, 1356, 1757, 1106, 23, 312, 197, 119, 1476, 206, 1773, 1781, 1473, 60, 380, 203, 61, 25, 26, 1076, 1539, 1471, 1705, 28, 1778, 1681]. Mor [771]. morphodynamic [339, 862]. morphology [338]. Morrison [215]. Morro [1228]. mortality [273, 1345]. mortar [77]. mortgage [984]. MOSFET [1733, 1108]. MOSFETs [892]. mosquito [1862]. mosquito-borne [1862]. moth [397]. Motion [550, 942, 1366, 76, 1230, 509, 268, 209, 208]. motor [1060, 686, 1297, 692, 188, 1628, 883, 189, 845, 1699, 691, 1295, 1726, 1594]. motors [699, 192, 932, 1301]. motorway [918]. mounted [699]. movement [296]. moving [1309, 1663, 1366, 64, 1102, 1872, 52, 388, 1884, 131]. MRAS [1060]. MSOGIs [721]. Multi [1734, 171, 1106, 884, 63, 1266, 1315, 1651, 634, 1589, 1600, 1272, 1755, 863, 1309, 1224, 874, 205, 1667, 204, 1735, 377, 970, 1628, 690, 1614, 727, 297, 206, 648, 717, 633, 1789]. Multi-almost [884]. multi-component [1614, 297]. multi-core [206]. multi-cracked [1589]. Multi-criteria [1266]. multi-dimension [1224]. multi-dimensional [205, 204]. Multi-drive [171]. multi-innovation [1309]. multi-layer [863]. multi-level [690]. multi-loop [727]. multi-market [1789]. multi-markets [970]. multi-objective [1755, 1628, 633]. Multi-physics [1106]. multi-populations [648]. Multi-port [1734]. multi-precision [634]. multi-resolution [377]. Multi-scale [63, 1315, 1600]. multi-source [717]. multi-stage [1600]. multi-stranded [1272]. multi-valued [1667]. Multi-vehicle [1651]. multi-wing [874]. multibody [474]. multicables [882]. multicell [1293]. multichannel [440]. multicomponent [1390, 1378, 1409]. multidimensional [1796, 226]. multidirectional [960]. multidomain [1601]. Multifractal [1215, 1111]. multigrid [78, 1080, 57]. multigrid-based [1080]. multilayer [1572]. multilayered [1690]. Multilevel [571, 1794, 1476, 1289, 992]. multiobjective [1130]. multioperators [1427, 1564]. multioperators-based [1427, 1564]. multipatch [1380]. multipath [311]. multiphase [1848, 1433, 97, 160]. Multiple [1636, 1599, 692, 995, 1376, 466, 1861, 1381, 138, 112, 975, 912, 710, 1189, 1797, 1582]. multiple-coupled [692]. Multiple-criteria [1636]. multiple-input [466]. multiple-zero [1376]. multiplicative [351]. multipole [1527]. multirate [466]. Multiresolution [342, 1423, 1091, 1573]. Multiscale [557, 1346, 860, 909, 1195, 1557, 1826]. Multistability [992]. multistage [101]. multistationarity [1333]. Multivariate [154, 340, 774, 1002, 920]. multiwavelets [342, 1027]. Murray [544]. muscle [39]. musical [34]. mutation [702, 1805]. mutual [1310].

Nagumo [452]. Nano [431]. Nano-GHz [431]. nanofluid [1570]. nanoparticle [1607]. nanoparticles [1686]. nanostructures [58]. nanowires [200]. NARMAX [832]. Nash [290]. National [1358]. Natural [1386, 1818, 1259, 358, 376, 1409]. nature [974]. Navier [1054]. navigation [653, 829]. near [799, 132]. near-singular [132]. nearly [247, 1262]. Nearwell [1110]. negative [993, 187]. neighbours [778]. NEM [677]. nested [1322]. net [435, 27, 717]. nets [1723]. network [457, 1710, 438, 187, 1152, 1714, 1286, 1069, 714, 104, 1299, 203, 1661, 1269, 1314, 1408, 923, 985, 1698, 718, 887, 1537, 651, 643, 960]. network-computer [457]. networks [232, 1662, 625, 626, 1646, 884, 1195, 1067, 144, 536, 591, 1666, 15, 1645, 678, 1001, 1656, 1010, 647, 650, 229, 648, 1679, 649, 1582, 735, 475, 371, 644, 436]. Neumann [219, 319, 1227, 1076]. Neural [436, 648, 232, 457, 1662, 1646, 884, 1067, 1152, 1286, 144, 536, 1069, 591, 1661, 1645, 1314, 1656, 647, 718, 650, 229, 1679, 649, 1582, 887, 735, 475, 651, 643, 371, 960, 644]. neuro [687, 97]. neuro-fuzzy [687, 97]. neurodynamic [1166]. neuron [1661]. neurons [1152, 452]. neutral [1132, 1377, 787, 732, 1067, 14, 624, 831]. neutral-type [1067]. neutron [127, 128, 132]. News [16, 48, 66, 91, 106, 120, 133, 149, 166, 193, 220, 235, 249, 262, 291, 305, 324, 362, 381, 393, 412, 426, 444, 462, 478, 499, 522, 537, 551, 578, 596, 609, 618, 628, 639, 655, 666, 681, 705, 728, 731, 760, 783, 804, 818, 839, 847, 865, 877, 888, 905, 913, 924, 938, 953, 966, 986, 1004, 1020, 1030, 1038, 1047, 1055, 1063, 1072, 1085, 1118, 1133, 1147, 1161, 1170, 1179, 1190, 1208, 1218, 1245, 1253, 1263, 1303, 1325, 1350, 1360, 1372, 1383, 1411, 1420, 1428, 1452]. News [1464, 1488, 1519, 1541, 1565, 1574, 1585, 1595, 1610, 1629, 1638, 1693, 1706, 1743, 1767, 1782, 1801, 1815, 1855, 1891]. Newton [1432, 574]. Newtonian [1449, 89, 84, 1570, 46]. Niasar [1752]. Nicolson [1531]. Nikkei [269]. Nikkei-225 [269]. nilpotent [1497, 385, 1257]. nine [686]. nine-phase [686]. no [1602, 870]. Nodal [128]. node [420]. nodes [949]. nodule [1581]. Noise [644, 653, 1487, 698, 546, 269, 114, 270, 1665]. noises [351]. Non [1077, 46, 1606, 1469, 1837, 1330, 857, 485, 1874, 335, 1018, 1702, 1397, 1547, 993, 964, 1691, 965, 1425, 671, 347, 567, 1166, 1762, 577, 506, 89, 1657, 241, 1477, 591, 1666, 1237, 1711, 1555, 84, 920, 1483, 1570, 1560, 387, 798, 28, 1345, 793, 1198, 1083, 1747, 1324]. non-Abelian [1547]. non-affine [1691]. non-autonomous [591, 1324]. non-boundary [1747]. non-conforming [965]. non-convex [1166]. non-destructive [857]. non-deterministic [1477]. non-diagonal [28]. Non-diminishing [1077]. Non-dominated [1606]. non-essential [1198]. non-fragile [1657, 1666]. non-isothermal [1702]. non-linear [1837, 1330, 964, 1691, 567, 1762, 1711, 1560, 1345, 1083]. non-locality [1483]. non-monotone [1425]. non-negative [993]. Non-Newtonian [46, 89, 84, 1570]. non-orthogonal [1555]. non-oscillating [1237]. non-periodic [506]. non-regular [335]. non-shrinkage [1874]. non-smooth [241, 798]. non-smoothnesses [793]. non-standard [1018, 347, 920]. non-stiff [485]. non-transitive [671]. noncharacteristic [1261]. nonconcave [982]. nonconforming [1398, 1053, 1054]. nonhomogeneous [592, 1117]. nonhyperbolic [799]. Nonlinear [1340, 1311, 1220, 188, 977, 1195, 1701, 1233, 1019, 1882, 517, 980, 1643, 831, 1789, 1627, 1329, 543, 1091, 154, 595, 504, 490, 572, 665, 1667, 1849, 1103, 1425, 1871, 497, 582, 1275, 461, 1229, 1677, 142, 885, 450, 1664, 555, 129, 680, 431, 244, 1156, 1590, 632, 352, 771, 1863, 1240, 26, 1076, 1539, 792, 521, 1482, 1126, 1053, 391, 1642, 470, 518, 1189, 959, 1860, 111, 1244, 1524, 961, 146, 1243]. nonlinearities [334, 488]. Nonlinearity [771, 1344]. Nonlinearity-aware [771]. Nonlocal [904, 1105, 836, 1865]. nonnegativity [1425]. nonoverlapping [860]. Nonparametric [590, 272, 425]. Nonsmooth [1509, 1051, 800, 1009]. Nonstandard [1202, 29, 396, 1154, 1451, 1349, 916]. nonstationary [1199, 871, 210]. nontransverse [799]. nonuniform [1396]. norm [1871]. normal [163]. Normalization [601]. normalized [859]. note [765, 556]. Novák [817]. Novel [1811, 990, 1649, 1687, 1879]. November [630, 926, 1074, 1255, 1622, 1834]. nozzle [406]. NSFD [1235]. nuclear [333, 119]. null [1109]. number [685, 81, 317, 829, 498, 1776, 876, 1704, 1714, 1477, 24]. numbers [1617, 1777, 1364, 90, 1528]. numeric [1886]. Numerical [837, 1089, 1386, 1818, 486, 1760, 1702, 1100, 881, 1101, 1438, 1399, 494, 1225, 1214, 79, 1226, 1756, 392, 994, 407, 165, 787, 75, 811, 1762, 1079, 506, 89, 507, 568, 85, 1518, 813, 1633, 794, 406, 207, 84, 90, 855, 164, 1601, 1859, 1451, 1434, 520, 1198, 1249, 1354, 1850, 146, 1388, 1617, 1837, 1390, 1391, 864, 1874, 1018, 1196, 1733, 602, 786, 389, 823, 1070, 533, 572, 1447, 1016, 238, 1007, 1437, 1800, 772, 1369, 569, 1158, 1827, 990, 1614, 1796, 1805, 662, 1793, 471, 679, 1884, 41, 351, 1234, 607, 244, 130, 353]. numerical [354, 118, 1838, 514, 1322, 792, 1705, 798, 1885, 1814, 599, 835, 793, 1865, 303, 1330, 319, 563, 42]. numerically [1432]. NURBS [547]. NWBUE [1059]. NXFEM [1398]. Nyström [1090, 768].

O [62, 1686]. obeys [1449]. object [1617]. objective [1755, 1628, 633]. Observations [808, 822]. observer [1046, 471, 688, 1605, 845, 470]. observers [1060]. obstacles [218, 829]. obstructive [943, 320]. obtain [1204]. occur [793]. occurring [1657]. ocean [855]. oceanic [496]. octal [1472]. Octave [1408]. October [414, 620, 1065, 1247, 1430, 1612]. OD [1644]. ODE [1068, 509]. ODEs [1625, 1016, 1370, 798]. off [278]. off-the-shelf [278]. OFFD [1746]. oil [1823, 861, 780, 323, 367, 828]. old [273]. old-age [273]. oligopolies [978]. oligopolistic [99]. Oligopoly [976, 1685, 625, 626]. Olkin [1534]. oncolytic [1578]. One [126, 796, 158, 1764, 1515, 234, 885, 794, 1877, 1450, 831, 1310, 1143, 1244, 1795]. one-dimensional [796, 158, 1515, 234, 885, 1877, 1450, 1244]. One-group [126]. one-leg [831]. one-parameter [1143]. one-predator [1310]. online [471]. Ono [519]. onset [726]. onto [1345]. OOA [27]. OOF3D [604]. OPAL [1719]. OPAL-RT [1719]. Open [1296, 685, 686, 823, 690, 1212, 1739, 1554, 1555, 1276]. Open-circuit [1296]. open-end [686, 690, 1276]. open-source [1739, 1554, 1555]. OpenComp3d [1739]. OpenFOAM [1747]. openings [1270]. operating [1265, 712]. operation [177, 1536, 1733, 182, 1293, 694, 633]. Operational [1450, 1762, 453, 1850]. operations [702, 749, 1003]. operator [1227, 601, 1231, 1169, 349, 1845, 1763, 514, 132]. operator-fitted [349]. operators [942, 1313, 1547, 1448, 407, 1682]. Optimal [211, 695, 685, 176, 702, 709, 1828, 213, 1151, 357, 1379, 1658, 846, 329, 1584, 1719, 875, 950, 1725, 247, 1774, 1376, 910, 1282, 1578, 1302, 214, 1156, 912, 1689, 792, 1126, 476, 1718, 733, 1524, 585, 330, 216]. optimality [945]. optimisation [192]. Optimization [405, 898, 1654, 1617, 573, 457, 1736, 715, 21, 1052, 1130, 180, 703, 1146, 702, 1377, 1870, 1139, 1675, 1648, 1786, 1720, 1724, 1880, 1166, 1628, 1343, 1714, 535, 1687, 555, 322, 697, 1002, 1732, 1450, 1841, 645]. Optimized [687, 1887, 1266, 1314]. optimum [874, 932]. Option [153, 3, 267, 572, 767, 1312, 1866]. options [824, 459, 205, 873, 1174, 117, 103]. OQBEM [1870]. orbits [1540, 389]. order [156, 1424, 232, 1445, 573, 1154, 1840, 1593, 124, 1165, 595, 558, 370, 1438, 995, 248, 1016, 1168, 1691, 1591, 495, 965, 433, 385, 1376, 932, 1786, 1044, 1677, 85, 664, 679, 1531, 129, 930, 454, 1635, 1359, 1791, 912, 1626, 795, 1029, 1493, 1863, 1844, 1240, 1560, 1126, 1053, 1027, 1427, 1564, 518, 1201, 132, 746, 675, 1809, 1746, 649]. ordering [935]. ordinal [1880]. ordinary [248]. organizations [1861]. organizing [437]. oriented [80, 1556]. original [1813, 1285]. orthogonal [1827, 617, 130, 1555, 1753]. oscillating [1237]. Oscillation [1367]. Oscillations [1792, 483, 874]. oscillator [791, 794, 675]. oscillators [533, 255, 541, 648]. oscillatory [1331, 565, 510]. Oseen [562]. other [1460, 1206, 277]. outbound [755]. outlier [102]. outliers [936]. outlook [800]. output [1286, 1276, 1319, 1746, 1123]. overdriven [1028]. overestimator [1166]. overlapping [77, 1885, 1071, 1354, 1686]. Overview [696, 741, 773, 1082, 1619]. Overvoltage [1726]. OWA [1169]. own [350].

Pacific [505]. package [370, 919]. packet [1647]. Padovan [1207]. Padovan-like [1207]. Pages [135, 222, 4, 108, 93, 151, 293, 383, 446, 251, 364, 307, 414, 524, 464, 580, 539, 630, 620, 611, 762, 668, 657, 730, 673, 841, 867, 879, 820, 926, 907, 955, 1032, 1040, 988, 1022, 1074, 1065, 1057, 1163, 1210, 1135, 1120, 1181, 1149, 1172, 1255, 1247, 1352, 1413, 1454, 1305, 1374, 1430, 1521, 1587, 1631, 1491, 1567, 1576, 1622, 1612, 1597, 1784, 1857, 1671, 1750, 1695, 1748, 1834, 1803]. paired [257]. Pakhtunkhwa [1847]. Pakistan [1847]. panel [1283, 271, 886, 112]. pantograph [1590]. papillomavirus [455]. Parabolic [334, 1627, 1388, 881, 73, 1425, 1457, 1845, 355, 26, 1539, 521, 87]. paradigm [399]. paradigms [808]. paradox [659, 811]. Parallel [205, 1675, 474, 337, 1323, 1725, 722, 1552, 616, 990, 1364, 189, 1717, 1512, 1293, 1637, 562, 1459]. parallelization [366]. Parallelizing [1408]. parameter [1332, 316, 1052, 279, 911, 242, 1515, 260, 1500, 101, 1108, 418, 466, 1509, 613, 352, 700, 44, 1852, 1143, 1746, 1599]. parameter-dependent [1515]. parameterized [1807]. parameters [405, 1283, 1752, 1097, 991, 38, 527, 317, 343, 179, 567, 898, 255, 843, 1436, 691, 1534, 1778, 1643, 1594, 936, 1177, 844]. parametric [917, 1370, 1377, 856, 226, 357, 470]. parametrically [541]. parametrized [1517]. parareal [589]. Pareto [1130, 369]. park [1358, 182]. part [1642, 625, 626, 1099, 1100, 1554, 1555]. Partial [1168, 1469, 1487, 1052, 246, 1762, 164, 1560, 1201, 1616]. particle [1589, 513, 1243, 208]. particular [1239]. partition [397]. partitioning [695]. partitions [344]. Passive [138, 1661]. passivity [1646]. past [197]. patch [1338]. path [1184, 1416, 1605, 203, 1001, 792, 998]. path-independent [998]. pathogen [1766]. paths [189]. pathways [1710]. Patient [320]. Patient-specific [320]. patients [373, 943]. Pattern [1062, 322]. patterned [1107]. patterning [1530]. patterns [1331, 1308, 565, 1846, 1000]. payment [749, 1003, 1201]. PCF [734]. PCM [1746]. PCR [96]. PDE [1425, 512, 1494]. peak [711]. PEM [1737, 1721]. penalty [1380]. penalty-free [1380]. penetrable [218]. penetration [1536]. perfect [1223]. Performance [584, 1035, 1673, 217, 1636, 259, 1237, 678, 1499, 141, 1143]. performances [1583]. period [1661, 638]. Periodic [96, 1540, 1752, 389, 1017, 493, 495, 142, 506, 1523, 519, 1419, 371, 960, 1650]. periodically [1759, 794, 475]. periodically-forced [794]. periodicity [884]. peristaltic [1570]. permanence [1167]. Permanency [1700]. permanent [1272, 699, 685, 1719, 1297, 176, 188, 1715, 932, 1274, 1628, 1271, 883, 1302, 700, 697, 1270, 696]. permanent-magnet [1274]. permeability [861]. permissible [1201]. persistence [1177]. persistent [809]. perspectives [214]. perturbation [450, 1882]. perturbations [627, 348, 736]. perturbed [349, 1851, 1845, 1763, 801, 1494]. Petri [1723, 435, 717]. Petviashvili [1175]. pH [1533]. pharmacophore [1336]. phase [837, 1389, 1390, 1819, 853, 1296, 1719, 686, 1400, 1675, 1715, 75, 885, 690, 727, 350, 1853, 568, 182, 60, 380, 680, 354, 725, 59, 1741, 356, 724, 1482, 56, 1409, 57, 1660]. phase-field [59, 56]. phases [685]. Phenomena [1220, 1737, 55]. phenomenon [1530, 633]. phi [1140]. phi-divergence [1140]. phonon [200]. photodiode [710]. photoinhibition [44]. photosynthesis [44]. photosynthetic [815, 575]. photovoltaic [1283, 1287, 709, 1794, 1286, 1712, 1742, 710, 1537]. photovoltaic/wind [1537]. physical [1733, 1561]. physically [164]. physics [1106, 1432]. PID [874]. Piecewise [788, 1678, 158, 448, 1078, 239, 797, 1682, 1010, 1878]. piezoelectric [192, 1300]. Pinning [647]. pipeline [594]. Piper [344]. piracy [1867]. pitfalls [278]. pivot [571]. placement [387]. planar [100, 881, 670, 1451]. Planck [854]. Plane [490, 634, 1203, 1709, 1849, 1212, 930]. Plane-wave [490]. plankton [602]. plankton-induced [602]. Plant [961, 1535, 1736, 1766, 828, 1436, 7, 476, 1618]. plants [1583, 961]. plasma [63, 513, 406]. plastic [772, 1798, 574]. plasticity [1028]. plate [1728]. plates [654, 1812, 562]. platforms [1461]. PMSM [687, 186, 182]. POD [679, 1531]. POD-based [1531]. Poincaré [373]. point [1445, 159, 1080, 385, 1777, 143, 419, 1468, 1556, 113, 1314, 1126, 999, 934]. points [1573, 799]. Poisson [1424, 1667, 1829, 230, 737]. Pol [790, 541, 675]. polar [1096, 1355, 1116]. polarization [1243]. polarized [517, 1778]. Pole [387, 1714]. policies [1692]. policy [983]. pollutant [419]. polluted [1722]. pollution [980]. Pólya [35]. polycrystalline [52]. polycyclic [1340]. Polynomial [1497, 997, 1757, 592, 1590, 35, 96, 999, 571, 468]. polynomially [403]. polynomials [1444, 1502, 1207, 1474, 1827, 920, 1250]. polynomiography [1752]. pools [438]. popular [769]. population [330, 1329, 1330, 1340, 810, 1446, 911, 811, 982, 1838, 896, 275, 397, 1517, 1528, 1529]. populations [752, 872, 1014, 273, 648, 1189, 1496]. pore [336, 1859]. pore-scale [336]. porosity [861]. porous [837, 1390, 1819, 1392, 857, 485, 1433, 1102, 341, 1870, 346, 1762, 350, 568, 1830, 160, 59, 557, 1315, 958, 361, 1354]. porous-saturated [1354]. port [1734]. posed [316]. posedness [521]. position [1297, 1764, 693]. Positive [129, 1419, 1878]. posteriori [80, 1516, 72, 576, 1875, 1830]. POT [781]. potential [1811, 1160, 278, 1533, 1878]. Poussin [1562]. Powell [343, 856, 859]. Power [1265, 1727, 717, 846, 1340, 1535, 1536, 720, 1733, 1736, 715, 1734, 1277, 1583, 1449, 1772, 716, 1794, 1720, 727, 711, 1291, 1281, 1286, 714, 1731, 1739, 1711, 1722, 1741, 1266, 724, 410, 1288, 1730, 1580, 719, 1716, 1537, 1709, 178]. Practical [1215]. practice [1145]. PrBa [62]. Pre [1848, 1269]. Pre-conditioning [1848]. pre-design [1269]. preassigned [949]. precision [634, 1777]. preconditioned [1080]. preconditioner [78]. Preconditioning [559, 570, 571, 566]. predation [1842]. predator [1700, 928, 1184, 1308, 1579, 1000, 1530, 1335, 636, 1079, 764, 1127, 1790, 830, 423, 1310, 934, 1795, 766]. predators [1184]. predict [1653]. Predicting [317, 726, 266]. Prediction [822, 699, 97, 1318, 833, 1787, 1144, 1580, 719, 2]. predictions [637]. Predictive [1649, 931, 1719, 188, 716, 861, 1741, 719]. predictor [1077]. Preface [327, 1088, 69, 554, 684, 708, 850, 891, 1280, 1268, 642, 1641, 36, 1050, 807, 265, 608, 1708]. premium [266]. prescribed [297]. presence [495]. preservation [1556, 1011]. preserving [1104, 952, 1012]. pressure [1388, 484, 219, 491, 772, 350, 623, 534]. pressurized [1400]. prestress [47]. prevention [1569]. prey [1700, 928, 1184, 1308, 1579, 1000, 1530, 1335, 636, 1079, 764, 1127, 1790, 830, 423, 1213, 1310, 1842, 934, 1795, 766]. prey-dependent [1530]. Price [416, 459, 780, 323]. prices [288, 112, 754, 747]. Pricing [103, 267, 572, 153, 205, 767, 873, 3, 1866, 1616]. prime [1772]. primitive [855]. principal [1376]. principle [1309, 73, 1682, 118, 1741, 87]. principles [62]. prismatic [87]. probabilistic [34, 1823, 1260, 1880]. probability [846, 1807, 311]. probes [1607]. problem [1089, 864, 901, 1668, 1094, 853, 895, 1761, 1888, 408, 1800, 218, 897, 1403, 1404, 1457, 127, 910, 577, 1406, 885, 1513, 1607, 1687, 1884, 118, 355, 559, 26, 1076, 1539, 360, 958, 1643, 586, 1875, 1865, 476, 1176, 424, 303, 1608, 1788, 733, 1261, 1571, 1538]. problematical [808]. problems [329, 1589, 1090, 1445, 1755, 1018, 316, 1395, 531, 212, 1099, 1100, 881, 1101, 224, 492, 319, 1370, 79, 665, 411, 1103, 1402, 228, 569, 1509, 1405, 1880, 1166, 1828, 566, 1614, 349, 80, 1518, 679, 1845, 1512, 1156, 454, 1763, 622, 1626, 1506, 1112, 1557, 603, 1510, 1527, 231, 77, 1516, 1482, 798, 574, 87, 887, 1674, 768, 361, 585]. procedure [548, 860, 259, 13, 397, 1143]. procedures [159, 1108]. process [224, 430, 437, 613, 873, 1337, 473, 304]. processes [776, 64, 1652, 584, 105, 1525, 557, 1117, 115, 1690]. processing [904, 857, 1091, 828]. processor [206]. procurement [1359]. Product [360, 1043, 1197, 1549, 1515, 584]. production [1823, 1720, 982, 710]. productivity [1146]. products [1311, 259, 1143]. profile [1636]. profiles [529]. profit [869, 1797]. profit-centered [1797]. programmable [1277]. programming [1755, 1736, 535, 1607, 1156, 887]. progress [1659]. progressive [743, 1143]. progressively [1393, 1799, 421, 1580, 526]. projected [1510]. Projection [1095, 290, 1051, 1482]. projective [680]. Projector [566]. proliferation [810]. proof [1556, 947]. propagated [961]. propagating [494]. propagation [95, 492, 505, 1235, 514, 1601, 516, 520]. Proper [549, 226, 1690]. properties [1756, 617, 737, 65, 1238, 745, 1034, 54, 576]. property [473]. proportional [1165, 1378, 526]. proportions [1140, 358]. prospects [756]. protein [509, 1348]. proteins [20]. proton [1299, 170]. protons [20]. providers [1731]. pseudo [1388]. pseudo-parabolic [1388]. pseudolocalized [1224]. pseudorandom [876, 1477]. pseudospectra [556]. pseudospectral [124, 492, 1601]. Publisher [765]. Pulsatile [1885, 1354]. Pulse [1459, 186, 1661, 440, 1177]. pulsed [1296, 996]. pumps [605]. pure [1563]. purpose [370, 186]. purposes [1727, 1269, 696]. PWM [172, 686, 725, 1726]. Pythagorean [550]. Pythagorean-hodograph [550].

QCQP [1510]. QMC [202]. QMLE [279]. Quad [592]. Quad/triangle [592]. quadratic [1081, 1344, 297, 1002, 887]. quadratic-cubic [1344]. quadrature [1017, 1019]. Quadrupole [483]. Qualitative [875, 1195, 1079, 930, 423, 766]. quality [1776, 1704, 1722, 1741, 357, 1890]. Quantification [1287, 861]. quantile [323]. quantitative [391]. quantities [297, 1863, 999]. quantity [1635, 1201]. quantum [233, 622]. quartic [1376]. quartic-order [1376]. Quasi [911, 204, 1781, 328, 332, 852, 1391, 1821, 1822, 1026, 1096, 1449, 347, 1827, 1853, 197, 25, 360, 1115, 1116, 1243, 198]. quasi-interpolant [332, 1822]. quasi-interpolants [328, 852, 1391, 1026, 347, 1115, 1116]. quasi-interpolation [1821, 1096, 1827, 360]. Quasi-Monte [911, 204, 1781, 197, 25]. quasi-Newtonian [1449]. quasi-particle [1243]. quasi-random [198]. quasilinear [901]. quasistatic [569, 1506]. queue [457]. queueing [678]. queues [1876]. quintic [504, 550, 1244].

R&D [625, 626]. radar [1188]. radial [1273, 700]. radiation [227, 1778]. radiative [607]. Radio [1644]. radiotherapy [1343, 1183]. rail [187]. railway [1266, 398, 793, 321, 1716]. RAMs [1827]. Random [1371, 563, 1364, 822, 964, 935, 458, 317, 616, 1776, 1779, 230, 1704, 1773, 617, 1042, 736, 809, 24, 575, 255, 1029, 1493, 1863, 422, 1471, 198, 1479, 633].

Randomization [422]. Randomized [21, 197]. randomly [1657, 1859]. Randomness [11]. range [1724]. rank [312, 422, 425]. Rankine [1564]. ranking [1122, 1636]. Rao [331, 851]. Raphson [1432]. Rapid [861]. rare [1475, 752]. rate [329, 153, 1334, 564, 1378, 3, 770, 982, 1554, 113, 1157, 526, 676, 1324]. Rates [161, 1340, 824, 1335, 769, 1485, 300, 754, 1189, 755]. ratings [780, 285]. ratio [1700, 1376, 1530, 1704, 114, 896, 96]. ratio-dependent [1700, 1530]. rational [1685, 1165, 1474, 1758, 425]. rationale [70]. rationality [869, 1789]. rationalized [1849]. Rayleigh [487]. rays [1185]. RBF [1548, 893]. Re [1617]. reactance [845]. reacting [212]. Reaction [587, 245, 1846, 1438, 408, 430, 497, 228, 1000, 1158, 349, 536, 1458, 591, 352, 417]. reactions [1471]. reactive [1390, 1819, 1401, 1794, 1731, 1409]. reactors [1034, 468]. reading [864]. Real [1297, 1723, 1281, 930, 105, 1683, 788, 720, 600, 949, 348, 443, 545, 1730, 735, 1719]. Real-time [1297, 1723, 105, 720, 348, 1730, 735, 1719]. real-world [600]. realization [569]. Realized [614, 267, 269, 270]. Rearranged [1105]. reboot [368]. receptor [1336]. recognition [1098]. recombinant [981]. reconfigurable [720]. reconfiguration [1742]. reconstruction [1443, 335, 951, 118, 139, 303, 1569, 1665]. record [1093, 1142, 259]. Recovering [858]. recovery [1426, 1324]. recruitment [1517]. rectangle [1549]. rectifier [1296, 176, 725]. rectifiers [727]. recurrent [976, 735, 643, 436]. Recursion [1547, 35]. recursive [466]. Reduced [573, 1389, 433, 679, 1531]. reduced-order [679, 1531]. reducing [81]. Reduction [241, 1478, 1165, 993, 824, 1591, 1628, 189, 138, 1029, 1493, 1863, 1436, 1839, 1718, 1665]. refinable [407, 213]. refinement [592, 1110, 606, 1875]. refinements [946]. refracting [835]. refuge [636]. regeneration [1811]. regime [834, 1185]. regimes [279, 899, 97]. region [973, 1237]. Regional [756]. regionally [276]. regions [789, 387, 1451, 403]. registration [941, 1548]. Regression [1294, 1478, 253, 99, 323, 643, 645, 1250]. regression-based [1478]. regular [335]. Regularization [635, 424, 316, 298, 256, 836, 586, 386, 733, 157, 1571]. regularized [994, 1494]. regulating [1123]. regulatory [1195, 1320]. reinforcement [1343]. rejection [1741]. related [1387, 671, 1619, 55, 402, 1241]. relating [1530]. relations [876]. relationship [1517]. relationships [584]. relative [1764, 1077]. relaxation [589]. Reliability [421, 526, 548, 337, 616, 1378, 402]. Reliable [111]. reluctance [185, 1714, 189, 701, 1729, 1269, 1718]. remarks [1547]. Remodelling [42, 1155]. remote [32, 1407, 442]. removal [1808]. renal [441]. Renewable [174, 979]. renewal [976]. repair [368]. repeated [947]. replacement [45]. representations [1252, 1493, 1863]. represented [399]. reproduction [592, 1528]. Requiem [33]. required [1731]. Requirements [1364]. resampling [198]. research [1636]. reservoir [1887, 1826, 1110, 1409, 1410, 825]. reservoirs [1402, 861]. reset [1827]. residence [1473]. Residual [411, 910]. Residuals [408]. residue [675]. resistance [1315]. resistive [1827]. resistivity [118]. Resolution [24, 377, 1237]. Resolution-stationary [24]. Resonance [187]. Resonant [1231]. resource [358]. resources [979, 1386, 1818]. respect [1642]. respiratory [11, 1754]. response [1700, 1501, 564, 1335, 142, 1079, 764, 1291, 1045, 1790, 830, 86, 1456]. responses [1127]. restarting [593]. restoration [948]. resultants [30, 1206]. results [1401, 1185, 1339]. retail [1357]. retailers [1201]. retention [359]. retrieval [1660]. retrieving [1482]. return [1710]. returns [287]. reversed [1378]. reversible [1774, 104]. reversion [112]. reverting [1866]. review [1369, 52, 72]. revised [1556]. revisited [789, 1321, 976]. Revisiting [919, 425, 1681, 1569]. rhizosphere [1618]. Richards [1196]. Riemann [37, 964, 920, 1242]. Riesz [1434, 1889]. right [901, 421, 1580]. right-hand [901]. rigid [1499]. ring [1474]. ripple [185, 1718]. Risk [775, 773, 1312, 285, 779, 748, 1634, 780, 777, 273, 367, 453, 529, 782, 775, 453, 781]. risks [1393, 1025]. river [558, 1725]. road [565, 1113]. roads [903]. Robin [1826, 1424]. robot [469]. robots [829, 208, 1499]. Robust [1284, 1646, 1691, 932, 9, 1738, 1525, 1656, 789, 777, 946, 1791, 1763, 14, 113, 410, 1153]. robustness [290, 1813, 203, 923, 391]. rock [1389]. rod [520]. Rodrigues [550]. role [1501]. root [1752, 271, 1376, 1129, 1045, 1127, 512, 1012, 425]. roots [995, 912]. Rosenbrock [803]. Rosenbrock-type [803]. rotating [1870, 1205, 1236, 173]. rotor [1297, 1715, 701, 1718]. rough [1591, 1787, 102]. RQMC [1485]. RT [1719]. rubber [287]. rule [1787, 1001]. rules [1685, 1197]. run [312]. Runge [929, 145, 768].

S [931, 651]. Saadatmandi [1752]. Sabin [343, 856, 859]. SABR [767]. saddles [788]. Sage [1242]. sags [171]. saltwater [1633]. sample [1617, 822, 1093, 616, 1406, 421, 114, 467]. sampled [1657]. sampled-data [1657]. samples [1393, 1318, 515, 1580, 526]. sampling [752, 186, 614, 198, 111]. sampling/importance [198]. samplings [214]. sandwich [832]. SARFIMA [843]. satisfying [945]. saturated [1701, 1157, 1354]. Saturation [340, 1324]. Saturn [670]. Saunders [116, 1532]. saving [1658]. scalar [1323, 964]. scale [1600, 336, 600, 1139, 1794, 1735, 1291, 43, 63, 44, 474, 1028, 1315]. scaled [1713]. scales [1238, 1419]. Scaling [1713, 1559]. scattered [1546, 893]. scattering [504, 493, 218, 411, 132, 1176]. scenario [1475, 789, 973]. scenarios [584]. schedules [1205]. scheduling [435, 1288, 672]. scheme [1423, 1819, 543, 95, 558, 1080, 572, 1433, 1871, 1403, 1786, 1872, 1851, 1235, 454, 1780, 1556, 105, 356, 603, 1877, 1494, 1563, 1564, 1674, 378]. schemes [1475, 941, 1091, 1573, 1154, 335, 1478, 1702, 872, 1014, 945, 929, 1229, 1187, 353, 1237, 1011, 952, 1322, 1699, 1560, 800, 1427, 1324]. Schrödinger [517, 1243, 1837, 504, 490, 1871, 1229, 1434, 518, 1479, 959, 1878, 146]. science [604]. Scientific [1544, 207, 404, 599]. Scirocco [406]. score [1129]. scorecard [1864]. scorecard-based [1864]. scrambled [201]. screen [1561]. Screening [1600, 933]. SDEs [1780]. SDS2014 [1193]. sea [602, 496]. Search [1062, 815, 322]. seasonal [627, 843, 1525]. Seasonality [318, 281]. seawater [1089, 355]. secant [1131]. Second [370, 1677, 852, 1165, 248, 1016, 215, 155, 1844]. Second-order [370, 1677, 1165, 1844]. second-sound [155]. secondary [1472]. sector [654]. secular [670]. security [1416, 289]. Sediment [1584, 338]. seeking [430, 297]. seen [1564]. seepage [677]. segmentation [1107, 1141, 386]. segmented [1718]. segmented-rotor [1718]. segmenting [1407]. segments [491]. SEIADR [1808]. SEIRS [1167]. SEIS [638]. Selecting [460]. Selection [1436, 1887, 1687, 560, 719]. Selective [723, 1692]. self [759, 437, 1607, 115, 734]. self-assembled [1607]. self-organizing [437]. self-similar [115, 734]. Semi [1114, 102, 948, 1433, 1403, 1404, 228, 884, 1458, 356, 1645, 1301, 637]. semi-analytical [1301]. semi-discrete [1433, 228, 356]. semi-discretization [637]. semi-discretizations [884]. Semi-hidden [1114, 1404]. semi-linear [1458]. semi-Markov [1645]. semi-stationary [1403]. Semi-supervised [102]. semi-transparent [948]. Semiclassical [489]. semigroups [1836]. semilinear [1457]. semiparametric [843]. semismooth [574]. sensing [1407, 651]. Sensitivity [1806, 1029, 199, 1252, 662, 886, 1288, 933, 1528]. sensor [653, 1296, 208, 1699]. sensorless [695, 1060, 1297, 186, 932]. sensors [440, 710]. Seo [386]. separable [1011, 1510]. separation [147]. SEPIC [727]. SEPRAN [370]. September [151, 611, 907, 1057, 1597, 1803]. septic [385]. SEQIJR [1528]. sequences [186, 1321, 1369, 1772, 1207, 1474, 1365, 1753, 1114, 201]. sequential [1390, 747]. Serial [33, 758]. series [232, 337, 1725, 833, 1759, 1787, 284, 272, 1215, 2]. series-parallel [1725]. series/parallel [337]. server [1655]. servers [1655]. service [1864, 1731]. sessile [1107]. set [1634, 143, 1677, 1468, 1787, 1112, 1113, 545, 719]. set-point [143]. SET50 [117]. sets [436, 591, 734]. settings [1775]. SEV [117]. seven [686]. seven- [686]. Seventh [1220]. several [1433, 912, 1674]. sexually [1862]. shallow [863, 1095, 1138, 1572, 603, 1322]. shallow-water [1572]. Shape [1446, 573, 1104, 555, 952]. shape-preserving [952]. shaped [1225]. shapes [70, 600]. sharing [1711, 545]. Sharp [1008, 959, 1399]. sharpening [1545]. shear [1800, 1756, 510]. shear-thinning [1756]. shedding [1536]. sheet [1764]. shelf [278]. shells [1380]. Shepard [1448]. Sherman [215]. shielding [1274]. shifting [284]. ship [1725, 672]. shocks [1223, 323, 1028, 276]. shop [622]. short [1733, 153, 3, 770]. short-circuit [1733]. short-term [770]. shortages [435, 1201]. shortest [1416, 203]. shortest-path [1416]. shrinkage [1874]. shuffled [322]. shunt [1722, 1741, 724]. SI [1154]. SiC [1733, 1726]. SiC-based [1726]. side [901]. sided [671]. sigmoid [1342]. sign [1323]. signal [1297, 1105, 1648, 114, 1649]. signal-to-noise [114]. signalized [1876]. signals [373]. signature [1715]. silos [1697]. similar [115, 734]. SIML [269]. simple [863, 1026, 295, 983, 690, 957, 1704, 791, 946, 117, 971, 47]. Simpler [593]. simplex [1116]. Simplified [1275, 1700, 692, 1582]. Simul [153, 937]. simulate [692, 1591]. simulated [1367, 460]. simulating [1366, 320]. Simulation [219, 441, 1490, 874, 1343, 473, 1619, 1003, 1697, 207, 376, 719, 1475, 921, 918, 1390, 290, 1600, 469, 1330, 741, 405, 457, 720, 862, 1188, 258, 1438, 527, 1146, 79, 64, 1122, 1756, 1680, 1870, 1368, 433, 1727, 753, 994, 1275, 861, 228, 1826, 418, 1872, 1735, 22, 1044, 1336, 990, 1079, 1853, 1110, 1174, 568, 1518, 206, 1572, 1781, 617, 1633, 1045, 886, 882, 679, 1111, 1869, 104, 130, 55, 354, 1712, 575, 1358, 84, 255, 1495, 275, 164, 557, 422, 1471, 1153, 1409, 184, 1499, 115, 599, 1754, 432, 520, 57]. simulation [468, 1599, 1878, 1071, 1354, 1145, 1569, 58, 652, 969]. simulation-analytical [1146]. Simulation-Based [1619, 1343, 719]. Simulations [1332, 1397, 1668, 615, 336, 943, 1848, 1225, 53, 811, 233, 1805, 119, 1476, 736, 1884, 63, 607, 59, 1814]. simulator [1195]. simulators [1730]. Simulink [718]. Simultaneous [1732, 340, 751, 747]. Sinc [1788]. sine [489]. sine-Gordon [489]. single [1819, 1283, 1296, 1647, 1624, 727, 354, 725]. single-linkage [1647]. single-phase [727, 354]. singular [1197, 242, 1527, 360, 132, 999, 1674]. singularities [1758]. singularity [812]. Singularly [349, 1851, 1845, 1763, 1494]. sinh [1758]. sinh-mapped [1758]. sink [1728]. sintering [60, 380]. SIR [1154, 827, 1157, 1379]. SIRS [309]. SIS [1498, 816]. situations [1262]. size [1092, 822, 1093, 616, 1528, 1713, 1496]. size-scaled [1713]. Sizing [1725, 1709, 1720, 1302, 1732, 1288]. slab [126]. Sleep [320, 943]. Sliding [845, 240, 1123, 1323, 137, 1691, 1008, 722, 932, 314, 1677, 688, 632, 1605, 725, 1009]. Sliding-mode [845, 932, 632]. slightly [958]. slip [1513, 562]. sliver [1545]. slot [1270]. slotless [176]. slotted [1301]. slow [1338, 1517]. slow-fast [1517]. Smagorinsky [855]. small [711, 1643]. Smart [1537, 1736, 702, 717]. smooth [158, 1078, 239, 797, 202, 1470, 241, 798]. smoothing [343, 1825, 892]. smoothnesses [793]. Sneyd [378]. Sobol’ [199, 1772, 662]. Sobolev [1531, 1053, 1765]. SoC [1742]. Social [289]. soft [176, 6, 43, 651]. software [695, 278, 1714, 1867, 1408, 704]. soil [1045]. soils [1702]. solar [227]. solid [100, 1761, 79]. solidification [65]. solids [1230, 516, 54]. Solitary [98, 518, 484, 1226, 507, 1344, 297, 516, 520]. solitary-wave-like [507]. soliton [495, 1249]. solitons [483, 493, 494, 1160, 1231, 1241, 517, 1244]. solute [336, 679]. solutes [1458]. Solution [1761, 634, 1617, 1329, 1445, 486, 1018, 1196, 1017, 881, 1070, 1447, 1016, 76, 1515, 1437, 1800, 218, 215, 411, 74, 1158, 1762, 1614, 244, 1156, 1883, 1763, 226, 231, 26, 1076, 1539, 1408, 963, 798, 1410, 562, 1809, 735, 1249, 1850, 146, 960]. solutions [1388, 329, 1392, 531, 490, 1224, 964, 1214, 239, 1758, 811, 577, 256, 129, 512, 1239, 90, 544, 1601, 1434, 98, 518, 519, 1419, 1529, 371]. solve [1687]. solved [87, 1788]. solver [366, 863, 37, 1080, 604, 81, 376, 45, 835, 57]. solvers [1068, 391, 28]. Solving [1424, 1876, 1130, 1849, 278, 417, 1860, 1571, 1469, 1821, 1888, 1786, 1880, 1166, 1513, 1314, 1494, 1844, 1126, 887, 1681]. SOM [718]. Some [1203, 899, 901, 1103, 802, 1342, 769, 745, 1185, 1482, 72, 574, 519, 1547, 238, 1334, 1485]. Song [995]. sorption [346]. Sorting [1485, 1606]. sound [155, 1213]. source [1387, 1535, 370, 1457, 419, 1654, 1739, 1554, 1555, 147, 635, 717]. sources [303, 713]. Space [1405, 477, 871, 1433, 125, 204, 185, 690, 1110, 859, 1234, 356, 103, 435, 603, 1665, 1660, 1561, 694, 635, 1889, 733, 1571, 1128]. space-fractional [733, 1571]. space-time [103]. spaces [1217, 1826, 1470]. spalling [82]. span [833, 2]. spanning [1538]. sparse [1252]. sparsification [1758]. Sparsified [422]. Spatial [1000, 510, 886, 1581, 361]. spatially [823, 872, 1014]. spatio [1308]. spatio-temporal [1308]. SPDE [1476]. SPDEs [351]. Special [652, 1386, 1818, 1770, 1202, 248, 969, 916, 1775, 29, 519]. species [627, 1624, 766]. specific [320]. specification [271]. speckled [590]. spectra [671, 243]. Spectral [281, 546, 157, 489, 233, 1759, 1889]. spectrum [1198]. speed [176, 1274, 1302, 694, 637]. sphere [1822, 79]. spheres [1366, 1471]. spherical [1313, 944, 515]. spheroids [1764]. spillover [377]. spillovers [778]. spindle [637]. Spline [838, 332, 852, 1391, 1821, 1822, 1026, 903, 1828, 664, 859, 1355, 1115, 1116, 1447]. splines [328, 343, 856, 1825, 1141, 892, 898, 1116, 835]. split [994, 1649]. Splitting [1614, 1011, 1094, 233, 28, 1765]. spot [287, 105]. Spread [1698]. spreads [1250]. Spurious [304]. square [254, 1197, 1887, 313, 1462, 1644, 929, 1129, 536, 1127, 512]. square-root [1129, 512]. squares [994, 466, 352, 376]. Stability [1308, 1013, 827, 1251, 1344, 1152, 144, 1615, 1157, 1831, 1417, 637, 542, 825, 371, 378, 1662, 1338, 1813, 254, 504, 1125, 899, 582, 726, 929, 636, 1067, 507, 536, 1664, 528, 243, 1019, 1590, 14, 624, 670, 1814, 1642, 1882, 1563, 1357, 229, 676, 831, 1177, 1419, 1679, 649, 162, 110, 1317]. stabilizability [528]. stabilization [895, 416, 9, 801, 534, 1524]. stabilized [1395, 1054]. stabilizers [410]. Stabilizing [983]. stable [1329, 1330, 1396, 572, 310, 298, 1062]. Stackelberg [1416]. stage [1600, 441, 337, 1624, 1127, 1790]. stage-structure [1127]. stages [1459]. stagnation [981]. stand [1537]. stand-alone [1537]. standard [1445, 1018, 347, 920]. standards [1890]. Standby [368]. standpoint [368]. standstill [179]. star [1481, 279, 416, 13]. STAR-GARCH [279]. starter [1271]. state [477, 871, 204, 126, 787, 1460, 1664, 1666, 1605, 1299, 435, 423, 1128]. state-dependent [787, 423]. state-space [477]. states [1329, 1330, 496, 348, 511]. Static [767, 654, 1583, 1737, 1509, 1722]. station [633]. stationarity [1356]. Stationary [1331, 210, 1403, 419, 112, 24, 304, 1054]. statistic [331]. Statistical [964, 1142, 1799, 418, 1776, 368, 1704]. statistics [156, 894, 851, 1109, 1140, 1746]. stator [692, 698, 176, 1302, 1270]. steadily [494]. steady [408, 126]. steel [910, 88]. Stefan [885]. Steiner [922]. stem [330]. stems [1155]. stenosed [1354, 1686]. step [1092, 310, 248, 1591, 1062, 1342, 1844]. step-size [1092]. stepping [461]. sticker [1538]. stiff [1092, 485, 1625, 310, 1460]. stiffness [1757]. Stochastic [314, 751, 809, 1159, 1482, 1348, 1199, 1487, 1094, 1760, 854, 824, 533, 1864, 1667, 937, 900, 902, 767, 1312, 418, 929, 909, 834, 390, 546, 261, 450, 1480, 1460, 769, 536, 1498, 1590, 1642, 583, 1674, 1850, 304]. stochasticity [1012]. stock [778, 780, 323, 112, 1787, 1517, 1866, 374, 274, 1084, 425]. stock-recruitment [1517]. stocks [614]. Stokes [1403, 1513, 1527, 1875, 534, 1054].

Storage [1709, 181, 333, 1735, 568, 714, 1697, 1288, 1537, 713]. straight [1418]. strain [510, 1234, 1186]. strains [82]. stranded [1272]. strategies [1540, 1052, 774, 1848, 1550, 769, 1359, 1276, 7, 178, 961]. strategy [318, 1722, 1678, 1379, 530]. Stratified [1478, 1368, 1399, 1159, 1237, 1705]. Stravinsky [33]. strength [548, 1378, 419]. stress [548, 1378, 510, 45]. strictly [1510]. Strong [113, 1624, 1381, 830, 1795]. strongly [1836]. Structural [1724, 758, 1081, 112, 276, 645, 238, 1007]. Structure [110, 810, 1514, 1624, 547, 1518, 1127, 1790, 1301, 747, 54]. structured [330, 1329, 1330, 1340, 455]. structures [1204, 951, 76, 1870, 1206, 1793, 1233, 1111, 515, 1859]. structuring [1507]. Student [274]. Studies [141, 874]. Study [1475, 1663, 180, 1024, 1091, 200, 1874, 1393, 216, 1551, 605, 1680, 1356, 660, 726, 1108, 817, 506, 1714, 1124, 886, 351, 721, 1883, 1729, 1554, 1555, 397, 1885, 1698, 650, 476, 1249]. Studying [509]. sub [1507, 771]. sub-model [771]. sub-structuring [1507]. subdifferential [1667]. subdiffusion [1438]. subdivision [335, 210, 592, 353, 1556, 952]. subject [458, 1605, 792, 1746]. submarine [1071]. submersible [605]. subspace [1415]. substation [1716]. substations [1266]. substitutable [1635]. substructuring [560]. subsurface [897, 1106, 568]. Subtraction [437]. subway [187]. success [1145]. sufficient [1008]. sum [1773]. summability [345]. sun [710, 670]. sun-tracking [710]. sunflower [476]. super [488, 859, 1605]. super-critical [488]. supercompact [342]. superconducting [701, 515]. superconductors [511]. Superconvergent [1391, 1090, 1820, 1096]. supersonic [165]. superstable [158]. supervised [102]. supervision [174, 1266, 713, 178]. supervisor [715]. supplier [584, 1687]. supply [846, 1551, 476]. support [1887, 643, 645]. supporting [70]. suppress [874, 644]. Suppressing [1244]. Suppression [487, 1168]. sure [1590]. surface [484, 485, 573, 699, 338, 602, 258, 1591, 1008, 79, 897, 910, 677, 379, 296, 209, 1261]. surfaces [1546, 343, 342, 856, 893, 898, 794]. surge [1322]. surgical [943]. surrender [824]. Surrogate [1823]. surrounded [1473]. surrounding [100, 1111, 1418]. survey [752, 430, 240, 713]. surveys [358]. survival [260]. sustainability [280]. SVE [601]. swarm [1658]. swell [505]. switched [189, 1729, 792]. switches [1740]. switching [172, 720, 368, 834, 1044, 528, 1010, 1177, 1842]. switchings [1642]. Symmetric [768, 1497, 507, 511, 35, 1115]. symmetries [1689]. symmetrized [1468]. symmetry [1437, 812]. symplectic [768]. Synchronisation [810]. Synchronization [140, 452, 595, 137, 1634, 314, 1044, 883, 1152, 1666, 680, 1645, 931, 647, 470, 299, 475]. synchronous [685, 1284, 1719, 1297, 176, 188, 932, 185, 1274, 179, 1724, 1628, 883, 1302, 689, 688, 1738, 697, 1718]. syncytial [11]. syndrome [943, 320]. synthesis [646]. System [1841, 181, 486, 695, 124, 1734, 1846, 171, 1037, 1526, 1046, 1691, 1132, 1377, 1277, 605, 722, 385, 497, 1634, 101, 616, 1735, 368, 1044, 757, 1282, 636, 764, 1316, 1572, 833, 9, 886, 1381, 1458, 191, 140, 1810, 1883, 1738, 749, 1003, 1382, 1418, 1689, 410, 398, 670, 1319, 545, 1295, 1882, 832, 423, 710, 1730, 460, 1713, 2, 717, 1310, 1419, 1797, 1083, 141, 1809, 1320, 1123, 1537, 110, 934, 1145, 1795, 766, 1813, 178]. system-on-chip [1730]. systematic [367]. systems [846, 469, 1536, 788, 1840, 928, 457, 337, 803, 1165, 895, 715, 786, 448, 1602, 1309, 1811, 310, 137, 1370, 1007, 1497, 239, 247, 430, 97, 1425, 663, 392, 1257, 1530, 929, 466, 732, 399, 962, 1062, 1735, 1677, 1655, 1307, 434, 217, 1129, 1657, 241, 1291, 509, 813, 1664, 1432, 375, 138, 1262, 812, 680, 721, 528, 243, 632, 1523, 1791, 1011, 1692, 795, 1029, 1493, 1863, 14, 624, 1844, 792, 402, 1241, 1561, 1194, 931, 1642, 1243, 1730, 470, 650, 299, 374, 1009, 1349, 999, 321, 1746, 111, 1659, 1716, 1524, 1709, 174, 238, 32, 1723, 1259]. Systems [1665, 713].

T [1225, 27, 931, 542, 651]. T-cells [542]. T-S [651]. T-shaped [1225]. table [751]. tactile [1300]. TAHA [1868]. tail [748, 671]. tailed [1084, 1128]. tails [1703]. Taiwan [288, 416]. Takeoff [981]. taking [1302, 398]. tandem [457]. tangential [148]. tangentially [388]. Tanner [1456]. target [1189, 315]. targeting [718]. targets [980]. Taylor [1445, 487, 1223]. TB [1847]. Team [1868]. technical [1292, 1731]. technique [1166, 1156, 828, 704, 1879]. techniques [216, 411, 601, 903, 1108, 30, 256, 1653, 227, 694]. Techno [1720, 1716]. Techno-economic [1716]. Techno-economical [1720]. technologies [1719, 979]. Technology [1797]. Telegrapher [1681]. temperature [1697, 1721, 1533]. temporal [373, 1308, 1195, 217, 1081]. Temporomandibular [40]. tend [947]. tension [485, 602, 664, 268]. tensor [1549, 1515, 1426]. tensor-product [1515]. tenth [1564]. tenth-order [1564]. term [770, 1654, 1053, 747, 1564, 386, 110, 378]. term-structure [747]. terminal [921, 632]. terminals [1726]. terms [1387, 937, 900, 902, 1355]. Ternary [952, 1556]. terrain [353]. terrorist [1861]. tessellations [826]. test [257, 186, 1277, 894, 1776, 112, 13, 1140, 425]. test-statistics [1140]. Testing [613, 1829, 331, 758, 851, 1143]. tests [1059, 154, 271, 1356, 660, 1776, 179, 1282, 1406, 253, 662, 354]. tetrahedral [344]. textile [1599]. th [1691, 1376, 1012]. their [1475, 1098, 238, 1376, 671, 984, 1472, 1863, 1027, 54, 576, 403]. theorem [922, 410]. theoretic [290, 1333]. Theoretical [449, 1793, 1015]. Theory [1868, 1333, 1228, 1787, 1185, 782, 1683, 1220, 786]. therapy [1836, 1604, 1501]. There [946, 257]. Thermal [1271, 1733, 1101, 1511, 1869, 700, 131, 899]. Thermo [1495, 461, 71]. thermo-elastodynamics [461]. Thermo-mechanical [1495]. thermodynamic [438, 54, 65]. thermodynamically [564]. thermoelastic [1884]. thermoviscous [1223, 1240]. theta [1242]. thick [654, 1812]. thickness [56]. thin [634, 1473]. thinking [1307]. thinning [1756]. third [1394, 495, 1844]. third-derivative [1844]. Thomas [544]. Three [258, 1437, 1854, 418, 354, 837, 1573, 1445, 366, 565, 973, 627, 605, 317, 385, 1776, 40, 1316, 182, 561, 794, 1741, 356, 1382, 1481, 724, 110]. Three-dimensional [258, 1437, 1854, 354, 366, 605, 40, 1316, 1382, 110]. three-level [1776]. three-order [385]. three-phase [837, 1741, 724]. three-species [627]. threshold [892, 1498, 1320]. thresholds [7]. tidal [823]. tide [1322]. tied [1712]. ties [257]. Tight [353]. Tikhonov [1571]. Time [1094, 600, 233, 1787, 921, 918, 232, 928, 720, 808, 1395, 1719, 862, 1702, 1184, 1297, 1646, 1876, 1438, 1046, 186, 1366, 1691, 869, 1132, 1870, 965, 74, 461, 1723, 1786, 732, 1229, 1735, 1405, 636, 348, 1067, 1110, 1460, 1473, 1281, 144, 751, 9, 1381, 591, 367, 1759, 975, 1653, 622, 300, 105, 801, 1112, 44, 103, 603, 1645, 1382, 1877, 284, 544, 1456, 1560, 1656, 272, 1882, 1215, 1730, 1778, 1713, 1879, 229, 1419, 1679, 1603, 638, 542, 111, 303, 735, 1459, 825, 1249, 1524, 530, 157, 1650]. time-accelerated [1713]. time-delay [1184, 111]. time-delays [732]. Time-dependent [233, 1395, 74, 1112, 303]. time-domain [1870, 965]. time-fractional [1702, 1877, 1879, 1603]. time-periodic [1650]. Time-scale [600, 1735, 44]. time-simultaneous [751]. Time-splitting [1094, 233]. time-stepping [461]. time-varying [1646, 1067, 144, 591, 367, 1645, 1656, 229, 735]. Timed [435]. times [1652]. Timestepping [800]. Timoshenko [1036]. TiO [1686]. tire [1647, 1495]. tissue [100, 1533]. tissues [43, 1418]. tolerant [1296, 1684, 1297, 1295]. tomato [1874]. tomography [835]. tool [172, 1366, 702, 433, 65]. Tools [404, 472]. Topology [1548, 1289, 321]. Torque [1274, 699, 685, 185, 1628, 845, 1270, 1718]. tortuosity [1859]. torus [1484]. Torvik [1314]. total [836, 1712, 1528]. totally [1271]. tourism [281, 277]. Tow [337]. Toxoplasmosis [1496]. trabecular [857]. traces [1113]. tracker [1691]. tracking [390, 711, 710, 315, 1145]. tract [1754]. traction [181]. trader [746]. trading [650]. Traffic [1668, 921, 918, 565, 1876, 1648, 1649, 147]. traffic-induced [147]. train [1732, 1569]. training [442]. trajectories [1869]. trajectory [771]. transfer [1036, 1614, 882, 607, 1885]. transform [493, 182]. transformation [370, 279, 566, 155, 1704]. transformations [1548]. transformer [1293]. Transient [173, 634, 1854, 354, 534, 585]. transients [172]. transistors [1108, 892]. transition [972, 748, 1888]. transitive [671]. transmission [11, 318, 1615, 311]. transmissions [1338]. transmitted [1862]. transonic [75]. transparent [948]. transport [1387, 1390, 1819, 336, 338, 125, 1401, 346, 128, 567, 80, 679, 1570, 1409, 132, 1249]. transposed [1275]. trap [1332]. trap-insect [1332]. Trapezoidal [1015]. Traub [1125]. travel [1653, 755]. Traveling [1529, 1175, 1223, 180, 497, 811, 512]. Travelling [245, 1329]. treating [1423]. treatment [875, 1101, 1808, 1013, 437, 1701, 1183, 42, 1151, 86, 1157, 1533, 1379, 1578]. tree [1345, 1538]. trees [286]. trend [304]. triangle [592, 947, 606]. triangular [148]. triangulation [859]. triangulations [1396]. triggered [1645]. trigonometric [664]. Trigonometrically [248]. trinomial [257]. triopoly [1311]. triplet [1807]. triplet-wise [1807]. trisection [947, 606]. Trivariate [1116, 344]. Troesch [1788]. truck [1717]. true [703]. Truncation [1775]. Tuberculosis [1847]. tumor [100, 1, 1490, 527, 41, 1533]. tumors [1604, 1418]. tumour [996]. Tuning [1813, 1001, 1774]. tunnel [406]. turbine [177, 1673, 711, 191, 712]. turbulence [899, 508, 855, 1754]. turbulent [1627, 1505]. Turing [1846, 497, 1000]. Tweedie [1829]. Twister [876, 1777]. twisting [1605]. Two [1684, 1217, 466, 1655, 182, 1216, 882, 1506, 1699, 1389, 1390, 788, 1338, 862, 448, 310, 248, 1591, 242, 1377, 1675, 1356, 260, 75, 40, 1138, 1406, 885, 350, 1853, 507, 568, 1458, 43, 679, 1127, 623, 1759, 794, 1810, 244, 44, 356, 1844, 843, 1483, 1276, 1409, 562, 541, 1310, 1417, 1186, 1459, 110, 1795, 1054, 378]. two-delay [244]. Two-dimensional [1217, 40, 679, 1483, 378]. two-grid [1458]. two-layer [1138, 507]. Two-level [1216, 1054]. two-output [1276]. two-parameter [242, 260]. two-patch [1338]. two-phase [1389, 1390, 1675, 75, 885, 350, 1853, 568, 356, 1409]. two-prey [1310]. two-sample [1406]. two-scale [43]. two-stage [1127]. two-step [310, 248, 1591, 1844]. two-strain [1186]. two-term [110]. twofold [633]. type [1175, 1700, 1311, 1393, 803, 1733, 254, 950, 1037, 494, 81, 1667, 1139, 418, 1474, 564, 1067, 1799, 421, 144, 1790, 1140, 1844, 515, 516, 1580, 936, 999, 1143, 649, 735, 1324, 12]. type-II [1393, 1799, 515]. type-IV [1700]. types [1389, 692]. Typical [973]. Tyson [817].

U.S. [780]. UK [300]. Ultimate [1634]. ultimately [390]. ultracapacitors [1285]. ultrahigh [694]. ultrasonic [192]. unbalanced [1794, 189, 1722]. unbounded [895, 146]. uncertain [1646, 319, 1677, 138, 1666, 14, 624, 8, 1656, 1866]. uncertainties [1423, 1657, 468, 1746]. Uncertainty [86, 997, 1691, 861, 1682, 118, 980]. unconstrained [1139, 322]. under-frequency [1536]. underestimator [1166]. understanding [957]. underwater [143]. underwriters [750]. Unfitted [336]. Unified [704, 1423, 1068, 1129, 795, 1319]. Uniform [664, 331, 70, 851]. Uniformly [1396, 390, 1753]. unilateral [801]. unique [577, 1316]. unit [271, 23, 1069, 425]. units [1797]. unity [397]. univariate [1115, 1144]. Universal [712, 1349]. university [1636]. unknown [271, 1457, 1062, 1034, 635, 1643]. unmatched [1605]. unmeasured [348]. unreliable [368]. unsaturated [991]. Unscented [314]. unstable [1062]. Unsteady [439, 945, 603, 90, 1689, 1686]. Unstructured [942, 615, 339]. unsupervised [1314]. unwinding [693]. updates [743]. updating [272]. upon [825]. upper [1142, 259]. upscaled [1399]. upscaling [1402]. Uranus [670]. urban [419]. urethra [37]. Use [1714, 645, 124, 215, 246]. used [100, 1060, 527, 1710, 1185]. user [1648]. Users [1461]. Using [1416, 459, 1836, 1424, 1199, 34, 918, 290, 748, 543, 19, 1332, 942, 487, 1287, 1313, 1719, 531, 1297, 1188, 1646, 993, 279, 774, 1323, 1309, 943, 493, 1130, 1333, 186, 1146, 1725, 205, 1227, 1400, 1870, 1581, 767, 343, 344, 856, 1824, 903, 1887, 797, 1257, 407, 1644, 418, 1723, 1141, 466, 1776, 314, 1827, 1628, 1762, 777, 233, 434, 419, 182, 1252, 1355, 509, 1682, 138, 555, 930, 688, 922, 1299, 1435, 1867, 103, 190, 226, 1322, 275, 828, 1689, 1408, 704, 410, 1699, 398, 919, 442, 376, 1436, 923, 1319, 1839, 1499, 1345]. using [832, 1730, 173, 637, 425, 361, 1747, 1850, 1569, 1250, 1128]. Utilizing [1653].

UV [1841].

vaccination [455, 875, 1792, 1808, 318, 1701, 1498, 816, 1459, 1379]. Vague [1839]. Validation [1843, 1719, 1399, 1286]. Vallée [1562]. valley [716]. valuation [824]. Value [285, 798, 779, 329, 1445, 1018, 319, 1370, 777, 273, 1844, 1539, 782, 1865, 1860, 162, 283, 399, 775, 453, 781]. Value-at-Risk [285, 775, 453, 781]. valued [1667, 919, 1679]. values [1093, 1142, 242, 1776, 259]. Valuing [1174]. VAR [276]. variability [1169, 282]. Variable [471, 290, 1392, 1092, 334, 935, 1102, 1786, 98, 637, 361]. variable-order [1786]. variables [1806, 1779, 230, 1773, 1042, 1642, 650, 1674]. variance [1478, 253, 270, 467, 998]. variate [459, 1775, 460]. variates [225]. variation [836, 1780]. Variational [344, 1092, 1593, 1888, 1051, 898, 1185, 1674]. various [70]. Varley [764]. VARMAX [477]. varying [1646, 823, 1067, 144, 591, 367, 1645, 1656, 1528, 1882, 229, 735, 1496]. vector [1175, 490, 1887, 690, 1615, 694, 835, 1852, 744, 643, 645]. vector-bias [1615]. vectors [114]. vegetatively [961]. vehicle [1684, 1714, 697, 793, 1651, 141, 645]. vehicles [1292, 143]. velocity [491, 458, 772, 838, 1594]. verification [687, 1554, 1555]. versus [1476]. vertex [55, 923]. vertical [318, 893]. VHDL [721, 704]. VHDL-AMS [704]. VI [1386]. VI-2015 [1386]. via [776, 1502, 1752, 948, 370, 472, 345, 671, 230, 1657, 1701, 1206, 1664, 367, 1666, 452, 1035, 1562, 1852, 1860, 475, 1145, 1524]. vibrating [1885]. Vibration [1729, 634, 189]. vibrations [147, 83]. vibro [1112]. vibro-impact [1112]. videos [1873]. view [185, 1300, 1841]. VII [1818, 207]. VII-2017 [1818]. viral [1341, 142, 1189]. virus [11, 1341, 814, 1862, 1578, 809, 1698, 676, 1417]. viscoelastic [634, 1756, 1870, 1812, 510, 130]. viscosity [1449]. viscous [615, 1604, 1397, 1214, 1437, 1800, 420, 514]. visibility [948]. Visualization [404]. vivo [38]. VIX [775]. volatile [1886]. Volatility [778, 267, 288, 287, 767, 909, 1872, 269, 10, 117, 105, 1144, 583, 1128]. Voltage [1711, 172, 1535, 1296, 1719, 171, 1715, 892, 1827, 714, 1731, 1741, 1294, 723, 178]. Volterra [1017, 1019, 831]. volume [1387, 1819, 333, 95, 338, 339, 863, 1095, 862, 1308, 944, 547, 1458, 679, 1531, 130, 417, 1843]. Voronoi [826]. Vortex [511, 409, 515, 1564]. vortices [504]. Voter [374]. vs [1655, 575, 981].

Wald [1140]. walk [1366, 575, 1471, 1479]. walks [1371]. wall [1754]. walls [388]. ward [1069]. waste [333]. wastewater [437]. water [338, 863, 1095, 439, 567, 1138, 1572, 1793, 296, 603, 1322, 476]. Wave [1220, 488, 512, 1329, 1024, 1175, 573, 124, 490, 492, 180, 994, 811, 1230, 507, 508, 509, 1883, 1601, 440, 98, 518, 1529]. wave-propagation [492]. waveform [589]. waveforms [685]. waveguides [233]. Wavelet [377, 1837, 1849, 1867, 451, 1809]. Wavelet-based [377]. wavelet-finite [1837]. wavelets [1313, 1396, 664]. waves [484, 245, 491, 1223, 1226, 506, 507, 1344, 297, 1240, 516, 520]. Weak [1487, 1774, 1042, 1478, 1476, 10]. weakening [1302, 845]. weakly [1197, 360]. wear [392]. web [757]. web-based [757]. Weibull [156, 1676, 317, 1500, 1318, 1799, 259, 737]. weight [1376, 671, 719, 1608]. weight-spectra [671]. Weighted [949, 1562, 260, 910, 460, 408]. Weighted-Residuals [408]. weighting [1645]. weights [1486]. well [558, 354]. well-balanced [558]. well-tests [354]. wells [1600, 1160]. WENO [1108]. WENO-based [1108]. Westervelt [514]. wetland [1358, 1618]. wheels [1558]. where [350]. white [351]. Whitney [70]. whose [1449, 1763]. wide [1648]. width [1296]. Wigner [963]. Wind [1535, 177, 1536, 1284, 1736, 1583, 1673, 1720, 711, 838, 191, 1738, 700, 406, 712, 1713, 1537]. winding [686, 693, 690, 1276]. windings [189]. window [1320]. wing [874]. wise [1807]. WITHDRAWN [902]. within [789, 772, 1798, 1417]. within-host [1417]. without [808, 703, 371]. workforce [978]. working [712]. works [1619]. world [600]. wormlike [1225]. Worsey [344]. worst [1775]. WTO [288].

XIV [1544].

YBa [62]. yearly [453]. yeast [817]. yield [1258]. yoke [176]. Yokeless [1273].

Zabusky [493]. Zealand [273]. Zero [1602, 1376, 947]. Zero-Hopf [1602]. Zeya [825]. Zhang [1145]. Zhang-gradient [1145]. Zhou [995]. Zika [1862, 1805, 1698]. zones [448, 1450].

References

BerguesCabrales:2008:MMT

[1] Luis Enrique Bergues Cabrales, Andrés Ramírez Aguilera, Rolando Placeres Jiménez, Manuel Verdecia Jarque, Héctor Manuel Camué Ciria, Juan Bory Reyes, Miguel Angel O’Farril Mateus, Fabiola Suárez Palencia, and Marisela González Ávila. Mathematical modeling of tumor growth in mice following low-level direct electric current. Mathematics and Computers in Simulation, 78(1):112– 120, June 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                              /www.sciencedirect.com/science/

article/pii/S037847540700208X. See

corrigendum [1490].

Wang:2009:LSP

[2] Lichun Wang and Xuan Wang. The life-span prediction of a system connected in series. Mathematics and Computers in Simulation, 79(5):1770– 1777, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408003157. See

comments [833].

Kung:2009:OPU

[3] James J. Kung and Lung-Sheng Lee. Option pricing under the Merton model of the short rate. Mathematics and Computers in Simulation, 80 (2):378–386, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002341. See

comment [153].

Anonymous:2010:PJa

[4] Anonymous. Pages 887–1044 (January 2010). Mathematics and Computers in Simulation, 80(5):??, January 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2010:EBa

[5] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 80(5):CO2, January 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000091.

Kalayathankal:2010:FSF

[6] Sunny Joseph Kalayathankal and G. Suresh Singh. A fuzzy soft flood alarm model. Mathematics and Computers in Simulation, 80(5): 887–893, January 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540900319X.

Tang:2010:DAP

[7] Sanyi Tang, Yanni Xiao, and Robert A. Cheke. Dynamical analysis of plant disease models with cultural control strategies and economic thresholds. Mathematics and Computers in Simulation, 80(5):894–921, January 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409003206.

Qin:2010:MDE

[8] Rui Qin and Yan-Kui Liu. Modeling data envelopment analysis by chance method in hybrid uncertain environments. Mathematics and Computers in Simulation, 80(5):922– 950, January 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409003218.

Liu:2010:RSL

[9] Leipo Liu, Zhengzhi Han, Xiushan Cai, and Jun Huang. Robust stabilization of linear differential inclusion system with time delay. Mathematics and Computers in Simulation, 80 (5):951–958, January 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540900322X.

Mackevicius:2010:WAC

[10] Vigirdas Mackevičius. On weak approximations of CIR equation with high volatility. Mathematics and Computers in Simulation, 80(5):959– 970, January 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409003358.

Arenas:2010:RMM

[11] Abraham J. Arenas, Gilberto González-Parra, and Lucas Jódar. Randomness in a mathematical model for the transmission of respiratory syncytial virus (RSV). Mathematics and Computers in Simulation, 80(5): 971–981, January 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409003553.

Zuoxiang:2010:AEL

[12] Peng Zuoxiang, Liu Miaomiao, and Saralees Nadarajah. Asymptotic expansions for the location invariant moment-type estimator. Mathematics and Computers in Simulation, 80 (5):982–998, January 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409003565.

Maki:2010:APT

[13] Daiki Maki. An alternative procedure to test for cointegration in STAR models. Mathematics and Computers in Simulation, 80(5):999– 1006, January 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409003577.

Qian:2010:LCR

[14] Wei Qian, Juan Liu, Youxian Sun, and Shumin Fei. A less conservative robust stability criteria for uncertain neutral systems with mixed delays. Mathematics and Computers in Simulation, 80 (5):1007–1017, January 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409003619.

Mei:2010:CAN

[15] Shan Mei, P. M. A. Sloot, Rick Quax, Yifan Zhu, and Weiping Wang. Complex agent networks explaining the HIV epidemic among homosexual men in Amsterdam. Mathematics and Computers in Simulation, 80 (5):1018–1030, January 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409003620.

Anonymous:2010:NIa

[16] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 80(5):1031–1032, January 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475410000133.

Anonymous:2010:ICEa

[17] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 80(5):1033– 1038, January 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410000145.

Anonymous:2010:EBb

[18] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 80(6):CO2, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000273.

Andrei:2010:AMA

[19] Petru Andrei and Ayodeji Adedoyin. Analysis of magnetic aftereffect by using Monte Carlo methods. Mathematics and Computers in Simulation, 80(6):1045–1055, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002869.

Bazavov:2010:ABM

[20] Alexei Bazavov, Bernd A. Berg, and Huan-Xiang Zhou. Application of Biased Metropolis Algorithms: From protons to proteins. Mathematics and Computers in Simulation, 80(6): 1056–1067, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409001475.

Calvin:2010:RAG

[21] James M. Calvin. Randomized algorithm for global optimization with bounded memory. Mathematics and Computers in Simulation, 80(6):1068– 1081, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408003406.

Horntrop:2010:CEM

[22] David J. Horntrop. Concentration effects in mesoscopic simulation of coarsening. Mathematics and Computers in Simulation, 80(6):1082– 1088, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409003188.

Hwang:2010:MCM

[23] Chi-Ok Hwang, Michael Mascagni, and Taeyoung Won. Monte Carlo methods for computing the capacitance of the unit cube. Mathematics and Computers in Simulation, 80(6): 1089–1095, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408001274.

Panneton:2010:RSR

[24] François Panneton and Pierre L’Ecuyer. Resolution-stationary random number generators. Mathematics and Computers in Simulation, 80(6):1096– 1103, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002625.

Pillards:2010:AAC

[25] Tim Pillards, Bart Vandewoestyne, and Ronald Cools. An adaptive approach to cube-based quasi-Monte Carlo integration on R^{d}. Mathematics and Computers in Simulation, 80(6): 1104–1117, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000032.

Rasulov:2010:MCS

[26] A. Rasulov, G. Raimova, and M. Mascagni. Monte Carlo solution of Cauchy problem for a nonlinear parabolic equation. Mathematics and Computers in Simulation, 80(6): 1118–1123, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409003632.

Schurer:2010:AAN

[27] Rudolf Schürer and Wolfgang Ch. Schmid. M in T — Architecture and applications of the ( t, m, s )- net and OOA database. Mathematics and Computers in Simulation, 80(6): 1124–1132, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002637.

Srinivasan:2010:MCL

[28] A. Srinivasan. Monte Carlo linear solvers with non-diagonal splitting. Mathematics and Computers in Simulation, 80(6):1133–1143, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409000962.

Roanes-Lozano:2010:FSI

[29] Eugenio Roanes-Lozano, Michael J. Wester, and Stanly Steinberg. Foreword to the special issue on “Nonstandard Applications of Computer Algebra”. Mathematics and Computers in Simulation, 80(6):1145, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410000194.

Lewis:2010:CAT

[30] Robert H. Lewis. Comparing acceleration techniques for the Dixon and Macaulay resultants. Mathematics and Computers in Simulation, 80(6): 1146–1152, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408001742.

Roanes-Lozano:2010:CDC

[31] Eugenio Roanes-Lozano, Nicolas van Labeke, and Eugenio Roanes-Macías. Connecting the 3D DGS Calques3D with the CAS Maple. Mathematics and Computers in Simulation, 80(6): 1153–1176, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002961.

Escribano:2010:ARC

[32] Jesús Escribano, Francisco Botana, and Miguel A. Abánades. Adding remote computational capabilities to Dynamic Geometry Systems. Mathematics and Computers in Simulation, 80(6): 1177–1184, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540800178X.

Lombardi:2010:SMS

[33] Paul Lombardi and Michael J. Wester. Serial mistakes in Stravinsky’s Requiem Canticles. Mathematics and Computers in Simulation, 80(6):1185– 1199, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002584.

Aguilera:2010:AGC

[34] Gabriel Aguilera, José Luis Galán, Rafael Madrid, Antonio Manuel Martínez, Yolanda Padilla, and Pedro Rodríguez. Automated generation of contrapuntal musical compositions using probabilistic logic in Derive. Mathematics and Computers in Simulation, 80(6):1200– 1211, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001165.

Pletsch:2010:RPP

[35] Bill Pletsch. A recursion of the Pólya polynomial for the symmetric group. Mathematics and Computers in Simulation, 80(6):1212– 1220, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408001791.

Danek:2010:P

[36] Josef Daněk and Eduard Rohan. Preface. Mathematics and Computers in Simulation, 80(6):1221, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410000200.

Brandner:2010:ARS

[37] Marek Brandner, Jiří Egermaier, and Hana Kopincová. Augmented Riemann solver for urethra flow modeling. Mathematics and Computers in Simulation, 80(6):1222–1231, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409002420.

Cimrman:2010:IAM

[38] Robert Cimrman and Eduard Rohan. On identification of the arterial model parameters from experiments applicable “in vivo”. Mathematics and Computers in Simulation, 80(6): 1232–1245, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000408.

Cihalova:2010:MFF

[39] Lenka Číhalová. The muscle fatigue fiber model. Mathematics and Computers in Simulation, 80(6):1246– 1255, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002286.

Hlinakova:2010:TJT

[40] P. Hliňáková, T. Dostálová, J. Daněk, J. Nedoma, and I. Hlaváček. Temporomandibular joint and its twodimensional and three-dimensional modelling. Mathematics and Computers in Simulation, 80(6):1256– 1268, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002432.

Mahmood:2010:NAA

[41] Mohammed Shuker Mahmood, Silvia Mahmood, and Dušan Dobrota. A numerical algorithm for avascular tumor growth model. Mathematics and Computers in Simulation, 80(6): 1269–1277, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409003127.

Marsik:2010:RLB

[42] František Maršík, Václav Klika, and Hynek Chlup. Remodelling of living bone induced by dynamic loading and drug delivery — Numerical modelling and clinical treatment. Mathematics and Computers in Simulation, 80(6): 1278–1288, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000664.

Lukes:2010:MBT

[43] Vladimír Lukeš and Eduard Rohan. Microstructure based two-scale modelling of soft tissues. Mathematics and Computers in Simulation, 80(6): 1289–1301, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000688.

Papacek:2010:EDP

[44] Štěpán Papáček, Sergej Čelikovský, Branislav Rehák, and Dalibor Štys. Experimental design for parameter estimation of two time-scale model of photosynthesis and photoinhibition in microalgae. Mathematics and Computers in Simulation, 80(6): 1302–1309, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002122.

Sistek:2010:BFS

[45] Jakub Šístek, Jaroslav Novotný, Jan Mandel, Marta Čertíková, and Pavel Burda. BDDC by a frontal solver and the stress computation in a hip joint replacement. Mathematics and Computers in Simulation, 80(6):1310– 1323, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409000044.

Vimmr:2010:NNE

[46] Jan Vimmr and Alena Jonášová. Non-Newtonian effects of blood flow in complete coronary and femoral bypasses. Mathematics and Computers in Simulation, 80(6):1324– 1336, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409000068.

Vychytil:2010:SMC

[47] Jan Vychytil and Miroslav Holeček. The simple model of cell prestress maintained by cell incompressibility. Mathematics and Computers in Simulation, 80(6):1337–1344, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409000391.

Anonymous:2010:NIb

[48] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 80(6):1345–1346, February 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475410000339.

Anonymous:2010:ICEb

[49] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 80(6):1347– 1352, February 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410000340.

Anonymous:2010:EBc

[50] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 80(7):CO2, March 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410000546.

Moelans:2010:E

[51] Nele Moelans and Giovanni Samaey. Editorial. Mathematics and Computers in Simulation, 80(7):1359– 1360, March 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410000698.

Janssens:2010:IRC

[52] K. G. F. Janssens. An introductory review of cellular automata modeling of moving grain boundaries in polycrystalline materials. Mathematics and Computers in Simulation, 80 (7):1361–1381, March 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409000470.

Hoyt:2010:FMD

[53] J. J. Hoyt, Z. T. Trautt, and M. Upmanyu. Fluctuations in molecular dynamics simulations. Mathematics and Computers in Simulation, 80(7): 1382–1392, March 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001001.

VanderVen:2010:LES

[54] A. Van der Ven, J. C. Thomas, Qingchuan Xu, and J. Bhattacharya. Linking the electronic structure of solids to their thermodynamic and kinetic properties. Mathematics and Computers in Simulation, 80(7):1393– 1410, March 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002456.

Mora:2010:VMS

[55] L. A. Barrales Mora. 2D vertex modeling for the simulation of grain growth and related phenomena. Mathematics and Computers in Simulation, 80 (7):1411–1427, March 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002407.

Selzer:2010:IPD

[56] Michael Selzer, Britta Nestler, and Denis Danilov. Influence of the phase diagram on the diffuse interface thickness and on the microstructure formation in a phase-field model for binary alloy. Mathematics and Computers in Simulation, 80(7):1428– 1437, March 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410000054.

Vanherpe:2010:MSP

[57] Liesbeth Vanherpe, Frank Wendler, Britta Nestler, and Stefan Vandewalle. A multigrid solver for phase field simulation of microstructure evolution. Mathematics and Computers in Simulation, 80(7):1438– 1448, March 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540900336X.

Hausser:2010:IEF

[58] Frank Haußer, Sandra Rasche, and Axel Voigt. The influence of electric fields on nanostructures — Simulation and control. Mathematics and Computers in Simulation, 80(7): 1449–1457, March 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409001657.

Nestler:2010:CLB

[59] Britta Nestler, Ali Aksi, and Michael Selzer. Combined Lattice Boltzmann and phase-field simulations for incompressible fluid flow in porous media. Mathematics and Computers in Simulation, 80(7):1458– 1468, March 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540900353X.

Luque:2010:GMC

[60] A. Luque, J. Aldazabal, J. M. Martínez-Esnaola, A. Martín-Meizoso, J. Gil Sevillano, and R. S. Farr.

Geometrical Monte Carlo model of liquid-phase sintering. Mathematics and Computers in Simulation, 80(7): 1469–1486, March 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409003152.

Pedersen:2010:DIA

[61] Andreas Pedersen and Hannes Jónsson. Distributed implementation of the adaptive kinetic Monte Carlo method. Mathematics and Computers in Simulation, 80(7):1487–1498, March 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475409000500.

Gomez:2010:FPC

[62] J. A. Gomez, I. Larkin, and U. Schwingenschlögl. First-principles calculations of the YBa_{2}Cu_{3}O_{7}/PrBa_{2}Cu_{3}O_{7} interface. Mathematics and Computers in Simulation, 80(7):1499– 1508, March 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001670.

Markidis:2010:MSS

[63] Stefano Markidis, Giovanni Lapenta, and Rizwan-uddin. Multi-scale simulations of plasma with iPIC3D. Mathematics and Computers in Simulation, 80(7):1509–1519, March 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002444.

Drapala:2010:CSD

[64] J. Drápala, P. Kubíček, and O. Vlach. Computer simulation of diffusion processes with moving interface boundary. Mathematics and Computers in Simulation, 80(7):1520– 1535, March 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410000066.

Miettinen:2010:ITK

[65] J. Miettinen, S. Louhenkilpi, H. Kytönen, and J. Laine. IDS: Thermodynamic– kinetic–empirical tool for modelling of solidification, microstructure and material properties. Mathematics and Computers in Simulation, 80(7): 1536–1550, March 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409003528.

Anonymous:2010:NIc

[66] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 80(7):1551, March 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410000613.

Anonymous:2010:ICEc

[67] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 80(7):1552– 1558, March 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410000625.

Anonymous:2010:EBd

[68] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 80(8):CO2, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000935.

Anonymous:2010:P

[69] Anonymous. Preface. Mathematics and Computers in Simulation, 80 (8):1565–1566, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001059.

Bossavit:2010:URW

[70] A. Bossavit. A uniform rationale for Whitney forms on various supporting shapes. Mathematics and Computers in Simulation, 80(8):1567– 1577, April 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408003807.

Kruis:2010:ECI

[71] Jaroslav Kruis, Tomáš Koudelka, and Tomáš Krejčí. Efficient computer implementation of coupled hydro-thermo-mechanical analysis. Mathematics and Computers in Simulation, 80(8): 1578–1588, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408003844.

Segeth:2010:RSP

[72] Karel Segeth. A review of some a posteriori error estimates for adaptive finite element methods. Mathematics and Computers in Simulation, 80 (8):1589–1600, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408004230.

Farago:2010:DMP

[73] István Faragó. Discrete maximum principle for finite element parabolic models in higher dimensions. Mathematics and Computers in Simulation, 80 (8):1601–1611, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000342.

Feistauer:2010:DGS

[74] M. Feistauer, V. Kučera, and J. Prokopová. Discontinuous Galerkin solution of compressible flow in time-dependent domains. Mathematics and Computers in Simulation, 80(8):1612– 1623, April 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540900038X.

Halama:2010:NMT

[75] Jan Halama, Fayssal Benkhaldoun, and Jaroslav Fořt. Numerical modeling of two-phase transonic flow. Mathematics and Computers in Simulation, 80 (8):1624–1635, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000421.

Dolezel:2010:IAS

[76] Ivo Doležel, Pavel Karban, Martina Donátová, and Pavel Šolín. Integrodifferential approach to solution of eddy currents in linear structures with motion. Mathematics and Computers in Simulation, 80(8): 1636–1646, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000457.

Rapetti:2010:OME

[77] Francesca Rapetti. An overlapping mortar element approach to coupled magneto-mechanical problems. Mathematics and Computers in Simulation, 80(8):1647–1656, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000469.

Banas:2010:EMP

[78] L’ubomír Baňas. An efficient multigrid preconditioner for Maxwell’s equations in micromagnetism. Mathematics and Computers in Simulation, 80 (8):1657–1663, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000494.

Do-Quang:2010:NSC

[79] Minh Do-Quang and Gustav Amberg. Numerical simulation of the coupling problems of a solid sphere impacting on a liquid free surface. Mathematics and Computers in Simulation, 80 (8):1664–1673, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000676.

Kuzmin:2010:GOP

[80] Dmitri Kuzmin and Sergey Korotov. Goal-oriented a posteriori error estimates for transport problems. Mathematics and Computers in Simulation, 80(8):1674–1683, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000755.

Dobrzynski:2010:FIN

[81] Michał Dobrzyński and Jagoda Plata. Fill-ins number reducing direct solver designed for FIT-type matrix. Mathematics and Computers in Simulation, 80(8):1684–1693, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001256.

Majorana:2010:AMC

[82] C. E. Majorana, V. A. Salomoni, G. Mazzucco, and G. A. Khoury. An approach for modelling concrete spalling in finite strains. Mathematics and Computers in Simulation, 80 (8):1694–1712, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001669.

Svacek:2010:ECF

[83] P. Sváček. On energy conservation for finite element approximation of flowinduced airfoil vibrations. Mathematics and Computers in Simulation, 80 (8):1713–1724, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001682.

Prokop:2010:NSN

[84] Vladimír Prokop and Karel Kozel. Numerical simulation of Newtonian and non-Newtonian flows in bypass. Mathematics and Computers in Simulation, 80(8):1725–1733, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001700.

Kozel:2010:NMH

[85] K. Kozel, P. Louda, and J. Příhoda. Numerical methods of higher order of accuracy for incompressible flows. Mathematics and Computers in Simulation, 80(8):1734– 1745, April 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002134.

Rahimi:2010:UTF

[86] Mohammad Rahimi and Saeed Ziaei-Rad. Uncertainty treatment in forced response calculation of mistuned bladed disk. Mathematics and Computers in Simulation, 80(8): 1746–1757, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409002298.

Vejchodsky:2010:DMP

[87] Tomáš Vejchodský, Sergey Korotov, and Antti Hannukainen. Discrete maximum principle for parabolic problems solved by prismatic finite elements. Mathematics and Computers in Simulation, 80(8):1758– 1770, April 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409003176.

Karban:2010:CIH

[88] Pavel Karban and Martina Donátová. Continual induction hardening of steel bodies. Mathematics and Computers in Simulation, 80(8):1771– 1782, April 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409003589.

Keslerova:2010:NMI

[89] Radka Keslerová and Karel Kozel. Numerical modelling of incompressible flows for Newtonian and non-Newtonian fluids. Mathematics and Computers in Simulation, 80(8):1783– 1794, April 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409003590.

Puncocharova-Porizkova:2010:NSU

[90] Petra Punčochářová-Pořízková, Jiří Fürst, Jaromír Horáček, and Karel Kozel. Numerical solutions of unsteady flows with low inlet Mach numbers. Mathematics and Computers in Simulation, 80(8):1795– 1805, April 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409003607.

Anonymous:2010:NId

[91] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 80(8):1806, April 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541000100X.

Anonymous:2010:ICEd

[92] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 80(8):1807– 1813, April 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001011.

Anonymous:2010:PM

[93] Anonymous. Pages 1821–1984 (May 2010). Mathematics and Computers in Simulation, 80(9):??, May 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2010:EBe

[94] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 80(9):CO2, May 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001278.

Bendahmane:2010:FVS

[95] Mostafa Bendahmane, Raimund Bürger, and Ricardo Ruiz-Baier. A finite volume scheme for cardiac propagation in media with isotropic conductivities. Mathematics and Computers in Simulation, 80(9):1821– 1840, May 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409003644.

Tavazoei:2010:PCR

[96] Mohammad Saleh Tavazoei and Mohammad Haeri. Periodic characteristic ratio (PCR) method: an alternative method to determine the characteristic polynomial. Mathematics and Computers in Simulation, 80 (9):1841–1853, May 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000029.

El-Sebakhy:2010:FRI

[97] Emad A. El-Sebakhy. Flow regimes identification and liquid-holdup prediction in horizontal multiphase flow based on neuro-fuzzy inference systems. Mathematics and Computers in Simulation, 80(9):1854– 1866, May 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410000030.

Triki:2010:SWS

[98] Houria Triki, Thiab R. Taha, and Abdul-Majid Wazwaz. Solitary wave solutions for a generalized KdV– mKdV equation with variable coefficients. Mathematics and Computers in Simulation, 80(9):1867– 1873, May 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410000248.

Kamalinejad:2010:CGL

[99] Howra Kamalinejad, Vahid Johari Majd, Hamed Kebriaei, and Ashkan Rahimi-Kian. Cournot games with linear regression expectations in oligopolistic markets. Mathematics and Computers in Simulation, 80 (9):1874–1885, May 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541000025X.

Aguilera:2010:ECD

[100] Andrés Ramírez Aguilera, Luis Enrique Bergues Cabrales, Héctor Manuel Camué Ciria, Yudelmis Soler Pérez, Fidel Gilart González, Maraelys Morales González, Lisset Ortíz Zamora, Fabiola Suárez Palencia, Miriam Fariñas Salas, Nailet Real Bestard, Gustavo Sierra González, and Idelisa Bergues Cabrales. Electric current density distribution in planar solid tumor and its surrounding healthy tissue generated by an electrode elliptic array used in electrotherapy. Mathematics and Computers in Simulation, 80 (9):1886–1902, May 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410000479.

Gong:2010:MSM

[101] Zhaohua Gong. A multistage system of microbial fed-batch fermentation and its parameter identification. Mathematics and Computers in Simulation, 80(9):1903–1910, May 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410000480.

Xue:2010:SSO

[102] Zhenxia Xue, Youlin Shang, and Aifen Feng. Semi-supervised outlier detection based on fuzzy rough C-means clustering. Mathematics and Computers in Simulation, 80 (9):1911–1921, May 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410000510.

Persson:2010:PAO

[103] Jonas Persson and Lina von Sydow. Pricing American options using a space-time adaptive finite difference method. Mathematics and Computers in Simulation, 80(9):1922– 1935, May 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410000522.

Mokdad:2010:FDM

[104] B. Mokdad, A. Ammar, M. Normandin, F. Chinesta, and J. R. Clermont. A fully deterministic micro–macro simulation of complex flows involving reversible network fluid models. Mathematics and Computers in Simulation, 80(9):1936– 1961, May 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541000073X.

Ogawa:2010:RTE

[105] Shigeyoshi Ogawa and Hoang-Long Ngo. Real-time estimation scheme for the spot cross volatility of jump diffusion processes. Mathematics and Computers in Simulation, 80(9):1962– 1976, May 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410000777.

Anonymous:2010:NIe

[106] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 80(9):1977, May 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001308.

Anonymous:2010:ICEe

[107] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 80(9):1978– 1984, May 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541000131X.

Anonymous:2010:PJb

[108] Anonymous. Pages 1985–2122 (June 2010). Mathematics and Computers in Simulation, 80(10):??, June 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2010:EBf

[109] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 80(10):CO2, June 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410001667.

Zhang:2010:SBS

[110] Xianxia Zhang, Shaoyuan Li, and Hanxiong Li. Structure and BIBO stability of a three-dimensional fuzzy two-term control system. Mathematics and Computers in Simulation, 80 (10):1985–2004, June 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410000728.

Yang:2010:RGC

[111] Dedong Yang and Kai-Yuan Cai. Reliable guaranteed cost sampling control for nonlinear time-delay systems. Mathematics and Computers in Simulation, 80(10):2005– 2018, June 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410000753.

Lu:2010:MRG

[112] Yang-Cheng Lu, Tsangyao Chang, Ken Hung, and Wen-Chi Liu. Mean reversion in G-7 stock prices: Further evidence from a panel stationary test with multiple structural breaks. Mathematics and Computers in Simulation, 80 (10):2019–2025, June 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000807.

Qin:2010:SCR

[113] Ruibing Qin, Zheng Tian, Hao Jin, and Xiaowei Zhang. Strong convergence rate of robust estimator of change point. Mathematics and Computers in Simulation, 80(10):2026– 2032, June 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410000820.

Kurita:2010:ESN

[114] Takamitsu Kurita. Effects of a signalto-noise ratio on finite sample inference for cointegrating vectors. Mathematics and Computers in Simulation, 80 (10):2033–2039, June 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000832.

Sousa-Vieira:2010:FSS

[115] M. E. Sousa-Vieira, A. Suárez-González, C. López-García, M. Fernández-Veiga, J. C. López-Ardao, and R. F. Rodríguez-Rubio. Fast simulation of self-similar and correlated processes. Mathematics and Computers in Simulation, 80(10):2040– 2061, June 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410000856.

Bhatti:2010:BSA

[116] Chad R. Bhatti. The Birnbaum– Saunders autoregressive conditional duration model. Mathematics and Computers in Simulation, 80(10):2062– 2078, June 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410000868.

McAleer:2010:SEV

[117] Michael McAleer and Chatayan Wiphatthanananthakul. A simple expected volatility (SEV) index: Application to SET50 index options. Mathematics and Computers in Simulation, 80(10):2079–2090, June 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541000090X.

Mukanova:2010:IRP

[118] Balgaisha Mukanova and Murat Orunkhanov. Inverse resistivity problem: Geoelectric uncertainty principle and numerical reconstruction method. Mathematics and Computers in Simulation, 80(10):2091– 2108, June 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410000911.

Lanconelli:2010:IMC

[119] Nico Lanconelli. The importance of Monte Carlo simulations in modeling detectors for Nuclear Medicine. Mathematics and Computers in Simulation, 80(10):2109–2114, June 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001084.

Anonymous:2010:NIf

[120] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 80(10):2115, June 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001692.

Anonymous:2010:ICEf

[121] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 80(10):2116– 2122, June 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001709.

Anonymous:2010:EBg

[122] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 80(11):CO2, July 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002077.

Deville:2010:F

[123] M. Deville. Foreword. Mathematics and Computers in Simulation, 80(11):2123, July 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001114.

Bazan:2010:CPM

[124] F. S. V. Bazán. Chebyshev pseudospectral method for wave equation with absorbing boundary conditions that does not use a first order hyperbolic system. Mathematics and Computers in Simulation, 80 (11):2124–2133, July 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001199.

Dulla:2010:TDT

[125] S. Dulla, S. Canepa, and P. Ravetto. From transport to diffusion through a space asymptotic approach. Mathematics and Computers in Simulation, 80 (11):2134–2141, July 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001126.

Ganapol:2010:OGS

[126] B. Ganapol. One-group steady state diffusion in a heterogeneous slab. Mathematics and Computers in Simulation, 80(11):2142–2158, July 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001138.

Guerin:2010:DDM

[127] Pierre Guérin, Anne-Marie Baudron, and Jean-Jacques Lautard. Domain decomposition methods for the neutron diffusion problem. Mathematics and Computers in Simulation, 80 (11):2159–2167, July 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541000114X.

Hennart:2010:NFE

[128] Jean-Pierre Hennart and Edmundo del Valle. Nodal finite element approximations for the neutron transport equation. Mathematics and Computers in Simulation, 80(11):2168– 2176, July 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001151.

Ma:2010:PSF

[129] To Fu Ma and André Luís Machado Martinez. Positive solutions for a fourth order equation with nonlinear boundary conditions. Mathematics and Computers in Simulation, 80 (11):2177–2184, July 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001163.

Mompean:2010:FVN

[130] G. Mompean and L. Thais. Finite volume numerical simulation of viscoelastic flows in general orthogonal coordinates. Mathematics and Computers in Simulation, 80(11):2185– 2199, July 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001175.

Rincon:2010:EAT

[131] M. A. Rincon and M. O. Lacerda. Error analysis of thermal equation with moving ends. Mathematics and Computers in Simulation, 80 (11):2200–2215, July 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410001187.

VandenEynde:2010:CNS

[132] Gert Van den Eynde, Robert Beauwens, and Ernest Mund. Calculating nearsingular eigenvalues of the neutron transport operator with arbitrary order anisotropic scattering. Mathematics and Computers in Simulation, 80 (11):2216–2230, July 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410001205.

Anonymous:2010:NIg

[133] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 80(11):2231, July 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410002144.

Anonymous:2010:ICEg

[134] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 80(11):2232– 2238, July 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410002156.

Anonymous:2010:PA

[135] Anonymous. Pages 2245–2416 (August 2010). Mathematics and Computers in Simulation, 80(12):??, August 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2010:EBh

[136] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 80(12):CO2, August 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002387.

Dadras:2010:ASM

[137] Sara Dadras and Hamid Reza Momeni. Adaptive sliding mode control of chaotic dynamical systems with application to synchronization. Mathematics and Computers in Simulation, 80 (12):2245–2257, August 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001096.

Lou:2010:PCU

[138] Xuyang Lou and Baotong Cui. Passive control of uncertain multiple input-delayed systems using reduction method. Mathematics and Computers in Simulation, 80(12):2258– 2271, August 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001382.

Popa:2010:HKC

[139] Constantin Popa. A hybrid Kaczmarz– Conjugate Gradient algorithm for image reconstruction. Mathematics and Computers in Simulation, 80(12):2272– 2285, August 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001412.

Mahmoud:2010:SCH

[140] Gamal M. Mahmoud and Emad E. Mahmoud. Synchronization and control of hyperchaotic complex Lorenz system. Mathematics and Computers in Simulation, 80(12):2286– 2296, August 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001448.

Wu:2010:SIV

[141] Jianyong Wu, Qingping Wang, Xue Wei, and Houjun Tang. Studies on improving vehicle handling and lane keeping performance of closedloop driver–vehicle system with integrated chassis control. Mathematics and Computers in Simulation, 80(12): 2297–2308, August 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001485.

Ji:2010:VIM

[142] Yu Ji, Lequan Min, Yu Zheng, and Yongmei Su. A viral infection model with periodic immune response and nonlinear CTL response. Mathematics and Computers in Simulation, 80 (12):2309–2316, August 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001539.

Herman:2010:MSP

[143] Przemysław Herman. Modified setpoint controller for underwater vehicles. Mathematics and Computers in Simulation, 80(12):2317– 2328, August 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001606.

Li:2010:SIC

[144] Kelin Li and Huanglin Zeng. Stability in impulsive Cohen–Grossbergtype BAM neural networks with time-varying delays: a general analysis. Mathematics and Computers in Simulation, 80(12):2329– 2349, August 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001618.

Xu:2010:ERK

[145] Y. Xu, J. J. Zhao, and Z. N. Sui. Exponential Runge–Kutta methods for delay differential equations. Mathematics and Computers in Simulation, 80 (12):2350–2361, August 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001795.

Zhou:2010:NSC

[146] Shenggao Zhou and Xiaoliang Cheng. Numerical solution to coupled nonlinear Schrödinger equations on unbounded domains. Mathematics and Computers in Simulation, 80(12):2362– 2373, August 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410001825.

Popescu:2010:ATI

[147] Th. D. Popescu. Analysis of trafficinduced vibrations by blind source separation with application in building monitoring. Mathematics and Computers in Simulation, 80(12):2374– 2385, August 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410001837.

Wu:2010:LTL

[148] Jyh-Yang Wu, Mei-Hsiu Chi, and Sheng-Gwo Chen. A local tangential lifting differential method for triangular meshes. Mathematics and Computers in Simulation, 80(12):2386– 2402, August 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410001898.

Anonymous:2010:NIh

[149] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 80(12):2403, August 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410002417.

Anonymous:2010:ICEh

[150] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 80(12):2404– 2410, August 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002429.

Anonymous:2010:PS

[151] Anonymous. Pages 1–156 (September 2010). Mathematics and Computers in Simulation, 81(1):??, September 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2010:EBi

[152] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 81(1):CO2, September 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002764.

Cui:2010:COP

[153] Zhenyu Cui and Don Mcleish. Comment on “Option pricing under the Merton model of the short rate” by Kung and Lee [Math. Comput. Simul. 80 (2009) 378–386]. Mathematics and Computers in Simulation, 81(1):1–4, September 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001953. See

[3].

Bai:2010:MLN

[154] Zhidong Bai, Wing-Keung Wong, and Bingzhi Zhang. Multivariate linear and nonlinear causality tests. Mathematics and Computers in Simulation, 81 (1):5–17, September 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410001977.

Jordan:2010:ACH

[155] P. M. Jordan. On the application of the Cole–Hopf transformation to hyperbolic equations based on second-sound models. Mathematics and Computers in Simulation, 81 (1):18–25, September 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410002004.

Aboeleneen:2010:IWD

[156] Z. A. Aboeleneen. Inference for Weibull distribution under generalized order statistics. Mathematics and Computers in Simulation, 81(1): 26–36, September 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410002028.

Zheng:2010:SRM

[157] G. H. Zheng and T. Wei. Spectral regularization method for the time fractional inverse advection– dispersion equation. Mathematics and Computers in Simulation, 81(1): 37–51, September 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410002181.

Brianzoni:2010:BCB

[158] Serena Brianzoni, Elisabetta Michetti, and Iryna Sushko. Border collision bifurcations of superstable cycles in a one-dimensional piecewise smooth map. Mathematics and Computers in Simulation, 81(1):52– 61, September 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002193.

Chen:2010:MPC

[159] Zhanshou Chen and Zheng Tian. Modified procedures for change point monitoring in linear models. Mathematics and Computers in Simulation, 81 (1):62–75, September 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002235.

Mroginski:2010:FEA

[160] Javier L. Mroginski, H. Ariel Di Rado, Pablo A. Beneyto, and Armando M. Awruch. A finite element approach for multiphase fluid flow in porous media. Mathematics and Computers in Simulation, 81 (1):76–91, September 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002247.

Zuoxiang:2010:RCE

[161] Peng Zuoxiang, Zhichao Weng, and Saralees Nadarajah. Rates of convergence of extremes for mixed exponential distributions. Mathematics and Computers in Simulation, 81 (1):92–99, September 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410002272.

Zhang:2010:ASB

[162] Chengjian Zhang and Hao Chen. Asymptotic stability of block boundary value methods for delay differentialalgebraic equations. Mathematics and Computers in Simulation, 81(1): 100–108, September 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410002491.

Pryce:2010:CNF

[163] J. D. Pryce, R. Khoshsiar Ghaziani, V. De Witte, and W. Govaerts. Computation of normal form coefficients of cycle bifurcations of maps by algorithmic differentiation. Mathematics and Computers in Simulation, 81 (1):109–119, September 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541000251X.

Respondek:2010:NSP

[164] Jerzy Stefan Respondek. Numerical simulation in the partial differential equation controllability analysis with physically meaningful constraints. Mathematics and Computers in Simulation, 81(1):120–132, September 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410002533.

Gross:2010:NIS

[165] Andreas Gross and Hermann F. Fasel. Numerical investigation of supersonic flow for axisymmetric cones. Mathematics and Computers in Simulation, 81(1):133–142, September 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002557.

Anonymous:2010:NIi

[166] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 81(1):143, September 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002806.

Anonymous:2010:ICEi

[167] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 81(1):144–150, September 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002818.

Anonymous:2010:EBj

[168] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 81(2):CO2, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003344.

Monmasson:2010:E

[169] Eric Monmasson, Benoit Robyns, Guy Olivier, Francis Labrique, Yves Perriard, and Ilhem Slama-Belkhodja. Editorial. Mathematics and Computers in Simulation, 81(2):157, October 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410003460.

Steiner:2010:MBD

[170] N. Yousfi Steiner, D. Candusso, D. Hissel, and P. Moçoteguy. Modelbased diagnosis for proton exchange membrane fuel cells. Mathematics and Computers in Simulation, 81(2): 158–170, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000509.

Cardenas:2010:MDM

[171] Alben Cardenas, Pierre Sicard, and Ahmed Chériti. Multi-drive management system to mitigate voltage sags. Mathematics and Computers in Simulation, 81(2):171– 179, October 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410000078.

Amarir:2010:NMT

[172] Saïd Amarir and Kamal Al-Haddad. A new mathematical tool to investigate the influence of cable characteristics and IGBT fast switching on voltage transients and differential mode currents for PWM drives. Mathematics and Computers in Simulation, 81 (2):180–193, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000467.

Tu:2010:TMI

[173] Xiaoping Tu, Louis-A. Dessaint, Roger Champagne, and Kamal Al-Haddad. Transient model of induction machine using rotating magnetic field approach. Mathematics and Computers in Simulation, 81(2): 194–207, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000492.

Courtecuisse:2010:MDF

[174] Vincent Courtecuisse, Jonathan Sprooten, Benoît Robyns, Marc Petit, Bruno François, and Jacques Deuse. A methodology to design a fuzzy logic based supervision of Hybrid Renewable Energy Systems. Mathematics and Computers in Simulation, 81 (2):208–224, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000741.

Patin:2010:MCC

[175] N. Patin, E. Monmasson, and J.- P. Louis. Modeling and control of a cascaded doubly-fed induction generator based on dynamical equivalent circuits. Mathematics and Computers in Simulation, 81(2):225– 238, October 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001217.

Chebak:2010:ODH

[176] Ahmed Chebak, Philippe Viarouge, and Jérôme Cros. Optimal design of a high-speed slotless permanent magnet synchronous generator with soft magnetic composite stator yoke and rectifier load. Mathematics and Computers in Simulation, 81 (2):239–251, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410001503.

Aguglia:2010:DFO

[177] Davide Aguglia, Philippe Viarouge, René Wamkeue, and Jérôme Cros. Determination of fault operation dynamical constraints for the design of wind turbine DFIG drives. Mathematics and Computers in Simulation, 81 (2):252–262, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410001497.

Zhang:2010:CDD

[178] He Zhang, Christophe Saudemont, Benoît Robyns, and Régis Meuret. Comparison of different DC voltage supervision strategies in a local Power Distribution System of More Electric Aircraft. Mathematics and Computers in Simulation, 81 (2):263–276, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410001588.

Hasni:2010:ESM

[179] M. Hasni, O. Touhami, R. Ibtiouen, M. Fadel, and S. Caux. Estimation of synchronous machine parameters by standstill tests. Mathematics and Computers in Simulation, 81 (2):277–289, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541000159X.

Dehez:2010:SOT

[180] Bruno Dehez, Christophe Vloebergh, and Francis Labrique. Study and optimization of traveling wave generation in finite-length beams. Mathematics and Computers in Simulation, 81 (2):290–301, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541000162X.

Allegre:2010:IML

[181] A.-L. Allègre, P. Delarue, P. Barrade, A. Bouscayrol, E. Chattot, and S. El-Fassi. Influence of the mechanical limitations of a traction system on energy storage design. Mathematics and Computers in Simulation, 81 (2):302–314, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541000203X.

Labbe:2010:TPO

[182] Thibaut Labbé, Bruno Dehez, and Francis Labrique. Two phase operation for a three phase PMSM using a control model based on a Concordia like transform associated to a classic Park transform. Mathematics and Computers in Simulation, 81 (2):315–326, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002326.

Menard:2010:DEM

[183] Laurianne Ménard, Guillaume Fontès, and Stéphan Astier. Dynamic energy model of a lithium-ion battery. Mathematics and Computers in Simulation, 81(2):327–339, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002636.

Sincero:2010:CMS

[184] G. C. R. Sincero, J. Ghannou, J. Cros, and P. Viarouge. Collector model for simulation of brush machines. Mathematics and Computers in Simulation, 81(2):340–353, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002624.

Hamiti:2010:MSR

[185] Tahar Hamiti, Thierry Lubin, Lotfi Baghli, and Abderrezak Rezzoug. Modeling of a synchronous reluctance machine accounting for space harmonics in view of torque ripple minimization. Mathematics and Computers in Simulation, 81(2):354– 366, October 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002612.

DeBelie:2010:SPD

[186] Frederik M. De Belie, Peter Sergeant, and Jan A. Melkebeek. A sensorless PMSM drive using modified high-frequency test pulse sequences for the purpose of a discrete-time current controller with fixed sampling frequency. Mathematics and Computers in Simulation, 81(2):367– 381, October 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002600.

Hamimi:2010:REI

[187] Amina N. Hamimi, François Ruelland, and Kamal Al-Haddad. Resonance effect of insulated negative current rail in a subway network. Mathematics and Computers in Simulation, 81 (2):382–393, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002740.

Errouissi:2010:NPC

[188] Rachid Errouissi and Mohand Ouhrouche. Nonlinear predictive controller for a permanent magnet synchronous motor drive. Mathematics and Computers in Simulation, 81(2):394– 406, October 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002855.

Li:2010:DRU

[189] Jian Li and Yunhyun Cho. Dynamic reduction of unbalanced magnetic force and vibration in switched reluctance motor by the parallel paths in windings. Mathematics and Computers in Simulation, 81(2):407– 419, October 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002879.

Piquet:2010:DLC

[190] Hubert Piquet, Rafael Díez, Jean-Marc Blaquière, Sounil Bhosle, and Nicolas Roux. DBD lamp converter design using an electrical model of the load. Mathematics and Computers in Simulation, 81(2): 420–432, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002867.

Machmoum:2010:FMD

[191] Mohamed Machmoum, Ahmad Hatoum, and Toufik Bouaouiche. Flicker mitigation in a doubly fed induction generator wind turbine system. Mathematics and Computers in Simulation, 81(2):433–445, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002910.

Flueckiger:2010:FEM

[192] Markus Flueckiger, José M. Fernandez, and Yves Perriard. Finite element method based design and optimisation methodology for piezoelectric ultrasonic motors. Mathematics and Computers in Simulation, 81 (2):446–459, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002909.

Anonymous:2010:NIj

[193] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 81(2):460, October 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003411.

Anonymous:2010:ICEj

[194] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 81(2):461– 467, October 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410003423.

Anonymous:2010:EBk

[195] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 81(3):CO2, November 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003526.

Alexandrov:2010:E

[196] Vassil Alexandrov. Editorial. Mathematics and Computers in Simulation, 81(3):475, November 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003885.

LEcuyer:2010:CPR

[197] P. L’Ecuyer and C. Sanvido. Coupling from the past with randomized quasi-Monte Carlo. Mathematics and Computers in Simulation, 81 (3):476–489, November 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002912.

Vandewoestyne:2010:CQR

[198] Bart Vandewoestyne and Ronald Cools. On the convergence of quasirandom sampling/importance resampling. Mathematics and Computers in Simulation, 81(3):490–505, November 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002924.

Dimov:2010:MCA

[199] I. Dimov and R. Georgieva. Monte Carlo algorithms for evaluating Sobol’ sensitivity indices. Mathematics and Computers in Simulation, 81(3):506– 514, November 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002936.

Atanassov:2010:EPI

[200] E. Atanassov, T. Gurov, A. Karaivanova, M. Nedjalkov, D. Vasileska, and K. Raleva. Electron–phonon interaction in nanowires: a Monte Carlo study of the effect of the field. Mathematics and Computers in Simulation, 81 (3):515–521, November 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002948.

Vandewoestyne:2010:CIS

[201] Bart Vandewoestyne, Hongmei Chi, and Ronald Cools. Computational investigations of scrambled Faure sequences. Mathematics and Computers in Simulation, 81(3):522–535, November 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540900295X.

Han:2010:SEM

[202] Chuan-Hsiang Han and Yongzeng Lai. A smooth estimator for MC/QMC methods in finance. Mathematics and Computers in Simulation, 81(3): 536–550, November 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002508.

Morohosi:2010:MNR

[203] Hozumi Morohosi. Measuring the network robustness by Monte Carlo estimation of shortest path length distribution. Mathematics and Computers in Simulation, 81(3):551–559, November 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002594.

ElHaddad:2010:QMC

[204] R. El Haddad, C. Lécot, P. L’Ecuyer, and N. Nassif. Quasi-Monte Carlo methods for Markov chains with continuous multi-dimensional state space. Mathematics and Computers in Simulation, 81(3):560–567, November 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002648.

Doan:2010:PPA

[205] Viet Dung Doan, Abhijeet Gaikwad, Mireille Bossy, Françoise Baude, and Ian Stokes-Rees. Parallel pricing algorithms for multi-dimensional Bermudan/American options using Monte Carlo methods. Mathematics and Computers in Simulation, 81(3):568– 577, November 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002739.

Larsson:2010:MCI

[206] Jonas Larsson. Monte Carlo implementation of financial simulation on Cell/ B.E. multi-core processor. Mathematics and Computers in Simulation, 81 (3):578–587, November 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002727.

Pistella:2010:ASC

[207] Francesca Pistella and Rosa Maria Spitaleri. Applied scientific computing VII. Forward numerical grid generation, approximation and simulation. Mathematics and Computers in Simulation, 81(3):589, November 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410003502.

Rigatos:2010:EKP

[208] Gerasimos G. Rigatos. Extended Kalman and Particle Filtering for sensor fusion in motion control of mobile robots. Mathematics and Computers in Simulation, 81(3):590– 607, November 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001515.

Quatember:2010:GMM

[209] Bernhard Quatember, Martin Mayr, Wolfgang Recheis, Stefanos Demertzis, Giampietro Allasia, Alessandra De Rossi, Roberto Cavoretto, and Ezio Venturino. Geometric modeling and motion analysis of the epicardial surface of the heart. Mathematics and Computers in Simulation, 81(3):608– 622, November 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002016.

Conti:2010:SNA

[210] Costanza Conti. Stationary and nonstationary affine combination of subdivision masks. Mathematics and Computers in Simulation, 81(3):623– 635, November 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409003541.

Anotaipaiboon:2010:OGF

[211] W. Anotaipaiboon and S. S. Makhanov. Optimal grids for five-axis machining. Mathematics and Computers in Simulation, 81(3):636–655, November 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410001527.

Bucchignani:2010:CFA

[212] Edoardo Bucchignani and Giuseppe Pezzella. Computational flowfield analyses of hypersonic problems with reacting boundary layer. Mathematics and Computers in Simulation, 81(3): 656–669, November 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410000042.

Pellegrino:2010:OLR

[213] E. Pellegrino and E. Santi. Optimal local refinable interpolants. Mathematics and Computers in Simulation, 81 (3):670–680, November 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410001102.

Lovison:2010:EOD

[214] A. Lovison and E. Rigoni. Extracting optimal datasets for metamodelling and perspectives for incremental samplings. Mathematics and Computers in Simulation, 81(3): 681–692, November 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000790.

Egidi:2010:USM

[215] N. Egidi and P. Maponi. The use of Sherman–Morrison formula in the solution of Fredholm integral equation of second kind. Mathematics and Computers in Simulation, 81(3): 693–704, November 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000789.

Bersani:2010:AOC

[216] Alberto Maria Bersani, Elisabetta Carlini, Piero Lanucara, Marco Rorro, and Vittorio Ruggiero. Application of Optimal Control techniques and Advanced Computing to the study of enzyme kinetics. Mathematics and Computers in Simulation, 81(3): 705–716, November 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003009.

Kervarc:2010:GLT

[217] Romain Kervarc, Jean Bourrely, and Claire Quillien. A generic logicaltemporal performance analysis method for complex systems. Mathematics and Computers in Simulation, 81(3): 717–730, November 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002971.

Egidi:2010:EMS

[218] N. Egidi and P. Maponi. An efficient method for the solution of the inverse scattering problem for penetrable obstacles. Mathematics and Computers in Simulation, 81(3): 731–741, November 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410003046.

Arbel:2010:SIF

[219] A. Arbel and A. Shklyar. Simulation of incompressible flow with alternate pressure Dirichlet and Neumann conditions. Mathematics and Computers in Simulation, 81(3):742– 756, November 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847541000306X.

Anonymous:2010:NIk

[220] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 81(3):757, November 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410003599.

Anonymous:2010:ICEk

[221] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 81(3):758–764, November 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410003605.

Anonymous:2010:PD

[222] Anonymous. Pages 771–928 (December 2010). Mathematics and Computers in Simulation, 81(4):??, December 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2010:EBl

[223] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 81(4):CO2, December 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003757.

Chen:2010:BCG

[224] Jianwei Chen, Kim-Hung Li, and Yeh Lam. Bayesian computation for geometric process in maintenance problems. Mathematics and Computers in Simulation, 81(4):771– 781, December 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541000193X.

Hormann:2010:CGC

[225] Wolfgang Hörmann and Halis Sak. t-Copula generation for control variates. Mathematics and Computers in Simulation, 81(4):782–790, December 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002296.

Pruliere:2010:DSM

[226] E. Pruliere, F. Chinesta, and A. Ammar. On the deterministic solution of multidimensional parametric models using the Proper Generalized Decomposition. Mathematics and Computers in Simulation, 81(4): 791–810, December 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002521.

Mora:2010:CDB

[227] Juan Mora and Llanos Mora-López. Comparing distributions with bootstrap techniques: an application to global solar radiation. Mathematics and Computers in Simulation, 81(4): 811–819, December 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002934.

Guias:2010:DSI

[228] Flavius Guiaş. Direct simulation of the infinitesimal dynamics of semi-discrete approximations for convection–diffusion–reaction problems. Mathematics and Computers in Simulation, 81(4):820–836, December 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002946.

Wang:2010:LCE

[229] Huiwei Wang, Qiankun Song, and Chengjun Duan. LMI criteria on exponential stability of BAM neural networks with both time-varying delays and general activation functions. Mathematics and Computers in Simulation, 81(4):837–850, December 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002958.

Jodra:2010:CGR

[230] P. Jodrá. Computer generation of random variables with Lindley or Poisson–Lindley distribution via the Lambert W function. Mathematics and Computers in Simulation, 81(4): 851–859, December 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541000296X.

Ramos:2010:NAS

[231] Airton Ramos. A new approach to the solution of electromagnetic problems with the impedance method. Mathematics and Computers in Simulation, 81(4):860–874, December 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475410002983.

Aladag:2010:HOF

[232] Cagdas Hakan Aladag, Ufuk Yolcu, and Erol Egrioglu. A high order fuzzy time series forecasting model based on adaptive expectation and artificial neural networks. Mathematics and Computers in Simulation, 81(4): 875–882, December 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003010.

Jungel:2010:TDS

[233] Ansgar Jüngel and Jan-Frederik Mennemann. Time-dependent simulations of quantum waveguides using a timesplitting spectral method. Mathematics and Computers in Simulation, 81 (4):883–898, December 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410003034.

Gardini:2010:BCB

[234] Laura Gardini, Fabio Tramontana, and Iryna Sushko. Border collision bifurcations in one-dimensional linear-hyperbolic maps. Mathematics and Computers in Simulation, 81(4): 899–914, December 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410003071.

Anonymous:2010:NIl

[235] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 81(4):915, December 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410003782.

Anonymous:2010:ICEl

[236] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 81(4):916–922, December 2010. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410003794.

Anonymous:2011:EBa

[237] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 81(5):CO2, January 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000036.

DelBuono:2011:GES

[238] Nicoletta Del Buono. Guest Editorial: Some important aspects on Structural Dynamical Systems and their numerical computation. Mathematics and Computers in Simulation, 81 (5):929–931, January 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003083.

Dieci:2011:FMS

[239] Luca Dieci and Luciano Lopez. Fundamental matrix solutions of piecewise smooth differential systems. Mathematics and Computers in Simulation, 81(5):932–953, January 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003186.

Pisano:2011:SMC

[240] Alessandro Pisano and Elio Usai. Sliding mode control: a survey with applications in math. Mathematics and Computers in Simulation, 81 (5):954–979, January 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003095.

Kupper:2011:RIC

[241] T. Küpper and H. A. Hosham. Reduction to invariant cones for nonsmooth systems. Mathematics and Computers in Simulation, 81(5):980– 995, January 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410003101.

Dieci:2011:LCS

[242] Luca Dieci, M. Grazia Gasparo, Alessandra Papini, and Alessandro Pugliese. Locating coalescing singular values of large two-parameter matrices. Mathematics and Computers in Simulation, 81(5):996– 1005, January 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410003113.

Menning:2011:EAS

[243] Melissa Menning and Erik S. Van Vleck. On the error in approximating stability spectra for discrete dynamical systems. Mathematics and Computers in Simulation, 81(5):1006– 1016, January 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410003125.

Messina:2011:CAN

[244] E. Messina, E. Russo, and A. Vecchio. Comparing analytical and numerical solution of a nonlinear twodelay integral equations. Mathematics and Computers in Simulation, 81 (5):1017–1026, January 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003137.

Bozzini:2011:TWR

[245] Benedetto Bozzini, Deborah Lacitignola, and Ivonne Sgura. Travelling waves in a reaction–diffusion model for electrodeposition. Mathematics and Computers in Simulation, 81(5): 1027–1044, January 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003149.

Garrappa:2011:UMF

[246] Roberto Garrappa and Marina Popolizio. On the use of matrix functions for fractional partial differential equations. Mathematics and Computers in Simulation, 81(5):1045– 1056, January 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410003150.

Diele:2011:ELI

[247] F. Diele, C. Marangi, and S. Ragni. Exponential Lawson integration for nearly Hamiltonian systems arising in optimal control. Mathematics and Computers in Simulation, 81(5): 1057–1067, January 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003162.

DAmbrosio:2011:TFT

[248] R. D’Ambrosio, M. Ferro, and B. Paternoster. Trigonometrically fitted two-step hybrid methods for special second order ordinary differential equations. Mathematics and Computers in Simulation, 81(5):1068– 1084, January 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410003174.

Anonymous:2011:NIa

[249] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 81(5):1085, January 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000097.

Anonymous:2011:ICEa

[250] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 81(5):1086– 1092, January 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411000103.

Anonymous:2011:PF

[251] Anonymous. Pages 1099–1232 (February 2011). Mathematics and Computers in Simulation, 81(6):??, February 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2011:EBb 81(6):CO2, February 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000164.

Jin:2011:MTV

[253] Hao Jin and Jinsuo Zhang. Modified tests for variance changes in autoregressive regression. Mathematics and Computers in Simulation, 81(6): 1099–1109, February 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000819.

Buckwar:2011:CLM

[254] Evelyn Buckwar and Thorsten Sickenberger. A comparative linear mean-square stability analysis of Maruyama- and Milstein-type methods. Mathematics and Computers in Simulation, 81(6):1110–1127, February 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410003058.

Pulch:2011:MSA

[255] Roland Pulch. Modelling and simulation of autonomous oscillators with random parameters. Mathematics and Computers in Simulation, 81(6): 1128–1143, February 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003721.

Lin:2011:NIR

[256] Ji Lin, Wen Chen, and Fuzhang Wang. A new investigation into regularization techniques for the method of fundamental solutions. Mathematics and Computers in Simulation, 81(6): 1144–1152, February 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410003836.

Bian:2011:TTP

[257] Guorui Bian, Michael McAleer, and Wing-Keung Wong. A trinomial test for paired data when there are many ties. Mathematics and Computers in Simulation, 81(6):1153– 1160, February 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410003848.

Chao:2011:TDF

[258] Chien-Chi Chao. Three-dimensional free surface simulation. Mathematics and Computers in Simulation, 81(6): 1161–1176, February 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541000385X.

Lee:2011:CPA

[259] Hsiu-Mei Lee, Wen-Chuan Lee, Chia-Ling Lei, and Jong-Wuu Wu. Computational procedure of assessing lifetime performance index of Weibull lifetime products with the upper record values. Mathematics and Computers in Simulation, 81(6):1177– 1189, February 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410003861.

Ghitany:2011:TPW

[260] M. E. Ghitany, F. Alqallaf, D. K. Al-Mutairi, and H. A. Husain. A two-parameter weighted Lindley distribution and its applications to survival data. Mathematics and Computers in Simulation, 81(6):1190– 1201, February 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410003873.

Kebriaei:2011:DMD

[261] Hamed Kebriaei and Ashkan Rahimi-Kian. Decision making in dynamic stochastic Cournot games. Mathematics and Computers in Simulation, 81(6):1202–1217, February 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475410003915.

Anonymous:2011:NIb

[262] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 81(6):1218, February 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000206.

Anonymous:2011:ICEb

[263] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 81(6):1219– 1225, February 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411000218.

Anonymous:2011:EBc

[264] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 81(7):CO2, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000383.

McAleer:2011:P

[265] Michael McAleer and Les Oxley. Preface. Mathematics and Computers in Simulation, 81(7):1233– 1234, March 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411000565.

Hsiao:2011:CFM

[266] Cheng Hsiao and Shui Ki Wan. Comparison of forecasting methods with an application to predicting excess equity premium. Mathematics and Computers in Simulation, 81(7):1235– 1246, March 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847541000087X.

Allen:2011:MCO

[267] David E. Allen, Michael McAleer, and Marcel Scharth. Monte Carlo option pricing with asymmetric realized volatility dynamics. Mathematics and Computers in Simulation, 81 (7):1247–1256, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001990.

Mullins:2011:MDM

[268] B. J. Mullins, R. D. Braddock, and I. E. Agranovski. Modelling droplet motion and interfacial tension in filters collecting liquid aerosols. Mathematics and Computers in Simulation, 81 (7):1257–1271, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001576.

Kunitomo:2011:SER

[269] Naoto Kunitomo and Seisho Sato. The SIML estimation of realized volatility of the Nikkei-225 Futures and hedging coefficient with micromarket noise. Mathematics and Computers in Simulation, 81(7):1272– 1289, March 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002715.

Oya:2011:BCR

[270] Kosuke Oya. Bias-corrected realized variance under dependent microstructure noise. Mathematics and Computers in Simulation, 81(7):1290– 1298, March 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001229.

Chan:2011:MSP

[271] Felix Chan and Laurent Pauwels. Model specification in panel data unit root tests with an unknown break. Mathematics and Computers in Simulation, 81(7):1299– 1309, March 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410001242.

Shang:2011:NTS

[272] Han Lin Shang and Rob. J. Hyndman. Nonparametric time series forecasting with dynamic updating. Mathematics and Computers in Simulation, 81 (7):1310–1324, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541000145X.

Li:2011:MOA

[273] J. S. H. Li, A. C. Y. Ng, and W. S. Chan. Modeling old-age mortality risk for the populations of Australia and New Zealand: an extreme value approach. Mathematics and Computers in Simulation, 81 (7):1325–1333, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410001424.

Wong:2011:MHK

[274] C. S. Wong. Modeling Hong Kong’s stock index with the Student t-mixture autoregressive model. Mathematics and Computers in Simulation, 81(7): 1334–1343, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001631.

Ramilan:2011:AEE

[275] Thiagarajah Ramilan, Frank Scrimgeour, and Dan Marsh. Analysis of environmental and economic efficiency using a farm population micro-simulation model. Mathematics and Computers in Simulation, 81 (7):1344–1352, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001230.

Sato:2011:ISR

[276] Kiyotaka Sato, Zhaoyong Zhang, and Michael McAleer. Identifying shocks in regionally integrated East Asian economies with structural VAR and block exogeneity. Mathematics and Computers in Simulation, 81(7):1353– 1364, March 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001989.

Yap:2011:IOL

[277] Ghialy Yap and David Allen. Investigating other leading indicators influencing Australian domestic tourism demand. Mathematics and Computers in Simulation, 81(7):1365– 1374, March 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001540.

Herbert:2011:SMM

[278] Ric D. Herbert and Peter J. Stemp. Solving macroeconomic models with “off-the-shelf” software: an example of potential pitfalls. Mathematics and Computers in Simulation, 81 (7):1375–1384, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410002697.

Chan:2011:ERS

[279] Felix Chan and Billy Theoharakis. Estimating m-regimes STAR-GARCH model using QMLE with parameter transformation. Mathematics and Computers in Simulation, 81(7):1385– 1396, March 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410001862.

Todorov:2011:MS

[280] Vladislav Todorov and Dora Marinova. Modelling sustainability. Mathematics and Computers in Simulation, 81 (7):1397–1408, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410001850.

Chan:2011:SAS

[281] Felix Chan and Christine Lim. Spectral analysis of seasonality in tourism demand. Mathematics and Computers in Simulation, 81(7):1409– 1418, March 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001941.

Kim:2011:AMC

[282] H. S. Kim, B. F. W. Croke, A. J. Jakeman, and F. H. S. Chiew. An assessment of modelling capacity to identify the impacts of climate variability on catchment hydrology. Mathematics and Computers in Simulation, 81 (7):1419–1429, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001564.

Zhao:2011:GDE

[283] Xin Zhao, Carl John Scarrott, Les Oxley, and Marco Reale. GARCH dependence in extreme value models with Bayesian inference. Mathematics and Computers in Simulation, 81 (7):1430–1440, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002703.

Rea:2011:LMS

[284] William Rea, Marco Reale, Jennifer Brown, and Les Oxley. Long memory or shifting means in geophysical time series? Mathematics and Computers in Simulation, 81(7):1441– 1453, March 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001965.

McAleer:2011:VRC

[285] Michael McAleer, Bernardo da Veiga, and Suhejla Hoti. Value-at-risk for country risk ratings. Mathematics and Computers in Simulation, 81(7): 1454–1463, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002053.

David:2011:GDB

[286] T. David, Thomas van Kempen, Huaxiong Huang, and Phillip Wilson. The geometry and dynamics of binary trees. Mathematics and Computers in Simulation, 81(7): 1464–1481, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001254.

Chang:2011:MCC

[287] Chia-Lin Chang, Thanchanok Khamkaew, Michael McAleer, and Roengchai Tansuchat. Modelling conditional correlations in the volatility of Asian rubber spot and futures returns. Mathematics and Computers in Simulation, 81 (7):1482–1490, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002284.

Chang:2011:MAV

[288] Chia-Lin Chang, Biing-Wen Huang, Meng-Gu Chen, and Michael McAleer. Modelling the asymmetric volatility in hog prices in Taiwan: the impact of joining the WTO. Mathematics and Computers in Simulation, 81 (7):1491–1506, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001928.

Sakata:2011:SSD

[289] K. Sakata and C. R. McKenzie. Social security and divorce in Japan. Mathematics and Computers in Simulation, 81(7):1507–1517, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001643.

Aiyoshi:2011:RNE

[290] Eitaro Aiyoshi, Atsushi Maki, and Takashi Okamoto. The robustness of a Nash equilibrium simulation model: Game-theoretic approach using variable metric projection method. Mathematics and Computers in Simulation, 81(7):1518–1526, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001904.

Anonymous:2011:NIc

[291] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 81(7):1527, March 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000449.

Anonymous:2011:ICEc

[292] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 81(7):1528– 1534, March 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411000450.

Anonymous:2011:PAa

[293] Anonymous. Pages 1541–1696 (April 2011). Mathematics and Computers in Simulation, 81(8):??, April 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2011:EBd

[294] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 81(8):CO2, April 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000632.

Collins:2011:SML

[295] Jeb Collins and Pierre Gremaud. A simple model for laser drilling. Mathematics and Computers in Simulation, 81(8):1541–1552, April 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410002351.

Oqielat:2011:MWD

[296] Moa’ath N. Oqielat, Ian W. Turner, John A. Belward, and Scott W. Mc-Cue. Modelling water droplet movement on a leaf surface. Mathematics and Computers in Simulation, 81 (8):1553–1571, April 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410002922.

Lakoba:2011:CCI

[297] T. I. Lakoba. Convergence conditions for iterative methods seeking multi-component solitary waves with prescribed quadratic conserved quantities. Mathematics and Computers in Simulation, 81(8):1572– 1592, April 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410003927.

Deng:2011:MRM

[298] Zhi-Liang Deng, Chu-Li Fu, Xiao-Li Feng, and Yuan-Xiang Zhang. A mollification regularization method for stable analytic continuation. Mathematics and Computers in Simulation, 81 (8):1593–1608, April 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003976.

Wan:2011:AIS

[299] Xiaojun Wan and Jitao Sun. Adaptive– impulsive synchronization of chaotic systems. Mathematics and Computers in Simulation, 81(8):1609– 1617, April 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410003988.

Nowman:2011:GEC

[300] K. Ben Nowman. Gaussian estimation of continuous time diffusions of UK interest rates. Mathematics and Computers in Simulation, 81 (8):1618–1624, April 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410004003.

Tramontana:2011:ECD

[301] Fabio Tramontana, Laura Gardini, and Anna Agliari. Endogenous cycles in discontinuous growth models. Mathematics and Computers in Simulation, 81(8):1625–1639, April 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410004027.

Foroni:2011:CDA

[302] Ilaria Foroni and Anna Agliari. Complex dynamics associated with the appearance/disappearance of invariant closed curves. Mathematics and Computers in Simulation, 81(8):1640– 1655, April 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847541000409X.

Yang:2011:IPT

[303] Liu Yang, Mehdi Dehghan, Jian-Ning Yu, and Guan-Wei Luo. Inverse problem of time-dependent heat sources numerical reconstruction. Mathematics and Computers in Simulation, 81 (8):1656–1672, April 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000280.

Zhang:2011:SGC

[304] Lingxiang Zhang and Xiaotong Zhang. Spurious Granger causality between a broken-trend stationary process and a stochastic trend process. Mathematics and Computers in Simulation, 81 (8):1673–1681, April 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541100036X.

Anonymous:2011:NId

[305] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 81(8):1682, April 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000668.

Anonymous:2011:ICEd

[306] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 81(8):1683– 1689, April 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541100067X.

Anonymous:2011:PM

[307] Anonymous. Pages 1697–1934 (May 2011). Mathematics and Computers in Simulation, 81(9):??, May 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2011:EBe

[308] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 81(9):CO2, May 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001017.

Zhang:2011:DAE

[309] Zhonghua Zhang, Jianhua Wu, Yaohong Suo, and Xinyu Song. The domain of attraction for the endemic equilibrium of an SIRS epidemic model. Mathematics and Computers in Simulation, 81(9):1697– 1706, May 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541000371X.

DAmbrosio:2011:CIH

[310] Raffaele D’Ambrosio and Zdzislaw Jackiewicz. Construction and implementation of highly stable twostep continuous methods for stiff differential systems. Mathematics and Computers in Simulation, 81(9):1707– 1728, May 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411000334.

Zhong:2011:AAE

[311] Sheng Zhong and Yuan Zhang. Approximation algorithm for estimating failure probability of multipath transmission. Mathematics and Computers in Simulation, 81(9):1729– 1732, May 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411000346.

Kurita:2011:LRE

[312] Takamitsu Kurita. Long-run exclusion and the determination of cointegrating rank: Monte Carlo evidence. Mathematics and Computers in Simulation, 81(9):1733–1740, May 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000358.

Garcia:2011:ACL

[313] Salvador Garcia. Aperiodic, chaotic lid-driven square cavity flows. Mathematics and Computers in Simulation, 81(9):1741–1769, May 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000516.

Heydari:2011:SCS

[314] Mahdi Heydari, Hassan Salarieh, and Mehdi Behzad. Stochastic chaos synchronization using Unscented Kalman–Bucy Filter and sliding mode control. Mathematics and Computers in Simulation, 81(9):1770– 1784, May 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541100053X.

Wu:2011:NEB

[315] Chun-Mu Wu, Ching-Kao Chang, and Tung-Te Chu. A new EP-based α–β– γ–δ filter for target tracking. Mathematics and Computers in Simulation, 81(9):1785–1794, May 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000590.

Bauer:2011:CPC

[316] Frank Bauer and Mark A. Lukas. Comparing parameter choice methods for regularization of ill-posed problems. Mathematics and Computers in Simulation, 81(9):1795– 1841, May 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411000607.

El-Adll:2011:PFL

[317] Magdy E. El-Adll. Predicting future lifetime based on random number of three parameters Weibull distribution. Mathematics and Computers in Simulation, 81(9):1842– 1854, May 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411000619.

Gao:2011:SMV

[318] Shujing Gao, Yujiang Liu, Juan J. Nieto, and Helena Andrade. Seasonality and mixed vaccination strategy in an epidemic model with vertical transmission. Mathematics and Computers in Simulation, 81(9):1855– 1868, May 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411000735.

Chleboun:2011:ADN

[319] Jan Chleboun. On an approach to deal with Neumann boundary value problems defined on uncertain domains: Numerical experiments. Mathematics and Computers in Simulation, 81 (9):1869–1875, May 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000759.

Ito:2011:PSG

[320] Yasushi Ito, Gary C. Cheng, Alan M. Shih, Roy P. Koomullil, Bharat K. Soni, Somsak Sittitavornwong, and Peter D. Waite. Patient-specific geometry modeling and mesh generation for simulating Obstructive Sleep Apnea Syndrome cases by Maxillomandibular Advancement. Mathematics and Computers in Simulation, 81 (9):1876–1891, May 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000760.

Xiangxian:2011:CBT

[321] Chen Xiangxian, He Yulin, and Huang hai. A component-based topology model for railway interlocking systems. Mathematics and Computers in Simulation, 81(9):1892– 1900, May 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411000772.

Mariani:2011:HSC

[322] Viviana Cocco Mariani and Leandro dos Santos Coelho. A hybrid shuffled complex evolution approach with pattern search for unconstrained optimization. Mathematics and Computers in Simulation, 81 (9):1901–1909, May 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000802.

Lee:2011:IOP

[323] Chien-Chiang Lee and Jhih-Hong Zeng. The impact of oil price shocks on stock market activities: Asymmetric effect with quantile regression. Mathematics and Computers in Simulation, 81(9):1910–1920, May 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000838.

Anonymous:2011:NIe

[324] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 81(9):1921, May 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411001042.

Anonymous:2011:ICEe

[325] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 81(9):1922– 1928, May 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411001054.

Anonymous:2011:EBf

[326] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 81(10):CO2, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541100125X.

Amaziane:2011:P

[327] Brahim Amaziane and Domingo Barrera. Preface. Mathematics and Computers in Simulation, 81(10): 1935, June 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411001376.

Abbadi:2011:CQI

[328] A. Abbadi, M. J. Ibáñez, and D. Sbibih. Computing quasi-interpolants from the B-form of B-splines. Mathematics and Computers in Simulation, 81 (10):1936–1948, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410004064.

Agouzal:2011:OCR

[329] Abdellatif Agouzal, Konstantin Lipnikov, and Yuri V. Vassilevski. On optimal convergence rate of finite element solutions of boundary value problems on adaptive anisotropic meshes. Mathematics and Computers in Simulation, 81(10):1949–1961, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410004283.

Ainseba:2011:CDO

[330] Bedr’Eddine Aïnseba and Chahrazed Benosman. CML dynamics: Optimal control of age-structured stem cell population. Mathematics and Computers in Simulation, 81(10):1962– 1977, June 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541000412X.

Alba-Fernandez:2011:ERS

[331] V. Alba-Fernández and M. D. Jiménez-Gamero. Estimating Rao’s statistic distribution for testing uniform association in cross-classifications. Mathematics and Computers in Simulation, 81 (10):1978–1990, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000541.

Allouch:2011:MKM

[332] C. Allouch, P. Sablonnière, and D. Sbibih. A modified Kulkarni’s method based on a discrete spline quasi-interpolant. Mathematics and Computers in Simulation, 81(10):1991– 2000, June 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410004180.

Angelini:2011:FVA

[333] O. Angelini, C. Chavant, E. Chénier, R. Eymard, and S. Granet. Finite volume approximation of a diffusion– dissolution model and application to nuclear waste storage. Mathematics and Computers in Simulation, 81 (10):2001–2017, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410004179.

Antontsev:2011:PED

[334] S. Antontsev and S. Shmarev. Parabolic equations with double variable nonlinearities. Mathematics and Computers in Simulation, 81(10):2018– 2032, June 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410004167.

Baccou:2011:KBS

[335] Jean Baccou and Jacques Liandrat. Kriging-based subdivision schemes: Application to the reconstruction of non-regular environmental data. Mathematics and Computers in Simulation, 81(10):2033–2050, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410004106.

Bastian:2011:UDG

[336] Peter Bastian, Christian Engwer, Jorrit Fahlke, and Olaf Ippisch. An Unfitted Discontinuous Galerkin method for pore-scale simulations of solute transport. Mathematics and Computers in Simulation, 81(10):2051– 2061, June 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410004258.

Benkamra:2011:TSD

[337] Zohra Benkamra, Mekki Terbeche, and Mounir Tlemcani. Tow stage design for estimating the reliability of series/parallel systems. Mathematics and Computers in Simulation, 81 (10):2062–2072, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000322.

Benkhaldoun:2011:CCF

[338] Fayssal Benkhaldoun and Mohammed Seaïd. Combined characteristics and finite volume methods for sediment transport and bed morphology in surface water flows. Mathematics and Computers in Simulation, 81 (10):2073–2086, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541000426X.

Benkhaldoun:2011:CUF

[339] Fayssal Benkhaldoun, Salah Daoudi, Imad Elmahi, and Mohammed Seaïd. Comparison of unstructured finitevolume morphodynamic models in contracting channel flows. Mathematics and Computers in Simulation, 81 (10):2087–2097, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410004349.

Cardenas-Morales:2011:SMS

[340] D. Cárdenas-Morales, A. Fraguela, and P. Garrancho. Saturation in multivariate simultaneous approximation. Mathematics and Computers in Simulation, 81(10):2098–2102, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410004143.

ElOssmani:2011:CMF

[341] Mustapha El Ossmani. A convergence of a MFE-FV method for immiscible compressible flow in heterogeneous porous media. Mathematics and Computers in Simulation, 81 (10):2103–2128, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410004088.

Fortes:2011:MAS

[342] M. A. Fortes and M. Moncayo. Multiresolution analysis and supercompact multiwavelets for surfaces. Mathematics and Computers in Simulation, 81 (10):2129–2149, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410004234.

Fortes:2011:HFM

[343] M. A. Fortes, P. González, M. Pasadas, and M. L. Rodríguez. A hole filling method for surfaces by using C^{1}- Powell–Sabin splines. Estimation of the smoothing parameters. Mathematics and Computers in Simulation, 81 (10):2150–2160, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410004040.

Fortes:2011:VTF

[344] M. A. Fortes, M. Pasadas, D. Sbibih, A. Serghini, and A. Tijini. Variational trivariate fitting using Worsey– Piper macro elements on tetrahedral partitions. Mathematics and Computers in Simulation, 81(10):2161– 2173, June 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411000929.

Garrancho:2011:AFS

[345] P. Garrancho, D. Cárdenas-Morales, and F. Aguilera. On asymptotic formulae via summability. Mathematics and Computers in Simulation, 81 (10):2174–2180, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410004076.

Golder:2011:MTS

[346] J. Golder, M. Joelson, and M. C. Néel. Mass transport with sorption in porous media. Mathematics and Computers in Simulation, 81 (10):2181–2189, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410004337.

Guessab:2011:EAN

[347] A. Guessab, M. J. Ibáñez, and O. Nouisser. Error analysis for a nonstandard class of differential quasiinterpolants. Mathematics and Computers in Simulation, 81(10):2190– 2200, June 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410004246.

Jean-Baptiste:2011:DAR

[348] Nelly Jean-Baptiste, Pierre-Olivier Malaterre, Christophe Dorée, and Jacques Sau. Data assimilation for real-time estimation of hydraulic states and unmeasured perturbations in a 1D hydrodynamic model. Mathematics and Computers in Simulation, 81 (10):2201–2214, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410004222.

Kavcic:2011:SPA

[349] I. Kavčič, M. Rogina, and T. Bosner. Singularly perturbed advection–diffusion– reaction problems: Comparison of operator-fitted methods. Mathematics and Computers in Simulation, 81 (10):2215–2224, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410004313.

Khalil:2011:DTP

[350] Ziad Khalil and Mazen Saad. Degenerate two-phase compressible immiscible flow in porous media: the case where the density of each phase depends on its own pressure. Mathematics and Computers in Simulation, 81 (10):2225–2233, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410004131.

Manouzi:2011:SDA

[351] Hassan Manouzi. SPDEs driven by additive and multiplicative white noises: a numerical study. Mathematics and Computers in Simulation, 81 (10):2234–2243, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410004210.

Mocenni:2011:LLS

[352] C. Mocenni, D. Madeo, and E. Sparacino. Linear least squares parameter estimation of nonlinear reaction diffusion equations. Mathematics and Computers in Simulation, 81 (10):2244–2257, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411001182.

Moncayo:2011:TNB

[353] M. Moncayo, J. F. Reinoso, and S. Amat. Tight numerical bounds for digital terrain modeling by interpolatory subdivision schemes. Mathematics and Computers in Simulation, 81 (10):2258–2269, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410004118.

Mourzenko:2011:TDN

[354] V. V. Mourzenko, I. I. Bogdanov, J.-F. Thovert, and P. M. Adler. Three-dimensional numerical simulation of single-phase transient compressible flows and well-tests in fractured formations. Mathematics and Computers in Simulation, 81(10):2270– 2281, June 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410004155.

Najib:2011:GED

[355] Khalid Najib and Carole Rosier. On the global existence for a degenerate elliptic–parabolic seawater intrusion problem. Mathematics and Computers in Simulation, 81(10):2282– 2295, June 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410004271.

Pereira:2011:SDC

[356] F. Pereira and A. Rahunanthan. A semi-discrete central scheme for the approximation of two-phase flows in three space dimensions. Mathematics and Computers in Simulation, 81 (10):2296–2306, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000528.

Reinoso:2011:OQI

[357] J. F. Reinoso and M. Moncayo. Optimal quality for image fusion with interpolatory parametric filters. Mathematics and Computers in Simulation, 81 (10):2307–2316, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410004052.

Rueda:2011:IEP

[358] M. Rueda, J. F. Muñoz, A. Arcos, and E. Álvarez. Indirect estimation of proportions in natural resource surveys. Mathematics and Computers in Simulation, 81(10):2317– 2325, June 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410004301.

Rust:2011:MMA

[359] Bert W. Rust. A mathematical model of atmospheric retention of manmade CO _{2} emissions. Mathematics and Computers in Simulation, 81 (10):2326–2336, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410004209.

Sablonniere:2011:PQI

[360] P. Sablonnière, D. Sbibih, and M. Tahrichi. Product quasi-interpolation method for weakly singular integral equation eigenvalue problem. Mathematics and Computers in Simulation, 81(10):2337–2345, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410004192.

Younes:2011:MVD

[361] A. Younes, M. Konz, M. Fahs, A. Zidane, and P. Huggenberger. Modelling variable density flow problems in heterogeneous porous media using the method of lines and advanced spatial discretization methods. Mathematics and Computers in Simulation, 81 (10):2346–2355, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000905.

Anonymous:2011:NIf

[362] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 81(10):2356, June 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001327.

Anonymous:2011:ICEf

[363] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 81(10):2357– 2363, June 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411001339.

Anonymous:2011:PJ

[364] Anonymous. Pages 2371–2594 (July 2011). Mathematics and Computers in Simulation, 81(11):??, July 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2011:EBg

[365] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 81(11):CO2, July 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001406.

Assous:2011:DDM

[366] Franck Assous, J. Segré, and E. Sonnendrücker. A domain decomposition method for the parallelization of a three-dimensional Maxwell solver based on a constrained formulation. Mathematics and Computers in Simulation, 81(11):2371– 2388, July 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411000784.

Lu:2011:EOF

[367] Jin-Ray Lu, Pei-Hsuan Lee, and I-Yuan Chuang. Estimation of oil firm’s systematic risk via composite time-varying models. Mathematics and Computers in Simulation, 81 (11):2389–2399, July 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000814.

Hsu:2011:SSG

[368] Ying-Lin Hsu, Jau-Chuan Ke, and Tzu-Hsin Liu. Standby system with general repair, reboot delay, switching failure and unreliable repair facility — a statistical standpoint. Mathematics and Computers in Simulation, 81 (11):2400–2413, July 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000826.

Mahmoudi:2011:BGP

[369] Eisa Mahmoudi. The beta generalized Pareto distribution with application to lifetime data. Mathematics and Computers in Simulation, 81 (11):2414–2430, July 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000851.

Bucker:2011:SOD

[370] H. Martin Bücker and Emil Slusanschi. Second-order derivatives of the general-purpose finite element package SEPRAN via source transformation. Mathematics and Computers in Simulation, 81(11):2431– 2439, July 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411000875.

Zhou:2011:SAA

[371] Jun Zhou, Weirui Zhao, Xiaohong Lv, and Huaping Zhu. Stability analysis of almost periodic solutions for delayed neural networks without global Lipschitz activation functions. Mathematics and Computers in Simulation, 81 (11):2440–2455, July 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000887.

Lanzendorfer:2011:MMJ

[372] Martin Lanzendörfer and Jan Stebel. On a mathematical model of journal bearing lubrication. Mathematics and Computers in Simulation, 81 (11):2456–2470, July 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000917.

Amiri:2011:BAP

[373] Mahmood Amiri, Esmaeil Davoodi-Bojd, Fariba Bahrami, and Mohsin Raza. Bifurcation analysis of the Poincaré map function of intracranial EEG signals in temporal lobe epilepsy patients. Mathematics and Computers in Simulation, 81(11):2471– 2491, July 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411000930.

Wang:2011:VIS

[374] Tiansong Wang, Jun Wang, Junhuan Zhang, and Wen Fang. Voter interacting systems applied to Chinese stock markets. Mathematics and Computers in Simulation, 81 (11):2492–2506, July 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000942.

Liu:2011:FMB

[375] Yang Liu and Shouwei Zhao. T – S fuzzy model-based impulsive control for chaotic systems and its application. Mathematics and Computers in Simulation, 81(11):2507– 2516, July 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411000978.

Ruspini:2011:SNC

[376] L. C. Ruspini, C. A. Dorao, and M. Fernandino. Simulation of a natural circulation loop using a least squares hp-adaptive solver. Mathematics and Computers in Simulation, 81 (11):2517–2528, July 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000991.

Huang:2011:WBM

[377] Shian-Chang Huang. Wavelet-based multi-resolution GARCH model for financial spillover effects. Mathematics and Computers in Simulation, 81 (11):2529–2539, July 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001091.

intHout:2011:SMC

[378] K. J. in ’t Hout and C. Mishra. Stability of the modified Craig– Sneyd scheme for two-dimensional convection–diffusion equations with mixed derivative term. Mathematics and Computers in Simulation, 81 (11):2540–2548, July 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001108.

Morel:2011:GCA

[379] C. Morel, R. Vlad, J.-Y. Morel, and D. Petreus. Generating chaotic attractors on a surface. Mathematics and Computers in Simulation, 81 (11):2549–2563, July 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001157.

Luque:2011:DMC

[380] A. Luque, J. Aldazabal, J. M. Martínez-Esnaola, A. Martín-Meizoso, J. Gil Sevillano, and R. S. Farr. Diffusional Monte Carlo model of liquid-phase sintering. Mathematics and Computers in Simulation, 81 (11):2564–2580, July 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001170.

Anonymous:2011:NIg

[381] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 81(11):2581, July 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411001431.

Anonymous:2011:ICEg

[382] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 81(11):2582– 2588, July 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411001443.

Anonymous:2011:PAb

[383] Anonymous. Pages 2595–2714 (August 2011). Mathematics and Computers in Simulation, 81(12):??, August 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2011:EBh

[384] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 81(12):CO2, August 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411001558.

Feng:2011:CCB

[385] Li Feng, Liu Yirong, and Li Hongwei. Center conditions and bifurcation of limit cycles at three-order nilpotent critical point in a septic Lyapunov system. Mathematics and Computers in Simulation, 81(12):

2595–2607, August 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541100111X.

Xin:2011:LSM

[386] Qiao Xin, Chunlai Mu, and Meng Li. The Lee–Seo model with regularization term for bimodal image segmentation. Mathematics and Computers in Simulation, 81(12):2608– 2616, August 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411001121.

Sari:2011:PPN

[387] B. Sari, O. Bachelier, J. Bosche, N. Maamri, and D. Mehdi. Pole placement in non connected regions for descriptor models. Mathematics and Computers in Simulation, 81(12): 2617–2631, August 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001145.

LeCoupanec:2011:MCB

[388] Erwan Le Coupanec and Joris C. G. Verschaeve. A mass conserving boundary condition for the lattice Boltzmann method for tangentially moving walls. Mathematics and Computers in Simulation, 81(12):2632– 2645, August 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411001169.

Calvo:2011:EGN

[389] M. Calvo, M. P. Laburta, J. I. Montijano, and L. Rández. Error growth in the numerical integration of periodic orbits. Mathematics and Computers in Simulation, 81(12):2646– 2661, August 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411001194.

Huang:2011:UUB

[390] Jun Huang, Zhengzhi Han, Xiushan Cai, and Leipo Liu. Uniformly ultimately bounded tracking control of linear differential inclusions with stochastic disturbance. Mathematics and Computers in Simulation, 81(12): 2662–2672, August 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411001200.

Sielemann:2011:QMR

[391] M. Sielemann and G. Schmitz. A quantitative metric for robustness of nonlinear algebraic equation solvers. Mathematics and Computers in Simulation, 81(12):2673– 2687, August 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411001224.

Frischmuth:2011:NCW

[392] Kurt Frischmuth and Dirk Langemann. Numerical calculation of wear in mechanical systems. Mathematics and Computers in Simulation, 81 (12):2688–2701, August 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001236.

Anonymous:2011:NIh

[393] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 81(12):2702, August 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001595.

Anonymous:2011:ICEh

[394] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 81(12):2703– 2708, August 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411001601.

Anonymous:2011:EBi

[395] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 82(1):CO2, September 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001984.

Roanes-Lozano:2011:NAC

[396] Eugenio Roanes-Lozano, Michael J. Wester, and Stanly Steinberg. Nonstandard applications of computer algebra. Mathematics and Computers in Simulation, 82(1):1, September 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475411001868.

See erratum [608].

Saenz-de-Cabezon:2011:CIP

[397] Eduardo Sáenz de Cabezón, L. Javier Hernández, M. Teresa Rivas, Esteban García-Ruiz, Vicente Marco, Ignacio Pérez-Moreno, and F. Javier Sáenz de Cabezón. A computer implementation of the partition of the unity procedure and its application to arthropod population dynamics. A case study on the European grape berry moth. Mathematics and Computers in Simulation, 82(1):2–14, September 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541000220X.

Roanes-Lozano:2011:LAD

[398] Eugenio Roanes-Lozano, Antonio Hernando, Jose Antonio Alonso, and Luis M. Laita. A logic approach to decision taking in a railway interlocking system using Maple. Mathematics and Computers in Simulation, 82 (1):15–28, September 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001874.

Hernando:2011:NAM

[399] Antonio Hernando. A new algebraic model for implementing expert systems represented under the ‘Concept-Attribute-Value’ paradigm. Mathematics and Computers in Simulation, 82 (1):29–43, September 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002211.

Garcia:2011:ACD

[400] R. E. Garcia, M. A. Lane, and B. M. Loft. Algebraic combinatorics of diametric magic circles. Mathematics and Computers in Simulation, 82 (1):44–53, September 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003198.

Recio:2011:ADH

[401] Tomas Recio, J. Rafael Sendra, Luis Felipe Tabera, and Carlos Villarino. Algorithmic detection of hypercircles. Mathematics and Computers in Simulation, 82(1):54–67, September 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002545.

Saenz-de-Cabezon:2011:CAA

[402] Eduardo Sáenz de Cabezón and Henry P. Wynn. Computational algebraic algorithms for the reliability of generalized k-out-of- n and related systems. Mathematics and Computers in Simulation, 82(1):68– 78, September 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541000265X.

Wester:2011:CIP

[403] Michael J. Wester, Yuzita Yaacob, and Stanly Steinberg. Computing integrals over polynomially defined regions and their boundaries in 2 and 3 dimensions. Mathematics and Computers in Simulation, 82(1): 79–101, September 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001522.

Spitaleri:2011:MAS

[404] Rosa Maria Spitaleri and Francesca Pistella. 8th meeting on Applied Scientific Computing and Tools Grid Generation, Approximation and Visualization. Mathematics and Computers in Simulation, 82(1):103, September 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541100187X.

Arbel:2011:OIP

[405] A. Arbel and A. Shklyar. Optimization of the iteration parameters of the simulation of incompressible flow. Mathematics and Computers in Simulation, 82(1):104– 117, September 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410003964.

Pezzella:2011:NAF

[406] Giuseppe Pezzella and Edoardo Bucchignani. Numerical assessment of the flowfield features at the exit of Scirocco plasma wind tunnel nozzle. Mathematics and Computers in Simulation, 82 (1):118–131, September 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000504.

Gori:2011:NEC

[407] L. Gori, E. Pellegrino, and E. Santi. Numerical evaluation of certain hypersingular integrals using refinable operators. Mathematics and Computers in Simulation, 82(1):132–143, September 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002302.

Deolmi:2011:BAW

[408] G. Deolmi, F. Marcuzzi, and M. Morandi Cecchi. The Best-Approximation Weighted-Residuals method for the steady convection–diffusion–reaction problem. Mathematics and Computers in Simulation, 82(1):144–162, September 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410003952.

Joubarne:2011:MBA

[409] Éric Joubarne and François Guibault. 3D Metric-based anisotropic mesh adaptation for vortex capture. Mathematics and Computers in Simulation, 82(1):163–180, September 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475411000747.

Rigatos:2011:DRE

[410] G. Rigatos and P. Siano. Design of robust electric power system stabilizers using Kharitonov’s theorem. Mathematics and Computers in Simulation, 82(1):181–191, September 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410002338.

Egidi:2011:RCT

[411] N. Egidi and P. Maponi. Residual correction techniques for the efficient solution of inverse scattering problems. Mathematics and Computers in Simulation, 82(1):192–204, September 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411000498.

Anonymous:2011:NIi

[412] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 82(1):205, September 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411002035.

Anonymous:2011:ICEi

[413] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 82(1):206–211, September 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411002047.

Anonymous:2011:PO

[414] Anonymous. Pages 213–364 (October 2011). Mathematics and Computers in Simulation, 82(2):??, October 2011.

CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2011:EBj

[415] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 82(2):CO2, October 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002242.

Hwang:2011:PST

[416] Tsorng-Chyi Hwang, Meng-Gu Chen, and Chia-Lin Chang. Price stabilization in the Taiwan hog and broiler industries: Evidence from a STAR approach. Mathematics and Computers in Simulation, 82(2):213– 219, October 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411000863.

Theeraek:2011:SCD

[417] P. Theeraek, S. Phongthanapanich, and P. Dechaumphai. Solving convection–diffusion-reaction equation by adaptive finite volume element method. Mathematics and Computers in Simulation, 82(2):220– 233, October 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411001510.

Gutierrez-Sanchez:2011:TPG

[418] R. Gutiérrez-Sánchez, A. Nafidi, A. Pascual, and E. Ramos-Ábalos. Three parameter gamma-type growth curve, using a stochastic gamma diffusion model: Computational statistical aspects and simulation. Mathematics and Computers in Simulation, 82 (2):234–243, October 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001662.

Kovalets:2011:ISL

[419] Ivan V. Kovalets, Spyros Andronopoulos, Alexander G. Venetsanos, and John G. Bartzis. Identification of strength and location of stationary point source of atmospheric pollutant in urban conditions using computational fluid dynamics model. Mathematics and Computers in Simulation, 82(2):244–257, October 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001686.

Li:2011:DMG

[420] Xiaolin Li. Development of a meshless Galerkin boundary node method for viscous fluid flows. Mathematics and Computers in Simulation, 82 (2):258–280, October 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001704.

Krishna:2011:REL

[421] Hare Krishna and Kapil Kumar. Reliability estimation in Lindley distribution with progressively type II right censored sample. Mathematics and Computers in Simulation, 82 (2):281–294, October 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001716.

Sabelfeld:2011:SRA

[422] K. K. Sabelfeld and N. S. Mozartova. Sparsified Randomization algorithms for low rank approximations and applications to integral equations and inhomogeneous random field simulation. Mathematics and Computers in Simulation, 82(2):295– 317, October 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411001777.

Tian:2011:MQA

[423] Yuan Tian, Kaibiao Sun, and Lansun Chen. Modelling and qualitative analysis of a predator–prey system with state-dependent impulsive effects. Mathematics and Computers in Simulation, 82(2):318– 331, October 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411001789.

Xiong:2011:RMP

[424] Xiangtuan Xiong, Liqin Zhu, and Ming Li. Regularization methods for a problem of analytic continuation. Mathematics and Computers in Simulation, 82(2):332–345, October 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001807.

Ye:2011:RRB

[425] Yonggang Ye, Tsangyao Chang, Ken Hung, and Yang-Cheng Lu. Revisiting rational bubbles in the G-7 stock markets using the Fourier unit root test and the nonparametric rank test for cointegration. Mathematics and Computers in Simulation, 82 (2):346–357, October 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411001856.

Anonymous:2011:NIj

[426] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 82(2):358, October 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541100228X.

Anonymous:2011:ICEj

[427] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 82(2):359– 364, October 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411002291.

Anonymous:2011:EBk

[428] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 82(3):CO2, November 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411002394.

Troch:2011:E

[429] Inge Troch and Felix Breitenecker. Editorial. Mathematics and Computers in Simulation, 82(3):365–368, November 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411002370.

Dochain:2011:ESC

[430] Denis Dochain, Michel Perrier, and Martin Guay. Extremum seeking control and its application to process and reaction systems: a survey. Mathematics and Computers in Simulation, 82 (3):369–380, November 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003290.

Mathis:2011:DNC

[431] W. Mathis and J.-K. Bremer. Design of nonlinear CMOS circuits in the nano-GHz Era and its mathematical challenges. Mathematics and Computers in Simulation, 82(3): 381–391, November 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003204.

Steinboeck:2011:FSM

[432] A. Steinboeck, D. Wild, T. Kiefer, and A. Kugi. A fast simulation method for 1D heat conduction. Mathematics and Computers in Simulation, 82 (3):392–403, November 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541000323X.

Fassbender:2011:MTS

[433] Heike Faßbender and Andreas Soppa. Machine tool simulation based on reduced order FE models. Mathematics and Computers in Simulation, 82 (3):404–413, November 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410003277.

Karer:2011:FCC

[434] Gorazd Karer, Gašper Mušič, Igor Škrjanc, and Borut Zupančič. Feedforward control of a class of hybrid systems using an inverse model. Mathematics and Computers in Simulation, 82(3):414–427, November 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475410003228.

Piera:2011:CPN

[435] M. A. Piera and G. Mušič. Coloured Petri net scheduling models: Timed state space exploration shortages. Mathematics and Computers in Simulation, 82(3):428–441, November 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475410003216.

Heidl:2011:LSR

[436] W. Heidl, C. Eitzinger, M. Gyimesi, and F. Breitenecker. Learning over sets with Recurrent Neural Networks: an empirical categorization of aggregation functions. Mathematics and Computers in Simulation, 82(3):442– 449, November 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410003253.

Heikkinen:2011:SAB

[437] M. Heikkinen, H. Poutiainen, M. Liukkonen, T. Heikkinen, and Y. Hiltunen. Subtraction analysis based on selforganizing maps for an industrial wastewater treatment process. Mathematics and Computers in Simulation, 82(3):450–459, November 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475410003289.

Hadlich:2011:DFD

[438] Frieder Hadlich, Katharina Nöh, and Wolfgang Wiechert. Determination of flux directions by thermodynamic network analysis: Computing informative metabolite pools. Mathematics and Computers in Simulation, 82(3): 460–470, November 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003320.

Dabnichki:2011:UFM

[439] Peter Dabnichki. Unsteady fluid mechanics effects in water based human locomotion. Mathematics and Computers in Simulation, 82(3):471– 482, November 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411002205.

Rosenkranz:2011:IMS

[440] S. Rosenkranz, C. Mayer, J. Kropf, and S. Wassertheurer. Intelligent multichannel sensors for pulse wave analysis. Mathematics and Computers in Simulation, 82(3):483–493, November 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410003307.

Atanasijevic-Kunc:2011:SAC

[441] M. Atanasijević-Kunc, J. Drinovec, S. Ručigaj, and A. Mrhar. Simulation analysis of coronary heart disease, congestive heart failure and end-stage renal disease economic burden. Mathematics and Computers in Simulation, 82(3):494–507, November 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475410003319.

Rojko:2011:TMU

[442] A. Rojko, D. Hercog, and K. Jezernik. E-training in mechatronics using innovative remote laboratory. Mathematics and Computers in Simulation, 82 (3):508–516, November 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410003241.

Logar:2011:ARL

[443] V. Logar, R. Karba, M. Papič, and M. Atanasijević-Kunc. Artificial and real laboratory environment in an elearning competition. Mathematics and Computers in Simulation, 82(3): 517–524, November 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003265.

Anonymous:2011:NIk

[444] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 82(3):525, November 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002461.

Anonymous:2011:ICEk

[445] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 82(3):526–531, November 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411002473.

Anonymous:2011:PD

[446] Anonymous. Pages 533–776 (December 2011). Mathematics and Computers in Simulation, 82(4):??, December 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2011:EBl

[447] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 82(4):CO2, December 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002710.

Buzzi:2011:LCC

[448] C. A. Buzzi, J. Llibre, and J. C. Medrado. On the limit cycles of a class of piecewise linear differential systems in R^{4} with two zones. Mathematics and Computers in Simulation, 82 (4):533–539, December 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001832.

Zhang:2011:TEI

[449] Hao Zhang, Xiao-Ping Yang, Xi-Kui Ma, and Feng Zheng. Theoretical and experimental investigation of bidirectional Hopf bifurcations in cascade DC–DC buck converters. Mathematics and Computers in Simulation, 82 (4):540–557, December 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002199.

Krstic:2011:ESP

[450] Marija Krstić. The effect of stochastic perturbation on a nonlinear delay malaria epidemic model. Mathematics and Computers in Simulation, 82 (4):558–569, December 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002187.

Shi:2011:LFW

[451] Yan Shi and Xiaoyuan Yang. The lifting factorization of wavelet bi-frames with arbitrary generators. Mathematics and Computers in Simulation, 82 (4):570–589, December 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002345.

Nguyen:2011:SCC

[452] Le Hoa Nguyen and Keum-Shik Hong. Synchronization of coupled chaotic FitzHugh–Nagumo neurons via Lyapunov functions. Mathematics and Computers in Simulation, 82(4):590– 603, December 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411002540.

Lu:2011:MYV

[453] Zhaoyang Lu. Modeling the yearly Value-at-Risk for operational risk in Chinese commercial banks. Mathematics and Computers in Simulation, 82 (4):604–616, December 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002527.

Mishra:2011:EEC

[454] Nachiketa Mishra and Sanyasiraju V. S. S. Yedida. Efficient exponential compact higher order difference scheme for convection dominated problems. Mathematics and Computers in Simulation, 82(4):617– 628, December 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411002539.

Al-arydah:2011:ASM

[455] Mo’tassem Al-arydah and Robert Smith. An age-structured model of human papillomavirus vaccination. Mathematics and Computers in Simulation, 82(4):629–652, December 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411002552.

Gao:2011:MAH

[456] Ting Gao, Wendi Wang, and Xianning Liu. Mathematical analysis of an HIV model with impulsive antiretroviral drug doses. Mathematics and Computers in Simulation, 82(4): 653–665, December 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002564.

Azadeh:2011:IAN

[457] A. Azadeh, Z. S. Faiz, S. M. Asadzadeh, and R. Tavakkoli-Moghaddam. An integrated artificial neural networkcomputer simulation for optimization of complex tandem queue systems. Mathematics and Computers in Simulation, 82(4):666–678, December 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411002588.

Dorini:2011:LAE

[458] F. A. Dorini and M. C. C. Cunha. On the linear advection equation subject to random velocity fields. Mathematics and Computers in Simulation, 82 (4):679–690, December 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541100259X.

Dingec:2011:UCP

[459] Kemal Dinçer Dingeç and Wolfgang Hörmann. Using the continuous price as control variate for discretely monitored options. Mathematics and Computers in Simulation, 82(4):691– 704, December 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411002606.

Tsai:2011:SBS

[460] Shing Chih Tsai. Selecting the best simulated system with weighted control-variate estimators. Mathematics and Computers in Simulation, 82 (4):705–717, December 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002618.

Gross:2011:GBE

[461] Michael Groß and Peter Betsch. Galerkin-based energy–momentum consistent time-stepping algorithms for classical nonlinear thermo-elastodynamics. Mathematics and Computers in Simulation, 82(4):718–770, December 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847541100262X.

Anonymous:2011:NIl

[462] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 82(4):771, December 2011. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411002758.

Anonymous:2011:ICEl

[463] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 82(4):772–776, December 2011. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847541100276X.

Anonymous:2012:PJa

[464] Anonymous. Pages 777–944 (January 2012). Mathematics and Computers in Simulation, 82(5):??, January 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2012:EBa

[465] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 82(5):CO2, January 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541200016X.

Han:2012:TRL

[466] Lili Han, Fangxiang Wu, Jie Sheng, and Feng Ding. Two recursive least squares parameter estimation algorithms for multirate multiple-input systems by using the auxiliary model. Mathematics and Computers in Simulation, 82 (5):777–789, January 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002576.

Vegas-Sanchez-Ferrero:2012:DCM

[467] Gonzalo Vegas-Sánchez-Ferrero, Santiago Aja-Fernández, Marcos Martín-Fernández, and César Palencia. A direct calculation of moments of the sample variance. Mathematics and Computers in Simulation, 82(5):790– 804, January 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411002631.

Villegas:2012:APC

[468] M. Villegas, F. Augustin, A. Gilg, A. Hmaidi, and U. Wever. Application of the Polynomial Chaos Expansion to the simulation of chemical reactors with uncertainties. Mathematics and Computers in Simulation, 82 (5):805–817, January 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002680.

Akdag:2012:IAS

[469] M. Akdağ, H. Karagülle, and L. Malgaca. An integrated approach for simulation of mechatronic systems applied to a hexapod robot. Mathematics and Computers in Simulation, 82 (5):818–835, January 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002813.

Torres:2012:ENO

[470] Lizeth Torres, Gildas Besançon, Didier Georges, and Cristina Verde. Exponential nonlinear observer for parametric identification and synchronization of chaotic systems. Mathematics and Computers in Simulation, 82 (5):836–846, January 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411002825.

Liu:2012:VHG

[471] Yan Liu and Dirk Söffker. Variable high-gain disturbance observer design with online adaption of observer gains embedded in numerical integration. Mathematics and Computers in Simulation, 82(5):847– 857, January 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847541200002X.

Ehlers:2012:CTC

[472] Ricardo S. Ehlers. Computational tools for comparing asymmetric GARCH models via Bayes factors. Mathematics and Computers in Simulation, 82 (5):858–867, January 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000043.

Lindsay:2012:SCP

[473] A. E. Lindsay and D. R. Brecher. Simulation of the CEV process and the local martingale property. Mathematics and Computers in Simulation, 82 (5):868–878, January 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000079.

Pazouki:2012:PCD

[474] A. Pazouki, H. Mazhar, and D. Negrut. Parallel collision detection of ellipsoids with applications in large scale multibody dynamics. Mathematics and Computers in Simulation, 82 (5):879–894, January 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000080.

Yu:2012:ELS

[475] Juan Yu, Cheng Hu, Haijun Jiang, and Zhidong Teng. Exponential lag synchronization for delayed fuzzy cellular neural networks via periodically intermittent control. Mathematics and Computers in Simulation, 82 (5):895–908, January 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000092.

Wu:2012:OCM

[476] Lin Wu, François-Xavier Le Dimet, Philippe de Reffye, Bao-Gang Hu, Paul-Henry Cournède, and Meng-Zhen Kang. An optimal control methodology for plant growth — Case study of a water supply problem of sunflower. Mathematics and Computers in Simulation, 82(5):909–923, January 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000109.

Casals:2012:GSS

[477] J. Casals, A. García-Hiernaux, and M. Jerez. From general state-space to VARMAX models. Mathematics and Computers in Simulation, 82 (5):924–936, January 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000110.

Anonymous:2012:NIa

[478] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 82(5):937, January 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000201.

Anonymous:2012:ICEa

[479] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 82(5):938– 943, January 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412000213.

Anonymous:2012:EBb

[480] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 82(6):CO2, February 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000316.

Anonymous:2012:Ca

[481] Anonymous. Contents. Mathematics and Computers in Simulation, 82 (6):v–vi, February 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000365.

Taha:2012:Fa

[482] Thiab R. Taha. Foreword. Mathematics and Computers in Simulation, 82(6):945, February 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000298.

Achilleos:2012:EBE

[483] V. A. Achilleos, T. P. Horikis, G. Theocharis, P. G. Kevrekidis, and D. J. Frantzeskakis. Excited Bose– Einstein condensates: Quadrupole oscillations and dark solitons. Mathematics and Computers in Simulation, 82 (6):946–957, February 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002569.

Akers:2012:GCG

[484] Benjamin Akers. The generation of capillary-gravity solitary waves by a surface pressure forcing. Mathematics and Computers in Simulation, 82 (6):958–967, February 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003022.

Ambrose:2012:NSB

[485] David M. Ambrose and Michael Siegel. A non-stiff boundary integral method for 3D porous media flow with surface tension. Mathematics and Computers in Simulation, 82(6): 968–983, February 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001813.

Antonopoulos:2012:NSC

[486] D. C. Antonopoulos and V. A. Dougalis. Numerical solution of the ‘classical’ Boussinesq system. Mathematics and Computers in Simulation, 82 (6):984–1007, February 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002515.

Barannyk:2012:SRT

[487] Lyudmyla L. Barannyk, Demetrios T. Papageorgiou, and Peter G. Petropoulos. Suppression of Rayleigh–Taylor instability using electric fields. Mathematics and Computers in Simulation, 82(6):1008–1016, February 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475410004039.

Bociu:2012:WES

[488] Lorena Bociu, Mohammad Rammaha, and Daniel Toundykov. Wave equations with super-critical interior and boundary nonlinearities. Mathematics and Computers in Simulation, 82(6): 1017–1029, February 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001133.

Buckingham:2012:SSC

[489] Robert Buckingham. Semiclassical spectral confinement for the sine-Gordon equation. Mathematics and Computers in Simulation, 82(6):1030– 1037, February 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002259.

Carter:2012:PWS

[490] John D. Carter. Plane-wave solutions of a dissipative generalization of the vector nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 82(6):1038– 1046, February 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002880.

Cascaval:2012:BMP

[491] Radu C. Cascaval. A Boussinesq model for pressure and flow velocity waves in arterial segments. Mathematics and Computers in Simulation, 82(6): 1047–1055, February 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000844.

Chen:2012:LCP

[492] Peter Y. P. Chen and Boris A. Malomed. Lanczos–Chebyshev pseudospectral methods for wave-propagation problems. Mathematics and Computers in Simulation, 82(6):1056– 1068, February 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411001388.

Christov:2012:HSZ

[493] Ivan C. Christov. Hidden solitons in the Zabusky–Kruskal experiment: Analysis using the periodic, inverse scattering transform. Mathematics and Computers in Simulation, 82(6): 1069–1078, February 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001849.

Christov:2012:NIA

[494] Christo I. Christov. Numerical implementation of the asymptotic boundary conditions for steadily propagating 2D solitons of Boussinesq type equations. Mathematics and Computers in Simulation, 82(6):1079– 1092, February 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002673.

Diaz-Otero:2012:EAL

[495] F. J. Diaz-Otero and P. Chamorro-Posada. Effects of the amplifier location in interchannel soliton collisions in periodic dispersion maps in the presence of third order dispersion. Mathematics and Computers in Simulation, 82(6):1093– 1101, February 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410003940.

Fedele:2012:ECO

[496] Francesco Fedele, Guillermo Gallego, Anthony Yezzi, Alvise Benetazzo, Luigi Cavaleri, Mauro Sclavo, and Mauro Bastianini. Euler characteristics of oceanic sea states. Mathematics and Computers in Simulation, 82(6): 1102–1111, February 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001212.

Gambino:2012:TIT

[497] G. Gambino, M. C. Lombardo, and M. Sammartino. Turing instability and traveling fronts for a nonlinear reaction–diffusion system with cross-diffusion. Mathematics and Computers in Simulation, 82(6):1112– 1132, February 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411002692.

Ghazaryan:2012:EHL

[498] Anna Ghazaryan and Christopher Jones. On the existence of high Lewis number combustion fronts. Mathematics and Computers in Simulation, 82(6):1133–1141, February 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475410001400.

Anonymous:2012:NIb

[499] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 82(6):1142, February 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000377.

Anonymous:2012:ICEb

[500] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 82(6):1143– 1148, February 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412000389.

Anonymous:2012:EBc

[501] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 82(7):CO2, March 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000493.

Anonymous:2012:Cb

[502] Anonymous. Contents. Mathematics and Computers in Simulation, 82(7):v–vi, March 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000547.

Taha:2012:Fb

[503] Thiab R. Taha. Foreword. Mathematics and Computers in Simulation, 82(7):1149, March 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541200047X.

Caplan:2012:ESS

[504] R. M. Caplan, R. Carretero-González, P. G. Kevrekidis, and B. A. Malomed. Existence, stability, and scattering of bright vortices in the cubic–quintic nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 82(7):1150–1171, March 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000250.

Henderson:2012:BFI

[505] Diane Henderson and Harvey Segur. The Benjamin–Feir instability and propagation of swell across the Pacific. Mathematics and Computers in Simulation, 82(7):1172– 1184, March 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002041.

Kao:2012:NSK

[506] Chiu-Yen Kao and Yuji Kodama. Numerical study of the KP equation for non-periodic waves. Mathematics and Computers in Simulation, 82 (7):1185–1218, March 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001886.

Kim:2012:NSS

[507] Hongjoong Kim, Won-Soung Bae, and Jeongwhan Choi. Numerical stability of symmetric solitary-wave-like waves of a two-layer fluid — Forced modified KdV equation. Mathematics and Computers in Simulation, 82(7): 1219–1227, March 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001765.

Korotkevich:2012:ICI

[508] A. O. Korotkevich. Influence of the condensate and inverse cascade on the direct cascade in wave turbulence. Mathematics and Computers in Simulation, 82(7):1228– 1238, March 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541000234X.

LeMesurier:2012:SDO

[509] Brenton LeMesurier. Studying Davydov’s ODE model of wave motion in α-helix protein using exactly energy– momentum conserving discretizations for Hamiltonian systems. Mathematics and Computers in Simulation, 82 (7):1239–1248, March 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410004325.

Lindley:2012:SSS

[510] Brandon S. Lindley, M. Gregory Forest, Breannan D. Smith, Sorin M. Mitran, and David B. Hill. Spatial stress and strain distributions of viscoelastic layers in oscillatory shear. Mathematics and Computers in Simulation, 82 (7):1249–1257, March 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002685.

Ludu:2012:VSA

[511] Andrei Ludu, Milorad V. Milošević, and François M. Peeters. Vortex states in axially symmetric superconductors in applied magnetic field. Mathematics and Computers in Simulation, 82 (7):1258–1270, March 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000249.

Mickens:2012:WFB

[512] Ronald E. Mickens. Wave front behavior of traveling wave solutions for a PDE having square-root dynamics. Mathematics and Computers in Simulation, 82(7):1271– 1277, March 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002892.

Pankavich:2012:PMC

[513] Stephen Pankavich. A particle method for a collisionless plasma with infinite mass. Mathematics and Computers in Simulation, 82(7):1278– 1286, March 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411001844.

Purrington:2012:NCW

[514] Robert D. Purrington and Guy V. Norton. A numerical comparison of the Westervelt equation with viscous attenuation and a causal propagation operator. Mathematics and Computers in Simulation, 82(7):1287– 1297, March 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001801.

Shevtsova:2012:GVS

[515] O. N. Shevtsova. Giant vortex structures in mesoscopic spherical type-II superconducting samples. Mathematics and Computers in Simulation, 82 (7):1298–1307, March 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002363.

Tamm:2012:PSW

[516] Kert Tamm and Andrus Salupere. On the propagation of 1D solitary waves in Mindlin-type microstructured solids. Mathematics and Computers in Simulation, 82(7):1308– 1320, March 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002260.

Todorov:2012:CDE

[517] M. D. Todorov and C. I. Christov. Collision dynamics of elliptically polarized solitons in Coupled Nonlinear Schrödinger Equations. Mathematics and Computers in Simulation, 82 (7):1321–1332, March 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410001394.

Triki:2012:SWS

[518] Houria Triki and Thiab R. Taha. Solitary wave solutions for a higher order nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 82(7):1333–1340, March 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002679.

Tutiya:2012:SSS

[519] Yohei Tutiya and Jun’ichi Shiraishi. On some special solutions to periodic Benjamin–Ono equation with discrete Laplacian. Mathematics and Computers in Simulation, 82(7):1341– 1347, March 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001552.

Vallikivi:2012:NSP

[520] Margit Vallikivi, Andrus Salupere, and Hui-Hui Dai. Numerical simulation of propagation of solitary deformation waves in a compressible hyperelastic rod. Mathematics and Computers in Simulation, 82(7):1348– 1362, March 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411001790.

Runzhang:2012:IPN

[521] Xu Runzhang and Liu Yacheng. Ill-posedness of nonlinear parabolic equation with critical initial condition. Mathematics and Computers in Simulation, 82(7):1363– 1374, March 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002995.

Anonymous:2012:NIc

[522] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 82(7):1375, March 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000559.

Anonymous:2012:ICEc

[523] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 82(7):1376– 1381, March 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412000560.

Anonymous:2012:PA

[524] Anonymous. Pages 1383–1550 (April 2012). Mathematics and Computers in Simulation, 82(8):??, April 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2012:EBd

[525] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 82(8):CO2, April 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000778.

Wang:2012:RAB

[526] Liang Wang and Yimin Shi. Reliability analysis based on progressively firstfailure-censored samples for the proportional hazard rate model. Mathematics and Computers in Simulation, 82(8):1383–1395, April 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000031.

Ciria:2012:IEA

[527] Héctor Manuel Camué Ciria, Luis Enrique Bergues Cabrales, Andrés Ramírez Aguilera, Juan José Godina Nava, Javier Antonio González Joa, Raudel Peña García, Gustavo Sierra González, Maraelys Morales González, Miriam Fariñas Salas, Manuel Verdecia Jarque, Tamara Rubio González, Miguel Angel O’Farril Mateus, Soraida Candida Acosta Brooks, Fabiola Suárez Palencia, Lisset Zamora Ortiz, María Cristina Céspedes Quevedo, Sarah Edward Seringe, Yadira Mesa Mariño, and Idelisa Bergues Cabrales. Influence of electrode array parameters used in electrotherapy on tumor growth kinetics: a mathematical simulation. Mathematics and Computers in Simulation, 82 (8):1396–1406, April 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000055.

Matcovschi:2012:DIE

[528] Mihaela-Hanako Matcovschi and Octavian Pastravanu. Diagonally invariant exponential stability and stabilizability of switching linear systems. Mathematics and Computers in Simulation, 82 (8):1407–1418, April 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000067.

Martel-Escobar:2012:IBR

[529] M. Martel-Escobar, A. Hernández-Bastida, and F. J. Vázquez-Polo. On the independence between risk profiles in the compound collective risk actuarial model. Mathematics and Computers in Simulation, 82 (8):1419–1431, April 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000134.

Zhao:2012:DEM

[530] Min Zhao, Xitao Wang, Hengguo Yu, and Jun Zhu. Dynamics of an ecological model with impulsive control strategy and distributed time delay. Mathematics and Computers in Simulation, 82(8):1432–1444, April 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000146.

Bin-Mohsin:2012:DIB

[531] B. Bin-Mohsin and D. Lesnic. Determination of inner boundaries in modified Helmholtz inverse geometric problems using the method of fundamental solutions. Mathematics and Computers in Simulation, 82 (8):1445–1458, April 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000262.

Spandl:2012:CCI

[532] Christoph Spandl. Computational complexity of iterated maps on the interval. Mathematics and Computers in Simulation, 82(8):1459– 1477, April 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412000274.

Cohen:2012:NDS

[533] David Cohen. On the numerical discretisation of stochastic oscillators. Mathematics and Computers in Simulation, 82(8):1478– 1495, April 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412000286.

Zhang:2012:FDF

[534] Tong Zhang and Yinnian He. Fully discrete finite element method based on pressure stabilization for the transient Stokes equations. Mathematics and Computers in Simulation, 82 (8):1496–1515, April 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000468.

Leyva:2012:DOB

[535] R. Leyva, U. Ribes-Mallada, P. Garces, and J. F. Reynaud. Design and optimization of buck and double buck converters by means of geometric programming. Mathematics and Computers in Simulation, 82(8): 1516–1530, April 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000626.

Li:2012:MSE

[536] Dingshi Li, Danhua He, and Daoyi Xu. Mean square exponential stability of impulsive stochastic reaction– diffusion Cohen–Grossberg neural networks with delays. Mathematics and Computers in Simulation, 82(8): 1531–1543, April 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000651.

Anonymous:2012:NId

[537] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 82(8):1544, April 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541200081X.

Anonymous:2012:ICEd

[538] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 82(8):1545– 1550, April 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412000821.

Anonymous:2012:PM

[539] Anonymous. Pages 1551–1718 (May 2012). Mathematics and Computers in Simulation, 82(9):??, May 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2012:EBe

[540] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 82(9):CO2, May 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541200095X.

Wang:2012:GDT

[541] Xia Wang and Fangqi Chen. Global dynamics of two coupled parametrically excited van der Pol oscillators. Mathematics and Computers in Simulation, 82(9):1551–1571, May 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000456.

Yan:2012:SAF

[542] Ye Yan and Chunhai Kou. Stability analysis for a fractional differential model of HIV infection of CD4 ^{+} T-cells with time delay. Mathematics and Computers in Simulation, 82 (9):1572–1585, May 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000663.

Amat:2012:AII

[543] S. Amat, J. Ruiz, and J. C. Trillo. Adaptive interpolation of images using a new nonlinear cell-average scheme. Mathematics and Computers in Simulation, 82(9):1586– 1596, May 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412000675.

Rionero:2012:LTB

[544] Salvatore Rionero and Maria Vitiello. Long-time behavior of the solutions of Murray–Thomas model for interacting chemicals. Mathematics and Computers in Simulation, 82 (9):1597–1614, May 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000687.

Singh:2012:LSS

[545] Bhupendra Singh and Puneet Kumar Gupta. Load-sharing system model and its application to the real data set. Mathematics and Computers in Simulation, 82(9):1615– 1629, May 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412000699.

Kamrani:2012:SCM

[546] Minoo Kamrani and S. Mohammad Hosseini. Spectral collocation method for stochastic Burgers equation driven by additive noise. Mathematics and Computers in Simulation, 82(9):1630– 1644, May 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412000729.

Heinrich:2012:FVM

[547] Ch. Heinrich, B. Simeon, and St. Boschert. A finite volume method on NURBS geometries and its application in isogeometric fluid–structure interaction. Mathematics and Computers in Simulation, 82(9):1645– 1666, May 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412000730.

Barbiero:2012:GDP

[548] Alessandro Barbiero. A general discretization procedure for reliability computation in complex stress– strength models. Mathematics and Computers in Simulation, 82(9):1667– 1676, May 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412000742.

Gonzalez:2012:PGD

[549] D. González, F. Masson, F. Poulhaon, A. Leygue, E. Cueto, and F. Chinesta. Proper Generalized Decomposition based dynamic data driven inverse identification. Mathematics and Computers in Simulation, 82(9):1677–1695, May 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000754.

Krajnc:2012:MDE

[550] Marjeta Krajnc and Vito Vitrih. Motion design with Euler–Rodrigues frames of quintic Pythagorean-hodograph curves. Mathematics and Computers in Simulation, 82(9):1696– 1711, May 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412000882.

Anonymous:2012:NIe

[551] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 82(9):1712, May 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000997.

Anonymous:2012:ICEe

[552] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 82(9):1713– 1718, May 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412001000.

Anonymous:2012:EBf

[553] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 82(10):CO2, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412001577.

Anonymous:2012:P

[554] Anonymous. Preface. Mathematics and Computers in Simulation, 82 (10):1719–1720, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541200153X.

Lukas:2012:SOM

[555] D. Lukáš, K. Postava, and O. Životský. A shape optimization method for nonlinear axisymmetric magnetostatics using a coupling of finite and boundary elements. Mathematics and Computers in Simulation, 82(10):1721– 1731, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411000553.

Janovska:2012:NCP

[556] Drahoslava Janovská, Vladimír Janovský, and Kunio Tanabe. A note on computation of pseudospectra. Mathematics and Computers in Simulation, 82 (10):1732–1743, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000596.

Rohan:2012:MFS

[557] E. Rohan and R. Cimrman. Multiscale FE simulation of diffusiondeformation processes in homogenized dual-porous media. Mathematics and Computers in Simulation, 82 (10):1744–1772, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000966.

Brandner:2012:HOW

[558] Marek Brandner, Jiří Egermaier, and Hana Kopincová. High order well-balanced scheme for river flow modeling. Mathematics and Computers in Simulation, 82(10):1773– 1787, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001255.

Pultarova:2012:PCP

[559] Ivana Pultarová. Preconditioning of the coarse problem in the method of balanced domain decomposition by constraints. Mathematics and Computers in Simulation, 82(10):1788– 1798, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411001819.

Sistek:2012:FBS

[560] Jakub Šístek, Marta Čertíková, Pavel Burda, and Jaroslav Novotný. Facebased selection of corners in 3D substructuring. Mathematics and Computers in Simulation, 82(10):1799– 1811, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411001820.

Mandel:2012:ABT

[561] Jan Mandel, Bedřich Sousedík, and Jakub Šístek. Adaptive BDDC in three dimensions. Mathematics and Computers in Simulation, 82(10):1812– 1831, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411000954.

Vimmr:2012:ASG

[562] Jan Vimmr, Hynek Klášterka, and Marek Hajžman. Analytical solution of gaseous slip flow between two parallel plates described by the Oseen equation. Mathematics and Computers in Simulation, 82(10):1832– 1840, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412000870.

Jodar:2012:RMH

[563] L. Jódar, J.-C. Cortés, and L. Villafuerte. Random mixed hyperbolic models: Numerical analysis and computing. Mathematics and Computers in Simulation, 82(10):1841– 1852, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411000309.

Hron:2012:TCR

[564] J. Hron, J. Kratochvíl, J. Málek, K. R. Rajagopal, and K. Tůma. A thermodynamically compatible rate type fluid to describe the response of asphalt. Mathematics and Computers in Simulation, 82(10):1853–1873, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000899.

Buric:2012:COP

[565] Lubor Buřič and Vladimír Janovský. Continuation of oscillatory patterns in a road traffic model of three cars. Mathematics and Computers in Simulation, 82(10):1874–1893, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001674.

Jarosova:2012:PPT

[566] Marta Jarošová, Axel Klawonn, and Oliver Rheinbach. Projector preconditioning and transformation of basis in FETI-DP algorithms for contact problems. Mathematics and Computers in Simulation, 82(10):1894– 1907, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410003939.

Hokr:2012:FCA

[567] Milan Hokr and Dalibor Frydrych. Fully coupled algorithm for heat and water transport – Estimation of nonlinear parameters based on the experimental data. Mathematics and Computers in Simulation, 82(10):1908– 1918, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001474.

Kolditz:2012:NST

[568] O. Kolditz, S. Bauer, N. Böttcher, D. Elsworth, U.-J. Görke, C.-I. Mc-Dermott, C.-H. Park, A. K. Singh, J. Taron, and W. Wang. Numerical simulation of two-phase flow in deformable porous media: Application to carbon dioxide storage in the subsurface. Mathematics and Computers in Simulation, 82(10):1919– 1935, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001450.

Haslinger:2012:ANR

[569] J. Haslinger, R. Kučera, O. Vlach, and C. C. Baniotopoulos. Approximation and numerical realization of 3D quasistatic contact problems with Coulomb friction. Mathematics and Computers in Simulation, 82 (10):1936–1951, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411000310.

Axelsson:2012:MEP

[570] Owe Axelsson. Macro-elementwise preconditioning methods. Mathematics and Computers in Simulation, 82 (10):1952–1963, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001401.

Boyanova:2012:MPG

[571] P. Boyanova, I. Georgiev, S. Margenov, and L. Zikatanov. Multilevel preconditioning of graph-Laplacians: Polynomial approximation of the pivot blocks inverses. Mathematics and Computers in Simulation, 82(10):1964– 1971, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001449.

Company:2012:CSN

[572] Rafael Company, Lucas Jódar, and José-Ramón Pintos. A consistent stable numerical scheme for a nonlinear option pricing model in illiquid markets. Mathematics and Computers in Simulation, 82(10):1972– 1985, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001436.

Antil:2012:ROM

[573] Harbir Antil, Matthias Heinkenschloss, Ronald H. W. Hoppe, Christopher Linsenmann, and Achim Wixforth. Reduced order modeling based shape optimization of surface acoustic wave driven microfluidic biochips. Mathematics and Computers in Simulation, 82(10):1986–2003, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003496.

Sysala:2012:AMS

[574] Stanislav Sysala. Application of a modified semismooth Newton method to some elasto-plastic problems. Mathematics and Computers in Simulation, 82(10):2004–2021, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001292.

Papacek:2012:MSP

[575] Štěpán Papáček, Ctirad Matonoha, Václav Štumbauer, and Dalibor Štys. Modelling and simulation of photosynthetic microorganism growth: random walk vs. finite difference method. Mathematics and Computers in Simulation, 82(10):2022–2032, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001728.

Vejchodsky:2012:CBP

[576] Tomáš Vejchodský. Complementarity based a posteriori error estimates and their properties. Mathematics and Computers in Simulation, 82 (10):2033–2046, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001509.

Janovsky:2012:CNU

[577] Vladimír Janovský and Tomáš Ligurský. Computing non unique solutions of the Coulomb friction problem. Mathematics and Computers in Simulation, 82(10):2047–2061, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410004295.

Anonymous:2012:NIf

[578] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 82(10):2062, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001632.

Anonymous:2012:ICEf

[579] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 82(10):2063– 2067, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001644.

Anonymous:2012:PJb

[580] Anonymous. Pages 2069–2288 (July 2012). Mathematics and Computers in Simulation, 82(11):??, July 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2012:EBg

[581] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 82(11):CO2, July 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412001875.

Gao:2012:ESN

[582] Xing Gao, Weijun Zhong, and Shue Mei. Equilibrium stability of a nonlinear heterogeneous duopoly game with extrapolative foresight. Mathematics and Computers in Simulation, 82 (11):2069–2078, July 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000912.

Wang:2012:AGS

[583] Joanna J. J. Wang. On asymmetric generalised t stochastic volatility models. Mathematics and Computers in Simulation, 82(11):2079– 2095, July 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412001073.

Lee:2012:PIN

[584] Yun-Huei Lee and Kung-Jeng Wang. Performance impact of new product development processes for distinct scenarios under different supplier– manufacturer relationships. Mathematics and Computers in Simulation, 82 (11):2096–2108, July 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001097.

Zhou:2012:CMF

[585] Zhaojie Zhou, Fengxin Chen, and Huanzhen Chen. Characteristic mixed finite element approximation of transient convection diffusion optimal control problems. Mathematics and Computers in Simulation, 82(11):2109– 2128, July 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001127.

Wei:2012:RMC

[586] T. Wei and Y. G. Chen. A regularization method for a Cauchy problem of Laplace’s equation in an annular domain. Mathematics and Computers in Simulation, 82(11):2129– 2144, July 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001188.

Kerl:2012:RDE

[587] Hermann-Johannes Kerl, Dirk Langemann, and Antje Vollrath. Reaction diffusion equations and the chronification of liver infections. Mathematics and Computers in Simulation, 82 (11):2145–2156, July 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541200122X.

Pandey:2012:MMF

[588] Anamika Pandey, Axel Klar, and Sudarshan Tiwari. Meshfree method for fluctuating hydrodynamics. Mathematics and Computers in Simulation, 82 (11):2157–2166, July 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412001310.

Liu:2012:PAB

[589] Jun Liu and Yao-Lin Jiang. A parareal algorithm based on waveform relaxation. Mathematics and Computers in Simulation, 82(11):2167– 2181, July 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412001358.

Giron:2012:NED

[590] Edwin Girón, Alejandro C. Frery, and Francisco Cribari-Neto. Nonparametric edge detection in speckled imagery. Mathematics and Computers in Simulation, 82(11):2182– 2198, July 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847541200136X.

Long:2012:AIS

[591] Shujun Long, Xiaohu Wang, and Dingshi Li. Attracting and invariant sets of non-autonomous reaction– diffusion neural networks with timevarying delays. Mathematics and Computers in Simulation, 82(11):2199– 2214, July 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001371.

Jiang:2012:QTS

[592] Qingtang Jiang and Baobin Li. Quad/ triangle subdivision, nonhomogeneous refinement equation and polynomial reproduction. Mathematics and Computers in Simulation, 82(11):2215– 2237, July 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001383.

Lin:2012:SGD

[593] Yiqin Lin, Liang Bao, and Qinghua Wu. Simpler GMRES with deflated restarting. Mathematics and Computers in Simulation, 82(11):2238– 2252, July 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001413.

Sun:2012:MMF

[594] Liang Sun. Mathematical modeling of the flow in a pipeline with a leak. Mathematics and Computers in Simulation, 82(11):2253–2267, July 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001425.

Boukabou:2012:GCC

[595] Abdelkrim Boukabou and Naim Mekircha. Generalized chaos control and synchronization by nonlinear high-order approach. Mathematics and Computers in Simulation, 82(11):2268– 2281, July 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412001553.

Anonymous:2012:NIg

[596] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 82(11):2282, July 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412001917.

Anonymous:2012:ICEg

[597] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 82(11):2283– 2287, July 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412001929.

Anonymous:2012:EBh

[598] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 82(12):CO2, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412002339.

Spitaleri:2012:ASC

[599] Rosa Maria Spitaleri. Applied scientific computing IX: Modeling, numerical methods, algorithms and simulation methodologies. Mathematics and Computers in Simulation, 82 (12):2889–2890, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002224.

Bruni:2012:TSE

[600] V. Bruni, D. De Canditiis, and D. Vitulano. Time-scale energy based analysis of contours of real-world shapes. Mathematics and Computers in Simulation, 82(12):2891–2907, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541000399X.

Egidi:2012:NTS

[601] Nadaniela Egidi and Pierluigi Maponi. Normalization techniques for the SVE of the Green function of Helmholtz operator. Mathematics and Computers in Simulation, 82(12):2908– 2915, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411001753.

Botte:2012:NEP

[602] Vincenzo Botte and Daniela Mansutti. A numerical estimate of the planktoninduced sea surface tension effects in a Langmuir circulation. Mathematics and Computers in Simulation, 82 (12):2916–2928, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001802.

Porta:2012:STA

[603] G. M. Porta, S. Perotto, and F. Ballio. A space–time adaptation scheme for unsteady shallow water problems. Mathematics and Computers in Simulation, 82(12):2929– 2950, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475411001534.

Coffman:2012:OIB

[604] Valerie R. Coffman, Andrew C. E. Reid, Stephen A. Langer, and Gunay Dogan. OOF3D: an imagebased finite element solver for materials science. Mathematics and Computers in Simulation, 82(12):2951– 2961, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412000602.

Egidi:2012:TDM

[605] N. Egidi, P. Maponi, L. Misici, and S. Rubino. A three-dimensional model for the study of the cooling system of submersible electric pumps. Mathematics and Computers in Simulation, 82 (12):2962–2970, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412001267.

Plaza:2012:LRA

[606] Ángel Plaza, Sergio Falcón, José P. Suárez, and Pilar Abad. A local refinement algorithm for the longest-edge trisection of triangle meshes. Mathematics and Computers in Simulation, 82(12):2971–2981, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411001698.

Melot:2012:CNM

[607] Matthieu Melot, Jean-Yves Trépanier, Ricardo Camarero, and Eddy Petro. Comparison of numerical models in radiative heat transfer with application to circuit-breaker simulations. Mathematics and Computers in Simulation, 82(12):2982–2996, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001516.

Pistella:2012:EPM

[608] Francesca Pistella, Rosa Spitaleri, and R. Beauwens. Erratum to “Preface [Math. Comput. 82 (1) (2011) 1–212]. Mathematics and Computers in Simulation, 82(12): 2997, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412002236. See

[396].

Anonymous:2012:NIh

[609] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 82(12):2998, June 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541200239X.

Anonymous:2012:ICEh

[610] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 82(12):2999– 3003, June 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412002406.

Anonymous:2012:PS

[611] Anonymous. Pages 1–86 (September 2012). Mathematics and Computers in Simulation, 83:1–86, September 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2012:EBi

[612] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 83:CO2, September 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412002480.

Huang:2012:TBC

[613] Jian Huang, Masahito Kobayashi, and Michael McAleer. Testing for the Box–Cox parameter for an integrated process. Mathematics and Computers in Simulation, 83:1–9, September 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408001766.

Nath:2012:RDB

[614] H. (Mindi) B. Nath, Jae H. Kim, and Robert D. Brooks. Realized dual-betas for leading Australian stocks: an evaluation of the estimation methods and the effect of the sampling interval. Mathematics and Computers in Simulation, 83: 10–22, September 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001694.

Bahrainian:2012:AUG

[615] Seyed Saied Bahrainian and Zahra Mehrdoost. An automatic unstructured grid generation method for viscous flow simulations. Mathematics and Computers in Simulation, 83: 23–43, September 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001760.

Gupta:2012:ERP

[616] Ramesh C. Gupta, M. E. Ghitany, and D. K. Al-Mutairi. Estimation of reliability in a parallel system with random sample size. Mathematics and Computers in Simulation, 83: 44–55, September 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001772.

Ledermann:2012:FPR

[617] Daniel Ledermann and Carol Alexander. Further properties of random orthogonal matrix simulation. Mathematics and Computers in Simulation, 83:56–79, September 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412001784.

Anonymous:2012:NIi

[618] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 83:80, September 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412002522.

Anonymous:2012:ICEi

[619] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 83:81–85, September 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475412002534.

Anonymous:2012:PO

[620] Anonymous. Pages 1–104 (October 2012). Mathematics and Computers in Simulation, 84:1–104, October 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2012:EBj

[621] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 84:CO2, October 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412002583.

Niu:2012:EQI

[622] Qun Niu, Taijin Zhou, Minrui Fei, and Bing Wang. An efficient quantum immune algorithm to minimize mean flow time for hybrid flow shop problems.

Mathematics and Computers in Simulation, 84:1–25, October 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001826.

Machado:2012:PCB

[623] Raúl Machado. On pressure and corner boundary conditions with two lattice Boltzmann construction approaches. Mathematics and Computers in Simulation, 84:26– 41, October 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001966.

Qian:2012:NAL

[624] Wei Qian, Juan Liu, and Shumin Fei. New augmented Lyapunov functional method for stability of uncertain neutral systems with equivalent delays. Mathematics and Computers in Simulation, 84:42– 50, October 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001991.

Bischi:2012:DMOa

[625] Gian Italo Bischi and Fabio Lamantia. A dynamic model of oligopoly with R&D externalities along networks. Part I. Mathematics and Computers in Simulation, 84:51– 65, October 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412002030.

Bischi:2012:DMOb

[626] Gian Italo Bischi and Fabio Lamantia. A dynamic model of oligopoly with R&D externalities along networks. Part II. Mathematics and Computers in Simulation, 84:66– 82, October 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412002078.

Dai:2012:CDB

[627] Chuanjun Dai, Min Zhao, and Lansun Chen. Complex dynamic behavior of three-species ecological model with impulse perturbations and seasonal disturbances. Mathematics and Computers in Simulation, 84:83– 97, October 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412002108.

Anonymous:2012:NIj

[628] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 84:98, October 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412002625.

Anonymous:2012:ICEj

[629] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 84:99–103, October 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412002637.

Anonymous:2012:PN

[630] Anonymous. Pages 1–102 (November 2012). Mathematics and Computers in Simulation, 85:1–102, November 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2012:EBk

[631] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 85:CO2, November 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002698.

Mobayen:2012:LBC

[632] Saleh Mobayen, Vahid Johari Majd, and Mahdi Sojoodi. An LMI-based composite nonlinear feedback terminal sliding-mode controller design for disturbed MIMO systems. Mathematics and Computers in Simulation, 85:1–10, November 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002133.

Xu:2012:CMO

[633] Jiuping Xu and Zhimiao Tao. A class of multi-objective equilibrium chance maximization model with twofold random phenomenon and its application to hydropower station operation. Mathematics and Computers in Simulation, 85:11–33, November 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002194.

Adamek:2012:AST

[634] V. Adámek and F. Valeš. Analytical solution of transient in-plane vibration of a thin viscoelastic disc and its multi-precision evaluation. Mathematics and Computers in Simulation, 85:34–44, November 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541200225X.

Tian:2012:RMU

[635] WenYi Tian, Can Li, Weihua Deng, and Yujiang Wu. Regularization methods for unknown source in space fractional diffusion equation. Mathematics and Computers in Simulation, 85:45–56, November 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002273.

Jana:2012:GSB

[636] Soovoojeet Jana, Milon Chakraborty, Kunal Chakraborty, and T. K. Kar. Global stability and bifurcation of time delayed prey–predator system incorporating prey refuge. Mathematics and Computers in Simulation, 85:57–77, November 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002285.

Xie:2012:SPM

[637] Qizhi Xie and Qichang Zhang. Stability predictions of milling with variable spindle speed using an improved semi-discretization method. Mathematics and Computers in Simulation, 85:78–89, November 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002303.

Xu:2012:GDS

[638] Rui Xu. Global dynamics of an SEIS epidemiological model with time delay describing a latent period. Mathematics and Computers in Simulation, 85:90–102, November 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002315.

Anonymous:2012:NIk

[639] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 85:I, November 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541200273X.

Anonymous:2012:ICEk

[640] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 85:II–VI, November 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475412002741.

Anonymous:2012:EBl

[641] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 86:CO2, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000190.

Cao:2012:P

[642] Jinde Cao, Jianquan Lu, Jinling Liang, and Qingshan Liu. Preface. Mathematics and Computers in Simulation, 86:1, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000177.

Zhao:2012:GRN

[643] Yan Zhao and Qingshan Liu. Generalized recurrent neural network for ϵ-insensitive support vector regression. Mathematics and Computers in Simulation, 86:2–9, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475412001796.

Zhu:2012:NSE

[644] Song Zhu, Yi Shen, and Guici Chen. Noise suppress exponential growth for hybrid Hopfield neural networks. Mathematics and Computers in Simulation, 86:10–20, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410004015.

Zhu:2012:USV

[645] Ping Zhu, Feng Pan, Wei Chen, and Siliang Zhang. Use of support vector regression in structural optimization: Application to vehicle crashworthiness design. Mathematics and Computers in Simulation, 86:21–31, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000936.

Zhang:2012:IHE

[646] Ying Zhang and Alice M. Agogino. Interactive hybrid evolutionary computation for MEMS design synthesis. Mathematics and Computers in Simulation, 86:32–38, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541100084X.

Song:2012:PSL

[647] Qiang Song, Jinde Cao, and Fang Liu. Pinning synchronization of linearly coupled delayed neural networks. Mathematics and Computers in Simulation, 86:39–51, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475411001741.

Wang:2012:NCN

[648] Rubin Wang, Zhikang Zhang, Chi K. Tse, Jingyi Qu, and Jianting Cao. Neural coding in networks of multipopulations of neural oscillators. Mathematics and Computers in Simulation, 86:52–66, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410003733.

Xu:2012:LBS

[649] Bingji Xu, Yuan Xu, and Linman He. LMI-based stability analysis of impulsive high-order Hopfieldtype neural networks. Mathematics and Computers in Simulation, 86:

67–77, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000796.

Vanstone:2012:CTS

[650] Bruce Vanstone, Gavin Finnie, and Tobias Hahn. Creating trading systems with fundamental variables and neural networks: the Aby case study. Mathematics and Computers in Simulation, 86:78–91, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000292.

Zhang:2012:AFN

[651] Ying Zhang. Application of T-S fuzzy neural network based on declination compensation in soft sensing. Mathematics and Computers in Simulation, 86:92–99, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475411000279.

Adali:2012:SII

[652] Sarp Adali, Ismail Kucuk, Evgeny V. Morozov, and Viktor E. Verijenko. Special ICMOSPS 2007 issue of Mathematics and Computers in Simulation. Mathematics and Computers in Simulation, 86:100, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475413000165.

Abdel-Hafez:2012:GIS

[653] Mamoun F. Abdel-Hafez. On the GPS/ IMU sensors’ noise estimation for enhanced navigation integrity. Mathematics and Computers in Simulation, 86:101–117, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410000765.

Aghdam:2012:EKM

[654] M. M. Aghdam, N. Shahmansouri, and M. Mohammadi. Extended Kantorovich method for static analysis of moderately thick functionally graded sector plates. Mathematics and Computers in Simulation, 86: 118–130, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475410002661.

Anonymous:2012:NIl

[655] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 86:131, December 2012. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000268.

Anonymous:2012:ICEl

[656] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 86:132–136, December 2012. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541300027X.

Anonymous:2013:PJa

[657] Anonymous. Pages 1–126 (January 2013). Mathematics and Computers in Simulation, 87:1–126, January 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2013:EBa

[658] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 87:CO2, January 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000517.

DalForno:2013:BCB

[659] Arianna Dal Forno and Ugo Merlone. Border-collision bifurcations in a model of Braess paradox. Mathematics and Computers in Simulation, 87:1–18, January 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000025.

Franco:2013:BTF

[660] G. C. Franco, V. A. Reisen, and F. A. Alves. Bootstrap tests for fractional integration and cointegration: a comparison study. Mathematics and Computers in Simulation, 87:19–29, January 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000098.

Sen:2013:MBC

[661] M. Sen, M. Banerjee, and E. Venturino. A model for biological control in agriculture. Mathematics and Computers in Simulation, 87:30– 44, January 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541300013X.

Lamboni:2013:DBG

[662] M. Lamboni, B. Iooss, A.-L. Popelin, and F. Gamboa. Derivative-based global sensitivity measures: General links with Sobol’ indices and numerical tests. Mathematics and Computers in Simulation, 87:45– 54, January 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413000141.

Fercec:2013:CCC

[663] Brigita Ferčec and Adam Mahdi. Center conditions and cyclicity for a family of cubic systems: Computer algebra approach. Mathematics and Computers in Simulation, 87:55–67, January 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000153.

Lamnii:2013:UTA

[664] A. Lamnii, H. Mraoui, D. Sbibih, and A. Zidna. Uniform tension algebraic trigonometric spline wavelets of class C ^{2} and order four. Mathematics and Computers in Simulation, 87:68–86, January 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002856.

Dolejsi:2013:HDN

[665] Vít Dolejší. hp-DGFEM for nonlinear convection–diffusion problems. Mathematics and Computers in Simulation, 87:87–118, January 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000402.

Anonymous:2013:NIa

[666] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 87:119, January 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000554.

Anonymous:2013:ICEa

[667] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 87:120–125, January 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413000566.

Anonymous:2013:PF

[668] Anonymous. Pages 1–46 (February 2013). Mathematics and Computers in Simulation, 88:1–46, February 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2013:EBb

[669] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 88:CO2, February 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541300116X.

Sansottera:2013:SSE

[670] M. Sansottera, U. Locatelli, and A. Giorgilli. On the stability of the secular evolution of the planar Sun– Jupiter–Saturn–Uranus system. Mathematics and Computers in Simulation, 88:1–14, February 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541100098X.

Gogin:2013:CMR

[671] N. Gogin and A. Mylläri. Conway matrices related to a non-transitive head-or-tail game with a q-sided die and their Hamming weight-spectra via DFT and the MacWilliams duality formula. Mathematics and Computers in Simulation, 88:15– 30, February 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412000614.

Wang:2013:ACD

[672] Xiaoping Wang, Yunliang Zhao, Peng Sun, and Xiaobin Wang. An analysis on convergence of data-driven approach to ship lock scheduling. Mathematics and Computers in Simulation, 88:31–38, February 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000463.

Anonymous:2013:PM

[673] Anonymous. Pages 1–90 (March 2013). Mathematics and Computers in Simulation, 89:1–90, March 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2013:EBc

[674] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 89:CO2, March 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000918.

Xiao:2013:AEF

[675] Min Xiao, Wei Xing Zheng, and Jinde Cao. Approximate expressions of a fractional order van der Pol oscillator by the residue harmonic balance method. Mathematics and Computers in Simulation, 89: 1–12, March 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413000384.

Wang:2013:GSA

[676] Tianlei Wang, Zhixing Hu, Fucheng Liao, and Wanbiao Ma. Global stability analysis for delayed virus infection model with general incidence rate and humoral immunity. Mathematics and Computers in Simulation, 89:13–22, March 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000451.

Jie:2013:ANS

[677] Yu xin Jie, Li zhen Liu, Wen jie Xu, and Guang xin Li. Application of NEM in seepage analysis with a free surface. Mathematics and Computers in Simulation, 89:23–37, March 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000475.

Rabta:2013:HMP

[678] Boualem Rabta. A hybrid method for performance analysis of G/G/m queueing networks. Mathematics and Computers in Simulation, 89: 38–49, March 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413000438.

Luo:2013:ROF

[679] Zhendong Luo, Hong Li, Ping Sun, Jing An, and Ionel Michael Navon. A reduced-order finite volume element formulation based on POD method and numerical simulation for twodimensional solute transport problems. Mathematics and Computers in Simulation, 89:50–68, March 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541300044X.

Mahmoud:2013:MPP

[680] Emad E. Mahmoud. Modified projective phase synchronization of chaotic complex nonlinear systems. Mathematics and Computers in Simulation, 89:69–85, March 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000426.

Anonymous:2013:NIb

[681] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 89:86, March 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000955.

Anonymous:2013:ICEb

[682] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 89:87–90, March 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475413000967.

Anonymous:2013:EBd

[683] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 90:CO2, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541300102X.

Anonymous:2013:Pa

[684] Anonymous. Preface. Mathematics and Computers in Simulation, 90:vii, April 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413001419.

Baudart:2013:OCW

[685] F. Baudart, E. Matagne, B. Dehez, and F. Labrique. Optimal current waveforms for torque control of permanent magnet synchronous machines with any number of phases in open circuit. Mathematics and Computers in Simulation, 90:1–14, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001176.

Bodo:2013:PMS

[686] N. Bodo, M. Jones, and E. Levi. A PWM method for seven- and ninephase open-end winding motor drives. Mathematics and Computers in Simulation, 90:15–27, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000894.

Chou:2013:OFD

[687] Hsin-Hung Chou, Ying-Shieh Kung, Nguyen Vu Quynh, and Stone Cheng. Optimized FPGA design, verification and implementation of a neuro-fuzzy controller for PMSM drives. Mathematics and Computers in Simulation, 90:28–44, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001759.

Meghnous:2013:DIS

[688] A. R. Meghnous, M. T. Pham, and X. Lin-Shi. Dynamic identification of a synchronous machine using an extended sliding mode observer. Mathematics and Computers in Simulation, 90:45–59, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001528.

Mbayed:2013:HES

[689] Rita Mbayed, Georges Salloum, Lionel Vido, Eric Monmasson, and Mohamed Gabsi. Hybrid excitation synchronous generator in embedded applications: Modeling and control. Mathematics and Computers in Simulation, 90:60–73, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001851.

Jones:2013:SML

[690] Martin Jones and I. Nyoman Wahyu Satiawan. A simple multi-level space vector modulation algorithm for fivephase open-end winding drives. Mathematics and Computers in Simulation, 90:74–85, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001152.

Sediki:2013:AAI

[691] Hamid Sediki, Ali Bechouche, Djaffar Ould Abdeslam, and Salah Haddad. ADALINE approach for induction motor mechanical parameters identification. Mathematics and Computers in Simulation, 90: 86–97, April 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001085.

Bouzid:2013:ESM

[692] M. Bouzid and G. Champenois. An efficient, simplified multiple-coupled circuit model of the induction motor aimed to simulate different types of stator faults. Mathematics and Computers in Simulation, 90:98– 115, April 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413000682.

Frechard:2013:MFP

[693] J. Frechard, D. Knittel, P. Dessagne, J. S. Pellé, G. Gaudiot, J. C. Caspar, and G. Heitz. Modelling and fast position control of a new unwinding– winding mechanism design. Mathematics and Computers in Simulation, 90:116–131, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541200208X.

Stumpf:2013:AIS

[694] Péter Stumpf, Rafael K. Járdán, and István Nagy. Analysis of the impact of space vector modulation techniques on the operation of ultrahigh speed induction machines. Mathematics and Computers in Simulation, 90:132–144, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002091.

Bahri:2013:OHS

[695] I. Bahri, L. Idkhajine, E. Monmasson, and M. E. A. Benkhelifa. Optimal hardware/software partitioning of a system on chip FPGA-based sensorless AC drive current controller. Mathematics and Computers in Simulation, 90:145–161, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412001437.

Tiegna:2013:OAM

[696] Huguette Tiegna, Yacine Amara, and Georges Barakat. Overview of analytical models of permanent magnet electrical machines for analysis and design purposes. Mathematics and Computers in Simulation, 90:162–177, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000049.

Nguyen:2013:BCO

[697] Phi Hung Nguyen, Emmanuel Hoang, and Mohamed Gabsi. Bi-criteria optimization design of an interior permanent magnet synchronous machine for a hybrid electric vehicle application. Mathematics and Computers in Simulation, 90:178– 191, April 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412002054.

Brudny:2013:IMM

[698] Jean-François Brudny, Cristian Demian, Lucian Petrea, and Thierry Belgrand. Induction machine magnetic noise: Impact of a new stator magnetic circuit design. Mathematics and Computers in Simulation, 90:192–204, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000104.

Baudart:2013:APC

[699] F. Baudart, E. Matagne, B. Dehez, and F. Labrique. Analytical prediction of cogging torque in surface mounted permanent magnet motors. Mathematics and Computers in Simulation, 90:205–217, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000633.

Nerg:2013:LPB

[700] Janne Nerg and Vesa Ruuskanen. Lumped-parameter-based thermal analysis of a doubly radial forced-air-cooled direct-driven permanent magnet wind generator. Mathematics and Computers in Simulation, 90: 218–229, April 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412002157.

Male:2013:ACF

[701] Gael Malé, Thierry Lubin, Smail Mezani, and Jean Lévêque. Analytical calculation of the flux density distribution in a superconducting reluctance machine with HTS bulks rotor. Mathematics and Computers in Simulation, 90:230–243, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000086.

Denies:2013:ODE

[702] J. Denies, H. Ben Ahmed, and B. Dehez. Optimal design of electromagnetic devices: Development of an efficient optimization tool based on smart mutation operations implemented in a genetic algorithm. Mathematics and Computers in Simulation, 90:244–255, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000074.

Delinchant:2013:MMO

[703] B. Delinchant, D. Lahaye, F. Wurtz, and J.-L. Coulomb. Manifold mapping optimization with or without true gradients. Mathematics and Computers in Simulation, 90:256– 265, April 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847541200211X.

Rezgui:2013:UMT

[704] A. Rezgui, L. Gerbaud, and B. Delinchant. Unified modeling technique using VHDL-AMS and software components. Mathematics and Computers in Simulation, 90:266– 276, April 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412002820.

Anonymous:2013:NIc

[705] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 90:277, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413001080.

Anonymous:2013:ICEc

[706] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 90:278–281, April 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475413001092.

Anonymous:2013:EBe

[707] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 91:CO2, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413001286.

Anonymous:2013:Pb

[708] Anonymous. Preface. Mathematics and Computers in Simulation, 91:viii, May 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413001420.

Femia:2013:OCP

[709] N. Femia, G. Petrone, G. Spagnuolo, and M. Vitelli. Optimal control of photovoltaic arrays. Mathematics and Computers in Simulation, 91:1–15, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001061.

Tina:2013:IST

[710] G. M. Tina, F. Arcidiacono, and A. Gagliano. Intelligent sun-tracking system based on multiple photodiode sensors for maximisation of photovoltaic energy production. Mathematics and Computers in Simulation, 91:16–28, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002066.

Kot:2013:CMP

[711] R. Kot, M. Rolak, and M. Malinowski. Comparison of maximum peak power tracking algorithms for a small wind turbine. Mathematics and Computers in Simulation, 91: 29–40, May 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413000657.

Rizo:2013:UWT

[712] Mario Rizo, Marco Liserre, Emilio Bueno, Francisco J. Rodríguez, and Francisco Huerta. Universal wind turbine working in grid-connected and island operating modes. Mathematics and Computers in Simulation, 91:41–51, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541200167X.

Robyns:2013:MSH

[713] Benoît Robyns, Arnaud Davigny, and Christophe Saudemont. Methodologies for supervision of Hybrid Energy Sources based on Storage Systems – a survey. Mathematics and Computers in Simulation, 91: 52–71, May 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001504.

Mariani:2013:DMV

[714] R. Rigo Mariani, F. Lacressonniere, G. Fontes, and X. Roboam. Design of a medium voltage power converter-storage devices embedded in a hybrid emergency network for more electrical aircraft. Mathematics and Computers in Simulation, 91: 72–90, May 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001723.

Breban:2013:EDG

[715] Stefan Breban, Christophe Saudemont, Sébastien Vieillard, and Benoît Robyns. Experimental design and genetic algorithm optimization of a fuzzy-logic supervisor for embedded electrical power systems. Mathematics and Computers in Simulation, 91:91–107, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001322.

Florez:2013:DCM

[716] J. Zárate Flórez, J. Martinez, G. Besançon, and D. Faille. Decentralizedcoordinated model predictive control for a hydro-power valley. Mathematics and Computers in Simulation, 91:108–118, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000900.

Wang:2013:PFP

[717] B. C. Wang, M. Sechilariu, and F. Locment. Power flow Petri Net modelling for building integrated multisource power system with smart grid interaction. Mathematics and Computers in Simulation, 91:119– 133, May 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413000128.

Tisan:2013:SNN

[718] A. Tisan and M. Cirstea. SOM neural network design – a new Simulink library based approach targeting FPGA implementation. Mathematics and Computers in Simulation, 91:134–149, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412001139.

Vyncke:2013:SBW

[719] Thomas J. Vyncke, Steven Thielemans, and Jan A. A. Melkebeek. Simulation-based weight factor selection and FPGA prediction core implementation for finite-set model based predictive control of power electronics. Mathematics and Computers in Simulation, 91:150–166, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000037.

Bachir:2013:FAR

[720] T. Ould Bachir, C. Dufour, J. Bélanger, J. Mahseredjian, and J. P. David. A fully automated reconfigurable calculation engine dedicated to the realtime simulation of high switching frequency power electronic circuits. Mathematics and Computers in Simulation, 91:167–177, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002121.

Martin:2013:FBA

[721] P. Martín, E. Bueno, Fco. J. Rodríguez, O. Machado, and B. Vuksanovic. An FPGA-based approach to the automatic generation of VHDL code for industrial control systems applications: a case study of MSOGIs implementation. Mathematics and Computers in Simulation, 91:178–192, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001541.

ElFadil:2013:ASM

[722] H. El Fadil, F. Giri, and Josep M. Guerrero. Adaptive sliding mode control of interleaved parallel boost converter for fuel cell energy generation system. Mathematics and Computers in Simulation, 91:193–210, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412001747.

Savaghebi:2013:SCV

[723] Mehdi Savaghebi, Juan C. Vasquez, Alireza Jalilian, Josep M. Guerrero, and Tzung-Lin Lee. Selective compensation of voltage harmonics in grid-connected microgrids. Mathematics and Computers in Simulation, 91:211–228, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412001279.

Rahmani:2013:DBI

[724] S. Rahmani, Ab. Hamadi, K. Al-Haddad, and A. I. Alolah. A DSPbased implementation of an instantaneous current control for a threephase shunt hybrid power filter. Mathematics and Computers in Simulation, 91:229–248, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412002248.

Naouar:2013:FBS

[725] M. W. Naouar, B. Ben Hania, I. Slama-Belkhodja, E. Monmasson, and A. A. Naassani. FPGA-based sliding mode direct control of single phase PWM boost rectifier. Mathematics and Computers in Simulation, 91:249–261, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412000924.

Gavagsaz-Ghoachani:2013:POB

[726] R. Gavagsaz-Ghoachani, M. Phattanasak, J.-P. Martin, S. Pierfederici, and B. Davat. Predicting the onset of bifurcation and stability study of a hybrid current controller for a boost converter. Mathematics and Computers in Simulation, 91:262–273, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000645.

Kanaan:2013:MML

[727] H. Y. Kanaan and K. Al-Haddad. Modeling and multi-loop feedback control design of a SEPIC power factor corrector in single-phase rectifiers. Mathematics and Computers in Simulation, 91:274–283, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001395.

Anonymous:2013:NId

[728] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 91:284, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413001341.

Anonymous:2013:ICEd

[729] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 91:285–288, May 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475413001353.

Anonymous:2013:PJb

[730] Anonymous. Pages 1–108 (June 2013). Mathematics and Computers in Simulation, 92:1–108, June 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2013:NIe

[731] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 92:CO2, June 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413001626.

He:2013:FCM

[732] Shuping He and Fei Liu. L _{2}– L _{∞} fuzzy control for Markov jump systems with neutral time-delays. Mathematics and Computers in Simulation, 92:1–13, June 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000499.

Zhang:2013:ORM

[733] Z. Q. Zhang and T. Wei. An optimal regularization method for spacefractional backward diffusion problem. Mathematics and Computers in Simulation, 92:14–27, June 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000712.

deAmo:2013:PSS

[734] Enrique de Amo, Manuel Díaz Carrillo, and Juan Fernández Sánchez. PCF selfsimilar sets and fractal interpolation.

Mathematics and Computers in Simulation, 92:28–39, June 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000852.

Yi:2013:NTR

[735] Chenfu Yi, Yunong Zhang, and Dongsheng Guo. A new type of recurrent neural networks for real-time solution of Lyapunov equation with timevarying coefficient matrices. Mathematics and Computers in Simulation, 92:40–52, June 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413001122.

Liu:2013:DSL

[736] Meng Liu and Ke Wang. Dynamics and simulations of a logistic model with impulsive perturbations in a random environment. Mathematics and Computers in Simulation, 92:53–75, June 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000773.

Mahmoudi:2013:EWP

[737] Eisa Mahmoudi and Afsaneh Sepahdar. Exponentiated Weibull– Poisson distribution: Model, properties and applications. Mathematics and Computers in Simulation, 92: 76–97, June 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413001134.

Anonymous:2013:ICEe

[738] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 92:98, June 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413001663.

Anonymous:2013:EBf

[739] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 92:99–102, June 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413001675.

Anonymous:2013:EBg

[740] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 93:CO2, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413001729.

Anonymous:2013:MSO

[741] Anonymous. Modelling and simulation: an overview. Mathematics and Computers in Simulation, 93: viii–xv, July 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413001833.

Franses:2013:IJA

[742] Philip Hans Franses. Improving judgmental adjustment of modelbased forecasts. Mathematics and Computers in Simulation, 93:1– 8, July 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412002868.

Chang:2013:FUP

[743] Chia-Lin Chang, Philip Hans Franses, and Michael McAleer. Are forecast updates progressive? Mathematics and Computers in Simulation, 93:9–18, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000487.

Wong:2013:CMV

[744] C. S. Wong. On a constrained mixture vector autoregressive model. Mathematics and Computers in Simulation, 93:19–28, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000761.

Rea:2013:AEB

[745] William Rea, Les Oxley, Marco Reale, and Jennifer Brown. Not all estimators are born equal: the empirical properties of some estimators of long memory. Mathematics and Computers in Simulation, 93:29–42, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002029.

Withanawasam:2013:CTM

[746] R. M. Withanawasam, P. A. Whigham, and T. F. Crack. Characterising trader manipulation in a limitorder driven market. Mathematics and Computers in Simulation, 93: 43–52, July 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412002212.

Suenaga:2013:MBT

[747] Hiroaki Suenaga. Measuring bias in a term-structure model of commodity prices through the comparison of simultaneous and sequential estimation. Mathematics and Computers in Simulation, 93:53–66, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000748.

Allen:2013:MTC

[748] D. E. Allen, A. R. Kramadibrata, R. J. Powell, and A. K. Singh. Modelling tail credit risk using transition matrices. Mathematics and Computers in Simulation, 93:67–75, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002200.

Nawata:2013:EDB

[749] K. Nawata and K. Kawabuchi. Evaluation of the DPC-based inclusive payment system in Japan for cataract operations by a new model. Mathematics and Computers in Simulation, 93:76–85, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/

article/pii/S0378475412002005. See

corrigendum [1003].

McKenzie:2013:MLU

[750] C. R. McKenzie and Sumiko Takaoka. The matching of lead underwriters and issuing firms in the Japanese corporate bond market. Mathematics and Computers in Simulation, 93: 86–97, July 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413000736.

Li:2013:SLT

[751] J. S. H. Li, A. C. Y. Ng, and W. S. Chan. Stochastic life table forecasting: a time-simultaneous fan chart application. Mathematics and Computers in Simulation, 93: 98–107, July 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412002042.

Brown:2013:ASD

[752] Jennifer A. Brown, Mohammad Salehi M., Mohammad Moradi, Bardia Panahbehagh, and David R. Smith. Adaptive survey designs for sampling rare and clustered populations. Mathematics and Computers in Simulation, 93:108–116, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002170.

Fukiharu:2013:IDI

[753] T. Fukiharu. Income distribution inequality, globalization, and innovation: a general equilibrium simulation. Mathematics and Computers in Simulation, 93:117–127, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001954.

Saha:2013:DER

[754] Shrabani Saha and Zhaoyong Zhang. Do exchange rates affect consumer prices? A comparative analysis for Australia, China and India. Mathematics and Computers in Simulation, 93:128–138, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002819.

Yap:2013:IER

[755] Ghialy Yap. The impacts of exchange rates on Australia’s domestic and outbound travel markets. Mathematics and Computers in Simulation, 93:139–150, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002182.

Hong:2013:CDM

[756] Jin Hong, Xiumei Guo, Dora Marinova, Fengli Yang, and Wentao Yu. Clean development mechanism in China: Regional distribution and prospects. Mathematics and Computers in Simulation, 93:151–163, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000116.

Iwanaga:2013:DIW

[757] Takuya Iwanaga, Sondoss El Sawah, and Anthony Jakeman. Design and implementation of a web-based groundwater data management system. Mathematics and Computers in Simulation, 93:164–174, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002881.

Chan:2013:ISC

[758] Felix Chan, Laurent L. Pauwels, and Johnathan Wongsosaputro. The impact of serial correlation on testing for structural change in binary choice model: Monte Carlo evidence. Mathematics and Computers in Simulation, 93:175–189, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002790.

Chang:2013:CJS

[759] Chia-Lin Chang, Michael McAleer, and Les Oxley. Coercive journal self citations, impact factor, Journal Influence and Article Influence. Mathematics and Computers in Simulation, 93:190–197, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000694.

Anonymous:2013:NIf

[760] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 93:198, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413001791.

Anonymous:2013:ICEf

[761] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 93:199–202, July 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475413001808.

Anonymous:2013:PAa

[762] Anonymous. Pages 1–332 (August 2013). Mathematics and Computers in Simulation, 94:1–332, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2013:EBh

[763] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 94:CO2, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541300236X.

Kim:2013:DCP

[764] Hye Kyung Kim and Hunki Baek. The dynamical complexity of a predator– prey system with Hassell–Varley functional response and impulsive effect. Mathematics and Computers in Simulation, 94:1–14, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413001559.

Anonymous:2013:PN

[765] Anonymous. Publisher’s note. Mathematics and Computers in Simulation, 94:15–32, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541300147X.

Zu:2013:GQA

[766] Jian Zu. Global qualitative analysis of a predator–prey system with Allee effect on the prey species. Mathematics and Computers in Simulation, 94:33–54, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413001523.

Fernandez:2013:SDS

[767] J. L. Fernández, A. M. Ferreiro, J. A. García-Rodríguez, A. Leitao, J. G. López-Salas, and C. Vázquez. Static and dynamic SABR stochastic volatility models: Calibration and option pricing using GPUs. Mathematics and Computers in Simulation, 94:55–75, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413001389.

You:2013:SSE

[768] Xiong You and Bingzhen Chen. Symmetric and symplectic exponentially fitted Runge–Kutta-Nyström methods for Hamiltonian problems. Mathematics and Computers in Simulation, 94:76–95, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413001535.

Kung:2013:ESP

[769] James J. Kung and E-Ching Wu. An evaluation of some popular investment strategies under stochastic interest rates. Mathematics and Computers in Simulation, 94:96– 108, August 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413001602.

Laurini:2013:IIF

[770] Márcio Poletti Laurini and Luiz Koodi Hotta. Indirect Inference in fractional short-term interest rate diffusions. Mathematics and Computers in Simulation, 94:109–126, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475413001870.

Nahvi:2013:NAS

[771] S. A. Nahvi, M. Nabi, and S. Janardhanan. Nonlinearity-aware submodel combination in trajectory based methods for nonlinear Mor. Mathematics and Computers in Simulation, 94:127–144, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413001936.

ElKhaoulani:2013:ENI

[772] R. El Khaoulani and P.-O. Bouchard. Efficient numerical integration of an elastic–plastic damage law within a mixed velocity–pressure formulation. Mathematics and Computers in Simulation, 94:145–158, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475413001882.

Chang:2013:RMM

[773] Chia-Lin Chang, David E. Allen, Michael McAleer, and Teodosio Perez Amaral. Risk modelling and management: an overview. Mathematics and Computers in Simulation, 94:159–163, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413001948.

Chang:2013:CHS

[774] Chia-Lin Chang, Lydia González-Serrano, and Juan-Angel Jimenez-Martin. Currency hedging strategies using dynamic multivariate GARCH. Mathematics and Computers in Simulation, 94:164–182, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475412000638.

Casarin:2013:RMR

[775] Roberto Casarin, Chia-Lin Chang, Juan-Angel Jimenez-Martin, Michael McAleer, and Teodosio Pérez-Amaral. Risk management of risk under the Basel Accord: a Bayesian approach to forecasting Value-at-Risk of VIX futures. Mathematics and Computers in Simulation, 94:183– 204, August 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412001498.

Aielli:2013:FCG

[776] Gian Piero Aielli and Massimiliano Caporin. Fast clustering of GARCH processes via Gaussian mixture models. Mathematics and Computers in Simulation, 94:205– 222, August 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412002261.

Jimenez-Martin:2013:GRR

[777] Juan-Angel Jimenez-Martin, Michael McAleer, Teodosio Pérez-Amaral, and Paulo Araújo Santos. GFCrobust risk management under the Basel Accord using extreme value methodologies. Mathematics and Computers in Simulation, 94:223– 237, August 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413002152.

Allen:2013:VSC

[778] David E. Allen, Ron Amram, and Michael McAleer. Volatility spillovers from the Chinese stock market to economic neighbours. Mathematics and Computers in Simulation, 94: 238–257, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000062.

Abad:2013:DCV

[779] Pilar Abad and Sonia Benito. A detailed comparison of value at risk estimates. Mathematics and Computers in Simulation, 94:258– 276, August 2013. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001218.

Hammoudeh:2013:DBC

[780] Shawkat Hammoudeh, Ramazan Sari, Mehmet Uzunkaya, and Tengdong Liu. The dynamics of BRICS’s country risk ratings and domestic stock markets, U.S. stock market and oil price. Mathematics and Computers in Simulation, 94:277–294, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412000122.

Santos:2013:FVR

[781] P. Araújo Santos and M. I. Fraga Alves. Forecasting Value-at-Risk with a duration-based POT method. Mathematics and Computers in Simulation, 94:295–309, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001838.

Singh:2013:EMR

[782] Abhay K. Singh, David E. Allen, and Powell J. Robert. Extreme market risk and extreme value theory. Mathematics and Computers in Simulation, 94:310–328, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412001206.

Anonymous:2013:NIg

[783] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 94:329, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413002437.

Anonymous:2013:ICEg

[784] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 94:330–332, August 2013. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475413002449.

Anonymous:2014:EBa

[785] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 95:CO2, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541300253X.

Brogliato:2014:DDS

[786] Bernard Brogliato, Luciano Lopez, Petri T. Piiroinen, and Tassilo Küpper. Discontinuous dynamical systems: Theory and numerical methods. Mathematics and Computers in Simulation, 95:1, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541300195X.

Guglielmi:2014:NAS

[787] Nicola Guglielmi and Ernst Hairer. Numerical approaches for statedependent neutral delay equations with discontinuities. Mathematics and Computers in Simulation, 95:2– 12, January 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475411002667.

Artes:2014:PLD

[788] Joan C. Artés, Jaume Llibre, Joao C. Medrado, and Marco A. Teixeira. Piecewise linear differential systems with two real saddles. Mathematics and Computers in Simulation, 95:13–22, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000396.

Avrutin:2014:BIS

[789] Viktor Avrutin, Bernd Eckstein, Michael Schanz, and Björn Schenke. Bandcount incrementing scenario revisited and floating regions within robust chaos. Mathematics and Computers in Simulation, 95:23– 38, January 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413001547.

Akhmet:2014:BDL

[790] Marat Akhmet and Mehmet Turan. Bifurcation of discontinuous limit cycles of the van der Pol equation. Mathematics and Computers in Simulation, 95:39–54, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000840.

Licsko:2014:CBS

[791] Gábor Licskó and Gábor Csernák. On the chaotic behaviour of a simple dry-friction oscillator. Mathematics and Computers in Simulation, 95:55–62, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000414.

Riedinger:2014:NFO

[792] Pierre Riedinger and Irinel-Constantin Morărescu. A numerical framework for optimal control of switched input affine nonlinear systems subject to path constraint. Mathematics and Computers in Simulation, 95:63–77, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000050.

True:2014:NCA

[793] H. True, A. P. Engsig-Karup, and D. Bigoni. On the numerical and computational aspects of nonsmoothnesses that occur in railway vehicle dynamics. Mathematics and Computers in Simulation, 95:78–97, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412002297.

Mason:2014:NAC

[794] J. F. Mason, N. Humphries, and P. T. Piiroinen. Numerical analysis of codimension-one, -two and - three bifurcations in a periodicallyforced impact oscillator with two discontinuity surfaces. Mathematics and Computers in Simulation, 95:98–110, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002169.

Pisano:2014:DDS

[795] Alessandro Pisano, Milan R. Rapaić, and Elio Usai. Discontinuous dynamical systems for fault detection. A unified approach including fractional and integer order dynamics. Mathematics and Computers in Simulation, 95:111–125, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002145.

Avrutin:2014:CCA

[796] Viktor Avrutin, Irina Sushko, and Laura Gardini. Cyclicity of chaotic attractors in one-dimensional discontinuous maps. Mathematics and Computers in Simulation, 95:126– 136, January 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541200198X.

Gardini:2014:BAI

[797] Laura Gardini, Fabio Tramontana, and Soumitro Banerjee. Bifurcation analysis of an inductorless chaos generator using 1D piecewise smooth map. Mathematics and Computers in Simulation, 95:137–145, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001346.

Sgura:2014:FDA

[798] Ivonne Sgura. A finite difference approach for the numerical solution of non-smooth problems for Boundary Value ODEs. Mathematics and Computers in Simulation, 95:146– 162, January 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001814.

Kryzhevich:2014:DNN

[799] Sergey Kryzhevich. Dynamics near nonhyperbolic fixed points or nontransverse homoclinic points. Mathematics and Computers in Simulation, 95:163–179, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541200170X.

Schindler:2014:TSN

[800] Thorsten Schindler and Vincent Acary. Timestepping schemes for nonsmooth dynamics based on discontinuous Galerkin methods: Definition and outlook. Mathematics and Computers in Simulation, 95: 180–199, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001231.

Oza:2014:FTS

[801] Harshal B. Oza, Yury V. Orlov, and Sarah K. Spurgeon. Finite time stabilization of a perturbed double integrator with unilateral constraints. Mathematics and Computers in Simulation, 95:200–212, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001164.

Garrappa:2014:SGI

[802] Roberto Garrappa. On some generalizations of the implicit Euler method for discontinuous fractional differential equations. Mathematics and Computers in Simulation, 95: 213–228, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001115.

Berardi:2014:RTM

[803] Marco Berardi. Rosenbrock-type methods applied to discontinuous differential systems. Mathematics and Computers in Simulation, 95:229– 243, January 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413001146.

Anonymous:2014:NIa

[804] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 95:244, January 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002607.

Anonymous:2014:ICEa

[805] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 95:245–247, January 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413002619.

Anonymous:2014:EBb

[806] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 96:CO2, February 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413002668.

Makroglou:2014:PD

[807] Athena Makroglou, Gennady Bocharov, Alistair Fitt, George Flessas, Yang Kuang, and Antonios Tsokaros. Preface — DIEBM 2010. Mathematics and Computers in Simulation, 96: 1–3, February 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413002140.

Baker:2014:OEM

[808] Christopher T. H. Baker. Observations on evolutionary models with (or without) time lag, and on problematical paradigms. Mathematics and Computers in Simulation, 96:4– 53, February 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847541200184X.

Luzyanina:2014:SMI

[809] T. Luzyanina and G. Bocharov. Stochastic modeling of the impact of random forcing on persistent hepatitis B virus infection. Mathematics and Computers in Simulation, 96:54–65, February 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475411002369.

Billy:2014:SCP

[810] Frédérique Billy, Jean Clairambaultt, Olivier Fercoq, Stéphane Gaubertt, Thomas Lepoutre, Thomas Ouillon, and Shoko Saito. Synchronisation and control of proliferation in cycling cell population models with age structure. Mathematics and Computers in Simulation, 96:66–94, February 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541200064X.

Jackiewicz:2014:NST

[811] Z. Jackiewicz, H. Liu, B. Li, and Y. Kuang. Numerical simulations of traveling wave solutions in a drift paradox inspired diffusive delay population model. Mathematics and Computers in Simulation, 96: 95–103, February 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001334.

Maharaj:2014:ASS

[812] A. Maharaj and P. G. L. Leach. Application of symmetry and singularity analyses to mathematical models of biological systems. Mathematics and Computers in Simulation, 96: 104–123, February 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413001894.

Lenz:2014:NCD

[813] Simon M. Lenz, Johannes P. Schlöder, and H. Georg Bock. Numerical computation of derivatives in systems of delay differential equations. Mathematics and Computers in Simulation, 96:124–156, February 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413001973.

Chen:2014:DAH

[814] Xiao Chen, Lequan Min, Yu Zheng, Yang Kuang, and Yongan Ye. Dynamics of acute hepatitis B virus infection in chimpanzees. Mathematics and Computers in Simulation, 96: 157–170, February 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000864.

Arnold:2014:SAM

[815] Anne Arnold and Zoran Nikoloski. In search for an accurate model of the photosynthetic carbon metabolism. Mathematics and Computers in Simulation, 96:171–194, February 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475412001280.

Safan:2014:MAS

[816] Muntaser Safan and Fathalla A. Rihan. Mathematical analysis of an SIS model with imperfect vaccination and backward bifurcation. Mathematics and Computers in Simulation, 96:195–206, February 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541100173X.

Govaerts:2014:CSB

[817] Willy Govaerts and Charlotte Sonck. Computational study of a budding yeast model of Tyson and Novák. Mathematics and Computers in Simulation, 96:207–223, February 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413000803.

Anonymous:2014:NIb

[818] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 96:224, February 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002735.

Anonymous:2014:ICEb

[819] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 96:225–227, February 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413002747.

Anonymous:2014:PM

[820] Anonymous. Pages 1–290 (March 2014). Mathematics and Computers in Simulation, 97:1–290, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2014:EBc

[821] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 97:CO2, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413002917.

Barakat:2014:PIF

[822] H. M. Barakat, Magdy E. El-Adll, and Amany E. Aly. Prediction intervals of future observations for a sample of random size from any continuous distribution. Mathematics and Computers in Simulation, 97: 1–13, March 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413002048.

Chen:2014:ESV

[823] Haibo Chen, Anzhou Cao, Jicai Zhang, Chunbao Miao, and Xianqing Lv. Estimation of spatially varying open boundary conditions for a numerical internal tidal model with adjoint method. Mathematics and Computers in Simulation, 97:14–38, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413002000.

Chang:2014:DRA

[824] Chih-Kai Chang. A dimensionreduction algorithm for the valuation of surrender options in EIA contracts with stochastic interest rates. Mathematics and Computers in Simulation, 97:39–52, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541300205X.

Yu:2014:SDA

[825] Hengguo Yu, Min Zhao, and Ravi P. Agarwal. Stability and dynamics analysis of time delayed eutrophication ecological model based upon the Zeya reservoir. Mathematics and Computers in Simulation, 97: 53–67, March 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413002061.

Jie:2014:MGF

[826] Yu xin Jie, Xu dong Fu, and Yan Liu. Mesh generation for FEM based on centroidal Voronoi tessellations. Mathematics and Computers in Simulation, 97:68–79, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002097.

Hu:2014:SBA

[827] Zengyun Hu, Zhidong Teng, and Long Zhang. Stability and bifurcation analysis in a discrete SIR epidemic model. Mathematics and Computers in Simulation, 97:80–93, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002103.

Rani:2014:CMB

[828] Monica Rani, Harish Garg, and S. P. Sharma. Cost minimization of butteroil processing plant using artificial bee colony technique. Mathematics and Computers in Simulation, 97:94–107, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002115.

Frank:2014:MMN

[829] T. D. Frank, T. D. Gifford, and S. Chiangga. Minimalistic model for navigation of mobile robots around obstacles based on complex-number calculus and inspired by human navigation behavior. Mathematics and Computers in Simulation, 97:108– 122, March 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413002188.

Pal:2014:BAM

[830] Pallav Jyoti Pal and Prashanta Kumar Mandal. Bifurcation analysis of a modified Leslie–Gower predator–prey model with Beddington–DeAngelis functional response and strong Allee effect. Mathematics and Computers in Simulation, 97:123–146, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002073.

Wang:2014:NSO

[831] Wansheng Wang. Nonlinear stability of one-leg methods for neutral Volterra delay-integro-differential equations. Mathematics and Computers in Simulation, 97:147–161, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475413001985.

Tan:2014:MDS

[832] Yonghong Tan and Liang Deng. Modeling the dynamic sandwich system with hysteresis using NAR-MAX model. Mathematics and Computers in Simulation, 97:162– 188, March 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413002206.

Liang:2014:CLS

[833] TaChen Liang and Ming-Chung Yang. Comments on “The life-span prediction of a system connected in series”. Mathematics and Computers in Simulation, 97:189–191, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002231. See

[2].

Hu:2014:IDS

[834] Guixin Hu. Invariant distribution of stochastic Gompertz equation under regime switching. Mathematics and Computers in Simulation, 97:192–206, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413002292.

Svetov:2014:NSB

[835] I. E. Svetov, E. Yu. Derevtsov, Yu. S. Volkov, and T. Schuster. A numerical solver based on B-splines for 2D vector field tomography in a refracting medium. Mathematics and Computers in Simulation, 97:207–223, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413002346.

Liu:2014:NNT

[836] Xinwu Liu and Lihong Huang. A new nonlocal total variation regularization algorithm for image denoising. Mathematics and Computers in Simulation, 97:224–233, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413002334.

Abreu:2014:NMT

[837] Eduardo Abreu. Numerical modelling of three-phase immiscible flow in heterogeneous porous media with gravitational effects. Mathematics and Computers in Simulation, 97: 234–259, March 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413002620.

LeGuyader:2014:SAG

[838] Carole Le Guyader, Dominique Apprato, and Christian Gout. Spline approximation of gradient fields: Applications to wind velocity fields. Mathematics and Computers in Simulation, 97:260–279, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002796.

Anonymous:2014:NIc

[839] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 97:280, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002954.

Anonymous:2014:ICEc

[840] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 97:281–283, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475413002966.

Anonymous:2014:PA

[841] Anonymous. Pages 1–72 (April 2014). Mathematics and Computers in Simulation, 98:1–72, April 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2014:EBd

[842] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 98:CO2, April 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000068.

Reisen:2014:SAE

[843] Valdério A. Reisen, Bartolomeu Zamprogno, Wilfredo Palma, and Josu Arteche. A semiparametric approach to estimate two seasonal fractional parameters in the SARFIMA model. Mathematics and Computers in Simulation, 98:1–17, April 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002863.

Yilmaz:2014:DLM

[844] Hulya Yilmaz and Hakan S. Sazak. Double-looped maximum likelihood estimation for the parameters of the generalized gamma distribution. Mathematics and Computers in Simulation, 98:18–30, April 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413003042.

Orlowska-Kowalska:2014:SMD

[845] T. Orlowska-Kowalska, G. Tarchala, and M. Dybkowski. Sliding-mode direct torque control and sliding-mode observer with a magnetizing reactance estimator for the field-weakening of the induction motor drive. Mathematics and Computers in Simulation, 98:31–45, April 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413001560.

Abbes:2014:LCC

[846] Dhaker Abbes, André Martinez, and Gérard Champenois. Life cycle cost, embodied energy and loss of power supply probability for the optimal design of hybrid power systems. Mathematics and Computers in Simulation, 98:46–62, April 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541300089X.

Anonymous:2014:NId

[847] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 98:63, April 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541400010X.

Anonymous:2014:ICEd

[848] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 98:64–66, April 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475414000111.

Anonymous:2014:EBe

[849] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 99:CO2, May 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000214.

Anonymous:2014:Pb

[850] Anonymous. Preface. Mathematics and Computers in Simulation, 99:vii, May 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414000329.

Jimenez-Gamero:2014:BRD

[851] M. D. Jiménez-Gamero, M. V. Alba-Fernández, and M. D. Estudillo-Martínez. Burbea–Rao divergence based statistics for testing uniform association. Mathematics and Computers in Simulation, 99: 1–18, May 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413000621.

Allouch:2014:IMF

[852] C. Allouch and P. Sablonnière. Iteration methods for Fredholm integral equations of the second kind based on spline quasi-interpolants. Mathematics and Computers in Simulation, 99:19–27, May 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000815.

BenGharbia:2014:GPA

[853] Ibtihel Ben Gharbia and Jérôme Jaffré. Gas phase appearance and disappearance as a problem with complementarity constraints. Mathematics and Computers in Simulation, 99: 28–36, May 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413001924.

Campillo:2014:AFP

[854] Fabien Campillo, Marc Joannides, and Irène Larramendy-Valverde. Approximation of the Fokker–Planck equation of the stochastic chemostat. Mathematics and Computers in Simulation, 99:37–53, May 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000785.

Rebollo:2014:NAS

[855] Tomás Chacón Rebollo, Frédéric Hecht, Macarena Gómez Mármol, Giordano Orzetti, and Samuele Rubino. Numerical approximation of the Smagorinsky turbulence model applied to the primitive equations of the ocean. Mathematics and Computers in Simulation, 99:54–70, May 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002012.

Fortes:2014:HFM

[856] M. A. Fortes, P. González, M. Pasadas, and M. L. Rodríguez. A hole filling method for explicit and parametric surfaces by using C^{1}-Powell Sabin splines. Mathematics and Computers in Simulation, 99:71–81, May 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000797.

Almhdie:2014:IPN

[857] Ahmad Almhdie, Olivier Rozenbaum, Eric Lespessailles, and Rachid Jennane. Image processing for the non-destructive characterization of porous media. Application to limestones and trabecular bones. Mathematics and Computers in Simulation, 99:82–94, May 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413002036.

Bozzini:2014:RFM

[858] Mira Bozzini and Licia Lenarduzzi. Recovering functions: a method based on domain decomposition. Mathematics and Computers in Simulation, 99:95–107, May 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413000670.

Lamnii:2014:NBC

[859] M. Lamnii, H. Mraoui, A. Tijini, and A. Zidna. A normalized basis for C^{1} cubic super spline space on Powell–Sabin triangulation. Mathematics and Computers in Simulation, 99:108–124, May 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413001511.

Francisco:2014:DIM

[860] A. Francisco, V. Ginting, F. Pereira, and J. Rigelo. Design and implementation of a multiscale mixed method based on a nonoverlapping domain decomposition procedure. Mathematics and Computers in Simulation, 99:125–138, May 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413001997.

Ginting:2014:RQU

[861] V. Ginting, F. Pereira, and A. Rahunanthan. Rapid quantification of uncertainty in permeability and porosity of oil reservoirs for enabling predictive simulation. Mathematics and Computers in Simulation, 99: 139–152, May 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413000827.

Bilanceri:2014:ITA

[862] M. Bilanceri, F. Beux, I. Elmahi, H. Guillard, and M. V. Salvetti. Implicit time advancing combined with two finite-volume methods in the simulation of morphodynamic flows. Mathematics and Computers in Simulation, 99:153–169, May 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002024.

Benkhaldoun:2014:SML

[863] Fayssal Benkhaldoun, Saida Sari, and Mohammed Seaid. A simple multilayer finite volume solver for densitydriven shallow water flows. Mathematics and Computers in Simulation, 99:170–189, May 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000839.

Arregui:2014:ANM

[864] I. Arregui, J. J. Cendán, and C. Vázquez. Adaptive numerical methods for an hydrodynamic problem arising in magnetic reading devices. Mathematics and Computers in Simulation, 99:190–202, May 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541300061X.

Anonymous:2014:NIe

[865] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 99:203, May 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414000275.

Anonymous:2014:ICEe

[866] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 99:204–206, May 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475414000287.

Anonymous:2014:PJa

[867] Anonymous. Pages 1–124 (June 2014). Mathematics and Computers in Simulation, 100:1–124, June 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2014:EBf

[868] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 100:CO2, June 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414000597.

Ding:2014:DCG

[869] Zhanwen Ding, Xiaofeng Zhu, and Shumin Jiang. Dynamical Cournot game with bounded rationality and time delay for marginal profit. Mathematics and Computers in Simulation, 100:1–12, June 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000032.

Wei:2014:NFE

[870] Zhouchao Wei, Rongrong Wang, and Anping Liu. A new finding of the existence of hidden hyperchaotic attractors with no equilibria. Mathematics and Computers in Simulation, 100:13–23, June 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000020.

Casals:2014:MCL

[871] J. Casals, S. Sotoca, and M. Jerez. Minimally conditioned likelihood for a nonstationary state space model. Mathematics and Computers in Simulation, 100:24–40, June 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000172.

Diele:2014:ISS

[872] Fasma Diele, Carmela Marangi, and Stefania Ragni. IMSP schemes for spatially explicit models of cyclic populations and metapopulation dynamics. Mathematics and Computers in Simulation, 100:41–53, June 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000366.

Imai:2014:CLD

[873] Junichi Imai. Comparison of low discrepancy mesh methods for pricing Bermudan options under a Lévy process. Mathematics and Computers in Simulation, 100:54– 71, June 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414000378.

Das:2014:SSD

[874] Saptarshi Das, Anish Acharya, and Indranil Pan. Simulation studies on the design of optimum PID controllers to suppress chaotic oscillations in a family of Lorenz-like multi-wing attractors. Mathematics and Computers in Simulation, 100:72–87, June 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000469.

Buonomo:2014:QAO

[875] Bruno Buonomo, Deborah Lacitignola, and Cruz Vargas-De-León. Qualitative analysis and optimal control of an epidemic model with vaccination and treatment. Mathematics and Computers in Simulation, 100: 88–102, June 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414000391.

Harase:2014:LRM

[876] Shin Harase. On the F_{2}-linear relations of Mersenne Twister pseudorandom number generators. Mathematics and Computers in Simulation, 100:103–113, June 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541400038X.

Anonymous:2014:NIf

[877] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 100:114, June 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000639.

Anonymous:2014:ICEf

[878] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 100:115–117, June 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475414000640.

Anonymous:2014:PJb

[879] Anonymous. Pages 1–122 (July 2014). Mathematics and Computers in Simulation, 101:1–122, July 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2014:EBg

[880] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 101:CO2, July 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000767.

Chapko:2014:NSP

[881] R. Chapko, B. T. Johansson, and V. Vavrychuk. Numerical solution of parabolic Cauchy problems in planar corner domains. Mathematics and Computers in Simulation, 101: 1–12, July 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414000457.

Loos:2014:TAH

[882] Florian Loos, Karl Dvorsky, and Hans-Dieter Liess. Two approaches for heat transfer simulation of current carrying multicables. Mathematics and Computers in Simulation, 101:13–30, July 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000500.

Kim:2014:ASM

[883] Seong-S. Kim and Han Ho Choi. Adaptive synchronization method for chaotic permanent magnet synchronous motor. Mathematics and Computers in Simulation, 101:31–42, July 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000536.

Huang:2014:MAP

[884] Zhenkun Huang, Sannay Mohamad, and Feng Gao. Multi-almost periodicity in semi-discretizations of a general class of neural networks. Mathematics and Computers in Simulation, 101:43–60, July 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000470.

Johansson:2014:MMI

[885] B. Tomas Johansson, Daniel Lesnic, and Thomas Reeve. A meshless method for an inverse two-phase onedimensional nonlinear Stefan problem. Mathematics and Computers in Simulation, 101:61–77, July 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000524.

Liu:2014:SSE

[886] Shuangzhe Liu, Tiefeng Ma, and Wolfgang Polasek. Spatial system estimators for panel models: a sensitivity and simulation study. Mathematics and Computers in Simulation, 101:78–102, July 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000512.

Yang:2014:NNN

[887] Yongqing Yang, Jinde Cao, Xianyun Xu, Manfeng Hu, and Yun Gao. A new neural network for solving quadratic programming problems with equality and inequality constraints. Mathematics and Computers in Simulation, 101:103–112, July 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541400055X.

Anonymous:2014:NIg

[888] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 101:113, July 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414000809.

Anonymous:2014:ICEg

[889] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 101:114–116, July 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475414000810.

Anonymous:2014:EBh

[890] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 102:ifc, August 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414000950.

Anonymous:2014:Pc

[891] Anonymous. Preface. Mathematics and Computers in Simulation, 102: vii, August 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414001062.

Ibanez:2014:CCT

[892] M. J. Ibáñez, J. B. Roldán, A. M. Roldán, and R. Yáñez. A comprehensive characterization of the threshold voltage extraction in MOSFETs transistors based on smoothing splines. Mathematics and Computers in Simulation, 102:1–10, August 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002085.

Izquierdo:2014:CRI

[893] Diego Izquierdo, María Cruz López de Silanes, María Cruz Parra, and Juan José Torrens. CS-RBF interpolation of surfaces with vertical faults from scattered data. Mathematics and Computers in Simulation, 102:11–23, August 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002127.

Fernandez:2014:ACG

[894] M. V. Alba Fernández, M. D. Jiménez Gamero, and S. Castillo Gutiérrez. Approximating a class of goodnessof-fit test statistics. Mathematics and Computers in Simulation, 102: 24–38, August 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413002139.

Boutoulout:2014:USP

[895] A. Boutoulout, R. El Ayadi, and M. Ouzahra. An unbounded stabilization problem of bilinear systems. Mathematics and Computers in Simulation, 102:39–50, August 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413002164.

Munoz:2014:NRD

[896] J. F. Muñoz, A. Arcos, E. Álvarez, and M. Rueda. New ratio and difference estimators of the finite population distribution function. Mathematics and Computers in Simulation, 102:51–61, August 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413002176.

Galiano:2014:FEA

[897] Gonzalo Galiano and Julián Velasco. Finite element approximation of a surface–subsurface coupled problem arising in forest dynamics. Mathematics and Computers in Simulation, 102:62–75, August 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541300219X.

Kouibia:2014:OPS

[898] A. Kouibia and M. Pasadas. Optimization of the parameters of surfaces by interpolating variational bicubic splines. Mathematics and Computers in Simulation, 102:76– 89, August 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413002267.

Durany:2014:SAL

[899] J. Durany, J. Pereira-Pérez, and F. Varas. Some aspects of lubrication in heavy regimes: Thermal effects, stability and turbulence. Mathematics and Computers in Simulation, 102:90–103, August 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002280.

ElMehdi:2014:ISF

[900] Rachida El Mehdi and Christian M. Hafner. Inference in stochastic frontier analysis with dependent error terms. Mathematics and Computers in Simulation, 102:104– 116, August 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413002310. See

erratum [937].

Azroul:2014:SQP

[901] E. Azroul, M. B. Benboubker, and M. Rhoudaf. On some p ( x )- quasilinear problem with right-hand side measure. Mathematics and Computers in Simulation, 102:117– 130, August 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413002322.

ElMehdi:2014:WIS

[902] Rachida El Mehdi and Christian M. Hafner. WITHDRAWN: Inference in stochastic frontier analysis with dependent error terms. Mathematics and Computers in Simulation, 102: 131–143, August 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413002632.

Garach:2014:DAE

[903] L. Garach, J. de Oña, and M. Pasadas. Determination of alignments in existing roads by using spline techniques. Mathematics and Computers in Simulation, 102:144–152, August 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475413003066.

Abderrahim:2014:NIL

[904] Elmoataz Abderrahim, Desquesnes Xavier, Lakhdari Zakaria, and Lézoray Olivier. Nonlocal infinity Laplacian equation on graphs with applications in image processing and machine learning. Mathematics and Computers in Simulation, 102:153– 163, August 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414000421.

Anonymous:2014:NIh

[905] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 102:164, August 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414001025.

Anonymous:2014:ICEh

[906] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 102:165–167, August 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475414001037.

Anonymous:2014:PS

[907] Anonymous. Pages 1–64 (September 2014). Mathematics and Computers in Simulation, 103:1–64, September 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2014:EBi

[908] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 103:ifc, September 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001128.

Han:2014:MEM

[909] Chuan-Hsiang Han, German Molina, and Jean-Pierre Fouque. McMC estimation of multiscale stochastic volatility models with applications. Mathematics and Computers in Simulation, 103:1–11, September 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000482.

Gupta:2014:DWR

[910] Nupur Gupta and Neela Nataraj. A dual weighted residual method for an optimal control problem of laser surface hardening of steel. Mathematics and Computers in Simulation, 103:12–32, September 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000494.

Chi:2014:QMC

[911] Hongmei Chi and Peter Beerli. Quasi-Monte Carlo method in population genetics parameter estimation. Mathematics and Computers in Simulation, 103:33–38, September 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414000548.

Neta:2014:BAS

[912] Beny Neta and Changbum Chun. Basins of attraction for several optimal fourth order methods for multiple roots. Mathematics and Computers in Simulation, 103:39–59, September 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414000688.

Anonymous:2014:NIi

[913] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 103:60, September 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414001165.

Anonymous:2014:ICEi

[914] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 103:61–63, September 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414001177.

Anonymous:2014:EBj

[915] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 104:ifc, October 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001293.

Hernando:2014:FSI

[916] Antonio Hernando, Francisco Botana, Eugenio Roanes-Lozano, and Michael J. Wester. Foreword to the special issue on “Nonstandard applications of computer algebra”. Mathematics and Computers in Simulation, 104: 1–2, October 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541400127X.

Botana:2014:PAD

[917] Francisco Botana. A parametric approach to 3D dynamic geometry. Mathematics and Computers in Simulation, 104:3–20, October 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413001390.

Aguilera:2014:ATS

[918] Gabriel Aguilera, José Luis Galán, José Manuel García, Enrique Mérida, and Pedro Rodríguez. An acceleratedtime simulation of car traffic on a motorway using a CAS. Mathematics and Computers in Simulation, 104:21–30, October 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412001140.

Roanes-Lozano:2014:RFV

[919] Eugenio Roanes-Lozano, Jose Antonio Alonso, and Antonio Hernando. Revisiting four-valued logics from Maple using the Logics Explorer package. Mathematics and Computers in Simulation, 104:31–42, October 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412001486.

Recio:2014:FAM

[920] Tomas Recio, J.-Rafael Sendra, Luis-Felipe Tabera, and Carlos Villarino. Factoring analytic multivariate polynomials and non-standard Cauchy– Riemann conditions. Mathematics and Computers in Simulation, 104: 43–57, October 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413001407.

Aguilera-Venegas:2014:ATS

[921] Gabriel Aguilera-Venegas, José L. Galán-García, Enrique Mérida-Casermeiro, and Pedro Rodríguez-Cielos. An accelerated-time simulation of baggage traffic in an airport terminal. Mathematics and Computers in Simulation, 104:58–66, October 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541400072X.

Montes:2014:GSL

[922] Antonio Montes and Tomás Recio. Generalizing the Steiner–Lehmus theorem using the Gröbner cover. Mathematics and Computers in Simulation, 104:67–81, October 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413001961.

Saenz-de-Cabezon:2014:MRN

[923] E. Sáenz de Cabezón and H. P. Wynn. Measuring the robustness of a network using minimal vertex covers. Mathematics and Computers in Simulation, 104:82–94, October 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000731.

Anonymous:2014:NIj

[924] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 104:95, October 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001360.

Anonymous:2014:ICEj

[925] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 104:96–98, October 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414001372.

Anonymous:2014:PN

[926] Anonymous. Pages 1–164 (November 2014). Mathematics and Computers in Simulation, 105:1–164, November 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2014:EBk

[927] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 105:ifc, November 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414001852.

Aulisa:2014:CTP

[928] Eugenio Aulisa and Sophia R.-J. Jang. Continuous-time predator– prey systems with Allee effects in the prey. Mathematics and Computers in Simulation, 105:1–16, November 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414000895.

Haghighi:2014:AAM

[929] A. Haghighi and S. M. Hosseini. Analysis of asymptotic mean-square stability of a class of Runge–Kutta schemes for linear systems of stochastic differential equations. Mathematics and Computers in Simulation, 105:17–48, November 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414001207.

Magrenan:2014:RQB

[930] Á. Alberto Magreñán, Alicia Cordero, José M. Gutiérrez, and Juan R. Torregrosa. Real qualitative behavior of a fourth-order family of iterative methods by using the convergence plane. Mathematics and Computers in Simulation, 105:49– 61, November 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414001086.

Senouci:2014:PCS

[931] Abdelkader Senouci and Abdelkrim Boukabou. Predictive control and synchronization of chaotic and hyperchaotic systems based on a T– S fuzzy model. Mathematics and Computers in Simulation, 105:62– 78, November 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414001268.

Hamida:2014:RAH

[932] Mohamed Assaad Hamida, Alain Glumineau, Jesus de Leon, and Luc Loron. Robust adaptive high order sliding-mode optimum controller for sensorless interior permanent magnet synchronous motors. Mathematics and Computers in Simulation, 105: 79–104, November 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001256.

Roustant:2014:CDB

[933] O. Roustant, J. Fruth, B. Iooss, and S. Kuhnt. Crossed-derivative based sensitivity measures for interaction screening. Mathematics and Computers in Simulation, 105:105– 118, November 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414001244.

Zhang:2014:DCD

[934] Limin Zhang, Chaofeng Zhang, and Min Zhao. Dynamic complexities in a discrete predator–prey system with lower critical point for the prey. Mathematics and Computers in Simulation, 105:119–131, November 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414001542.

Cui:2014:OIM

[935] Lijie Cui, Zhenzhou Lu, Pan Wang, and Weihu Wang. The ordering importance measure of random variable and its estimation. Mathematics and Computers in Simulation, 105: 132–143, November 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414001566.

Wang:2014:EEB

[936] Fu-Kwun Wang and Chih-Wen Lee. Mestimator for estimating the Burr type III parameters with outliers. Mathematics and Computers in Simulation, 105:144–159, November 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001505.

ElMehdi:2014:EIS

[937] Rachida El Mehdi and Christian M. Hafner. Erratum to “Inference in stochastic frontier analysis with dependent error terms” [Math. Comp. Simul. 102C (2014) 131–143]. Mathematics and Computers in Simulation, 105:160, November 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001591. See

[900].

Anonymous:2014:NIk

[938] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 105:161, November 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541400189X.

Anonymous:2014:ICEk

[939] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 105:162–163, November 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414001906.

Anonymous:2014:EBl

[940] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 106:ifc, December 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414002146.

Allasia:2014:LIS

[941] Giampietro Allasia, Roberto Cavoretto, and Alessandra De Rossi. Local interpolation schemes for landmarkbased image registration: a comparison. Mathematics and Computers in Simulation, 106:1–25, December 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475414001554.

Arabi:2014:UML

[942] Sina Arabi, Ricardo Camarero, and François Guibault. Unstructured meshes for large body motion using mapping operators. Mathematics and Computers in Simulation, 106: 26–43, December 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001103.

Cheng:2014:ASE

[943] Gary C. Cheng, Roy P. Koomullil, Yasushi Ito, Alan M. Shih, Somsak Sittitavornwong, and Peter D. Waite. Assessment of surgical effects on patients with obstructive sleep apnea syndrome using computational fluid dynamics simulations. Mathematics and Computers in Simulation, 106:44–59, December 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541200287X.

DeSantis:2014:ECF

[944] D. De Santis, G. Geraci, and A. Guardone. Equivalence conditions for the finite volume and finite element methods in spherical coordinates. Mathematics and Computers in Simulation, 106:60– 75, December 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541200119X.

Dominguez-Mota:2014:FDS

[945] Francisco J. Domínguez-Mota, Sanzon Mendoza Armenta, G. Tinoco-Guerrero, and J. G. Tinoco-Ruiz. Finite difference schemes satisfying an optimality condition for the unsteady heat equation. Mathematics and Computers in Simulation, 106: 76–83, December 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000718.

Marquez:2014:TSR

[946] Alberto Márquez, Auxiliadora Moreno-González, Ángel Plaza, and José P. Suárez. There are simple and robust refinements (almost) as good as Delaunay. Mathematics and Computers in Simulation, 106:84– 94, December 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001309.

Perdomo:2014:MPH

[947] Francisco Perdomo, Ángel Plaza, Eduardo Quevedo, and José P. Suárez.

A mathematical proof of how fast the diameters of a triangle mesh tend to zero after repeated trisection. Mathematics and Computers in Simulation, 106:95–108, December 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001979.

Bruni:2014:ARS

[948] V. Bruni, E. Rossi, and D. Vitulano. Automated restoration of semitransparent degradation via Lie groups and visibility laws. Mathematics and Computers in Simulation, 106: 109–123, December 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000724.

Capobianco:2014:WLI

[949] M. R. Capobianco and G. Criscuolo. Weighted Lagrange interpolation with preassigned nodes on the real line. Mathematics and Computers in Simulation, 106:124–132, December 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475414000196.

Cefalo:2014:OMG

[950] Massimo Cefalo and Maria Rosaria Lancia. An optimal mesh generation algorithm for domains with Koch type boundaries. Mathematics and Computers in Simulation, 106: 133–162, December 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001487.

Dassi:2014:EGR

[951] F. Dassi, S. Perotto, L. Formaggia, and P. Ruffo. Efficient geometric reconstruction of complex geological structures. Mathematics and Computers in Simulation, 106:163– 184, December 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414000184.

Pitolli:2014:TSP

[952] Francesca Pitolli. Ternary shapepreserving subdivision schemes. Mathematics and Computers in Simulation, 106:185–194, December 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000669.

Anonymous:2014:NIl

[953] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 106:195, December 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414002201.

Anonymous:2014:ICEl

[954] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 106:196–197, December 2014. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414002213.

Anonymous:2015:PJa

[955] Anonymous. Pages 1–160 (January 2015). Mathematics and Computers in Simulation, 107:1–160, January 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2015:EBa

[956] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 107:ifc, January 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414002493.

Khadraoui:2015:SMI

[957] Khader Khadraoui. A simple Markovian individual-based model as a means of understanding forest dynamics. Mathematics and Computers in Simulation, 107:1– 23, January 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847541400158X.

Sun:2015:MFE

[958] Tongjun Sun and Yirang Yuan. Mixed finite element method and the characteristics-mixed finite element method for a slightly compressible miscible displacement problem in porous media. Mathematics and Computers in Simulation, 107:24–45, January 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414001645.

Xu:2015:SCG

[959] Runzhang Xu, Xuemei Wang, Yi Niu, Mingyou Zhang, and Jie Liu. Sharp criterion of global existence for a class of nonlinear Schrödinger equations with critical exponent. Mathematics and Computers in Simulation, 107:46–51, January 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001578.

Zhou:2015:APS

[960] Tiejun Zhou, Min Wang, and Chen Li. Almost periodic solution for multidirectional associative memory neural network with distributed delays. Mathematics and Computers in Simulation, 107:52–60, January 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001931.

Zhao:2015:PDM

[961] Tingting Zhao and Yanni Xiao. Plant disease models with nonlinear impulsive cultural control strategies for vegetatively propagated plants. Mathematics and Computers in Simulation, 107:61–91, January 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001475.

Hernando:2015:AMI

[962] Antonio Hernando and Eugenio Roanes-Lozano. An algebraic model for implementing expert systems based on the knowledge of different experts. Mathematics and Computers in Simulation, 107:92– 107, January 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414001220.

Sellier:2015:CAS

[963] J. M. Sellier, M. Nedjalkov, I. Dimov, and S. Selberherr. A comparison of approaches for the solution of the Wigner equation. Mathematics and Computers in Simulation, 107:108–119, January 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001530.

Conceicao:2015:SMS

[964] D. Conceição and W. Lambert. Statistical moments for solutions of nonlinear scalar equations with random Riemann data. Mathematics and Computers in Simulation, 107:120– 133, January 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414001232.

Fahs:2015:HOD

[965] Hassan Fahs. High-order discontinuous Galerkin method for time-domain electromagnetics on non-conforming hybrid meshes. Mathematics and Computers in Simulation, 107:134– 156, January 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414000871.

Anonymous:2015:NIa

[966] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 107:157, January 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414002547.

Anonymous:2015:ICEa

[967] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 107:158–159, January 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414002559.

Anonymous:2015:EBb

[968] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 108:ifc, February 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414002717.

Gardini:2015:FSI

[969] Laura Gardini, Fabio Lamantia, and Fabio Tramontana. Foreword to the special issue of Mathematics and Computers in Simulation on complex dynamics in economics and finance. Mathematics and Computers in Simulation, 108:1–2, February 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414002365.

Huang:2015:HAM

[970] Weihong Huang and Zhenxi Chen. Heterogeneous agents in multi-markets: a coupled map lattices approach. Mathematics and Computers in Simulation, 108:3–15, February 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002814.

Tramontana:2015:SFM

[971] Fabio Tramontana, Frank Westerhoff, and Laura Gardini. A simple financial market model with chartists and fundamentalists: Market entry levels and discontinuities. Mathematics and Computers in Simulation, 108:16–40, February 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413001869.

Agliari:2015:CMI

[972] Anna Agliari, Francesco Rillosi, and George Vachadze. Credit market imperfection, financial market globalization, and catastrophic transition. Mathematics and Computers in Simulation, 108:41–62, February 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414001499.

Commendatore:2015:TBS

[973] Pasquale Commendatore, Ingrid Kubin, and Iryna Sushko. Typical bifurcation scenario in a three region identical New Economic Geography model. Mathematics and Computers in Simulation, 108:63– 80, February 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414000160.

Agliari:2015:ADF

[974] Anna Agliari, Pasquale Commendatore, Ilaria Foroni, and Ingrid Kubin. Agglomeration dynamics and first nature asymmetries. Mathematics and Computers in Simulation, 108:81–98, February 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001098.

Matsumoto:2015:DMM

[975] Akio Matsumoto and Ferenc Szidarovszky. Dynamic monopoly with multiple continuously distributed time delays. Mathematics and Computers in Simulation, 108:99–118, February 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414000159.

Panchuk:2015:OMR

[976] A. Panchuk and T. Puu. Oligopoly model with recurrent renewal of capital revisited. Mathematics and Computers in Simulation, 108:119– 128, February 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413002309.

Fanti:2015:NDC

[977] Luciano Fanti, Luca Gori, and Mauro Sodini. Nonlinear dynamics in a Cournot duopoly with isoelastic demand. Mathematics and Computers in Simulation, 108:129– 143, February 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413002279.

Merlone:2015:DOC

[978] Ugo Merlone and Ferenc Szidarovszky. Dynamic oligopolies with contingent workforce and investment costs. Mathematics and Computers in Simulation, 108:144–154, February 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414000408.

Lamantia:2015:ERR

[979] F. Lamantia and D. Radi. Exploitation of renewable resources with differentiated technologies: an evolutionary analysis. Mathematics and Computers in Simulation, 108:155– 174, February 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413002838.

Travaglini:2015:NDP

[980] Giuseppe Travaglini. Nonlinear dynamic pollution under uncertainty and binding targets. Mathematics and Computers in Simulation, 108: 175–183, February 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001621.

Privileggi:2015:TVS

[981] Fabio Privileggi. Takeoff vs. stagnation in endogenous recombinant growth models. Mathematics and Computers in Simulation, 108:184– 214, February 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541400161X.

Michetti:2015:CAB

[982] Elisabetta Michetti. Complex attractors and basins in a growth model with nonconcave production function and logistic population growth rate. Mathematics and Computers in Simulation, 108:215–232, February 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002243.

Colucci:2015:SIS

[983] Domenico Colucci and Vincenzo Valori. Stabilizing inflation in a simple monetary policy model with heterogeneous agents. Mathematics and Computers in Simulation, 108:233– 244, February 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413002255.

Landini:2015:IMM

[984] Simone Landini, Mariacristina Uberti, and Simone Casellina. Italian mortgage markets and their dynamics. Mathematics and Computers in Simulation, 108:245–259, February 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414001074.

Sgrignoli:2015:INE

[985] P. Sgrignoli, E. Agliari, R. Burioni, and A. Schianchi. Instability and network effects in innovative markets. Mathematics and Computers in Simulation, 108:260–271, February 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414001633.

Anonymous:2015:NIb

[986] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 108:272, February 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414002778.

Anonymous:2015:ICEb

[987] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 108:273–274, February 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847541400278X.

Anonymous:2015:PMa

[988] Anonymous. Pages 1–218 (March 2015). Mathematics and Computers in Simulation, 109:1–218, March 2015.

CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2015:EBc

[989] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 109:ifc, March 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414002882.

Jiang:2015:NGB

[990] Beini Jiang, Weizhong Dai, Abdul Khaliq, Michelle Carey, Xiaobo Zhou, and Le Zhang. Novel 3D gpu based numerical parallel diffusion algorithms in cylindrical coordinates for health care simulation. Mathematics and Computers in Simulation, 109:1–19, March 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001608.

Cariaga:2015:EHP

[991] Emilio Cariaga, Rubén Martínez, and Mauricio Sepúlveda. Estimation of hydraulic parameters under unsaturated flow conditions in heap leaching. Mathematics and Computers in Simulation, 109:20–31, March 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001943.

Zhusubaliyev:2015:MHA

[992] Zhanybai T. Zhusubaliyev and Erik Mosekilde. Multistability and hidden attractors in a multilevel DC/ DC converter. Mathematics and Computers in Simulation, 109:32– 45, March 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414001967.

Ch:2015:KRF

[993] Aswani Kumar Ch., Sérgio M. Dias, and Newton J. Vieira. Knowledge reduction in formal contexts using nonnegative matrix factorization. Mathematics and Computers in Simulation, 109:46–63, March 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001992.

Gao:2015:NSM

[994] Fuzheng Gao, Feng Qiao, and Hongxing Rui. Numerical simulation of the modified regularized long wave equation by split least-squares mixed finite element method. Mathematics and Computers in Simulation, 109:64–73, March 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414002006.

Chun:2015:BAZ

[995] Changbum Chun and Beny Neta. Basins of attraction for Zhou–Chen– Song fourth order family of methods for multiple roots. Mathematics and Computers in Simulation, 109:74–91, March 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414002031.

Yang:2015:MPI

[996] Jin Yang, Sanyi Tang, and Robert A. Cheke. Modelling pulsed immunotherapy of tumour–immune interaction. Mathematics and Computers in Simulation, 109:92–112, March 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475414002341.

Chen-Charpentier:2015:CAG

[997] B.-M. Chen-Charpentier, J.-C. Cortés, J.-A. Licea, J.-V. Romero, M.-D. Roselló, Francisco-J. Santonja, and Rafael-J. Villanueva. Constructing adaptive generalized polynomial chaos method to measure the uncertainty in continuous models: a computational approach. Mathematics and Computers in Simulation, 109:113– 129, March 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414002353.

Wang:2015:PIA

[998] Hengxu Wang, John G. O’Hara, and Nick Constantinou. A pathindependent approach to integrated variance under the CEV model. Mathematics and Computers in Simulation, 109:130–152, March 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414002377.

Wu:2015:CSP

[999] Yusen Wu, Peiluan Li, and Haibo Chen. Calculation of singular point quantities at infinity for a type of polynomial differential systems. Mathematics and Computers in Simulation, 109:153–173, March 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541400247X.

Guin:2015:SPT

[1000] Lakshmi Narayan Guin. Spatial patterns through Turing instability in a reaction–diffusion predator– prey model. Mathematics and Computers in Simulation, 109:174– 185, March 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414002468.

Reppas:2015:TAP

[1001] Andreas I. Reppas, Konstantinos Spiliotis, and Constantinos I. Siettos. Tuning the average path length of complex networks and its influence to the emergent dynamics of the majority-rule model. Mathematics and Computers in Simulation, 109: 186–196, March 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414002456.

Ouanes:2015:NQL

[1002] Mohand Ouanes, Hoai An Le Thi, Trong Phuc Nguyen, and Ahmed Zidna. New quadratic lower bound for multivariate functions in global optimization. Mathematics and Computers in Simulation, 109:197– 211, March 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414002444.

Nawata:2015:CED

[1003] K. Nawata and K. Kawabuchi. Corrigendum to “Evaluation of the DPCbased inclusive payment system in Japan for cataract operations by a new model” [Math. Comput. Simulation 93 (2013) 76–85]. Mathematics and Computers in Simulation, 109:212–214, March 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001980. See

[749].

Anonymous:2015:NIc

[1004] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 109:215, March 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414002936.

Anonymous:2015:ICEc

[1005] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 109:217–218, March 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475414002948.

Anonymous:2015:EBd

[1006] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 110:ifc, April 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415000427.

DelBuono:2015:GES

[1007] Nicoletta Del Buono. Guest editorial: Structural dynamical systems, discontinuity and numerical methods. Mathematics and Computers in Simulation, 110:1–2, April 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415000026.

Dieci:2015:SSA

[1008] L. Dieci, C. Elia, and L. Lopez. Sharp sufficient attractivity conditions for sliding on a co-dimension 2 discontinuity surface. Mathematics and Computers in Simulation, 110: 3–14, April 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414000354.

Weiss:2015:IMN

[1009] D. Weiss, T. Küpper, and H. A. Hosham. Invariant manifolds for nonsmooth systems with sliding mode. Mathematics and Computers in Simulation, 110:15–32, April 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541400041X.

Shahrear:2015:DPL

[1010] Pabel Shahrear, Leon Glass, Roy Wilds, and Rod Edwards. Dynamics in piecewise linear and continuous models of complex switching networks.

Mathematics and Computers in Simulation, 110:33–39, April 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413003054.

Pace:2015:SSE

[1011] Brigida Pace, Fasma Diele, and Carmela Marangi. Splitting schemes and energy preservation for separable Hamiltonian systems. Mathematics and Computers in Simulation, 110:40–52, April 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002875.

Politi:2015:SPM

[1012] Tiziano Politi and Marina Popolizio.

On stochasticity preserving methods for the computation of the matrix p th root. Mathematics and Computers in Simulation, 110:53– 68, April 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414000044.

Halanay:2015:SAE

[1013] A. Halanay, D. Cândea, and I. R.

Rădulescu. Stability analysis of equilibria in a delay differentialequations model of CML including asymmetric division and treatment. Mathematics and Computers in Simulation, 110:69–82, April 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001219.

Diele:2015:ISS

[1014] Fasma Diele, Carmela Marangi, and Stefania Ragni. IMSP schemes for spatially explicit models of cyclic populations and metapopulation dynamics. Mathematics and Computers in Simulation, 110:83–95, April 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415000099.

Garrappa:2015:TMF

[1015] Roberto Garrappa. Trapezoidal methods for fractional differential equations: Theoretical and computational aspects. Mathematics and Computers in Simulation, 110:96– 112, April 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413002802.

DAmbrosio:2015:GFN

[1016] Raffaele D’Ambrosio and Beatrice Paternoster. A general framework for the numerical solution of second order ODEs. Mathematics and Computers in Simulation, 110:113– 124, April 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414001104.

Cardone:2015:EGD

[1017] A. Cardone, L. Gr. Ixaru, B. Paternoster, and G. Santomauro. Ef-Gaussian direct quadrature methods for Volterra integral equations with periodic solution. Mathematics and Computers in Simulation, 110:125–143, April 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413003029.

Basile:2015:CNM

[1018] M. Basile, E. Messina, W. Themistoclakis, and A. Vecchio. Convergence of a numerical method for the solution of non-standard integro-differential boundary value problems. Mathematics and Computers in Simulation, 110:144–154, April 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413003030.

Messina:2015:NSD

[1019] E. Messina and A. Vecchio. Nonlinear stability of direct quadrature methods for Volterra integral equations. Mathematics and Computers in Simulation, 110:155–164, April 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413000700.

Anonymous:2015:NId

[1020] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 110:165, April 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541500049X.

Anonymous:2015:ICEd

[1021] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 110:166–167, April 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475415000506.

Anonymous:2015:PMb

[1022] Anonymous. Pages 1–98 (May 2015). Mathematics and Computers in Simulation, 111:1–98, May 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2015:EBe

[1023] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 111:ifc, May 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415000142.

Alonso-Mallo:2015:ABC

[1024] I. Alonso-Mallo and A. M. Portillo. Absorbing boundary conditions and geometric integration: a case study for the wave equation. Mathematics and Computers in Simulation, 111:1–16, May 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414003218.

Reboredo:2015:DRE

[1025] Juan C. Reboredo and Mikel Ugando. Downside risks in EU carbon and fossil fuel markets. Mathematics and Computers in Simulation, 111: 17–35, May 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414003255.

Boujraf:2015:SMC

[1026] A. Boujraf, D. Sbibih, M. Tahrichi, and A. Tijini. A simple method for constructing integro spline quasiinterpolants. Mathematics and Computers in Simulation, 111:36– 47, May 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414003188.

Sun:2015:BBM

[1027] Jianjun Sun, Weixing Wang, Lina Zhao, and Lihong Cui. Biorthogonal balanced multiwavelets with high armlets order and their application in image denoising. Mathematics and Computers in Simulation, 111: 48–62, May 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414003322.

Runnels:2015:CPE

[1028] Scott R. Runnels. Capturing plasticity effects in overdriven shocks on the finite scale. Mathematics and Computers in Simulation, 111: 63–79, May 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414003334.

Pulch:2015:SAM

[1029] Roland Pulch, E. Jan W. ter Maten, and Florian Augustin. Sensitivity analysis and model order reduction for random linear dynamical systems. Mathematics and Computers in Simulation, 111:80–95, May 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541500004X.

Anonymous:2015:NIe

[1030] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 111:96, May 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415000191.

Anonymous:2015:ICEe

[1031] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 111:97–98, May 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475415000208.

Anonymous:2015:PJb

[1032] Anonymous. Pages 1–58 (June 2015). Mathematics and Computers in Simulation, 112:1–58, June 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2015:EBf

[1033] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 112:ifc, June 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415000282.

Rincon:2015:NAC

[1034] A. Rincón, D. Piarpuzán, and F. Angulo. A new adaptive controller for bio-reactors with unknown kinetics and biomass concentration: Guarantees for the boundedness and convergence properties. Mathematics and Computers in Simulation, 112: 1–13, June 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415000063.

Soriano-Sanchez:2015:PIC

[1035] A. G. Soriano-Sánchez, C. Posadas-Castillo, M. A. Platas-Garza, and D. A. Diaz-Romero. Performance improvement of chaotic encryption via energy and frequency location criteria. Mathematics and Computers in Simulation, 112:14–27, June 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415000087.

Ghayesh:2015:IET

[1036] Mergen H. Ghayesh and Hamed Farokhi. Internal energy transfer in dynamical behaviour of Timoshenko microarches. Mathematics and Computers in Simulation, 112: 28–39, June 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415000075.

Chen:2015:NLT

[1037] Yuming Chen and Qigui Yang. A new Lorenz-type hyperchaotic system with a curve of equilibria. Mathematics and Computers in Simulation, 112:40–55, June 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541400305X.

Anonymous:2015:NIf

[1038] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 112:56, June 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415000336.

Anonymous:2015:ICEf

[1039] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 112:57–58, June 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475415000348.

Anonymous:2015:PJc

[1040] Anonymous. Pages 1–76 (July 2015). Mathematics and Computers in Simulation, 113:1–76, July 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2015:EBg

[1041] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 113:ifc, July 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415000555.

Lenksas:2015:WAH

[1042] A. Lenkšas and V. Mackevičius. Weak approximation of Heston model by discrete random variables. Mathematics and Computers in Simulation, 113:1–15, July 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415000129.

Andaluz:2015:DEG

[1043] Joaquín Andaluz and Gloria Jarne.

On the dynamics of economic games based on product differentiation. Mathematics and Computers in Simulation, 113:16–27, July 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415000385.

Huang:2015:CSC

[1044] Li-Lian Huang, Juan Zhang, and Shuai-Shuai Shi. Circuit simulation on control and synchronization of fractional order switching chaotic system. Mathematics and Computers in Simulation, 113:28–39, July 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415000531.

Li:2015:DRS

[1045] Songyang Li, Jixun Gao, Qingsheng Zhu, Lingqiu Zeng, and Ji Liu. A dynamic root simulation model in response to soil moisture heterogeneity. Mathematics and Computers in Simulation, 113:40–50, July 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415000233.

Dang:2015:AOA

[1046] Quoc Viet Dang, Laurent Vermeiren, Antoine Dequidt, and Michel Dambrine. Augmented observer approach for high-impedance haptic system with time delay. Mathematics and Computers in Simulation, 113: 51–68, July 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415000117.

Anonymous:2015:NIg

[1047] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 113:69, July 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415000609.

Anonymous:2015:ICEg

[1048] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 113:71–72, July 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475415000610.

Anonymous:2015:EBh

[1049] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 114:CO2, August 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415000865.

Feng:2015:P

[1050] Xiao-Bin Feng, Ai-Xiang Huang, Jinyun Yuan, and Nelson Maculan. Preface. Mathematics and Computers in Simulation, 114:

1, August 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001081.

Cruz:2015:FCA

[1051] J. Y. Bello Cruz and A. N. Iusem. Full convergence of an approximate projection method for nonsmooth variational inequalities. Mathematics and Computers in Simulation, 114: 2–13, August 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410001916.

Carvalho:2015:OSP

[1052] Esdras P. Carvalho, Julian Martínez, J. M. Martínez, and Feodor Pisnitchenko. On optimization strategies for parameter estimation in models governed by partial differential equations. Mathematics and Computers in Simulation, 114:14– 24, August 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475410002570.

Shi:2015:LOC

[1053] Dongyang Shi, Qili Tang, and Wei Gong. A low order characteristicnonconforming finite element method for nonlinear Sobolev equation with convection-dominated term. Mathematics and Computers in Simulation, 114:25–36, August 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414000743.

Zhu:2015:TLS

[1054] Liping Zhu and Zhangxin Chen. A two-level stabilized nonconforming finite element method for the stationary Navier–Stokes equations. Mathematics and Computers in Simulation, 114:37–48, August 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412001978.

Anonymous:2015:NIh

[1055] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 114:49, August 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415000932.

Anonymous:2015:ICEh

[1056] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 114:50–51, August 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475415000944.

Anonymous:2015:PS

[1057] Anonymous. Pages 1–52 (September 2015). Mathematics and Computers in Simulation, 115:1–52, September 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2015:EBi 115:ifc, September 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415000737.

Anis:2015:MCC

[1059] M. Z. Anis and Abhik Ghosh.

Monte Carlo comparison of tests of exponentiality against NWBUE alternatives. Mathematics and Computers in Simulation, 115:1–11, September 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415000671.

Bensiali:2015:CAB

[1060] N. Bensiali, E. Etien, and N. Benalia. Convergence analysis of back-EMF MRAS observers used in sensorless control of induction motor drives. Mathematics and Computers in Simulation, 115:12–23, September 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415000658.

Singh:2015:CCF

[1061] Anuraj Singh and Sunita Gakkhar.

Controlling chaos in a food chain model. Mathematics and Computers in Simulation, 115:24–36, September 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415000646.

Herrera:2015:GPS

[1062] J. Herrera, A. Ibeas, M. de la Sen, E. Rivera, and J. Peláez. Generalized Pattern Search Methods for control of stable, unstable and integrating systems with unknown delay under step input. Mathematics and Computers in Simulation, 115: 37–48, September 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415000269.

Anonymous:2015:NIi

[1063] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 115:49, September 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415000786.

Anonymous:2015:ICEi

[1064] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 115:50–51, September 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415000798.

Anonymous:2015:PO

[1065] Anonymous. Pages 1–92 (October 2015). Mathematics and Computers in Simulation, 116:1–92, October 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2015:EBj

[1066] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 116:ifc, October 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415000981.

Jian:2015:GLS

[1067] Jigui Jian and Baoxian Wang.

Global Lagrange stability for neutraltype Cohen–Grossberg BAM neural networks with mixed timevarying delays. Mathematics and Computers in Simulation, 116:1– 25, October 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415000683.

Andersson:2015:AUF

[1068] Christian Andersson, Claus Führer, and Johan Åkesson. Assimulo: a unified framework for ODE solvers. Mathematics and Computers in Simulation, 116:26–43, October 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415000701.

LoSchiavo:2015:ANN

[1069] Mauro Lo Schiavo, Barbara Prinari, Jessica A. Gronski, and Angelo V. Serio. An artificial neural network approach for modeling the ward atmosphere in a medical unit. Mathematics and Computers in Simulation, 116:44–58, October 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415000695.

Chumakov:2015:EGE

[1070] G. A. Chumakov, E. A. Lashina, and N. A. Chumakova. On estimation of the global error of numerical solution on canard-cycles. Mathematics and Computers in Simulation, 116:59–74, October 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414002596.

Zaghi:2015:ADO

[1071] S. Zaghi, A. Di Mascio, R. Broglia, and R. Muscari. Application of dynamic overlapping grids to the simulation of the flow around a fullyappended submarine. Mathematics and Computers in Simulation, 116: 75–88, October 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414003000.

Anonymous:2015:NIj

[1072] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 116:89, October 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415001032.

Anonymous:2015:ICEj

[1073] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 116:90–91, October 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415001044.

Anonymous:2015:PN

[1074] Anonymous. Pages 1–156 (November 2015). Mathematics and Computers in Simulation, 117:1–156, November 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2015:EBk

[1075] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 117:ifc, November 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541500138X.

Rasulov:2015:MCS

[1076] Abdujabar Rasulov and Gulnora Raimova. Monte Carlo solution of the Neumann problem for the nonlinear Helmholtz equation. Mathematics and Computers in Simulation, 117:1–9, November 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541500083X.

Liu:2015:NDR

[1077] Q. X. Liu, J. K. Liu, and Y. M.

Chen. Non-diminishing relative error of the predictor–corrector algorithm for certain fractional differential equations. Mathematics and Computers in Simulation, 117:10– 19, November 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415000828.

Canovas:2015:ADP

[1078] Jose S. Cánovas, Anastasiia Panchuk, and Tönu Puu. Asymptotic dynamics of a piecewise smooth map modelling a competitive market. Mathematics and Computers in Simulation, 117:20–38, November 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541500107X.

Jiang:2015:NSQ

[1079] Hongling Jiang, Lijuan Wang, and Ruofei Yao. Numerical simulation and qualitative analysis for a predator–prey model with B–D functional response. Mathematics and Computers in Simulation, 117:39– 53, November 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001184.

Cheng:2015:MBP

[1080] Dongsheng Cheng, Zhiyong Liu, and Tingting Wu. A multigrid-based preconditioned solver for the Helmholtz equation with a discretization by 25point difference scheme. Mathematics and Computers in Simulation, 117: 54–67, November 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001160.

Kim:2015:QTF

[1081] Jinkyu Kim and Dongkeon Kim. A quadratic temporal finite element method for linear elastic structural dynamics. Mathematics and Computers in Simulation, 117:68– 88, November 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001226.

Nguyen:2015:IAO

[1082] Vinh Phu Nguyen, Cosmin Anitescu, Stéphane P. A. Bordas, and Timon Rabczuk. Isogeometric analysis: an overview and computer implementation aspects. Mathematics and Computers in Simulation, 117: 89–116, November 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001214.

Wongvanich:2015:MMM

[1083] N. Wongvanich, C. E. Hann, and H. R.

Sirisena. Minimal modeling methodology to characterize non-linear damping in an electromechanical system. Mathematics and Computers in Simulation, 117:117–140, November 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001202.

Xi:2015:BAH

[1084] Yanhui Xi, Hui Peng, Yemei Qin, Wenbiao Xie, and Xiaohong Chen. Bayesian analysis of heavy-tailed market microstructure model and its application in stock markets. Mathematics and Computers in Simulation, 117: 141–153, November 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415001251.

Anonymous:2015:NIk

[1085] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 117:154, November 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415001433.

Anonymous:2015:ICEk

[1086] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 117:155–156, November 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415001445.

Anonymous:2015:EBl

[1087] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 118:ifc, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415001615.

Amaziane:2015:P

[1088] Brahim Amaziane, Domingo Barrera, and Driss Sbibih. Preface. Mathematics and Computers in Simulation, 118:1, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415001597.

Abudawia:2015:NAS

[1089] A. Abudawia and C. Rosier. Numerical analysis for a seawater intrusion problem in a confined aquifer. Mathematics and Computers in Simulation, 118:2–16, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414003152.

Allouch:2015:DSN

[1090] C. Allouch and M. Tahrichi. Discrete superconvergent Nyström method for integral equations and eigenvalue problems. Mathematics and Computers in Simulation, 118:17– 29, December 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414003097.

Amat:2015:FNC

[1091] S. Amat, J. Liandrat, J. Ruiz, and J. C. Trillo. On a family of nonlinear cell-average multiresolution schemes for image processing: an experimental study. Mathematics and Computers in Simulation, 118: 30–48, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415000038.

Amat:2015:VSS

[1092] Sergio Amat, M. José Legaz, and Pablo Pedregal. A variable step-size implementation of a variational method for stiff differential equations. Mathematics and Computers in Simulation, 118:49–57, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414003139.

Barranco-Chamorro:2015:ESS

[1093] I. Barranco-Chamorro, J. L. Moreno-Rebollo, M. D. Jiménez-Gamero, and M. V. Alba-Fernández. Estimation of the sample size n based on record values. Mathematics and Computers in Simulation, 118:58– 72, December 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541400336X.

Bauzet:2015:TSA

[1094] Caroline Bauzet. Time-splitting approximation of the Cauchy problem for a stochastic conservation law. Mathematics and Computers in Simulation, 118:73–86, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414003115.

Benkhaldoun:2015:PFV

[1095] Fayssal Benkhaldoun, Saida Sari, and Mohammed Seaid. Projection finite volume method for shallow water flows. Mathematics and Computers in Simulation, 118:87– 101, December 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414003280.

Boujraf:2015:SQI

[1096] A. Boujraf, M. Tahrichi, and A. Tijini.

C^{1} Superconvergent quasi-interpolation based on polar forms. Mathematics and Computers in Simulation, 118: 102–115, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414003310.

Boularouk:2015:NAA

[1097] Y. Boularouk and K. Djeddour. New approximation for ARMA parameters estimate. Mathematics and Computers in Simulation, 118:116– 122, December 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415000051.

Bouyachcher:2015:CFM

[1098] Fadoua Bouyachcher, Bouchta Bouali, and M’barek Nasri. Clifford Fourier– Mellin moments and their invariants for color image recognition. Mathematics and Computers in Simulation, 118:123–129, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414003048.

Buscaglia:2015:MCCa

[1099] Gustavo C. Buscaglia, Mohamed El Alaoui Talibi, and Mohammed Jai. Mass-conserving cavitation model for dynamical lubrication problems. Part I: Mathematical analysis. Mathematics and Computers in Simulation, 118: 130–145, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414003279.

Buscaglia:2015:MCCb

[1100] Gustavo C. Buscaglia, Mohamed El Alaoui Talibi, and Mohammed Jai. Mass-conserving cavitation model for dynamical lubrication problems. Part II: Numerical analysis. Mathematics and Computers in Simulation, 118: 146–162, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414003243.

Chau:2015:NTC

[1101] Oanh Chau and Rachid Oujja. Numerical treatment of a class of thermal contact problems. Mathematics and Computers in Simulation, 118: 163–176, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414003346.

Duque:2015:AMM

[1102] José C. M. Duque, Rui M. P. Almeida, and Stanislav N. Antontsev. Application of the moving mesh method to the porous medium equation with variable exponent. Mathematics and Computers in Simulation, 118:177– 185, December 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414003267.

Eymard:2015:MFE

[1103] Robert Eymard, Thierry Gallouët, and Raphaèle Herbin. RT_{k} mixed finite elements for some nonlinear problems. Mathematics and Computers in Simulation, 118:186–197, December 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414003127.

Fortes:2015:FHS

[1104] M. A. Fortes, P. González, A. Palomares, and M. Pasadas. Filling holes with shape preserving conditions. Mathematics and Computers in Simulation, 118:198–212, December 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414003371.

Galiano:2015:RNF

[1105] G. Galiano and J. Velasco. Rearranged nonlocal filters for signal denoising. Mathematics and Computers in Simulation, 118:213–223, December 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541400319X.

Ginting:2015:MPM

[1106] V. Ginting, F. Pereira, and A. Rahunanthan. Multi-physics Markov chain Monte Carlo methods for subsurface flows. Mathematics and Computers in Simulation, 118:224–238, December 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414003231.

Gomez-Lopera:2015:EIS

[1107] J. F. Gómez-Lopera, J. Martínez-Aroza, M. A. Rodríguez-Valverde, M. A. Cabrerizo-Vílchez, and F. J. Montes-Ruíz-Cabello. Entropic image segmentation of sessile drops over patterned acetate. Mathematics and Computers in Simulation, 118: 239–247, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414003061.

Gonzalez:2015:DSW

[1108] P. González, M. J. Ibáñez, A. M. Roldán, and J. B. Roldán. An in-depth study on WENO-based techniques to improve parameter extraction procedures in MOSFET transistors. Mathematics and Computers in Simulation, 118:248–257, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414003176.

Jimenez-Gamero:2015:AND

[1109] M. D. Jiménez-Gamero, M. V. Alba-Fernández, P. Jodrá, and I. Barranco-Chamorro. An approximation to the null distribution of a class of Cramér– von Mises statistics. Mathematics and Computers in Simulation, 118: 258–272, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414003103.

Kheriji:2015:NLS

[1110] Walid Kheriji, Roland Masson, and Arthur Moncorgé. Nearwell local space and time refinement in reservoir simulation. Mathematics and Computers in Simulation, 118:273– 292, December 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541400322X.

Martin:2015:CSI

[1111] Miguel Angel Martín, Miguel Reyes, and Francisco Javier Taguas. Computer simulation of the interplay between fractal structures and surrounding heterogeneous multifractal distributions. Applications. Mathematics and Computers in Simulation, 118: 293–301, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414003164.

Paoli:2015:EAV

[1112] Laetitia Paoli. Existence and approximation for vibro-impact problems with a time-dependent set of constraints. Mathematics and Computers in Simulation, 118:302–309, December 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414003140.

Reinoso:2015:NIA

[1113] J. F. Reinoso, M. Moncayo, and F. J. Ariza-López. A new iterative algorithm for creating a mean 3D axis of a road from a set of GNSS traces. Mathematics and Computers in Simulation, 118:310–319, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414003309.

Roman-Galvez:2015:SHM

[1114] R. Román-Gálvez, R. Román-Roldán, J. Martínez-Aroza, and J. F. Gómez-Lopera. Semi-hidden Markov models for generation and analysis of sequences. Mathematics and Computers in Simulation, 118:320–328, December 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414003085.

Serghini:2015:CUS

[1115] A. Serghini and A. Tijini. Construction of univariate spline quasi-interpolants with symmetric functions. Mathematics and Computers in Simulation, 118: 329–342, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414003358.

Serghini:2015:TSQ

[1116] A. Serghini and A. Tijini. Trivariate spline quasi-interpolants based on simplex splines and polar forms. Mathematics and Computers in Simulation, 118:343–359, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414003073.

Shmarev:2015:IDA

[1117] Sergey Shmarev, Viktor Vdovin, and Alexey Vlasov. Interfaces in diffusion– absorption processes in nonhomogeneous media. Mathematics and Computers in Simulation, 118:360– 378, December 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414003036.

Anonymous:2015:NIl

[1118] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 118:379, December 2015. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001688.

Anonymous:2015:ICEl

[1119] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 118:380–381, December 2015. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541500169X.

Anonymous:2016:PJa

[1120] Anonymous. Pages 1–184 (January 2016). Mathematics and Computers in Simulation, 119:1–184, January 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2016:EBa

[1121] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 119:ifc, January 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002001.

Ebrahimnejad:2016:IDE

[1122] Ali Ebrahimnejad, Madjid Tavana, and Francisco J. Santos-Arteaga. An integrated data envelopment analysis and simulation method for group consensus ranking. Mathematics and Computers in Simulation, 119: 1–17, January 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001871.

Yuan:2016:SMC

[1123] Xiaohui Yuan, Zhihuan Chen, Yanbin Yuan, Yuehua Huang, Xianshan Li, and Wenwu Li. Sliding mode controller of hydraulic generator regulating system based on the input/output feedback linearization method. Mathematics and Computers in Simulation, 119:18–34, January 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001858.

Lee:2016:CSC

[1124] Dongsun Lee and Junseok Kim. Comparison study of the conservative Allen–Cahn and the Cahn– Hilliard equations. Mathematics and Computers in Simulation, 119:35– 56, January 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001834.

Cordero:2016:DTM

[1125] Alicia Cordero, Alfredo Ferrero, and Juan R. Torregrosa. Damped Traub’s method: Convergence and stability. Mathematics and Computers in Simulation, 119:57–68, January 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001779.

Sharifi:2016:NCO

[1126] Somayeh Sharifi, Mehdi Salimi, Stefan Siegmund, and Taher Lotfi. A new class of optimal four-point methods with convergence order 16 for solving nonlinear equations. Mathematics and Computers in Simulation, 119:69–90, January 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001767.

Ma:2016:ITS

[1127] Xiangmin Ma, Yuanfu Shao, Zhen Wang, Mengzhuo Luo, Xianjia Fang, and Zhixiang Ju. An impulsive two-stage predator–prey model with stage-structure and square root functional responses. Mathematics and Computers in Simulation, 119:91– 107, January 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415001743.

dePinho:2016:MVU

[1128] Frank M. de Pinho, Glaura C. Franco, and Ralph S. Silva. Modeling volatility using state space models with heavy tailed distributions. Mathematics and Computers in Simulation, 119:108–127, January 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415001573.

Kulikova:2016:USR

[1129] M. V. Kulikova and J. V. Tsyganova. A unified square-root approach for the score and Fisher information matrix computation in linear dynamic systems. Mathematics and Computers in Simulation, 119:128– 141, January 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415001561.

Clempner:2016:SPF

[1130] Julio B. Clempner and Alexander S. Poznyak. Solving the Pareto front for multiobjective Markov chains using the minimum Euclidean distance gradientbased optimization method. Mathematics and Computers in Simulation, 119:142–160, January 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541500155X.

Magrenan:2016:NIC

[1131] Á. Alberto Magreñán and Ioannis K.

Argyros. New improved convergence analysis for the secant method. Mathematics and Computers in Simulation, 119:161–170, January 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001524.

Duda:2016:LFN

[1132] Jozef Duda. A Lyapunov functional for a neutral system with a distributed time delay. Mathematics and Computers in Simulation, 119: 171–181, January 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001512.

Anonymous:2016:NIa

[1133] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 119:182, January 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002050.

Anonymous:2016:ICEa

[1134] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 119:183–184, January 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415002062.

Anonymous:2016:PF

[1135] Anonymous. Pages 1–132 (February 2016). Mathematics and Computers in Simulation, 120:1–132, February 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2016:EBb

[1136] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 120:ifc, February 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415002244.

Llibre:2016:LCB

[1137] Jaume Llibre and Chara Pantazi. Limit cycles bifurcating from a degenerate center. Mathematics and Computers in Simulation, 120:1– 11, February 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415001147.

Izem:2016:DGM

[1138] Nouh Izem, Mohammed Seaid, and Mohamed Wakrim. A discontinuous Galerkin method for two-layer shallow water equations. Mathematics and Computers in Simulation, 120:12–23, February 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415001196.

Farid:2016:ADG

[1139] Mahboubeh Farid. Accelerated diagonal gradient-type method for large-scale unconstrained optimization. Mathematics and Computers in Simulation, 120:24–30, February 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001172.

Martin:2016:WTP

[1140] Nirian Martin, Raquel Mata, and Leandro Pardo. Wald type and phidivergence based test-statistics for isotonic binomial proportions. Mathematics and Computers in Simulation, 120:31–49, February 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001275.

Hadrich:2016:BEM

[1141] Atizez Hadrich, Mourad Zribi, and Afif Masmoudi. Bayesian expectation maximization algorithm by using B-splines functions: Application in image segmentation. Mathematics and Computers in Simulation, 120:50–63, February 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001263.

Dey:2016:SIG

[1142] Sanku Dey, Tanujit Dey, and Daniel J.

Luckett. Statistical inference for the generalized inverted exponential distribution based on upper record values. Mathematics and Computers in Simulation, 120:64– 78, February 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415001317.

Wu:2016:CTA

[1143] Shu-Fei Wu and Ying-Po Lin. Computational testing algorithmic procedure of assessment for lifetime performance index of products with one-parameter exponential distribution under progressive type I interval censoring. Mathematics and Computers in Simulation, 120:79–90, February 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415001329.

Trucios:2016:BPU

[1144] Carlos Trucíos and Luiz K. Hotta. Bootstrap prediction in univariate volatility models with leverage effect. Mathematics and Computers in Simulation, 120:91–103, February 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415001330.

Zhang:2016:CSP

[1145] Yunong Zhang, Keke Zhai, Dechao Chen, Long Jin, and Chaowei Hu. Challenging simulation practice (failure and success) on implicit tracking control of double-integrator system via Zhang-gradient method. Mathematics and Computers in Simulation, 120:104–119, February 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415001342.

Dengiz:2016:MMB

[1146] Berna Dengiz, Yusuf Tansel İç, and Onder Belgin. A meta-model based simulation optimization using hybrid simulation-analytical modeling to increase the productivity in automotive industry. Mathematics and Computers in Simulation, 120:120– 128, February 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001494.

Anonymous:2016:NIb

[1147] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 120:129, February 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002293.

Anonymous:2016:ICEb

[1148] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 120:131–132, February 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541500230X.

Anonymous:2016:PMa

[1149] Anonymous. Pages 1–182 (March 2016). Mathematics and Computers in Simulation, 121:1–182, March 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2016:EBc

[1150] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 121:ifc, March 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002402.

Radulescu:2016:OCA

[1151] I. R. Rădulescu, D. Cândea, and A. Halanay. Optimal control analysis of a leukemia model under imatinib treatment. Mathematics and Computers in Simulation, 121: 1–11, March 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001135.

Kundu:2016:SBS

[1152] Amitava Kundu, Pritha Das, and A. B. Roy. Stability, bifurcations and synchronization in a delayed neural network model of nidentical neurons. Mathematics and Computers in Simulation, 121:12– 33, March 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001500.

Sarkar:2016:MSR

[1153] M. Sarkar, S. Nandy, S. R. K. Vadali, S. Roy, and S. N. Shome. Modelling and simulation of a robust energy efficient AUV controller. Mathematics and Computers in Simulation, 121:34–47, March 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541500186X.

Arenas:2016:CNF

[1154] Abraham J. Arenas, Gilberto González-Parra, and Benito M. Chen-Charpentier. Construction of nonstandard finite difference schemes for the SI and SIR epidemic models of fractional order. Mathematics and Computers in Simulation, 121:48–63, March 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001883.

Belinha:2016:ABR

[1155] J. Belinha, L. M. J. S. Dinis, and R. M. Natal Jorge. The analysis of the bone remodelling around femoral stems: a meshless approach. Mathematics and Computers in Simulation, 121:64–94, March 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001895.

Mezzadri:2016:CTS

[1156] Francesco Mezzadri and Emanuele Galligani. A Chebyshev technique for the solution of optimal control problems with nonlinear programming methods. Mathematics and Computers in Simulation, 121:95– 108, March 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001913.

Rivero-Esquivel:2016:SBA

[1157] Erika Rivero-Esquivel, Eric Avila-Vales, and Gerardo García-Almeida. Stability and bifurcation analysis of a SIR model with saturated incidence rate and saturated treatment. Mathematics and Computers in Simulation, 121:109–132, March 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415001937.

Hernandez-Martinez:2016:GFA

[1158] Eliseo Hernandez-Martinez, Francisco Valdés-Parada, Jose Alvarez-Ramirez, Hector Puebla, and Epifanio Morales-Zarate. A Green’s function approach for the numerical solution of a class of fractional reaction– diffusion equations. Mathematics and Computers in Simulation, 121:133– 145, March 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415001925.

Maire:2016:SFD

[1159] Sylvain Maire and Giang Nguyen. Stochastic finite differences for elliptic diffusion equations in stratified domains. Mathematics and Computers in Simulation, 121:146– 165, March 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415001962.

Gerdjikov:2016:ABM

[1160] V. S. Gerdjikov, M. D. Todorov, and A. V. Kyuldjiev. Asymptotic behavior of Manakov solitons: Effects of potential wells and humps. Mathematics and Computers in Simulation, 121:166–178, March 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002189.

Anonymous:2016:NIc

[1161] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 121:179, March 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002451.

Anonymous:2016:ICEc

[1162] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 121:181–182, March 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475415002463.

Anonymous:2016:PAa

[1163] Anonymous. Pages 1–84 (April 2016).

Mathematics and Computers in Simulation, 122:1–84, April 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2016:EBd

[1164] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 122:ifc, April 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002621.

Bonin:2016:FAR

[1165] Thomas Bonin, Heike Faßbender, Andreas Soppa, and Michael Zaeh. A fully adaptive rational global Arnoldi method for the model-order reduction of second-order MIMO systems with proportional damping. Mathematics and Computers in Simulation, 122:1–19, April 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001822.

Hosseinipour-Mahani:2016:NOT

[1166] N. Hosseinipour-Mahani and A. Malek. A neurodynamic optimization technique based on overestimator and underestimator functions for solving a class of non-convex optimization problems. Mathematics and Computers in Simulation, 122:20– 34, April 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541500213X.

Jiang:2016:GHB

[1167] Zhichao Jiang, Wanbiao Ma, and Junjie Wei. Global Hopf bifurcation and permanence of a delayed SEIRS epidemic model. Mathematics and Computers in Simulation, 122: 35–54, April 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415002517.

David:2016:PCS

[1168] S. A. David, J. A. T. Machado, D. D. Quintino, and J. M. Balthazar. Partial chaos suppression in a fractional order macroeconomic model. Mathematics and Computers in Simulation, 122:55–68, April 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002542.

Kaur:2016:MVO

[1169] Gurbinder Kaur, Joydip Dhar, and Rangan Kumar Guha. Minimal variability OWA operator combining ANFIS and fuzzy c-means for forecasting BSE index. Mathematics and Computers in Simulation, 122: 69–80, April 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415002530.

Anonymous:2016:NId

[1170] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 122:81, April 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002670.

Anonymous:2016:ICEd

[1171] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 122:83–84, April 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475415002682.

Anonymous:2016:PMb

[1172] Anonymous. Pages 1–82 (May 2016).

Mathematics and Computers in Simulation, 123:1–82, May 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2016:EBe

[1173] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 123:ifc, May 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416000185.

Klimsiak:2016:VAO

[1174] Tomasz Klimsiak, Andrzej Rozkosz, and Bartosz Ziemkiewicz. Valuing American options by simulation: a BSDEs approach. Mathematics and Computers in Simulation, 123:1–18, May 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415002724.

Alvarez:2016:PTM

[1175] J. Álvarez and A. Durán. Petviashvili type methods for traveling wave computations: II. Acceleration with vector extrapolation methods. Mathematics and Computers in Simulation, 123:19–36, May 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415002785.

Wu:2016:MBM

[1176] Qinghua Wu and Guozheng Yan. Mathematical basis for a mixed inverse scattering problem. Mathematics and Computers in Simulation, 123:37–52, May 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002773.

Wang:2016:GSP

[1177] Xiying Wang, Xinzhi Liu, Wei chau Xie, Wei Xu, and Yong Xu. Global stability and persistence of HIV models with switching parameters and pulse control. Mathematics and Computers in Simulation, 123: 53–67, May 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416000021.

Langemann:2016:CMM

[1178] Dirk Langemann, Igor Nesteruk, and Jürgen Prestin. Comparison of mathematical models for the dynamics of the Chernivtsi children disease. Mathematics and Computers in Simulation, 123:68–79, May 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416000069.

Anonymous:2016:NIe

[1179] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 123:80, May 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416000239.

Anonymous:2016:ICEe

[1180] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 123:81–82, May 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475416000240.

Anonymous:2016:PJb

[1181] Anonymous. Pages 1–106 (June 2016). Mathematics and Computers in Simulation, 124:1–106, June 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2016:EBf

[1182] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 124:ifc, June 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416000434.

Liu:2016:MMC

[1183] Zijian Liu and Chenxue Yang. A mathematical model of cancer treatment by radiotherapy followed by chemotherapy. Mathematics and Computers in Simulation, 124:1– 15, June 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415002761.

Boonrangsiman:2016:BPC

[1184] S. Boonrangsiman, K. Bunwong, and Elvin J. Moore. A bifurcation path to chaos in a timedelay fisheries predator–prey model with prey consumption by immature and mature predators. Mathematics and Computers in Simulation, 124:16–29, June 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416000033.

Riou:2016:SRA

[1185] Hervé Riou. On some results about the variational theory of complex rays used close to the high frequency regime. Mathematics and Computers in Simulation, 124:30–43, June 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416000057.

Xu:2016:GDT

[1186] Zhiting Xu, Liangcheng Qu, and Yehui Huang. Global dynamics of a twostrain flu model with delay. Mathematics and Computers in Simulation, 124:44–59, June 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416000070.

Jones:2016:MFD

[1187] Don A. Jones. Modified finite difference schemes for geophysical flows. Mathematics and Computers in Simulation, 124:60–68, June 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416000112.

Bozagac:2016:DSM

[1188] Doruk Bozağaç, İnci Batmaz, and Halit Oğuztüzün. Dynamic simulation metamodeling using MARS: a case of radar simulation. Mathematics and Computers in Simulation, 124:69–86, June 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416000100.

Wang:2016:AHM

[1189] Xia Wang, Xinyu Song, Sanyi Tang, and Libin Rong. Analysis of HIV models with multiple target cell populations and general nonlinear rates of viral infection and cell death. Mathematics and Computers in Simulation, 124:87–103, June 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416000148.

Anonymous:2016:NIf

[1190] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 124:104, June 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416000483.

Anonymous:2016:ICEf

[1191] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 124:105–106, June 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475416000495.

Anonymous:2016:EBg

[1192] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 125:ifc, July 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300088.

DelBuono:2016:SGE

[1193] Nicoletta Del Buono, Roberto Garrappa, and Giulia Spaletta. SDS2014 Guest Editorial. Mathematics and Computers in Simulation, 125:1– 2, July 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300015.

Schonlein:2016:CEL

[1194] Michael Schönlein and Uwe Helmke.

Controllability of ensembles of linear dynamical systems. Mathematics and Computers in Simulation, 125:3–14, July 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415002207.

Ironi:2016:NTM

[1195] Liliana Ironi and Diana X. Tran. Nonlinear and temporal multiscale dynamics of gene regulatory networks: a qualitative simulator. Mathematics and Computers in Simulation, 125: 15–37, July 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415002578.

Berardi:2016:NSR

[1196] Marco Berardi and Michele Vurro.

The numerical solution of Richards’ equation by means of method of lines and ensemble Kalman filter. Mathematics and Computers in Simulation, 125:38–47, July 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001846.

Capobianco:2016:APR

[1197] M. R. Capobianco and G. Criscuolo. On the approximation by product rules of weakly singular double integrals over the square. Mathematics and Computers in Simulation, 125:48–55, July 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002128.

Vermiglio:2016:NAN

[1198] Rossana Vermiglio. Numerical approximation of the non-essential spectrum of abstract delay differential equations. Mathematics and Computers in Simulation, 125:56–69, July 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002335.

Acciani:2016:IAC

[1199] Giuseppe Acciani, Filomena Di Modugno, Angelamaria Abrescia, and Giuseppe C. Marano. Integration algorithm for covariance nonstationary dynamic analysis using equivalent stochastic linearization. Mathematics and Computers in Simulation, 125:70–82, July 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002566.

Orlando:2016:DMM

[1200] Giuseppe Orlando. A discrete mathematical model for chaotic dynamics in economics: Kaldor’s model on business cycle. Mathematics and Computers in Simulation, 125: 83–98, July 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416000045.

Vandana:2016:EMR

[1201] Vandana and B. K. Sharma. An EOQ model for retailers partial permissible delay in payment linked to order quantity with shortages. Mathematics and Computers in Simulation, 125:99–112, July 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541500258X.

Botana:2016:FSI

[1202] Francisco Botana, Antonio Hernando, Eugenio Roanes-Lozano, and Michael J. Wester. Foreword to the special issue on “Nonstandard Applications of Computer Algebra (ACA’2013)”. Mathematics and Computers in Simulation, 125:113– 114, July 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300027.

Botana:2016:SIA

[1203] Francisco Botana and Tomas Recio.

Some issues on the automatic computation of plane envelopes in interactive environments. Mathematics and Computers in Simulation, 125:115–125, July 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414001529.

Ceballos:2016:AMO

[1204] Manuel Ceballos, Juan Núñez, and Ángel F. Tenorio. Algorithmic method to obtain combinatorial structures associated with Leibniz algebras. Mathematics and Computers in Simulation, 125:126–138, July 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414002845.

Falcon:2016:CEF

[1205] Raúl Falcón, Eva Barrena, David Canca, and Gilbert Laporte. Counting and enumerating feasible rotating schedules by means of Gröbner bases. Mathematics and Computers in Simulation, 125:139–151, July 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475414003292.

Lewis:2016:FBL

[1206] Robert H. Lewis and Evangelos A. Coutsias. Flexibility of Bricard’s linkages and other structures via resultants and computer algebra. Mathematics and Computers in Simulation, 125:152–167, July 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414002857.

Gogin:2016:PLS

[1207] N. Gogin and A. Mylläri. Padovan-like sequences and Bell polynomials. Mathematics and Computers in Simulation, 125:168–177, July 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001731.

Anonymous:2016:NIg

[1208] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 125:178, July 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300155.

Anonymous:2016:ICEg

[1209] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 125:179–180, July 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475416300167.

Anonymous:2016:PAb

[1210] Anonymous. Pages 1–126 (August 2016). Mathematics and Computers in Simulation, 126:1–126, August 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2016:EBh

[1211] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 126:ifc, August 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300398.

Kolar:2016:CAC

[1212] Miroslav Kolář, Michal Beneš, and Daniel Ševčovič. Computational analysis of the conserved curvature driven flow for open curves in the plane. Mathematics and Computers in Simulation, 126:1–13, August 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416000318.

Venturino:2016:MMS

[1213] Ezio Venturino. Metaecoepidemic models with sound and infected prey migration. Mathematics and Computers in Simulation, 126:14– 44, August 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416000550.

Crnjaric-Zic:2016:NAS

[1214] Nelida Črnjarić-Žic and Nermina Mujaković. Numerical analysis of the solutions for 1d compressible viscous micropolar fluid flow with different boundary conditions. Mathematics and Computers in Simulation, 126:45–62, August 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416000574.

Thompson:2016:MDF

[1215] James R. Thompson and James R. Wilson. Multifractal detrended fluctuation analysis: Practical applications to financial time series. Mathematics and Computers in Simulation, 126:63–88, August 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416000598.

Liu:2016:TLM

[1216] Qingfang Liu, Yanren Hou, Zhiheng Wang, and Jiakun Zhao. Two-level methods for the Cahn– Hilliard equation. Mathematics and Computers in Simulation, 126:89– 103, August 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416000604.

Devloo:2016:TDH

[1217] Philippe R. B. Devloo, Agnaldo M.

Farias, Sônia M. Gomes, and Denise de Siqueira. Two-dimensional hp adaptive finite element spaces for mixed formulations. Mathematics and Computers in Simulation, 126: 104–122, August 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541630026X.

Anonymous:2016:NIh

[1218] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 126:123, August 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541630043X.

Anonymous:2016:ICEh

[1219] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 126:124–125, August 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475416300441.

Anonymous:2016:SII

[1220] Anonymous. Seventh IMACS International Conference on Nonlinear Evolution Equations and Wave Phenomena: Computation and Theory. Mathematics and Computers in Simulation, 127:??, September 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2016:EBi

[1221] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 127:ifc, September 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416300532.

Taha:2016:F

[1222] Thiab R. Taha. Foreword. Mathematics and Computers in Simulation, 127:1, September 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416300490.

Christov:2016:ATW

[1223] Ivan C. Christov, P. M. Jordan, S. A. Chin-Bing, and A. Warn-Varnas. Acoustic traveling waves in thermoviscous perfect gases: Kinks, acceleration waves, and shocks under the Taylor–Lighthill balance. Mathematics and Computers in Simulation, 127:2–18, September 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541300075X.

Christov:2016:PSM

[1224] C. I. Christov. On the pseudolocalized solutions in multi-dimension of Boussinesq equation. Mathematics and Computers in Simulation, 127: 19–27, September 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412001681.

Colin:2016:NSW

[1225] M. Colin, T. Colin, and J. Dambrine.

Numerical simulations of wormlike micelles flows in micro-fluidic Tshaped junctions. Mathematics and Computers in Simulation, 127:28– 55, September 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414000433.

Dougalis:2016:NAS

[1226] V. A. Dougalis, A. Durán, and D. E.

Mitsotakis. Numerical approximation of solitary waves of the Benjamin equation. Mathematics and Computers in Simulation, 127:56– 79, September 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001711.

Dutykh:2016:DDB

[1227] Denys Dutykh and Olivier Goubet. Derivation of dissipative Boussinesq equations using the Dirichlet-to-Neumann operator approach. Mathematics and Computers in Simulation, 127:80–93, September 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000706.

Gentile:2016:HDC

[1228] M. Gentile and B. Straughan. Hyperbolic diffusion with Christov– Morro theory. Mathematics and Computers in Simulation, 127:94– 100, September 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412001735.

Hederi:2016:EET

[1229] M. Hederi, A. L. Islas, K. Reger, and C. M. Schober. Efficiency of exponential time differencing schemes for nonlinear Schrödinger equations. Mathematics and Computers in Simulation, 127:101–113, September 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413001572.

Joumaa:2016:DWM

[1230] Hady Joumaa and Martin Ostoja-Starzewski. On the dilatational wave motion in anisotropic fractal solids. Mathematics and Computers in Simulation, 127:114–130, September 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475413000876.

Kaup:2016:RSO

[1231] D. J. Kaup and Robert A. Van Gorder. Resonant solitons from the 3 × 3 operator. Mathematics and Computers in Simulation, 127:131– 150, September 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475412002832.

Li:2016:ILM

[1232] K. Li, P. G. Kevrekidis, H. Susanto, and V. Rothos. Intrinsic localized modes in coupled DNLS equations from the anti-continuum limit. Mathematics and Computers in Simulation, 127: 151–165, September 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475412002807.

Ma:2016:NBI

[1233] Wen-Xiu Ma, Jinghan Meng, and Mengshu Zhang. Nonlinear biintegrable couplings with Hamiltonian structures. Mathematics and Computers in Simulation, 127:166– 177, September 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541400069X.

Margolin:2016:SSF

[1234] L. G. Margolin. A strain space framework for numerical hyperplasticity. Mathematics and Computers in Simulation, 127:178–188, September 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475412001693.

Mickens:2016:NSA

[1235] Ronald E. Mickens. NSFD scheme for acoustic propagation with the linearized Euler equations. Mathematics and Computers in Simulation, 127: 189–193, September 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475414001955.

Mladenov:2016:GRL

[1236] Ivaïlo M. Mladenov and John Oprea. On the geometry of the rotating liquid drop. Mathematics and Computers in Simulation, 127:194–202, September 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414000883.

Moshkin:2016:PHR

[1237] N. P. Moshkin, G. G. Chernykh, and Kridsada Narong. On the performance of high resolution non-oscillating advection schemes in the context of the flow generated by a mixed region in a stratified fluid. Mathematics and Computers in Simulation, 127: 203–219, September 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412002844.

Peets:2016:ISD

[1238] Tanel Peets. Internal scales and dispersive properties of microstructured materials. Mathematics and Computers in Simulation, 127:220– 228, September 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414000561.

Porubov:2016:DKE

[1239] A. V. Porubov. Description of kink evolution by means of particular analytical solutions. Mathematics and Computers in Simulation, 127:229– 235, September 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414000445.

Rasmussen:2016:CWH

[1240] Anders Rønne Rasmussen, Mads Peter Sørensen, Yuri B. Gaididei, and Peter Leth Christiansen. Compound waves in a higher order nonlinear model of thermoviscous fluids. Mathematics and Computers in Simulation, 127: 236–251, September 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414000573.

Salupere:2016:HSK

[1241] Andrus Salupere. On hidden solitons in KdV related systems. Mathematics and Computers in Simulation, 127: 252–262, September 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541400202X.

Swierczewski:2016:CRT

[1242] Christopher Swierczewski and Bernard Deconinck. Computing Riemann theta functions in Sage with applications. Mathematics and Computers in Simulation, 127:263–272, September 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475413000888.

Todorov:2016:EEP

[1243] M. D. Todorov. The effect of the elliptic polarization on the quasiparticle dynamics of linearly coupled systems of Nonlinear Schrödinger Equations. Mathematics and Computers in Simulation, 127:273–286, September 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414001657.

Zeng:2016:SCC

[1244] Jianhua Zeng and Boris A. Malomed. Suppressing the critical collapse of solitons by one-dimensional quintic nonlinear lattices. Mathematics and Computers in Simulation, 127: 287–296, September 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475412002017.

Anonymous:2016:NIi

[1245] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 127:297, September 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541630060X.

Anonymous:2016:ICEi

[1246] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 127:298–299, September 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300611.

Anonymous:2016:PO

[1247] Anonymous. Pages 1–58 (October 2016). Mathematics and Computers in Simulation, 128:1–58, October 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2016:EBj

[1248] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 128:ifc, October 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300684.

Yu:2016:NSL

[1249] C. H. Yu and Tony W. H. Sheu. Numerical study of long-time Camassa– Holm solution behavior for soliton transport. Mathematics and Computers in Simulation, 128:1– 12, October 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300246.

deHierro:2016:FRA

[1250] Antonio Francisco Roldán López de Hierro, Juan Martínez-Moreno, Concepción Aguilar Peña, and Concepción Roldán López de Hierro. A fuzzy regression approach using Bernstein polynomials for the spreads: Computational aspects and applications to economic models. Mathematics and Computers in Simulation, 128: 13–25, October 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300295.

Hu:2016:SHB

[1251] Qing Hu, Zhixing Hu, and Fucheng Liao. Stability and Hopf bifurcation in a HIV-1 infection model with delays and logistic growth. Mathematics and Computers in Simulation, 128:26–41, October 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300325.

Lambert:2016:GSA

[1252] Romain S. C. Lambert, Frank Lemke, Sergei S. Kucherenko, Shufang Song, and Nilay Shah. Global sensitivity analysis using sparse high dimensional model representations generated by the group method of data handling. Mathematics and Computers in Simulation, 128:42–54, October 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300349.

Anonymous:2016:NIj

[1253] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 128:55, October 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300738.

Anonymous:2016:ICEj

[1254] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 128:56–57, October 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541630074X.

Anonymous:2016:PN

[1255] Anonymous. Pages 1–98 (November 2016). Mathematics and Computers in Simulation, 129:1–98, November 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2016:EBk

[1256] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 129:ifc, November 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300994.

Gine:2016:CCN

[1257] Jaume Giné. Center conditions for nilpotent cubic systems using the Cherkas method. Mathematics and Computers in Simulation, 129:1–9, November 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300313.

Maciel:2016:DEA

[1258] Leandro Maciel, Fernando Gomide, and Rosangela Ballini. A differential evolution algorithm for yield curve estimation. Mathematics and Computers in Simulation, 129:10– 30, November 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300337.

Hernando:2016:NLI

[1259] Antonio Hernando, Roberto Maestre-Martínez, and Eugenio Roanes-Lozano. A natural language for implementing algebraically Expert Systems. Mathematics and Computers in Simulation, 129:31–49, November 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300350.

Chaskalovic:2016:DMP

[1260] Joël Chaskalovic and Franck Assous. Data mining and probabilistic models for error estimate analysis of finite element method. Mathematics and Computers in Simulation, 129: 50–68, November 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300374.

Zhao:2016:DSH

[1261] Jing jun Zhao, Song shu Liu, and Tao Liu. Determining surface heat flux for noncharacteristic Cauchy problem for Laplace equation. Mathematics and Computers in Simulation, 129:69–80, November 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300647.

Maestre-Martinez:2016:AAD

[1262] Roberto Maestre-Martínez, Antonio Hernando, and Eugenio Roanes-Lozano. An algebraic approach for detecting nearly dangerous situations in expert systems. Mathematics and Computers in Simulation, 129: 81–93, November 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300301.

Anonymous:2016:NIk

[1263] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 129:95, November 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301045.

Anonymous:2016:ICEk

[1264] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 129:96–97, November 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416301057.

ElAchkar:2016:POD

[1265] Maria El Achkar, Rita Mbayed, Georges Salloum, Nicolas Patin, Sandrine Le Ballois, and Eric Monmasson. Power operating domain of a cascaded doubly fed induction machine. Mathematics and Computers in Simulation, 130:142–154, December 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415000105.

Pankovits:2016:MCF

[1266] Petronela Pankovits, Dhaker Abbes, Christophe Saudemont, Stephane Brisset, Julien Pouget, and Benoit Robyns. Multi-criteria fuzzy-logic optimized supervision for hybrid railway power substations. Mathematics and Computers in Simulation, 130:236–250, December 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475416300799.

Anonymous:2016:EBl

[1267] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 130:ifc, December 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541630129X.

Blasco-Gimenez:2016:P

[1268] Ramon Blasco-Gimenez. Preface. Mathematics and Computers in Simulation, 130:1–2, December 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475416301136.

Ouagued:2016:CHA

[1269] Sofiane Ouagued, Yacine Amara, and Georges Barakat. Comparison of hybrid analytical modelling and reluctance network modelling for predesign purposes. Mathematics and Computers in Simulation, 130:3– 21, December 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300660.

Sergeant:2016:ISS

[1270] Peter Sergeant, Hendrik Vansompel, and Luc Dupré. Influence of stator slot openings on losses and torque in axial flux permanent magnet machines. Mathematics and Computers in Simulation, 130:22– 31, December 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416000586.

Khlissa:2016:TMA

[1271] Radhouane Khlissa, Stephane Vivier, Guy Friedrich, Khadija El Kadri Benkara, and Bassel Assaad. Thermal modeling of an asymmetrical totally enclosed permanent magnet integrated starter generator. Mathematics and Computers in Simulation, 130:32–47, December 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416000094.

Arumugam:2016:MAE

[1272] Puvan Arumugam, Jiri Dusek, Smail Mezani, Tahar Hamiti, and Chris Gerada. Modeling and analysis of eddy current losses in permanent magnet machines with multi-stranded bundle conductors. Mathematics and Computers in Simulation, 130:48– 56, December 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415002347.

Dumont:2016:YRE

[1273] C. Dumont, V. Kluyskens, and B. Dehez. Yokeless radial electrodynamic bearing. Mathematics and Computers in Simulation, 130:57– 69, December 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415002177.

Hannon:2016:TTC

[1274] Bert Hannon, Peter Sergeant, and Luc Dupré. Torque and torque components in high-speed permanentmagnet synchronous machines with a shielding cylinder. Mathematics and Computers in Simulation, 130: 70–80, December 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415001949.

Geissler:2016:SSM

[1275] D. Geissler, C. Geissler, and T. Leibfried. Simplified simulation model of continuously transposed cable for linear and nonlinear buckling analysis. Mathematics and Computers in Simulation, 130:81–94, December 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541500124X.

Riedemann:2016:MSO

[1276] Javier Riedemann, Iván Andrade, Rubén Peña, Ramón Blasco-Gimenez, Jon Clare, Pedro Melín, and Marco Rivera. Modulation strategies for an open-end winding induction machine fed by a two-output indirect matrix converter. Mathematics and Computers in Simulation, 130:95– 111, December 2016. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300842.

Dufour:2016:IMP

[1277] Christian Dufour, Sébastien Cense, and Jean Bélanger. An induction machine and power electronic test system on a field-programmable gate array. Mathematics and Computers in Simulation, 130:112–123, December 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300659.

Ouddah:2016:ICE

[1278] N. Ouddah, M. Boukhnifer, A. Chaibet, and E. Monmasson. IMACS calendar of events. Mathematics and Computers in Simulation, 130: 252–253, December 2016. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416301379.

Anonymous:2017:EBa

[1279] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 131:ifc, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416301562.

Anonymous:2017:P

[1280] Anonymous. Preface. Mathematics and Computers in Simulation, 131: 1–2, January 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475416301537.

Lesel:2017:RTE

[1281] J. Lesel, D. Bourdon, G. Claisse, P. Debay, and B. Robyns. Real time electrical power estimation for the energy management of automatic metro lines. Mathematics and Computers in Simulation, 131: 3–20, January 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475416300891.

Jaafar:2017:EMH

[1282] A. Jaafar, C. Turpin, X. Roboam, E. Bru, and O. Rallieres. Energy management of a hybrid system based on a fuel cell and a Lithium Ion battery: Experimental tests and integrated optimal design. Mathematics and Computers in Simulation, 131:21–37, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416000136.

Bastidas-Rodriguez:2017:GAI

[1283] J. D. Bastidas-Rodriguez, G. Petrone, C. A. Ramos-Paja, and G. Spagnuolo. A genetic algorithm for identifying the single diode model parameters of a photovoltaic panel. Mathematics and Computers in Simulation, 131:38–54, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002220.

Berkoune:2017:RCH

[1284] Karima Berkoune, Emna Ben Sedrine, Lionel Vido, and Sandrine Le Ballois. Robust control of hybrid excitation synchronous generator for wind applications. Mathematics and Computers in Simulation, 131:55– 75, January 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415002165.

Turpin:2017:MAO

[1285] C. Turpin, D. Van Laethem, B. Morin, O. Rallières, X. Roboam, O. Verdu, and V. Chaudron. Modelling and analysis of an original direct hybridization of fuel cells and ultracapacitors. Mathematics and Computers in Simulation, 131:76–87, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001780.

Leva:2017:AVH

[1286] S. Leva, A. Dolara, F. Grimaccia, M. Mussetta, and E. Ogliari. Analysis and validation of 24 hours ahead neural network forecasting of photovoltaic output power. Mathematics and Computers in Simulation, 131:88–100, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415001238.

Bastidas-Rodriguez:2017:QPM

[1287] J. D. Bastidas-Rodriguez, E. Franco, G. Petrone, C. A. Ramos-Paja, and G. Spagnuolo. Quantification of photovoltaic module degradation using model based indicators. Mathematics and Computers in Simulation, 131: 101–113, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541500066X.

Rigo-Mariani:2017:FPF

[1288] R. Rigo-Mariani, B. Sareni, and X. Roboam. Fast power flow scheduling and sensitivity analysis for sizing a microgrid with storage. Mathematics and Computers in Simulation, 131:114–127, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416000124.

Vidal:2017:MMD

[1289] Ricardo Vidal, Diego Soto, Iván Andrade, Javier Riedemann, Cristián Pesce, Enrique Belenguer, Ruben Pena, and Ramon Blasco-Gimenez. A multilevel modular DC–DC converter topology. Mathematics and Computers in Simulation, 131:128– 141, January 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415002712.

Ghzaiel:2017:GIE

[1290] Walid Ghzaiel, Manel Jebali-Ben Ghorbal, Ilhem Slama-Belkhodja, and Josep M. Guerrero. Grid impedance estimation based hybrid islanding detection method for AC microgrids. Mathematics and Computers in Simulation, 131:142–156, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002219.

Le:2017:AFR

[1291] Thi-Tinh-Minh Le and Nicolas Retiere.

Approximation of the frequency response of power systems based on scale invariance. Mathematics and Computers in Simulation, 131:157– 171, January 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001809.

Bouallaga:2017:MTE

[1292] Anouar Bouallaga, Arnaud Davigny, Vincent Courtecuisse, and Benoit Robyns. Methodology for technical and economic assessment of electric vehicles integration in distribution grid. Mathematics and Computers in Simulation, 131:172–189, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416300805.

Sanchez:2017:DFO

[1293] S. Sanchez, F. Richardeau, and D. Risaletto. Design and faultoperation analysis of a modular cyclic cascade inter-cell transformer (ICT) for parallel multicell converters. Mathematics and Computers in Simulation, 131:190–199, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416000082.

Salameh:2017:RMI

[1294] F. Salameh, A. Picot, M. Chabert, and P. Maussion. Regression methods for improved lifespan modeling of low voltage machine insulation. Mathematics and Computers in Simulation, 131:200–216, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415002372.

Sobanski:2017:FDC

[1295] P. Sobanski and T. Orlowska-Kowalska. Faults diagnosis and control in a low-cost fault-tolerant induction motor drive system. Mathematics and Computers in Simulation, 131:217– 233, January 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415002360.

BenYoussef:2017:OCF

[1296] Ahlem Ben Youssef, Sejir Khojet El Khil, and Ilhem Slama Belkhodja. Open-circuit fault diagnosis and voltage sensor fault tolerant control of a single phase pulsed width modulated rectifier. Mathematics and Computers in Simulation, 131:234– 252, January 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415002190.

Bourogaoui:2017:RTE

[1297] M. Bourogaoui, H. Ben Attia Sethom, and I. Slama Belkhodja. Real-time encoder faults detection and rotor position estimation for permanent magnet synchronous motor drives fault tolerant sensorless control using digital signal controller. Mathematics and Computers in Simulation, 131: 253–267, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001986.

Cousseau:2017:IEM

[1298] R. Cousseau, N. Patin, C. Forgez, E. Monmasson, and L. Idkhajine. Improved electrical model of aluminum electrolytic capacitor with anomalous diffusion for health monitoring. Mathematics and Computers in Simulation, 131:268–282, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001792.

Morando:2017:AMA

[1299] S. Morando, S. Jemei, D. Hissel, R. Gouriveau, and N. Zerhouni. ANOVA method applied to proton exchange membrane fuel cell ageing forecasting using an echo state network. Mathematics and Computers in Simulation, 131:283– 294, January 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415001287.

Nadal:2017:MBE

[1300] Clément Nadal, Christophe Giraud-Audine, Frédéric Giraud, Michel Amberg, and Betty Lemaire-Semail. Modelling of a beam excited by piezoelectric actuators in view of tactile applications. Mathematics and Computers in Simulation, 131:295– 315, January 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475415002359.

Sprangers:2017:FSA

[1301] R. L. J. Sprangers, J. J. H. Paulides, B. L. J. Gysen, and E. A. Lomonova. A fast semi-analytical model for the slotted structure of induction motors. Mathematics and Computers in Simulation, 131:316– 327, January 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475416300908.

Kuttler:2017:FIL

[1302] Sulivan Kuttler, Khadija El Kadri Benkara, Guy Friedrich, Abdenour Abdelli, and Franck Vangraefschepe. Fast iron losses model of stator taking into account the flux weakening mode for the optimal sizing of high speed permanent internal magnet synchronous machine. Mathematics and Computers in Simulation, 131:328–343, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300957.

Anonymous:2017:NIa

[1303] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 131:344, January 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541630163X.

Anonymous:2017:ICEa

[1304] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 131:345–346, January 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416301641.

Anonymous:2017:PF

[1305] Anonymous. Pages 1–310 (February 2017). Mathematics and Computers in Simulation, 132:1–310, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2017:EBb

[1306] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 132:ifc, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301835.

Joly:2017:FCB

[1307] Marcel Joly and Patrícia H. C. Rondó. The future of computational biomedicine: Complex systems thinking. Mathematics and Computers in Simulation, 132:1– 27, February 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001299.

Burger:2017:SAF

[1308] Raimund Bürger, Ricardo Ruiz-Baier, and Canrong Tian. Stability analysis and finite volume element discretization for delay-driven spatiotemporal patterns in a predator– prey model. Mathematics and Computers in Simulation, 132:28– 52, February 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541630088X.

Chen:2017:DFB

[1309] Feiyan Chen, Feng Ding, Ahmed Alsaedi, and Tasawar Hayat. Data filtering based multi-innovation extended gradient method for controlled autoregressive autoregressive moving average systems using the maximum likelihood principle. Mathematics and Computers in Simulation, 132:53– 67, February 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300921.

Wang:2017:DBO

[1310] Zhen Wang, Yuanfu Shao, Xianjia Fang, and Xiangmin Ma. The dynamic behaviors of one-predator twoprey system with mutual interference and impulsive control. Mathematics and Computers in Simulation, 132:68–85, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300933.

Andaluz:2017:NCB

[1311] J. Andaluz, A. A. Elsadany, and G. Jarne. Nonlinear Cournot and Bertrand-type dynamic triopoly with differentiated products and heterogeneous expectations. Mathematics and Computers in Simulation, 132:86–99, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300969.

Gao:2017:RMG

[1312] Huan Gao, Rogemar Mamon, and Xiaoming Liu. Risk measurement of a guaranteed annuity option under a stochastic modelling framework. Mathematics and Computers in Simulation, 132:100–119, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416301100.

Behera:2017:ADO

[1313] Ratikanta Behera and Mani Mehra. Approximation of the differential operators on an adaptive spherical geodesic grid using spherical wavelets. Mathematics and Computers in Simulation, 132:120–138, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416301148.

Raja:2017:DUF

[1314] Muhammad Asif Zahoor Raja, Raza Samar, Muhammad Anwar Manzar, and Syed Muslim Shah. Design of unsupervised fractional neural network model optimized with interior point algorithm for solving Bagley– Torvik equation. Mathematics and Computers in Simulation, 132:139– 158, February 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475416301276.

Sobieski:2017:MSM

[1315] Wojciech Sobieski and Qiang Zhang. Multi-scale modeling of flow resistance in granular porous media. Mathematics and Computers in Simulation, 132:159–171, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416301422.

Kingni:2017:CAU

[1316] Sifeu Takougang Kingni, Sajad Jafari, Viet-Thanh Pham, and Paul Woafo. Constructing and analyzing of a unique three-dimensional chaotic autonomous system exhibiting three families of hidden attractors. Mathematics and Computers in Simulation, 132:172–182, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301434.

Zhang:2017:GES

[1317] Yu Zhang. Global exponential stability of delay difference equations with delayed impulses. Mathematics and Computers in Simulation, 132: 183–194, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301458.

Kotb:2017:IPM

[1318] M. S. Kotb and M. Z. Raqab. Inference and prediction for modified Weibull distribution based on doubly censored samples. Mathematics and Computers in Simulation, 132: 195–207, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541630146X.

Shen:2017:CCU

[1319] Zhiping Shen and Juntao Li. Chaos control for a unified chaotic system using output feedback controllers. Mathematics and Computers in Simulation, 132:208–219, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301471.

Yang:2017:MRS

[1320] Jin Yang, Guangyao Tang, and Sanyi Tang. Modelling the regulatory system of a chemostat model with a threshold window. Mathematics and Computers in Simulation, 132:220– 235, February 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416301495.

Faure:2017:LDS

[1321] Henri Faure and Christiane Lemieux. Low-discrepancy sequences: Atanassov’s methods revisited. Mathematics and Computers in Simulation, 132:236– 256, February 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416301513.

Rahman:2017:SWM

[1322] M. Mizanur Rahman, Gour Chandra Paul, and Ashabul Hoque. A shallow water model for computing water level due to tide and surge along the coast of Bangladesh using nested numerical schemes. Mathematics and Computers in Simulation, 132:257–276, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301525.

Chaouech:2017:DNF

[1323] Lotfi Chaouech, Moêz Soltani, Slim Dhahri, and Abdelkader Chaari. Design of new fuzzy sliding mode controller based on parallel distributed compensation controller and using the scalar sign function. Mathematics and Computers in Simulation, 132: 277–288, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301549.

Zhang:2017:NNA

[1324] Yan Zhang, Shihua Chen, Shujing Gao, Kuangang Fan, and Qingyun Wang. A new non-autonomous model for migratory birds with Leslie–Gower Holling-type II schemes and saturation recovery rate. Mathematics and Computers in Simulation, 132: 289–306, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301719.

Anonymous:2017:NIb

[1325] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 132:307, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301884.

Anonymous:2017:ICEb

[1326] Anonymous. IMACS calendar events.

Mathematics and Computers in Simulation, 132:308–309, February 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475416301896.

Anonymous:2017:EBc

[1327] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 133:ifc, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416302129.

Alt:2017:BB

[1328] René Alt and Svetoslav Markov. Biomath 2014 and Biomath 2015. Mathematics and Computers in Simulation, 133:1, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416302105.

Akimenko:2017:ASSa

[1329] Vitalii V. Akimenko. Asymptotically stable states of nonlinear age-structured monocyclic population model I. Travelling wave solution. Mathematics and Computers in Simulation, 133:2–23, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415001159.

Akimenko:2017:ASSb

[1330] Vitalii V. Akimenko. Asymptotically stable states of non-linear age-structured monocyclic population model II. Numerical simulation. Mathematics and Computers in Simulation, 133:24–38, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001123.

Anguelov:2017:SOP

[1331] Roumen Anguelov and Stephanus Marnus Stoltz. Stationary and oscillatory patterns in a coupled Brusselator model. Mathematics and Computers in Simulation, 133:39– 46, March 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001111.

Anguelov:2017:SPE

[1332] Roumen Anguelov, Claire Dufourd, and Yves Dumont. Simulations and parameter estimation of a trapinsect model using a finite element approach. Mathematics and Computers in Simulation, 133:47– 75, March 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001536.

Conradi:2017:GTA

[1333] Carsten Conradi and Maya Mincheva.

Graph-theoretic analysis of multistationarity using degree theory. Mathematics and Computers in Simulation, 133:76–90, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001755.

Dimitrov:2017:MRC

[1334] Stanko Dimitrov and Svetoslav Markov.

Metabolic rate constants: Some computational aspects. Mathematics and Computers in Simulation, 133:91– 110, March 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415002529.

Ivanov:2017:PPM

[1335] Tihomir Ivanov and Neli Dimitrova. A predator–prey model with generic birth and death rates for the predator and Beddington–DeAngelis functional response. Mathematics and Computers in Simulation, 133:111– 123, March 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001548.

Jereva:2017:MDS

[1336] Dessislava Jereva, Filip Fratev, Ivanka Tsakovska, Petko Alov, Tania Pencheva, and Ilza Pajeva. Molecular dynamics simulation of the human estrogen receptor alpha: contribution to the pharmacophore of the agonists. Mathematics and Computers in Simulation, 133:124–134, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001354.

Lazarova:2017:DPA

[1337] M. D. Lazarova and L. D. Minkova. I-Delaporte process and applications. Mathematics and Computers in Simulation, 133:135–141, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475415002608.

Berge:2017:GST

[1338] Tsanou Berge, Samuel Bowong, and Jean M.-S. Lubuma. Global stability of a two-patch cholera model with fast and slow transmissions. Mathematics and Computers in Simulation, 133:142–164, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002384.

Timofte:2017:HRC

[1339] Claudia Timofte. Homogenization results for the calcium dynamics in living cells. Mathematics and Computers in Simulation, 133:165– 174, March 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001305.

Akimenko:2017:NAS

[1340] Vitalii V. Akimenko. Nonlinear age-structured models of polycyclic population dynamics with death rates as power functions with exponent n. Mathematics and Computers in Simulation, 133:175– 205, March 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416301483.

Arenas:2017:LGM

[1341] Abbiana R. Arenas, Neil B. Thackar, and Evan C. Haskell. The logistic growth model as an approximating model for viral load measurements of influenza A virus. Mathematics and Computers in Simulation, 133:206–222, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301951.

Iliev:2017:ASF

[1342] A. Iliev, N. Kyurkchiev, and S. Markov. On the approximation of the step function by some sigmoid functions. Mathematics and Computers in Simulation, 133:223–234, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002554.

Jalalimanesh:2017:SBO

[1343] Ammar Jalalimanesh, Hamidreza Shahabi Haghighi, Abbas Ahmadi, and Madjid Soltani. Simulation-based optimization of radiotherapy: Agentbased modeling and reinforcement learning. Mathematics and Computers in Simulation, 133:235–248, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475416300878.

Kolkovska:2017:SIS

[1344] N. Kolkovska, M. Dimova, and N. Kutev. Stability or instability of solitary waves to double dispersion equation with quadraticcubic nonlinearity. Mathematics and Computers in Simulation, 133:249– 264, March 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300271.

Tamen:2017:MMT

[1345] A. Tchuinté Tamen, Y. Dumont, J. J.

Tewa, S. Bowong, and P. Couteron. A minimalistic model of tree–grass interactions using impulsive differential equations and non-linear feedback functions of grass biomass onto fire-induced tree mortality. Mathematics and Computers in Simulation, 133:265–297, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300258.

Timofte:2017:MAC

[1346] Claudia Timofte. Multiscale analysis of a carcinogenesis model. Mathematics and Computers in Simulation, 133:298–310, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300945.

Usaini:2017:DBE

[1347] Salisu Usaini, Alun L. Lloyd, Roumen Anguelov, and Salisu M. Garba. Dynamical behavior of an epidemiological model with a demographic Allee effect. Mathematics and Computers in Simulation, 133:311– 325, March 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300866.

Wadhwa:2017:SKC

[1348] R. R. Wadhwa, L. Zalányi, J. Szente, L. Négyessy, and P. Érdi. Stochastic kinetics of the circular gene hypothesis: Feedback effects and protein fluctuations. Mathematics and Computers in Simulation, 133:326– 336, March 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001585.

Wood:2017:UAA

[1349] Daniel T. Wood, Hristo V. Kojouharov, and Dobromir T. Dimitrov. Universal approaches to approximate biological systems with nonstandard finite difference methods. Mathematics and Computers in Simulation, 133:337–350, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300362.

Anonymous:2017:NIc

[1350] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 133:351, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416302191.

Anonymous:2017:ICEc

[1351] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 133:352–353, March 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475416302208.

Anonymous:2017:PAa

[1352] Anonymous. Pages 1–102 (April 2017).

Mathematics and Computers in Simulation, 134:1–102, April 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2017:EBd

[1353] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 134:ifc, April 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416302348.

Zaman:2017:NSP

[1354] A. Zaman, N. Ali, and M. Sajid. Numerical simulation of pulsatile flow of blood in a porous-saturated overlapping stenosed artery. Mathematics and Computers in Simulation, 134: 1–16, April 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416301756.

Lamnii:2017:CHI

[1355] A. Lamnii, M. Lamnii, and F. Oumellal. Computation of Hermite interpolation in terms of B-spline basis using polar forms. Mathematics and Computers in Simulation, 134: 17–27, April 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541630177X.

Ferrer-Perez:2017:CTM

[1356] H. Ferrer-Pérez, M. I. Ayuda, and A. Aznar. A comparison of two modified stationarity tests. A Monte Carlo study. Mathematics and Computers in Simulation, 134:28– 36, April 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416301781.

Villena:2017:DSR

[1357] Marcelo J. Villena and Axel A. Araneda. Dynamics and stability in retail competition. Mathematics and Computers in Simulation, 134: 37–53, April 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416301793.

Patra:2017:MSW

[1358] Atasi Patra, Jai Tushar, and B. Dubey. Modeling and simulation of a wetland park: an application to Keoladeo National Park, India. Mathematics and Computers in Simulation, 134:54–78, April 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301811.

Montanari:2017:MAP

[1359] Roberto Montanari, Eleonora Bottani, Ehsan Shekarian, and Nima Kazemi. A model for the analysis of procurement strategies in the economic order interval environment. Mathematics and Computers in Simulation, 134:79–98, April 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416302002.

Anonymous:2017:NId

[1360] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 134:99, April 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416302397.

Anonymous:2017:ICEd

[1361] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 134:100–101, April 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475416302403.

Anonymous:2017:EBe

[1362] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 135:ifc, May 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300137.

Anonymous:2017:E

[1363] Anonymous.  Editorial.  Mathe-

matics and Computers in Simulation, 135:1–2, May 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300046.

LEcuyer:2017:RNP

[1364] Pierre L’Ecuyer, David Munger, Boris Oreshkin, and Richard Simard. Random numbers for parallel computers: Requirements and methods, with emphasis on GPUs. Mathematics and Computers in Simulation, 135: 3–17, May 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300829.

Niederreiter:2017:RCL

[1365] Harald Niederreiter. Recent constructions of low-discrepancy sequences. Mathematics and Computers in Simulation, 135:18–27, May 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414002432.

Deaconu:2017:WMS

[1366] M. Deaconu, S. Herrmann, and S. Maire. The walk on moving spheres: a new tool for simulating Brownian motion’s exit time from a domain. Mathematics and Computers in Simulation, 135:28–38, May 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415001366.

Dermoune:2017:OMH

[1367] Azzouz Dermoune, Daoud Ounaissi, and Nadji Rahmania. Oscillation of Metropolis–Hastings and simulated annealing algorithms around LASSO estimator. Mathematics and Computers in Simulation, 135: 39–50, May 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415001901.

Fakhereddine:2017:SMC

[1368] Rana Fakhereddine, Rami El Haddad, Christian Lécot, and Joseph El Maalouf. Stratified Monte Carlo simulation of Markov chains. Mathematics and Computers in Simulation, 135:51–62, May 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416302531.

Faure:2017:RDB

[1369] Henri Faure and Christiane Lemieux.

A review of discrepancy bounds for (t, s) and (t, e, s)-sequences with numerical comparisons. Mathematics and Computers in Simulation, 135: 63–71, May 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475414002419.

Daun:2017:CPI

[1370] Thomas Daun and Stefan Heinrich.

Complexity of parametric initial value problems for systems of ODEs. Mathematics and Computers in Simulation, 135:72–85, May 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415000713.

Cheng:2017:RWG

[1371] Wen-Ju Cheng, Jim Cox, and Paula Whitlock. Random walks on graphs and Monte Carlo methods. Mathematics and Computers in Simulation, 135:86–94, May 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415002748.

Anonymous:2017:NIe

[1372] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 135:95, May 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417300204.

Anonymous:2017:ICEe

[1373] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 135:96–97, May 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475417300216.

Anonymous:2017:PJ

[1374] Anonymous. Pages 1–144 (June 2017). Mathematics and Computers in Simulation, 136:1–144, June 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2017:EBf

[1375] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 136:ifc, June 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417300459.

Geum:2017:FOQ

[1376] Young Hee Geum, Young Ik Kim, and Beny Neta. A family of optimal quartic-order multiple-zero finders with a weight function of the principal k-th root of a derivative-to-derivative ratio and their basins of attraction. Mathematics and Computers in Simulation, 136:1–21, June 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416302014.

Duda:2017:LMA

[1377] Jozef Duda. Lyapunov matrices approach to the parametric optimization of a neutral system with two delays. Mathematics and Computers in Simulation, 136:22–35, June 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416302026.

Kizilaslan:2017:CBE

[1378] Fatih Kızılaslan. Classical and Bayesian estimation of reliability in a multicomponent stress–strength model based on the proportional reversed hazard rate mode. Mathematics and Computers in Simulation, 136: 36–62, June 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416302063.

Zaman:2017:OSV

[1379] Gul Zaman, Yong Han Kang, Giphil Cho, and Il Hyo Jung. Optimal strategy of vaccination & treatment in an sir epidemic model. Mathematics and Computers in Simulation, 136:63–77, June 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541630249X.

Goyal:2017:PFF

[1380] Anmol Goyal and Bernd Simeon. On penalty-free formulations for multipatch isogeometric Kirchhoff– Love shells. Mathematics and Computers in Simulation, 136:78– 103, June 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475416302506.

Liu:2017:DAH

[1381] Chao Liu, Na Lu, and Qingling Zhang. Dynamical analysis in a hybrid bioeconomic system with multiple time delays and strong Allee effect. Mathematics and Computers in Simulation, 136:104–131, June 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417300010.

Prants:2017:CDT

[1382] Fabiola G. Prants and Paulo C. Rech. Complex dynamics of a three-dimensional continuous-time autonomous system. Mathematics and Computers in Simulation, 136:132– 139, June 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475417300022.

Anonymous:2017:NIf

[1383] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 136:140, June 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300502.

Anonymous:2017:ICEf

[1384] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 136:141–143, June 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475417300514.

Anonymous:2017:EBg

[1385] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 137:ifc, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300630.

Amaziane:2017:MSI

[1386] Brahim Amaziane, Domingo Barrera, and Driss Sbibih. MATCOM Special Issue MAMERN VI-2015: 6th International Conference on Approximation Methods and Numerical Modeling in Environment and Natural Resources. Mathematics and Computers in Simulation, 137:1, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300617.

Abreu:2017:NFV

[1387] E. Abreu, W. Lambert, J. Perez, and A. Santo. A new finite volume approach for transport models and related applications with balancing source terms.

Mathematics and Computers in Simulation, 137:2–28, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300289.

Abreu:2017:CNS

[1388] Eduardo Abreu and Jardel Vieira. Computing numerical solutions of the pseudo-parabolic Buckley–Leverett equation with dynamic capillary pressure. Mathematics and Computers in Simulation, 137:29–48, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475416301999.

Ahmed:2017:RFM

[1389] Elyes Ahmed, Jérôme Jaffré, and Jean E. Roberts. A reduced fracture model for two-phase flow with different rock types. Mathematics and Computers in Simulation, 137: 49–70, July 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416301987.

Ahusborde:2017:SAN

[1390] E. Ahusborde and M. El Ossmani. A sequential approach for numerical simulation of two-phase multicomponent flow with reactive transport in porous media. Mathematics and Computers in Simulation, 137: 71–89, July 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416302452.

Allouch:2017:SSQ

[1391] C. Allouch, A. Boujraf, and M. Tahrichi.

Superconvergent spline quasi-interpolants and an application to numerical integration. Mathematics and Computers in Simulation, 137:90–108, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847541630204X.

Almeida:2017:DSP

[1392] Rui M. P. Almeida, Stanislav N. Antontsev, and José C. M. Duque. Discrete solutions for the porous medium equation with absorption and variable exponents. Mathematics and Computers in Simulation, 137:109– 129, July 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417300071.

Barranco-Chamorro:2017:SRB

[1393] I. Barranco-Chamorro, P. L. Luque-Calvo, M. D. Jiménez-Gamero, and M. V. Alba-Fernández. A study of risks of Bayes estimators in the generalized half-logistic distribution for progressively type-II censored samples. Mathematics and Computers in Simulation, 137:130–147, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301690.

Bellour:2017:CCM

[1394] A. Bellour, E. A. Rawashdeh, and A. Zidna. Continuous collocation method for third kind integral– algebraic equations. Mathematics and Computers in Simulation, 137: 148–158, July 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416302038.

Benkhaldoun:2017:SMM

[1395] Fayssal Benkhaldoun, A. Halassi, Driss Ouazar, Mohammed Seaid, and Ahmed Taik. A stabilized meshless method for time-dependent convectiondominated flow problems. Mathematics and Computers in Simulation, 137:159–176, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416302099.

Bruvoll:2017:USW

[1396] Solveig Bruvoll, Tom Lyche, and Knut Mørken. Uniformly stable wavelets on nonuniform triangulations. Mathematics and Computers in Simulation, 137:177–200, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301732.

Calgaro:2017:SNH

[1397] Caterina Calgaro, Emmanuel Creusé, Thierry Goudon, and Stella Krell. Simulations of non homogeneous viscous flows with incompressibility constraints. Mathematics and Computers in Simulation, 137:201– 225, July 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416302324.

Capatina:2017:ENN

[1398] D. Capatina, H. El-Otmany, D. Graebling, and R. Luce. Extension of NXFEM to nonconforming finite elements. Mathematics and Computers in Simulation, 137:226– 245, July 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417300083.

Cherfils:2017:NVU

[1399] L. Cherfils, C. Choquet, and M. M.

Diédhiou. Numerical validation of an upscaled sharp–diffuse interface model for stratified miscible flows. Mathematics and Computers in Simulation, 137:246–265, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416302051.

Engwer:2017:PFA

[1400] Christian Engwer and Liesel Schumacher. A phase field approach to pressurized fractures using discontinuous Galerkin methods. Mathematics and Computers in Simulation, 137:266–285, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416302075.

Erhel:2017:AGR

[1401] Jocelyne Erhel and Souhila Sabit.

Analysis of a global reactive transport model and results for the Mo-MaS benchmark. Mathematics and Computers in Simulation, 137:286– 298, July 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416302464.

Fumagalli:2017:ACL

[1402] Alessio Fumagalli, Stefano Zonca, and Luca Formaggia. Advances in computation of local problems for a flow-based upscaling in fractured reservoirs. Mathematics and Computers in Simulation, 137:299– 324, July 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417300320.

Gallouet:2017:CMC

[1403] T. Gallouët, R. Herbin, D. Maltese, and A. Novotny. Convergence of the marker-and-cell scheme for the semi-stationary compressible Stokes problem. Mathematics and Computers in Simulation, 137:325– 349, July 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416301963.

Gomez-Lopera:2017:EPD

[1404] J. F. Gómez-Lopera, J. Martínez-Aroza, R. Román-Roldán, R. Román-Gálvez, and D. Blanco-Navarro. The evaluation problem in discrete semihidden Markov models. Mathematics and Computers in Simulation, 137:350–365, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416302518.

Hoang:2017:STD

[1405] Thi-Thao-Phuong Hoang, Caroline Japhet, Michel Kern, and Jean E. Roberts. Space–time domain decomposition for advection–diffusion problems in mixed formulations. Mathematics and Computers in Simulation, 137:366–389, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416302087.

Jimenez-Gamero:2017:FTT

[1406] M. D. Jiménez-Gamero, M. V. Alba-Fernández, P. Jodrá, and I. Barranco-Chamorro. Fast tests for the twosample problem based on the empirical characteristic function. Mathematics and Computers in Simulation, 137:390–410, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301744.

Lenarduzzi:2017:ESR

[1407] Licia Lenarduzzi and Monica Pepe. Enhancing and segmenting a remote sensing image of a glacier body. Mathematics and Computers in Simulation, 137:411–423, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301720.

Reinoso-Gordo:2017:PDN

[1408] J. F. Reinoso-Gordo, M. J. Ibáñez, and R. Romero-Zaliz. Parallelizing drainage network algorithm using free software: Octave as a solution. Mathematics and Computers in Simulation, 137:424–430, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301707.

Sin:2017:SNG

[1409] Irina Sin, Vincent Lagneau, Laurent De Windt, and Jérôme Corvisier. 2D simulation of natural gas reservoir by two-phase multicomponent reactive flow and transport — Description of a benchmarking exercise. Mathematics and Computers in Simulation, 137:431–447, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541630252X.

Touzani:2017:FES

[1410] R. Touzani, L. Alessio, D. Kuzmichev, and R. Buoy. Finite element solution for a coal-bed methane reservoir model. Mathematics and Computers in Simulation, 137:448– 461, July 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417300307.

Anonymous:2017:NIg

[1411] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 137:462, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300708.

Anonymous:2017:ICEg

[1412] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 137:463–465, July 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847541730071X.

Anonymous:2017:PAb

[1413] Anonymous. Pages 1–82 (August 2017). Mathematics and Computers in Simulation, 138:1–82, August 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2017:EBh

[1414] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 138:ifc, August 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300903.

Dinh:2017:AIK

[1415] Khanh N. Dinh and Roger B. Sidje.

Analysis of inexact Krylov subspace methods for approximating the matrix exponential. Mathematics and Computers in Simulation, 138:1– 13, August 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417300034.

Clempner:2017:UEM

[1416] Julio B. Clempner and Alexander S.

Poznyak. Using the extraproximal method for computing the shortestpath mixed Lyapunov equilibrium in Stackelberg security games. Mathematics and Computers in Simulation, 138:14–30, August 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300095.

Wang:2017:SHB

[1417] Yan Wang and Xianning Liu. Stability and Hopf bifurcation of a within-host Chikungunya virus infection model with two delays. Mathematics and Computers in Simulation, 138:31–48, August 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300101.

Pupo:2017:CDT

[1418] Ana Elisa Bergues Pupo, Maraelys Morales González, Luis Enrique Bergues Cabrales, Juan Bory Reyes, Eduardo José Roca Oria, Juan José Godina Nava, Rolando Placeres Jiménez, Francisco Martínez Sánchez, Héctor Manuel Camué Ciria, and Jesús Manuel Bergues Cabrales. 3d current density in tumors and surrounding healthy tissues generated by a system of straight electrode arrays. Mathematics and Computers in Simulation, 138:49–64, August 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300113.

Wang:2017:ESP

[1419] Qinglong Wang and Zhijun Liu. Existence and stability of positive almost periodic solutions for a competitive system on time scales. Mathematics and Computers in Simulation, 138:65–77, August 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300265.

Anonymous:2017:NIh

[1420] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 138:78, August 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300952.

Anonymous:2017:ICEh

[1421] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 138:79–81, August 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475417300964.

Anonymous:2017:EBi

[1422] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 139:ifc, September 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417301714.

Abgrall:2017:TUM

[1423] R. Abgrall, P. M. Congedo, and G. Geraci. Towards a unified multiresolution scheme for treating discontinuities in differential equations with uncertainties. Mathematics and Computers in Simulation, 139: 1–22, September 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416000161.

Abouali:2017:SPE

[1424] Mohammad Abouali and Jose E. Castillo. Solving Poisson equation with Robin boundary condition on a curvilinear mesh using high order mimetic discretization methods. Mathematics and Computers in Simulation, 139:23–36, September 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475414003206.

Farago:2017:DNN

[1425] István Faragó, János Karátson, and Sergey Korotov. Discrete nonnegativity for nonlinear cooperative parabolic PDE systems with nonmonotone coupling. Mathematics and Computers in Simulation, 139: 37–53, September 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301264.

Laug:2017:MTR

[1426] P. Laug and H. Borouchaki. Metric tensor recovery for adaptive meshing. Mathematics and Computers in Simulation, 139:54–66, September 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475415000373.

Tolstykh:2017:HOM

[1427] Andrei I. Tolstykh, Michael V.

Lipavskii, and Dmitrii A. Shirobokov. High-order multioperatorsbased schemes: Developments and applications. Mathematics and Computers in Simulation, 139:67–80, September 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417300794.

Anonymous:2017:NIi

[1428] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 139:81, September 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417301787.

Anonymous:2017:ICEi

[1429] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 139:82–84, September 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417301799.

Anonymous:2017:PO

[1430] Anonymous. Pages 1–142 (October 2017). Mathematics and Computers in Simulation, 140:1–142, October 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2017:EBj

[1431] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 140:ifc, October 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417301945.

Liang:2017:HFN

[1432] Yu Liang, Zhenjun Shi, and Peter W. Chung. A Hessian-free Newton– Raphson method for the configuration of physics systems featured by numerically asymmetric force field. Mathematics and Computers in Simulation, 140:1–23, October 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300290.

Correa:2017:SDC

[1433] Maicon R. Correa. A semi-discrete central scheme for incompressible multiphase flow in porous media in several space dimensions. Mathematics and Computers in Simulation, 140:24–52, October 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300332.

Sweilam:2017:NSD

[1434] N. H. Sweilam and M. M. Abou Hasan. Numerical solutions for 2-D fractional Schrödinger equation with the Riesz–Feller derivative. Mathematics and Computers in Simulation, 140:53–68, October 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300605.

Muroi:2017:CGJ

[1435] Yoshifumi Muroi and Shintaro Suda.

Computation of Greeks in jumpdiffusion models using discrete Malliavin calculus. Mathematics and Computers in Simulation, 140:69– 93, October 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417300782.

Rydel:2017:SRP

[1436] Marek Rydel and Włodzimierz Stanisławski. Selection of reduction parameters for complex plant MIMO LTI models using the evolutionary algorithm. Mathematics and Computers in Simulation, 140:94– 106, October 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417300812.

Drazic:2017:TDC

[1437] Ivan Dražić, Nermina Mujaković, and Nelida Črnjarić-Žic. Threedimensional compressible viscous micropolar fluid with cylindrical symmetry: Derivation of the model and a numerical solution. Mathematics and Computers in Simulation, 140: 107–124, October 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300824.

Chen:2017:NSH

[1438] Y. Chen and Chang-Ming Chen. Numerical simulation with high order accuracy for the time fractional reaction– subdiffusion equation. Mathematics and Computers in Simulation, 140: 125–138, October 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417300848.

Anonymous:2017:EBk

[1439] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 140:139, October 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417301994.

Anonymous:2017:EBl

[1440] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 140:140–142, October 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417302008.

Anonymous:2017:EBm

[1441] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 141:ifc, November 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417302586.

DellAccio:2017:E

[1442] Francesco Dell’Accio, Filomena Di Tommaso, Maria Italia Gualtieri, Gennaro Infante, and Anna Napoli. Editorial. Mathematics and Computers in Simulation, 141:1–2, November 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475417302562.

Ala:2017:AMA

[1443] Guido Ala, Gregory E. Fasshauer, Elisa Francomano, Salvatore Ganci, and Michael J. McCourt. An augmented MFS approach for brain activity reconstruction. Mathematics and Computers in Simulation, 141:3– 15, November 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416302488.

Altomare:2017:NDF

[1444] M. Altomare and F. A. Costabile.

A new determinant form of Bessel polynomials and applications. Mathematics and Computers in Simulation, 141:16–23, November 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301112.

Amodio:2017:GTM

[1445] P. Amodio, F. Iavernaro, F. Mazzia, M. S. Mukhametzhanov, and Ya. D. Sergeyev. A generalized Taylor method of order three for the solution of initial value problems in standard and infinity floating-point arithmetic. Mathematics and Computers in Simulation, 141:24–39, November 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300234.

Bulai:2017:SEH

[1446] Iulia Martina Bulai and Ezio Venturino. Shape effects on herd behavior in ecological interacting population models. Mathematics and Computers in Simulation, 141:40– 55, November 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475417301696.

Costabile:2017:CLS

[1447] Francesco A. Costabile, Maria Italia Gualtieri, and Giada Serafini. Cubic Lidstone-Spline for numerical solution of BVPs. Mathematics and Computers in Simulation, 141:56– 64, November 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475417300319.

DellAccio:2017:BSB

[1448] F. Dell’Accio and F. Di Tommaso. Bivariate Shepard–Bernoulli operators. Mathematics and Computers in Simulation, 141:65–82, November 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475417302732.

Farhloul:2017:DMF

[1449] Mohamed Farhloul and Abdelmalek Zine. A dual-mixed finite element method for quasi-Newtonian flows whose viscosity obeys a power law or the Carreau law. Mathematics and Computers in Simulation, 141:

83–95, November 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416302476.

Sergeyev:2017:OZC

[1450] Yaroslav D. Sergeyev, Dmitri E.

Kvasov, and Marat S. Mukhametzhanov. Operational zones for comparing metaheuristic and deterministic one-dimensional global optimization algorithms. Mathematics and Computers in Simulation, 141:96– 109, November 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300830.

Sommariva:2017:NHN

[1451] Alvise Sommariva and Marco Vianello.

Numerical hyperinterpolation over nonstandard planar regions. Mathematics and Computers in Simulation, 141:110–120, November 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301197.

Anonymous:2017:NIj

[1452] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 141:121, November 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417302653.

Anonymous:2017:ICEj

[1453] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 141:122–124, November 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475417302665.

Anonymous:2017:PD

[1454] Anonymous. Pages 1–134 (December 2017). Mathematics and Computers in Simulation, 142:1–134, December 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2017:EBn

[1455] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 142:ifc, December 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417302860.

Roy:2017:HTM

[1456] Banani Roy, Sankar Kumar Roy, and Dil Bahadur Gurung. Holling–Tanner model with Beddington–DeAngelis functional response and time delay introducing harvesting. Mathematics and Computers in Simulation, 142:1–14, December 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417300861.

Grimmonprez:2017:EEF

[1457] Marijke Grimmonprez and Marián Slodička. Error estimates for the full discretization of a semilinear parabolic problem with an unknown source. Mathematics and Computers in Simulation, 142:15– 33, December 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417300873.

Liu:2017:TGM

[1458] Wei Liu. A two-grid method for the semi-linear reaction–diffusion system of the solutes in the groundwater flow by finite volume element. Mathematics and Computers in Simulation, 142:34–50, December 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541730157X.

Yongzhen:2017:PVE

[1459] Pei Yongzhen, Li Shuping, Gao Shujing, and Zhong Min. Pulse vaccination of an epidemic model with two parallel infectious stages and time delays. Mathematics and Computers in Simulation, 142:51– 61, December 2017. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417301659.

Kulikov:2017:ASE

[1460] G. Yu. Kulikov and M. V. Kulikova.

Accurate state estimation of stiff continuous-time stochastic models in chemical and other engineering. Mathematics and Computers in Simulation, 142:62–81, December 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417301660.

Rayskin:2017:UDD

[1461] Victoria Rayskin. Users’ dynamics on digital platforms. Mathematics and Computers in Simulation, 142: 82–97, December 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417301672.

Garcia:2017:CLD

[1462] Salvador Garcia. Chaos in the liddriven square cavity. Mathematics and Computers in Simulation, 142: 98–112, December 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417301866.

Ndanguza:2017:ABE

[1463] Denis Ndanguza, Isambi S. Mbalawata, Heikki Haario, and Jean M. Tchuenche. Analysis of bias in an Ebola epidemic model by extended Kalman filter approach. Mathematics and Computers in Simulation, 142: 113–129, December 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541730191X.

Anonymous:2017:NIk

[1464] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 142:130, December 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417302914.

Anonymous:2017:EBo

[1465] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 142:131–133, December 2017. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417302926.

Anonymous:2018:EBa

[1466] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 143:ifc, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417303002.

Buckwar:2018:E

[1467] Evelyn Buckwar, Peter Kritzer, Gunther Leobacher, Friedrich Pillichshammer, and Arne Winterhof. Editorial. Mathematics and Computers in Simulation, 143:1–2, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417302987.

Kritzinger:2018:EFD

[1468] Ralph Kritzinger. An exact formula for the L_{2} discrepancy of the symmetrized Hammersley point set. Mathematics and Computers in Simulation, 143:3–13, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475415002591.

Acevedo:2018:NLA

[1469] José Infante Acevedo and Tony Lelièvre. A non linear approximation method for solving high dimensional partial differential equations: Application in finance. Mathematics and Computers in Simulation, 143:14–34, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301240.

Irrgeher:2018:IAC

[1470] Christian Irrgeher, Peter Kritzer, and Friedrich Pillichshammer. Integration and approximation in cosine spaces of smooth functions. Mathematics and Computers in Simulation, 143:35–45, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416000549.

Sabelfeld:2018:RWS

[1471] Karl K. Sabelfeld. A random walk on spheres based kinetic Monte Carlo method for simulation of the fluctuation-limited bimolecular reactions. Mathematics and Computers in Simulation, 143:46–56, January 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300283.

Meidl:2018:SCB

[1472] Wilfried Meidl. A secondary construction of bent functions, octal gbent functions and their duals. Mathematics and Computers in Simulation, 143:57–64, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416000306.

Lejay:2018:MCE

[1473] Antoine Lejay. A Monte Carlo estimation of the mean residence time in cells surrounded by thin layers. Mathematics and Computers in Simulation, 143:65–77, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417302069.

Hofer:2018:HTS

[1474] Roswitha Hofer. Halton-type sequences in rational bases in the ring of rational integers and in the ring of polynomials over a finite field. Mathematics and Computers in Simulation, 143:78–88, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301124.

Agarwal:2018:SNR

[1475] Ankush Agarwal, Stefano De Marco, Emmanuel Gobet, and Gang Liu. Study of new rare event simulation schemes and their application to extreme scenario generation. Mathematics and Computers in Simulation, 143:89–98, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417301908.

Lang:2018:MCV

[1476] Annika Lang and Andreas Petersson. Monte Carlo versus multilevel Monte Carlo in weak error simulations of SPDE approximations. Mathematics and Computers in Simulation, 143:99–113, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541730188X.

Li:2018:EDN

[1477] Jie Li, Jianliang Zheng, and Paula Whitlock. Efficient deterministic and non-deterministic pseudorandom number generation. Mathematics and Computers in Simulation, 143:114– 124, January 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416301227.

Belomestny:2018:SRB

[1478] D. Belomestny, S. Häfner, and M. Urusov. Stratified regression-based variance reduction approach for weak approximation schemes. Mathematics and Computers in Simulation, 143: 125–137, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417301891.

Wagner:2018:RWM

[1479] Wolfgang Wagner. A random walk model for the Schrödinger equation. Mathematics and Computers in Simulation, 143:138–148, January 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416301239.

Kruse:2018:DSG

[1480] Raphael Kruse and Michael Scheutzow. A discrete stochastic Gronwall lemma. Mathematics and Computers in Simulation, 143:149– 157, January 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416300970.

Puchhammer:2018:ELB

[1481] Florian Puchhammer. On an explicit lower bound for the star discrepancy in three dimensions. Mathematics and Computers in Simulation, 143:158–168, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301501.

Sabelfeld:2018:SPM

[1482] Karl K. Sabelfeld. Stochastic projection methods and applications to some nonlinear inverse problems of phase retrieving. Mathematics and Computers in Simulation, 143:169– 175, January 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416301252.

Reisinger:2018:NLM

[1483] C. Reisinger. The non-locality of Markov chain approximations to twodimensional diffusions. Mathematics and Computers in Simulation, 143: 176–185, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416300854.

Ullrich:2018:LBD

[1484] Mario Ullrich. A lower bound for the dispersion on the torus. Mathematics and Computers in Simulation, 143:186–190, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475415002736.

LEcuyer:2018:SMC

[1485] Pierre L’Ecuyer, David Munger, Christian Lécot, and Bruno Tuffin. Sorting methods and convergence rates for Array-RQMC: Some empirical comparisons. Mathematics and Computers in Simulation, 143: 191–201, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416301203.

Gilbert:2018:HWC

[1486] Alexander D. Gilbert, Frances Y. Kuo, and Ian H. Sloan. Hiding the weights — CBC black box algorithms with a guaranteed error bound. Mathematics and Computers in Simulation, 143:202–214, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541630091X.

Barth:2018:WCG

[1487] Andrea Barth and Tobias Stüwe. Weak convergence of Galerkin approximations of stochastic partial differential equations driven by additive Lévy noise. Mathematics and Computers in Simulation, 143:215– 225, January 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417300836.

Anonymous:2018:NIa

[1488] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 143:226, January 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417303075.

Anonymous:2018:ICEa

[1489] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 143:227–229, January 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417303087.

BerguesCabrales:2018:CMM

[1490] Luis Enrique Bergues Cabrales, Andrés Ramírez Aguilera, Rolando Placeres Jiménez, Manuel Verdecia Jarque, Héctor Manuel Camué Ciria and Juan Bory Reyes, Miguel Angel O Farril Mateus, Fabiola Suárez Palencia, and Marisela González Ávila. Corrigendum to “Mathematical modeling of tumor growth in mice following low-level direct electric current” [Math. Comput. Simulation 78 (2008) 112–120]. Mathematics and Computers in Simulation, 144:266, February 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541730085X. See

[1].

Anonymous:2018:PF

[1491] Anonymous. Pages 1–270 (February 2018). Mathematics and Computers in Simulation, 144:1–270, February 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2018:EBb

[1492] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 144:ifc, February 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417303257.

Pulch:2018:MOR

[1493] Roland Pulch. Model order reduction and low-dimensional representations for random linear dynamical systems. Mathematics and Computers in Simulation, 144:1– 20, February 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475417302057.

Rajan:2018:RIS

[1494] M. P. Rajan and G. D. Reddy. A regularized iterative scheme for solving singularly perturbed elliptic PDE. Mathematics and Computers in Simulation, 144:21–34, February 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417302082.

Rafei:2018:TMC

[1495] Mehdi Rafei, Mir Hamid Reza Ghoreishy, and Ghasem Naderi. Thermomechanical coupled finite element simulation of tire cornering characteristics — Effect of complex material models and friction law. Mathematics and Computers in Simulation, 144:35–51, February 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417302094.

Yongzhen:2018:DMT

[1496] Pei Yongzhen, Ji Xuehui, Li Changguo, and Gao Shujing. Dynamics of a model of Toxoplasmosis disease in cat and human with varying size populations. Mathematics and Computers in Simulation, 144: 52–59, February 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417302550.

Dias:2018:PHS

[1497] Fabio Scalco Dias, Jaume Llibre, and Claudia Valls. Polynomial Hamiltonian systems of degree 3 with symmetric nilpotent centers. Mathematics and Computers in Simulation, 144:60–77, February 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417302306.

Liu:2018:TSS

[1498] Qun Liu, Daqing Jiang, Ningzhong Shi, Tasawar Hayat, and Ahmed Alsaedi. The threshold of a stochastic SIS epidemic model with imperfect vaccination. Mathematics and Computers in Simulation, 144:78– 90, February 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541730232X.

Sobehart:2018:HPR

[1499] Lionel Sobehart and Hiroyuki Harada. High performance rigid body simulation of modularized robots using constraint-based models. Mathematics and Computers in Simulation, 144: 91–107, February 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417302744.

Gomez-Deniz:2018:APE

[1500] E. Gómez-Déniz. Adding a parameter to the exponential and Weibull distributions with applications. Mathematics and Computers in Simulation, 144:108–119, February 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417302756.

Dubey:2018:MRA

[1501] Preeti Dubey, Uma S. Dubey, and Balram Dubey. Modeling the role of acquired immune response and antiretroviral therapy in the dynamics of HIV infection. Mathematics and Computers in Simulation, 144: 120–137, February 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541730277X.

Belhaj:2018:BID

[1502] Skander Belhaj, Haithem Ben Kahla, Marwa Dridi, and Maher Moakher. Blind image deconvolution via Hankel based method for computing the GCD of polynomials. Mathematics and Computers in Simulation, 144: 138–152, February 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417302793.

Anonymous:2018:EBc

[1503] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 145:ifc, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417303580.

Blaheta:2018:E

[1504] Radim Blaheta, Jiří Starý, Zdeněk Dostál, and Tomáš Kozubek. Editorial. Mathematics and Computers in Simulation, 145:1–2, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541730349X.

Bastl:2018:IAT

[1505] Bohumír Bastl, Marek Brandner, Jiří Egermaier, Kristýna Michálková, and Eva Turnerová. Isogeometric analysis for turbulent flow. Mathematics and Computers in Simulation, 145: 3–17, March 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475416301161.

Panagiotopoulos:2018:TAC

[1506] Christos G. Panagiotopoulos, Vladislav Mantič, and Tomáš Roubíček. Two adhesive-contact models for quasistatic mixed-mode delamination problems. Mathematics and Computers in Simulation, 145:18–33, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475416301975.

Magoules:2018:AIS

[1507] Frédéric Magoulès and Cédric Venet. Asynchronous iterative sub-structuring methods. Mathematics and Computers in Simulation, 145:34–49, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S037847541630115X.

Kraus:2018:IFL

[1508] Johannes Kraus and Maria Lymbery. Incomplete factorization by local exact factorization (ILUE). Mathematics and Computers in Simulation, 145:50–61, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417303452.

Haslinger:2018:NCP

[1509] Jaroslav Haslinger, Vladimír Janovský, Radek Kučera, and Kristina Motyčková. Nonsmooth continuation of parameter dependent static contact problems with Coulomb friction. Mathematics and Computers in Simulation, 145:62–78, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417302975.

Pospisil:2018:PBB

[1510] Lukáš Pospíšil and Zdeněk Dostál. The projected Barzilai–Borwein method with fall-back for strictly convex QCQP problems with separable constraints. Mathematics and Computers in Simulation, 145:79–89, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475417303415.

Kruzik:2018:CMM

[1511] Martin Kružík and Jan Valdman. Computational modeling of magnetic hysteresis with thermal effects. Mathematics and Computers in Simulation, 145:90–105, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300800.

Merta:2018:PLB

[1512] M. Merta and J. Zapletal. A parallel library for boundary element discretization of engineering problems. Mathematics and Computers in Simulation, 145:106–113, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301446.

Kucera:2018:EMS

[1513] Radek Kučera, Jaroslav Haslinger, Václav Šátek, and Marta Jarošová. Efficient methods for solving the Stokes problem with slip boundary conditions. Mathematics and Computers in Simulation, 145:114– 124, March 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475416301215.

Cimrman:2018:IAE

[1514] Robert Cimrman, Matyáš Novák, Radek Kolman, Miroslav Tůma, and Jiří Vackář. Isogeometric analysis in electronic structure calculations. Mathematics and Computers in Simulation, 145:125–135, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301173.

Dolgov:2018:DTP

[1515] Sergey V. Dolgov, Vladimir A. Kazeev, and Boris N. Khoromskij. Direct tensor-product solution of one-dimensional elliptic equations with parameter-dependent coefficients. Mathematics and Computers in Simulation, 145:136–155, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475417303476.

Repin:2018:LFL

[1516] S. Repin. Localized forms of the LBB condition and a posteriori estimates for incompressible media problems. Mathematics and Computers in Simulation, 145:156–170, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416300817.

Schaarschmidt:2018:PSR

[1517] Ute Schaarschmidt, Trond Steihaug, and Sam Subbey. A parametrized stock-recruitment relationship derived from a slow-fast population dynamic model. Mathematics and Computers in Simulation, 145:171– 185, March 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417303464.

Langer:2018:NSF

[1518] Ulrich Langer and Huidong Yang. Numerical simulation of fluid–structure interaction problems with hyperelastic models: a monolithic approach. Mathematics and Computers in Simulation, 145:186–208, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301185.

Anonymous:2018:NIb

[1519] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 145:209, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417303658.

Anonymous:2018:ICEb

[1520] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 145:210–212, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S037847541730366X.

Anonymous:2018:PAa

[1521] Anonymous. Pages 1–266 (April 2018). Mathematics and Computers in Simulation, 146:1–266, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2018:EBd

[1522] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 146:ifc, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417303816.

Morel:2018:GFM

[1523] C. Morel, J.-Y. Morel, and M. F. Danca. Generalization of the Filippov method for systems with a large periodic input. Mathematics and Computers in Simulation, 146: 1–13, April 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475417303233.

Zhang:2018:FTS

[1524] Yu Zhang, Zhi Guo Feng, Xinsong Yang, Fuad E. Alsaadi, and Bashir Ahmad. Finite-time stabilization for a class of nonlinear systems via optimal control. Mathematics and Computers in Simulation, 146:14–26, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417303221.

Reisen:2018:REF

[1525] Valdério Anselmo Reisen, Edson Zambon Monte, Glaura da Conceição Franco, Adriano Marcio Sgrancio, Fábio Alexander Fajardo Molinares, Pascal Bondon, Flávio Augusto Ziegelmann, and Bovas Abraham. Robust estimation of fractional seasonal processes: Modeling and forecasting daily average SO _{2} concentrations. Mathematics and Computers in Simulation, 146:27–43, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417303427.

Chen:2018:NHS

[1526] Lijuan Chen, Song Tang, Qingdu Li, and Shouming Zhong. A new 4D hyperchaotic system with high complexity. Mathematics and Computers in Simulation, 146:44–56, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417303403.

Qu:2018:FMS

[1527] Wenzhen Qu, Wen Chen, Zhuojia Fu, and Yan Gu. Fast multipole singular boundary method for Stokes flow problems. Mathematics and Computers in Simulation, 146: 57–69, April 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417303397.

Siriprapaiwan:2018:GRN

[1528] Supatcha Siriprapaiwan, Elvin J. Moore, and Sanoe Koonprasert. Generalized reproduction numbers, sensitivity analysis and critical immunity levels of an SEQIJR disease model with immunization and varying total population size. Mathematics and Computers in Simulation, 146:70–89, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417303440.

Yamashita:2018:TWS

[1529] William M. S. Yamashita, Lucy T. Takahashi, and Grigori Chapiro. Traveling wave solutions for the dispersive models describing population dynamics of aedes aegypti. Mathematics and Computers in Simulation, 146:90–99, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417303506.

Guin:2018:CAB

[1530] Lakshmi Narayan Guin and Hunki Baek. Comparative analysis between prey-dependent and ratio-dependent predator–prey systems relating to patterning phenomenon. Mathematics and Computers in Simulation, 146: 100–117, April 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475417303531.

Luo:2018:PBR

[1531] Zhendong Luo, Fei Teng, and Jing Chen. A POD-based reduced-order Crank–Nicolson finite volume element extrapolating algorithm for 2D Sobolev equations. Mathematics and Computers in Simulation, 146:118– 133, April 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417303567.

Fonseca:2018:IBB

[1532] Rodney V. Fonseca and Francisco Cribari-Neto. Inference in a bimodal Birnbaum–Saunders model. Mathematics and Computers in Simulation, 146:134–159, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417303725.

Soba:2018:IAP

[1533] Alejandro Soba, Cecilia Suárez, Maraelys Morales González, Luis Enrique Bergues Cabrales, Ana Elisa Bergues Pupo, Juan Bory Reyes, and José Pablo Martínez Tassé. Integrated analysis of the potential, electric field, temperature, pH and tissue damage generated by different electrode arrays in a tumor under electrochemical treatment. Mathematics and Computers in Simulation, 146:160–176, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417303749.

Torabi:2018:EPM

[1534] Hamzeh Torabi, F. L. Bagheri, and E. Mahmoudi. Estimation of parameters for the Marshall–Olkin generalized exponential distribution based on complete data. Mathematics and Computers in Simulation, 146:177– 185, April 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475417303737.

Ano-Villalba:2018:WPP

[1535] S. Añó-Villalba, R. Blasco-Gimenez, S. Bernal-Perez, and E. Belenguer. Wind power plant integration in voltage source converter HVdc grids with voltage droop control. Mathematics and Computers in Simulation, 146:186–199, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541730006X.

Aparicio:2018:AUF

[1536] Néstor Aparicio, Salvador Añó-Villalba, Enrique Belenguer, and Ramon Blasco-Gimenez. Automatic under-frequency load shedding maloperation in power systems with high wind power penetration. Mathematics and Computers in Simulation, 146:200–209, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417300058.

Zaibi:2018:SPM

[1537] Malek Zaibi, Gérard Champenois, Xavier Roboam, Jamel Belhadj, and Bruno Sareni. Smart power management of a hybrid photovoltaic/wind stand-alone system coupling battery storage and hydraulic network. Mathematics and Computers in Simulation, 146:210–228, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475416301768.

Zhou:2018:MAS

[1538] Kang Zhou, Yufang Huang, Zhen Cheng, and Yafei Dong. Making all spanning tree problem based on sticker model. Mathematics and Computers in Simulation, 146:229– 239, April 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417300885.

Rasulov:2018:MCM

[1539] Abdujabar Rasulov and Gulnora Raimova. Monte Carlo method for solution of initial–boundary value problem for nonlinear parabolic equations. Mathematics and Computers in Simulation, 146:240–250, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475417301052.

Abad:2018:ESC

[1540] A. Abad and A. Elipe. Evolution strategies for computing periodic orbits. Mathematics and Computers in Simulation, 146:251– 261, April 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475414002018.

Anonymous:2018:NIc

[1541] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 146:262, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417303865.

Anonymous:2018:ICEc

[1542] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 146:263–265, April 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475417303877.

Anonymous:2018:EBe

[1543] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 147:ifc, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418300107.

Mansutti:2018:ASC

[1544] Daniela Mansutti and Rosa Maria Spitaleri. On Applied Scientific Computing XIV for Challenging Applications. Mathematics and Computers in Simulation, 147:1, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418300016.

Belokrys-Fedotov:2018:DMI

[1545] A. I. Belokrys-Fedotov, V. A. Garanzha, and L. N. Kudryavtseva. Delaunay meshing of implicit domains with boundary edge sharpening and sliver elimination. Mathematics and Computers in Simulation, 147: 2–26, May 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417302781.

Brakhage:2018:AID

[1546] Karl-Heinz Brakhage. Analytical investigations for the design of fast approximation methods for fitting curves and surfaces to scattered data. Mathematics and Computers in Simulation, 147:27–39, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417303993.

Carillo:2018:ROA

[1547] Sandra Carillo, Mauro Lo Schiavo, and Cornelia Schiebold. Recursion operators admitted by non-Abelian Burgers equations: Some remarks. Mathematics and Computers in Simulation, 147:40–51, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300551.

Cavoretto:2018:TAG

[1548] Roberto Cavoretto, Alessandra De Rossi, and Hanli Qiao. Topology analysis of global and local RBF transformations for image registration. Mathematics and Computers in Simulation, 147:52–72, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417303488.

Costabile:2018:MLB

[1549] F. A. Costabile, F. Di Tommaso, and E. Longo. A mixed Lagrange– Bernoulli tensor product expansion on the rectangle with applications. Mathematics and Computers in Simulation, 147:73–89, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417302525.

DeCanditiis:2018:DSG

[1550] D. De Canditiis, M. Pensky, and P. J. Wolfe. Denoising strategies for general finite frames. Mathematics and Computers in Simulation, 147:90–99, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300599.

Dellino:2018:MMF

[1551] G. Dellino, T. Laudadio, R. Mari, N. Mastronardi, and C. Meloni. Microforecasting methods for fresh food supply chain management: a computational study. Mathematics and Computers in Simulation, 147:100– 120, May 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417303981.

Festa:2018:DDB

[1552] Adriano Festa. Domain decomposition based parallel Howard’s algorithm. Mathematics and Computers in Simulation, 147:121–139, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475417301684.

Kervarc:2018:ACI

[1553] Romain Kervarc, Ariane Piel, Stéphanie Lala, and Jean Bourrely. Assessing component impairing at mission level. Mathematics and Computers in Simulation, 147:140–151, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475417301921.

Noriega:2018:CSOa

[1554] H. Noriega, F. Guibault, M. Reggio, and R. Magnan. A case-study in opensource CFD code verification, Part I: Convergence rate loss diagnosis. Mathematics and Computers in Simulation, 147:152–171, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417303774.

Noriega:2018:CSOb

[1555] H. Noriega, F. Guibault, M. Reggio, and R. Magnan. A case-study in open-source CFD code verification. Part II: Boundary condition non-orthogonal correction. Mathematics and Computers in Simulation, 147:172–193, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417303762.

Novara:2018:IPT

[1556] Paola Novara and Lucia Romani. On the interpolating 5-point ternary subdivision scheme: a revised proof of convexity-preservation and an application-oriented extension. Mathematics and Computers in Simulation, 147:194–209, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541630180X.

Pezza:2018:MCM

[1557] L. Pezza and F. Pitolli. A multiscale collocation method for fractional differential problems. Mathematics and Computers in Simulation, 147:210–219, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417302768.

Rom:2018:MMC

[1558] M. Rom, K.-H. Brakhage, S. Barth, C. Wrobel, P. Mattfeld, and F. Klocke. Mathematical modeling of ceramic bond bridges in grinding wheels. Mathematics and Computers in Simulation, 147:220–236, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417300563.

Rossini:2018:DSM

[1559] Milvia Rossini and Elena Volontè. On directional scaling matrices in dimension d = 2. Mathematics and Computers in Simulation, 147: 237–249, May 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417301878.

Sahu:2018:HOF

[1560] Smita Sahu. High-order filtered schemes for first order time dependent linear and non-linear partial differential equations. Mathematics and Computers in Simulation, 147: 250–263, May 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417302070.

Santana:2018:PMS

[1561] José M. Santana, Agustín Trujillo, and José P. Suárez. A physical model for screen space distribution of 3D marks on geographical information systems. Mathematics and Computers in Simulation, 147:264–278, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417302318.

Themistoclakis:2018:WAD

[1562] Woula Themistoclakis. Weighted L^{1} approximation on [−1, 1] via discrete de la Vallée Poussin means. Mathematics and Computers in Simulation, 147:279–292, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417302331.

Tinoco-Guerrero:2018:SAG

[1563] G. Tinoco-Guerrero, F. J. Domínguez-Mota, A. Gaona-Arias, M. L. Ruiz-Zavala, and J. G. Tinoco-Ruiz. A stability analysis for a generalized finite-difference scheme applied to the pure advection equation. Mathematics and Computers in Simulation, 147:293–300, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541730215X.

Tolstykh:2018:IAF

[1564] Andrei I. Tolstykh and Michael V. Lipavskii. Instability and acoustic fields of the Rankine vortex as seen from long-term calculations with the tenth-order multioperatorsbased scheme. Mathematics and Computers in Simulation, 147:301– 320, May 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475417303208.

Anonymous:2018:NId

[1565] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 147:321, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541830017X.

Anonymous:2018:EBf

[1566] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 147:322–325, May 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300181.

Anonymous:2018:PJa

[1567] Anonymous. Pages 1–98 (June 2018).

Mathematics and Computers in Simulation, 148:1–98, June 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2018:EBg

[1568] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 148:ifc, June 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300296.

Zhou:2018:RTA

[1569] Yonghua Zhou, Xin Tao, Lei Luan, and Jingjie Ning. Revisiting the 7/23 train accident using computer reconstruction simulation for causation and prevention analysis. Mathematics and Computers in Simulation, 148:1–15, June 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300053.

Sadaf:2018:IMF

[1570] Hina Sadaf, Muhammad Usman Akbar, and S. Nadeem. Induced magnetic field analysis for the peristaltic transport of non-Newtonian nanofluid in an annulus. Mathematics and Computers in Simulation, 148:16–36, June 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300028.

Zheng:2018:SBP

[1571] Guang-Hui Zheng and Quan-Guo Zhang. Solving the backward problem for space-fractional diffusion equation by a fractional Tikhonov regularization method. Mathematics and Computers in Simulation, 148: 37–47, June 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541730397X.

Lastra:2018:EMS

[1572] Miguel Lastra, Manuel J. Castro Díaz, Carlos Ureña, and Marc de la Asunción. Efficient multilayer shallow-water simulation system based on GPUs. Mathematics and Computers in Simulation, 148: 48–65, June 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417303786.

Amat:2018:CTP

[1573] S. Amat, J. Liandrat, M. Moncayo, J. Ruiz, and J. C. Trillo. On a class of three points cell-average multiresolution schemes. Mathematics and Computers in Simulation, 148: 66–93, June 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417303750.

Anonymous:2018:NIe

[1574] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 148:94, June 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541830034X.

Anonymous:2018:ICEd

[1575] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 148:95–98, June 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475418300351.

Anonymous:2018:PJb

[1576] Anonymous. Pages 1–128 (July 2018).

Mathematics and Computers in Simulation, 149:1–128, July 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2018:EBh

[1577] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 149:ifc, July 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541830051X.

Kim:2018:HBA

[1578] Kwang Su Kim, Sangil Kim, and Il Hyo Jung. Hopf bifurcation analysis and optimal control of Treatment in a delayed oncolytic virus dynamics. Mathematics and Computers in Simulation, 149:1–16, July 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300077.

Capone:2018:DIP

[1579] F. Capone, M. F. Carfora, R. De Luca, and I. Torcicollo. On the dynamics of an intraguild predator–prey model. Mathematics and Computers in Simulation, 149:17–31, July 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300089.

Valiollahi:2018:EPP

[1580] R. Valiollahi, Mohammad Z. Raqab, A. Asgharzadeh, and F. A. Alqallaf. Estimation and prediction for power Lindley distribution under progressively type II right censored samples. Mathematics and Computers in Simulation, 149:32–47, July 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300259.

Farahani:2018:HIA

[1581] Farzad Vasheghani Farahani, Abbas Ahmadi, and Mohammad Hossein Fazel Zarandi. Hybrid intelligent approach for diagnosis of the lung nodule from CT images using spatial kernelized fuzzy c-means and ensemble learning. Mathematics and Computers in Simulation, 149: 48–68, July 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418300272.

Xu:2018:LGH

[1582] Changjin Xu. Local and global Hopf bifurcation analysis on simplified bidirectional associative memory neural networks with multiple delays. Mathematics and Computers in Simulation, 149:69–90, July 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300405.

El-Shimy:2018:ILM

[1583] M. El-Shimy, N. Mostafa, A. N.

Afandi, A. M. Sharaf, and Mahmoud A. Attia. Impact of load models on the static and dynamic performances of grid-connected wind power plants: a comparative analysis. Mathematics and Computers in Simulation, 149:91–108, July 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300417.

Alvarez-Vazquez:2018:SMC

[1584] L. J. Alvarez-Vázquez, A. Martínez, C. Rodríguez, and M. E. Vázquez-Méndez. Sediment minimization in canals: an optimal control approach. Mathematics and Computers in Simulation, 149:109–122, July 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475418300478.

Anonymous:2018:NIf

[1585] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 149:123, July 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418300569.

Anonymous:2018:ICEe

[1586] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 149:124–127, July 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475418300570.

Anonymous:2018:PAb

[1587] Anonymous. Pages 1–144 (August 2018). Mathematics and Computers in Simulation, 150:1–144, August 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2018:EBi

[1588] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 150:ifc, August 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418300764.

Ai:2018:MCP

[1589] Weilong Ai and Charles E. Augarde. A multi-cracked particle method for complex fracture problems in 2D. Mathematics and Computers in Simulation, 150:1–24, August 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418300454.

Milosevic:2018:CAS

[1590] Marija Milošević. Convergence and almost sure polynomial stability of the backward and forward–backward Euler methods for highly nonlinear pantograph stochastic differential equations. Mathematics and Computers in Simulation, 150:25– 48, August 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418300466.

Deolmi:2018:TSM

[1591] G. Deolmi and S. Müller. A twostep model order reduction method to simulate a compressible flow over an extended rough surface. Mathematics and Computers in Simulation, 150:49–65, August 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541830048X.

Liu:2018:CDB

[1592] Yan Liu, Shengzhong Xiao, and Yiwu Lin. Continuous dependence for the Brinkman–Forchheimer fluid interfacing with a Darcy fluid in a bounded domain. Mathematics and Computers in Simulation, 150:66– 82, August 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418300491.

Bartelt:2018:EIE

[1593] M. Bartelt, J. Dietzsch, and M. Groß.

Efficient implementation of energy conservation for higher order finite elements with variational integrators. Mathematics and Computers in Simulation, 150:83–121, August 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475418300636.

Villarreal-Cervantes:2018:DEB

[1594] Miguel G. Villarreal-Cervantes, Efrén Mezura-Montes, and José Yaír Guzmán-Gaspar. Differential evolution based adaptation for the direct current motor velocity control parameters. Mathematics and Computers in Simulation, 150:122–141, August 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418300685.

Anonymous:2018:NIg

[1595] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 150:142, August 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418300806.

Anonymous:2018:ICEf

[1596] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 150:143–144, August 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475418300818.

Anonymous:2018:PS

[1597] Anonymous. Pages 1–172 (September 2018). Mathematics and Computers in Simulation, 151:1–172, September 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2018:EBj

[1598] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 151:ifc, September 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301009.

Xu:2018:MPD

[1599] Dinghua Xu, Yangao He, Yue Yu, and Qifeng Zhang. Multiple parameter determination in textile material design: a Bayesian inference approach based on simulation. Mathematics and Computers in Simulation, 151:1–14, September 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300879.

Akbari:2018:SWM

[1600] Hani Akbari and Allan P. Engsig-Karup. Screening wells by multi-scale grids for multi-stage Markov Chain Monte Carlo simulation. Mathematics and Computers in Simulation, 151: 15–28, September 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300867.

Rodriguez-Padilla:2018:NSE

[1601] Jairo Rodríguez-Padilla and Daniel Olmos-Liceaga. Numerical solutions of equations of cardiac wave propagation based on Chebyshev multidomain pseudospectral methods. Mathematics and Computers in Simulation, 151:29–53, September 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300740.

Candido:2018:ZHB

[1602] Murilo R. Cândido and Jaume Llibre. Zero-Hopf bifurcations in 3dimensional differential systems with no equilibria. Mathematics and Computers in Simulation, 151:54– 76, September 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418300697.

Wei:2018:IDC

[1603] T. Wei and Y. S. Li. Identifying a diffusion coefficient in a timefractional diffusion equation. Mathematics and Computers in Simulation, 151:77–95, September 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300673.

Cabrales:2018:VMG

[1604] Luis Enrique Bergues Cabrales, Juan I. Montijano, Maria Schonbek, and Antonio Rafael Selva Castañeda. A viscous modified Gompertz model for the analysis of the kinetics of tumors under electrochemical therapy. Mathematics and Computers in Simulation, 151:96–110, September 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300661.

Mohamed:2018:AST

[1605] Guermouche Mohamed, Ahmed Ali Sofiane, and Langlois Nicolas. Adaptive super twisting extended state observer based sliding mode control for diesel engine air path subject to matched and unmatched disturbance. Mathematics and Computers in Simulation, 151:111–130, September 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541830065X.

Wang:2018:IND

[1606] Yanfeng Wang, Yongpeng Shen, Xuncai Zhang, Guangzhao Cui, and Junwei Sun. An Improved Nondominated Sorting Genetic Algorithm-II (INSGA-II) applied to the design of DNA codewords. Mathematics and Computers in Simulation, 151: 131–139, September 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300727.

Li:2018:DCB

[1607] Fei Li, Jingming Liu, and Zheng Li. DNA computation based on selfassembled nanoparticle probes for 0–1 integer programming problem. Mathematics and Computers in Simulation, 151:140–146, September 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417300587.

Yin:2018:MBC

[1608] Zhixiang Yin, Jianzhong Cui, and Chen Zhen. Molecular beacon computing model for maximum weight clique problem. Mathematics and Computers in Simulation, 151:147– 155, September 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475417300575.

Siam:2018:MMH

[1609] F. M. Siam, M. Grinfeld, A. Bahar, H. A. Rahman, H. Ahmad, and F. Johar. A mechanistic model of high dose irradiation damage. Mathematics and Computers in Simulation, 151: 156–168, September 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475416000562.

Anonymous:2018:NIh

[1610] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 151:169, September 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418301046.

Anonymous:2018:ICEg

[1611] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 151:170–171, September 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475418301058.

Anonymous:2018:PO

[1612] Anonymous. Pages 1–100 (October 2018). Mathematics and Computers in Simulation, 152:1–100, October 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2018:EBk

[1613] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 152:ifc, October 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301289.

Juncu:2018:SMN

[1614] Gheorghe Juncu, Aurelian Nicola, and Constantin Popa. Splitting methods for the numerical solution of multi-component mass transfer problems. Mathematics and Computers in Simulation, 152:1– 14, October 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301137.

Li:2018:SBV

[1615] Jinhui Li, Zhidong Teng, and Long Zhang. Stability and bifurcation in a vector-bias model of malaria transmission with delay. Mathematics and Computers in Simulation, 152:15–34, October 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300958.

Yang:2018:CBP

[1616] Xiaofeng Yang, Jinping Yu, Mengna Xu, and Wenjing Fan. Convertible bond pricing with partial integrodifferential equation model. Mathematics and Computers in Simulation, 152:35–50, October 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300910.

Afshari:2018:NMO

[1617] Faraz Afshari, Hadi Ghasemi Zavaragh, Bayram Sahin, Roberta Cocci Grifoni, Francesco Corvaro, Barbara Marchetti, and Fabio Polonara. On numerical methods; optimization of CFD solution to evaluate fluid flow around a sample object at low Re numbers. Mathematics and Computers in Simulation, 152:51–68, October 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300909.

Zhao:2018:ERD

[1618] Zhong Zhao, Qiuying Li, and Lansun Chen. Effect of rhizosphere dispersal and impulsive input on the growth of wetland plant. Mathematics and Computers in Simulation, 152:69–80, October 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300892.

Mefteh:2018:SBD

[1619] Wafa Mefteh. Simulation-based Design: Overview about related works. Mathematics and Computers in Simulation, 152:81–97, October 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418300739.

Anonymous:2018:EBl

[1620] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 152:98, October 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301332.

Anonymous:2018:EBm

[1621] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 152:99–100, October 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301344.

Anonymous:2018:PN

[1622] Anonymous. Pages 1–98 (November 2018). Mathematics and Computers in Simulation, 153:1–98, November 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2018:EBn

[1623] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 153:ifc, November 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301551.

Hao:2018:HBA

[1624] Pengmiao Hao, Xuechen Wang, and Junjie Wei. Hopf bifurcation analysis of a diffusive single species model with stage structure and strong Allee effect. Mathematics and Computers in Simulation, 153:1– 14, November 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301162.

Coudiere:2018:EAB

[1625] Yves Coudière, Charlie Douanla-Lontsi, and Charles Pierre. Exponential Adams–Bashforth integrators for stiff ODEs, application to cardiac electrophysiology. Mathematics and Computers in Simulation, 153: 15–34, November 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300922.

Paipuri:2018:CHO

[1626] Mahendra Paipuri, Sonia Fernández-Méndez, and Carlos Tiago. Comparison of high-order continuous and hybridizable discontinuous Galerkin methods for incompressible fluid flow problems. Mathematics and Computers in Simulation, 153:35– 58, November 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301241.

Abdulla:2018:EIN

[1627] Ugur G. Abdulla, Jian Du, Adam Prinkey, Chloe Ondracek, and Suneil Parimoo. Evolution of interfaces for the nonlinear double degenerate parabolic equation of turbulent filtration with absorption. Mathematics and Computers in Simulation, 153: 59–82, November 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301411.

Ilka:2018:CTR

[1628] Reza Ilka, Yousef Alinejad-Beromi, and Hamid Yaghobi. Cogging torque reduction of permanent magnet synchronous motor using multiobjective optimization. Mathematics and Computers in Simulation, 153: 83–95, November 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301423.

Anonymous:2018:NIi

[1629] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 153:96, November 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301605.

Anonymous:2018:ICEh

[1630] Anonymous. IMACS calender of events. Mathematics and Computers in Simulation, 153:97–98, November 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301617.

Anonymous:2018:PD

[1631] Anonymous. Pages 1–96 (December 2018). Mathematics and Computers in Simulation, 154:1–96, December 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2018:EBo

[1632] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 154:ifc, December 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418301770.

Leye:2018:NSS

[1633] Babacar Lèye, Jonas Koko, Soulèye Kane, and Mamadou Sy. Numerical simulation of saltwater intrusion in coastal aquifers with anisotropic mesh adaptation. Mathematics and Computers in Simulation, 154:1– 18, December 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475418301538.

Gao:2018:UBE

[1634] Wei Gao, Li Yan, Mohammadhossein Saeedi, and Hassan Saberi Nik. Ultimate bound estimation set and chaos synchronization for a financial risk system. Mathematics and Computers in Simulation, 154:19– 33, December 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475418301514.

Mokhtari:2018:EOQ

[1635] Hadi Mokhtari. Economic order quantity for joint complementary and substitutable items. Mathematics and Computers in Simulation, 154:34– 47, December 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301496.

Kunsch:2018:MCP

[1636] Pierre L. Kunsch and Alessio Ishizaka.

Multiple-criteria performance ranking based on profile distributions: an application to university research evaluations. Mathematics and Computers in Simulation, 154:48– 64, December 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301459.

Silva:2018:PIC

[1637] Alisson R. Silva, Maury M. Gouvêa, Luís F. W. Góes, and Carlos A. P. S. Martins. A parallel implementation of a cloud dynamics model with cellular automaton. Mathematics and Computers in Simulation, 154:65–93, December 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301447.

Anonymous:2018:NIj

[1638] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 154:94, December 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418301824.

Anonymous:2018:ICEi

[1639] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 154:95–96, December 2018. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475418301836.

Anonymous:2019:EBa

[1640] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 155:ifc, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418302210.

Cao:2019:P

[1641] Jinde Cao, Raja Ramachandran, and Quanxin Zhu. Preface. Mathematics and Computers in Simulation, 155:1, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541830212X.

Socha:2019:ESR

[1642] Leslaw Socha and Quanxin Zhu. Exponential stability with respect to part of the variables for a class of nonlinear stochastic systems with Markovian switchings. Mathematics and Computers in Simulation, 155:2–14, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541730321X.

Vadivoo:2019:NID

[1643] B. Sundara Vadivoo, R. Raja, R. Aly Seadawy, and G. Rajchakit. Nonlinear integro-differential equations with small unknown parameters: a controllability analysis problem. Mathematics and Computers in Simulation, 155:15–26, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417303439.

Guo:2019:ELS

[1644] Jianhua Guo, Yu Liu, Xiugang Li, Wei Huang, Jinde Cao, and Yun Wei. Enhanced least square based dynamic OD matrix estimation using Radio Frequency Identification data. Mathematics and Computers in Simulation, 155:27–40, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541730352X.

Pradeep:2019:ETS

[1645] C. Pradeep, Yang Cao, R. Murugesu, and R. Rakkiyappan. An event-triggered synchronization of semi-Markov jump neural networks with time-varying delays based on generalized free-weighting-matrix approach. Mathematics and Computers in Simulation, 155:41–56, January 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475417303555.

Cao:2019:RPA

[1646] Yang Cao, R. Samidurai, and R. Sriraman. Robust passivity analysis for uncertain neural networks with leakage delay and additive timevarying delays by using general activation function. Mathematics and Computers in Simulation, 155:57– 77, January 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475417303543.

Bi:2019:BTP

[1647] Xia an Bi, Xianhao Luo, and Qi Sun. Branch tire packet classification algorithm based on singlelinkage clustering. Mathematics and Computers in Simulation, 155:78– 91, January 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475417303713.

Guo:2019:MGA

[1648] Jianhua Guo, Ye Kong, Zongzhi Li, Wei Huang, Jinde Cao, and Yun Wei. A model and genetic algorithm for areawide intersection signal optimization under user equilibrium traffic. Mathematics and Computers in Simulation, 155:92–104, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475417303798.

Lu:2019:NTS

[1649] Ke Lu, Pingping Du, Jinde Cao, Qiming Zou, Tianjia He, and Wei Huang. A novel traffic signal split approach based on Explicit Model Predictive Control. Mathematics and Computers in Simulation, 155:105– 114, January 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475417303920.

Zhu:2019:TPD

[1650] Min Zhu and Yong Xu. A timeperiodic dengue fever model in a heterogeneous environment. Mathematics and Computers in Simulation, 155: 115–129, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475417304007.

Wei:2019:MVD

[1651] Yun Wei, Qing Tian, Jianhua Guo, Wei Huang, and Jinde Cao. Multi-vehicle detection algorithm through combining Harr and HOG features. Mathematics and Computers in Simulation, 155:130–145, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300041.

Hu:2019:MDL

[1652] Wei Hu, Quanxin Zhu, and Yi Lu.

Moments and distributions of the last exit times for a class of Markov processes. Mathematics and Computers in Simulation, 155:146– 153, January 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541830003X.

Moonam:2019:UDM

[1653] Hasan M. Moonam, Xiao Qin, and Jun Zhang. Utilizing data mining techniques to predict expected freeway travel time from experienced travel time. Mathematics and Computers in Simulation, 155:154– 167, January 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418300260.

Ma:2019:OMD

[1654] Yong-Ki Ma, P. Prakash, and A. Deiveegan. Optimization method for determining the source term in fractional diffusion equation. Mathematics and Computers in Simulation, 155:168–176, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300648.

Jeganathan:2019:TSM

[1655] K. Jeganathan, M. Abdul Reiyas, K. Prasanna Lakshmi, and S. Saravanan. Two server Markovian inventory systems with server interruptions: Heterogeneous vs. homogeneous servers. Mathematics and Computers in Simulation, 155:177– 200, January 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418300624.

Samidurai:2019:RDA

[1656] R. Samidurai and R. Sriraman. Robust dissipativity analysis for uncertain neural networks with additive time-varying delays and general activation functions. Mathematics and Computers in Simulation, 155: 201–216, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300715.

Kumar:2019:DAA

[1657] S. Vimal Kumar, S. Marshal Anthoni, and R. Raja. Dissipative analysis for aircraft flight control systems with randomly occurring uncertainties via nonfragile sampled-data control. Mathematics and Computers in Simulation, 155:217–226, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300703.

Zheng:2019:OCL

[1658] Zhi xin Zheng, Jun qing Li, and Pei yong Duan. Optimal chiller loading by improved artificial fish swarm algorithm for energy saving. Mathematics and Computers in Simulation, 155:227–243, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301083.

Yang:2019:RPI

[1659] Xueyan Yang, Dongxue Peng, Xiaoxiao Lv, and Xiaodi Li. Recent progress in impulsive control systems. Mathematics and Computers in Simulation, 155:244–268, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418301150.

Rajupillai:2019:PRF

[1660] K. Rajupillai and S. Palaniammal. α-Phase retrieval frame in Hilbert space and its application. Mathematics and Computers in Simulation, 155:269–276, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418301174.

Nie:2019:APP

[1661] Rencan Nie, Jinde Cao, Dongming Zhou, and Wenhua Qian. Analysis of pulse period for passive neuron in pulse coupled neural network. Mathematics and Computers in Simulation, 155:277–289, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418301216.

Balasundaram:2019:IEC

[1662] K. Balasundaram, R. Raja, A. Pratap, and S. Chandrasekaran. Impulsive effects on competitive neural networks with mixed delays: Existence and exponential stability analysis. Mathematics and Computers in Simulation, 155:290–302, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301204.

Cheng:2019:SIB

[1663] Huimei Cheng, Huan Huang, and Kai Zhang. Study on the influence of bus front-end intrusion-free distance to the bus moving characteristics. Mathematics and Computers in Simulation, 155:303–313, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301435.

Li:2019:ISS

[1664] Peng Li and Xiaodi Li. Input-to-state stability of nonlinear impulsive systems via Lyapunov method involving indefinite derivative. Mathematics and Computers in Simulation, 155: 314–323, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301678.

Rajupillai:2019:FRN

[1665] K. Rajupillai and S. Palaniammal.

Frame Reconstruction with Noise Reduction in Hilbert space and Application in Communication Systems. Mathematics and Computers in Simulation, 155:324–334, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541830171X.

Luo:2019:LSU

[1666] Yiping Luo, Fei Deng, Zhaomin Ling, and Zifeng Cheng. Local H_{∞} synchronization of uncertain complex networks via non-fragile state feedback control. Mathematics and Computers in Simulation, 155:335– 346, January 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301939.

Durga:2019:EEB

[1667] N. Durga and P. Muthukumar. Existence and exponential behavior of multi-valued nonlinear fractional stochastic integro-differential equations with Poisson jumps of Clarke’s subdifferential type. Mathematics and Computers in Simulation, 155:347– 359, January 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301964.

Bagdasar:2019:TAM

[1668] Ovidiu Bagdasar, Stuart Berry, Sam O’Neill, Nicolae Popovici, and Ramachandran Raja. Traffic assignment: Methods and simulations for an alternative formulation of the fixed demand problem. Mathematics and Computers in Simulation, 155:360– 373, January 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418302027.

Anonymous:2019:EBb

[1669] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 155:374, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302283.

Anonymous:2019:EBc

[1670] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 155:375–376, January 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302295.

Anonymous:2019:PF

[1671] Anonymous. Pages 1–398 (February 2019). Mathematics and Computers in Simulation, 156:1–398, February 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2019:EBd

[1672] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 156:ifc, February 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302568.

Fekry:2019:CPA

[1673] M. Fekry, Abdelfatah M. Mohamed, Mohamed Fanni, and S. Yoshida. A Comprehensive performance assessment of the integration of magnetic bearings with horizontal axis wind turbine. Mathematics and Computers in Simulation, 156:1– 39, February 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541830168X.

Yoshioka:2019:FDS

[1674] Hidekazu Yoshioka and Yuta Yaegashi. A finite difference scheme for variational inequalities arising in stochastic control problems with several singular control variables. Mathematics and Computers in Simulation, 156:40–66, February 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301708.

Ferreiro:2019:PTP

[1675] Ana M. Ferreiro, José Antonio García-Rodríguez, Carlos Vázquez, E. Costa e Silva, and A. Correia. Parallel two-phase methods for global optimization on GPU. Mathematics and Computers in Simulation, 156: 67–90, February 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301502.

Barbiero:2019:BCM

[1676] A. Barbiero. A bivariate count model with discrete Weibull margins. Mathematics and Computers in Simulation, 156:91–109, February 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301757.

Jafari:2019:SOS

[1677] Maryam Jafari and Saleh Mobayen.

Second-order sliding set design for a class of uncertain nonlinear systems with disturbances: an LMI approach. Mathematics and Computers in Simulation, 156:110–125, February 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301733.

Yang:2019:PCM

[1678] Jin Yang and Guangyao Tang. Piecewise chemostat model with control strategy. Mathematics and Computers in Simulation, 156:126– 142, February 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301885.

Wang:2019:ESI

[1679] Zengyun Wang and Xinzhi Liu.

Exponential stability of impulsive complex-valued neural networks with time delay. Mathematics and Computers in Simulation, 156:143–157, February 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301903.

Ezzatneshan:2019:CSL

[1680] Eslam Ezzatneshan. Comparative study of the lattice Boltzmann collision models for simulation of incompressible fluid flows. Mathematics and Computers in Simulation, 156:

158–177, February 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301976.

Zhang:2019:RKM

[1681] Bolong Zhang, Wenjian Yu, and Michael Mascagni. Revisiting Kac’s method: a Monte Carlo algorithm for solving the Telegrapher’s equations. Mathematics and Computers in Simulation, 156:178–193, February 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418302052.

Long:2019:DDD

[1682] Sarah D. Long, Somayyeh Sheikholeslami, James V. Lambers, and Carley Walker. Diagonalization of 1-D differential operators with piecewise constant coefficients using the uncertainty principle. Mathematics and Computers in Simulation, 156: 194–226, February 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302003.

Agahi:2019:MLG

[1683] Hamzeh Agahi and Mohsen Alipour. Mittag-Leffler-Gaussian distribution: Theory and application to real data. Mathematics and Computers in Simulation, 156:227–235, February 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418302015.

Boukhari:2019:TLF

[1684] Mohamed Riad Boukhari, Ahmed Chaibet, Moussa Boukhnifer, and Sébastien Glaser. Two longitudinal fault tolerant control architectures for an autonomous vehicle. Mathematics and Computers in Simulation, 156: 236–253, February 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541830199X.

Baiardi:2019:OMR

[1685] Lorenzo Cerboni Baiardi and Ahmad K. Naimzada. An oligopoly model with rational and imitation rules. Mathematics and Computers in Simulation, 156:254–278, February 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418302155.

Zaman:2019:ENC

[1686] Akbar Zaman, Nasir Ali, and Mazhar Sajjad. Effects of nanoparticles (Cu, TiO_{2}, Al_{2}O_{3}) on unsteady blood flow through a curved overlapping stenosed channel. Mathematics and Computers in Simulation, 156(??): 279–293, February 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302106.

Luan:2019:NMS

[1687] Jing Luan, Zhong Yao, Futao Zhao, and Xin Song. A novel method to solve supplier selection problem: Hybrid algorithm of genetic algorithm and ant colony optimization. Mathematics and Computers in Simulation, 156 (??):294–309, February 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541830209X.

Castillo:2019:IAL

[1688] Paul Castillo, Sergio Gómez, and Sergio Manzanarez. Improving the accuracy of LDG approximations on coarse meshes. Mathematics and Computers in Simulation, 156(??): 310–326, February 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418302088.

Rashed:2019:ADU

[1689] A. S. Rashed. Analysis of (3 + 1)- dimensional unsteady gas flow using optimal system of Lie symmetries. Mathematics and Computers in Simulation, 156(??):327– 346, February 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475418302064.

Tertrais:2019:PGD

[1690] H. Tertrais, R. Ibañez, A. Barasinski, Ch. Ghnatios, and F. Chinesta. On the Proper Generalized Decomposition applied to microwave processes involving multilayered components. Mathematics and Computers in Simulation, 156 (??):347–363, February 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302180.

Deepika:2019:RFT

[1691] Deepika Deepika, Shiv Narayan, and Sandeep Kaur. Robust finite time integral sliding mode tracker for nth-order non-affine non-linear system with uncertainty and disturbance estimator. Mathematics and Computers in Simulation, 156(??):364– 376, February 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418302167.

Pei:2019:MBF

[1692] Yongzhen Pei, Miaomiao Chen, Xiyin Liang, and Changguo Li. Model-based on fishery management systems with selective harvest policies. Mathematics and Computers in Simulation, 156 (??):377–395, February 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302076.

Anonymous:2019:NIa

[1693] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 156(??):396, February 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302611.

Anonymous:2019:ICEa

[1694] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 156(??):397– 398, February 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475418302623.

Anonymous:2019:PMa

[1695] Anonymous. Pages 1–162 (March 2019). Mathematics and Computers in Simulation, 157:1–162, March 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2019:EBe

[1696] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 157:ifc, March 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418302866.

Novoa-Munoz:2019:STB

[1697] Francisco Novoa-Muñoz. Simulation of the temperature of barley during its storage in cylindrical silos. Mathematics and Computers in Simulation, 157:1–14, March 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418302143.

Srivastav:2019:SZV

[1698] Akhil Kumar Srivastav, Junyuan Yang, XiaoFeng Luo, and Mini Ghosh. Spread of Zika virus disease on complex network — a mathematical study. Mathematics and Computers in Simulation, 157:15–38, March 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302465.

Rkhissi-Kammoun:2019:TCS

[1699] Yosra Rkhissi-Kammoun, Jawhar Ghommam, Moussa Boukhnifer, and Faisal Mnif. Two current sensor fault detection and isolation schemes for induction motor drives using algebraic estimation approach. Mathematics and Computers in Simulation, 157:39–62, March 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302428.

Amirabad:2019:PPP

[1700] H. Qolizadeh Amirabad, O. RabieiMotlagh, and H. M. MohammadiNejad. Permanency in predator–prey models of Leslie type with ratiodependent simplified Holling type-IV functional response. Mathematics and Computers in Simulation, 157:63–76, March 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302660.

Kumar:2019:NDI

[1701] Anuj Kumar, Prashant K. Srivastava, and R. P. Gupta. Nonlinear dynamics of infectious diseases via information-induced vaccination and saturated treatment. Mathematics and Computers in Simulation, 157:77–99, March 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418302696.

Bohaienko:2019:NSM

[1702] V. O. Bohaienko. Numerical schemes for modelling time-fractional dynamics of non-isothermal diffusion in soils. Mathematics and Computers in Simulation, 157:100–114, March 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475418302738.

Gomez:2019:GDH

[1703] Yolanda M. Gómez, Heleno Bolfarine, and Héctor W. Gómez. Gumbel distribution with heavy tails and applications to environmental data. Mathematics and Computers in Simulation, 157:115–129, March 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418302726.

Kolonko:2019:ISQ

[1704] Michael Kolonko, Feng Gu, and Zijun Wu. Improving the statistical quality of random number generators by applying a simple ratio transformation. Mathematics and Computers in Simulation, 157:130–142, March 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418302714.

Sayah:2019:ASM

[1705] Toni Sayah. Adaptive stratified Monte Carlo algorithm for numerical computation of integrals. Mathematics and Computers in Simulation, 157:143–158, March 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541830274X.

Anonymous:2019:NIb

[1706] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 157:159, March 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541830291X.

Anonymous:2019:ICEb

[1707] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 157:160–161, March 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475418302921.

Roboam:2019:P

[1708] Xavier Roboam, Eric Monmasson, Ramon Blasco-Gimenez, Benoit Robyns, Giovanni Spagnuolo, Nicolas Patin, and Sareni Bruno. Preface. Mathematics and Computers in Simulation, 158:1, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418303240.

Cabello:2019:DSC

[1709] Javier M. Cabello, Xavier Roboam, Sergio Junco, and Christophe Turpin. Direct sizing and characterization of Energy Storage Systems in the Energy-Power plane. Mathematics and Computers in Simulation, 158: 2–17, April 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418300880.

Goddet:2019:MAC

[1710] Etienne Goddet, Nicolas Retière, Vojislav Stojanović, Anca Dieudonné, Jérôme Genoulaz, and Jean-Michel Guichon. Maximizing the algebraic connectivity of meshed electrical pathways used as current return network. Mathematics and Computers in Simulation, 158:18–31, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301149.

Moussa:2019:VHD

[1711] Hassan Moussa, Jean-Philippe Martin, Serge Pierfederici, Farid Meibody-Tabar, and Nazih Moubayed. Voltage harmonic distortion compensation with non-linear load power sharing in low-voltage islanded microgrid. Mathematics and Computers in Simulation, 158:32–48, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301198.

Orozco-Gutierrez:2019:EST

[1712] M. L. Orozco-Gutierrez, G. Spagnuolo, C. A. Ramos-Paja, J. M. Ramirez-Scarpetta, and B. Ospina-Agudelo. Enhanced simulation of total cross tied photovoltaic arrays. Mathematics and Computers in Simulation, 158:49–64, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541830140X.

Varais:2019:SWE

[1713] A. Varais, X. Roboam, F. Lacressonnière, C. Turpin, J. M. Cabello, E. Bru, and J. Pulido. Scaling of wind energy conversion system for time-accelerated and sizescaled experiments. Mathematics and Computers in Simulation, 158: 65–78, April 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301393.

LeGuyadec:2019:URN

[1714] M. Le Guyadec, L. Gerbaud, E. Vinot, V. Reinbold, and C. Dumont. Use of reluctance network modelling and software component to study the influence of electrical machine pole number on hybrid electric vehicle global optimization. Mathematics and Computers in Simulation, 158: 79–90, April 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301460.

Gritli:2019:ACV

[1715] Y. Gritli, A. Tani, C. Rossi, and D. Casadei. Assessment of current and voltage signature analysis for the diagnosis of rotor magnet demagnetization in five-phase AC permanent magnet generator drives. Mathematics and Computers in Simulation, 158:91–106, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418301472.

Yang:2019:TED

[1716] Nanfang Yang, Julien Pouget, Tony Letrouvé, Cristian Jecu, and Loic Joseph-Auguste. Techno-economic design methodology of hybrid energy systems connected to electrical grid: an application of hybrid railway power substation. Mathematics and Computers in Simulation, 158:107–119, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418301691.

Mayet:2019:ICF

[1717] C. Mayet, J. Welles, A. Bouscayrol, T. Hofman, and B. Lemaire-Semail. Influence of a CVT on the fuel consumption of a parallel mediumduty electric hybrid truck. Mathematics and Computers in Simulation, 158:120–129, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418301745.

Wu:2019:ATR

[1718] Hailong Wu, Daniel Depernet, and Vincent Lanfranchi. Analysis of torque ripple reduction in a segmentedrotor synchronous reluctance machine by optimal currents. Mathematics and Computers in Simulation, 158: 130–147, April 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301721.

Bermudez:2019:MPO

[1719] M. Bermudez, O. Gomozov, X. Kestelyn, F. Barrero, N. K. Nguyen, and E. Semail. Model predictive optimal control considering current and voltage limitations: Real-time validation using OPAL-RT technologies and five-phase permanent magnet synchronous machines. Mathematics and Computers in Simulation, 158:148–161, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301897.

Hernandez-Torres:2019:TEO

[1720] David Hernández-Torres, Christophe Turpin, Xavier Roboam, and Bruno Sareni. Techno-economical optimization of wind power production including lithium and/or hydrogen sizing in the context of the day ahead market in island grids. Mathematics and Computers in Simulation, 158:162–178, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301940.

Pessot:2019:CML

[1721] Alexandra Pessot, Christophe Turpin, Amine Jaafar, Emilie Soyez, Olivier Rallières, Guillaume Gager, and Julien d Arbigny. Contribution to the modelling of a low temperature PEM fuel cell in aeronautical conditions by design of experiments. Mathematics and Computers in Simulation, 158:179–198, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418301927.

Nohra:2019:DDC

[1722] Antoine F. Hanna Nohra, Maurice Fadel, and Hadi Y. Kanaan. Design of a direct control strategy for a static shunt compensator to improve power quality in polluted and unbalanced grids. Mathematics and Computers in Simulation, 158:199– 215, April 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475418301915.

Gutierrez:2019:RTE

[1723] A. Gutierrez, M. Bressan, J. F. Jimenez, and C. Alonso. Realtime emulation of boost inverter using the Systems Modeling Language and Petri nets. Mathematics and Computers in Simulation, 158:216– 234, April 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475418301952.

Hoang:2019:SOM

[1724] Trung-Kien Hoang, Lionel Vido, Frederic Gillon, and Mohamed Gabsi. Structural optimization to maximize the flux control range of a double excitation synchronous machine. Mathematics and Computers in Simulation, 158:235–247, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302519.

Derollepot:2019:SCS

[1725] Romain Derollepot and Emmanuel Vinot. Sizing of a combined series-parallel hybrid architecture for river ship application using genetic algorithm and optimal energy management. Mathematics and Computers in Simulation, 158:248– 263, April 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418302441.

Taghia:2019:OMT

[1726] Bouazza Taghia, Bernardo Cougo, Hubert Piquet, David Malec, Antoine Belinger, and Jean-Pierre Carayon. Overvoltage at motor terminals in SiC-based PWM drives. Mathematics and Computers in Simulation, 158:264–280, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302192.

Filleau:2019:PAS

[1727] Clément Filleau, Jacques Saint-Michel, Émile Mouni, Antoine Picot, and Pascal Maussion. Power alternator simulation models for diagnostic purposes. Mathematics and Computers in Simulation, 158:281– 295, April 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418302453.

Castelan:2019:AMP

[1728] A. Castelan, B. Cougo, S. Dutour, and T. Meynard. 3D analytical modelling of plate fin heat sink on forced convection. Mathematics and Computers in Simulation, 158:296–307, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541830243X.

Nguyen:2019:VSI

[1729] Duy-Minh Nguyen, Imen Bahri, Guillaume Krebs, Eric Berthelot, and Claude Marchand. Vibration study of the intermittent control for a switched reluctance machine. Mathematics and Computers in Simulation, 158:308–325, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418302477.

Tormo:2019:ERT

[1730] Daniel Tormo, Lahoucine Idkhajine, Eric Monmasson, Ricardo Vidal-Albalate, and Ramon Blasco-Gimenez. Embedded real-time simulators for electromechanical and power electronic systems using system-on-chip devices. Mathematics and Computers in Simulation, 158:326– 343, April 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475418302490.

Merai:2019:GCC

[1731] M. Merai, M. W. Naouar, I. Slama-Belkhodja, and E. Monmasson. Grid connected converters as reactive power ancillary service providers: Technical analysis for minimum required DC-link voltage. Mathematics and Computers in Simulation, 158:344– 354, April 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418302489.

Poline:2019:SOS

[1732] Marie Poline, Laurent Gerbaud, Julien Pouget, and Frédéric Chauvet. Simultaneous optimization of sizing and energy management — Application to hybrid train. Mathematics and Computers in Simulation, 158:355–374, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302544.

Boige:2019:SPM

[1733] F. Boige, F. Richardeau, S. Lefebvre, and M. Cousineau. SiC power MOSFET in short-circuit operation: Electro-thermal macromodelling combining physical and numerical approaches with circuit-type implementation. Mathematics and Computers in Simulation, 158:375– 386, April 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418302532.

Buticchi:2019:MPD

[1734] Giampaolo Buticchi, Levy Ferreira Costa, and Marco Liserre. Multiport DC/DC converter for the electrical power distribution system of the more electric aircraft. Mathematics and Computers in Simulation, 158:387–402, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418302520.

Hmam:2019:CBF

[1735] S. Hmam, J.-C. Olivier, S. Bourguet, L. Loron, N. Bernard, and E. Schaeffer. A cycle-based formulation for the simulation of multi time-scale systems — Application to the modeling of the storage system of a fully electric ferry. Mathematics and Computers in Simulation, 158:403–417, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418302799.

Bourbon:2019:EMO

[1736] R. Bourbon, S. U. Ngueveu, X. Roboam, B. Sareni, C. Turpin, and D. Hernandez-Torres. Energy management optimization of a smart wind power plant comparing heuristic and linear programming methods. Mathematics and Computers in Simulation, 158:418–431, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418302659.

ElAabid:2019:GAC

[1737] S. El Aabid, J. Regnier, C. Turpin, O. Rallières, E. Soyez, N. Chadourne, J. d Arbigny, T. Hordé, and E. Foch. A global approach for a consistent identification of static and dynamic phenomena in a PEM Fuel Cell. Mathematics and Computers in Simulation, 158:432–452, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302787.

Mseddi:2019:RCW

[1738] Amina Mseddi, Sandrine Le Ballois, Helmi Aloui, and Lionel Vido. Robust control of a wind conversion system based on a hybrid excitation synchronous generator: a comparison between H_{∞} and CRONE controllers. Mathematics and Computers in Simulation, 158: 453–476, April 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418302994.

Morentin:2019:OOS

[1739] A. Morentin, G. Fontes, M. Mannes Hillesheim, T. Meynard, D. Flumian, J. Bourdon, and H. Piquet. OpenComp3d: an opensource framework dedicated to design in power electronics. Mathematics and Computers in Simulation, 158:477–489, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541830301X.

Vermeersch:2019:EDS

[1740] Pierre Vermeersch, François Gruson, Xavier Guillaud, Michael M. C. Merlin, and Philippe Egrot. Energy and director switches commutation controls for the alternate arm converter. Mathematics and Computers in Simulation, 158:490–505, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541830315X.

Ouchen:2019:EQI

[1741] Sabir Ouchen, Jean-Paul Gaubert, Heinrich Steinhart, and Achour Betka. Energy quality improvement of threephase shunt active power filter under different voltage conditions based on predictive direct power control with disturbance rejection principle. Mathematics and Computers in Simulation, 158:506–519, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418303203.

Petrone:2019:SIP

[1742] G. Petrone, F. Serra, G. Spagnuolo, and E. Monmasson. SoC implementation of a photovoltaic reconfiguration algorithm by exploiting a HLS-based architecture. Mathematics and Computers in Simulation, 158:520–537, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418303331.

Anonymous:2019:NIc

[1743] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 158:538, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541930014X.

Anonymous:2019:ICEc

[1744] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 158:539–540, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475419300151.

Anonymous:2019:EBf

[1745] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 158:ifc, April 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300072.

Xie:2019:MPO

[1746] Junfei Xie, Yan Wan, Kevin Mills, James J. Filliben, Yu Lei, and Zongli Lin. M-PCM-OFFD: an effective output statistics estimation method for systems of high dimensional uncertainties subject to low-order parameter interactions. Mathematics and Computers in Simulation, 159: 93–118, May 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418302817.

Zakaria:2019:CNB

[1747] Mohamad Shukri Zakaria, Farzad Ismail, Masaaki Tamagawa, Ahmad Fazli Abdul Aziz, Surjatin Wiriadidjaja, Adi Azrif Basri, and Kamarul Arifin Ahmad. A Cartesian non-boundary fitted grid method on complex geometries and its application to the blood flow in the aorta using OpenFOAM. Mathematics and Computers in Simulation, 159:220–250, May 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418303094.

Anonymous:2019:PMb

[1748] Anonymous. Pages 1–254 (May 2019). Mathematics and Computers in Simulation, 159:1–254, May 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2019:EBg

[1749] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 159:ifc, May 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419300369.

Anonymous:2019:PJ

[1750] Anonymous. Pages 1–204 (June 2019). Mathematics and Computers in Simulation, 160:1–204, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2019:EBh

[1751] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 160:ifc, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419300588.

Bisheh-Niasar:2019:BNS

[1752] Morteza Bisheh-Niasar and Krzysztof Gdawiec. Bisheh-niasar–saadatmandi root finding method via the Siteration with periodic parameters and its polynomiography. Mathematics and Computers in Simulation, 160:1–12, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418303070.

Pausinger:2019:UDS

[1753] Florian Pausinger. Uniformly distributed sequences in the orthogonal group and on the Grassmannian manifold. Mathematics and Computers in Simulation, 160:13– 22, June 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418303173.

Srivastav:2019:CWT

[1754] Vivek Kumar Srivastav, Akshoy R.

Paul, and Anuj Jain. Capturing the wall turbulence in CFD simulation of human respiratory tract. Mathematics and Computers in Simulation, 160:23–38, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418303148.

Arya:2019:FEI

[1755] Rubi Arya and Pitam Singh. Fuzzy efficient iterative method for multiobjective linear fractional programming problems. Mathematics and Computers in Simulation, 160:39– 54, June 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418303082.

Elhanafy:2019:NSB

[1756] Ahmed Elhanafy, Amr Guaily, and Ahmed Elsaid. Numerical simulation of blood flow in abdominal aortic aneurysms: Effects of blood shearthinning and viscoelastic properties. Mathematics and Computers in Simulation, 160:55–71, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418303215.

Galvis:2019:MCA

[1757] Juan Galvis and O. Andrés Cuervo. A Monte Carlo approach to computing stiffness matrices arising in polynomial chaos approximations. Mathematics and Computers in Simulation, 160:72–81, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418303033.

Huang:2019:WIS

[1758] Zhu Huang and John P. Boyd. When integration sparsification fails: Banded Galerkin discretizations for Hermite functions, rational Chebyshev functions and sinh-mapped Fourier functions on an infinite domain, and Chebyshev methods for solutions with C^{∞} endpoint singularities. Mathematics and Computers in Simulation, 160:82–102, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418303227.

Mahmoudi:2019:NMC

[1759] Mohammad Reza Mahmoudi, Mohammad Hossein Heydari, and Reza Roohi. A new method to compare the spectral densities of two independent periodically correlated time series. Mathematics and Computers in Simulation, 160:103–110, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418303276.

Beznea:2019:NAS

[1760] Lucian Beznea, Madalina Deaconu, and Oana Lupaşcu-Stamate. Numerical approach for stochastic differential equations of fragmentation; application to avalanches. Mathematics and Computers in Simulation, 160:111–125, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418303239.

Canelas:2019:SEP

[1761] Alfredo Canelas, Antonio Pereira, Jean R. Roche, and Jean P. Brancher. Solution of the equilibrium problem in electromagnetic casting considering a solid inclusion in the melt. Mathematics and Computers in Simulation, 160:126–137, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418303288.

Jaiswal:2019:NSN

[1762] Shubham Jaiswal, Manish Chopra, and S. Das. Numerical solution of non-linear partial differential equation for porous media using operational matrices. Mathematics and Computers in Simulation, 160:138– 154, June 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475418303264.

Munyakazi:2019:RFO

[1763] Justin B. Munyakazi, Kailash C. Patidar, and Mbani T. Sayi. A robust fitted operator finite difference method for singularly perturbed problems whose solution has an interior layer. Mathematics and Computers in Simulation, 160:155–167, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541830329X.

Castro:2019:ADR

[1764] Paula M. Castro, Adriana Dapena, María J. Souto-Salorio, and Ana D. Tarrío-Tobar. Algorithms for determining relative position between spheroids and hyperboloids with one sheet. Mathematics and Computers in Simulation, 160:168–179, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418303252.

Zhang:2019:MCC

[1765] Jiansong Zhang, Yuezhi Zhang, Hui Guo, and Hongfei Fu. A massconservative characteristic splitting mixed finite element method for convection-dominated Sobolev equation. Mathematics and Computers in Simulation, 160:180–191, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475419300047.

Mukherjee:2019:ECI

[1766] Debasis Mukherjee. Effect of constant immigration in plant–pathogen– herbivore interactions. Mathematics and Computers in Simulation, 160:192–200, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300035.

Anonymous:2019:NId

[1767] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 160:201, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300631.

Anonymous:2019:ICEd

[1768] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 160:202–203, June 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475419300643.

Anonymous:2019:EBi

[1769] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 161:ifc, July 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419300849.

LEcuyer:2019:EIS

[1770] Pierre L’Ecuyer. Editorial introduction to the special issue on MCM 2017. Mathematics and Computers in Simulation, 161:1, July 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419300710.

Keller:2019:IEM

[1771] Alexander Keller and Ken Dahm. Integral equations and machine learning. Mathematics and Computers in Simulation, 161:2–12, July 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419300308.

Faure:2019:IIS

[1772] Henri Faure and Christiane Lemieux. Implementation of irreducible Sobol’ sequences in prime power bases. Mathematics and Computers in Simulation, 161:13–22, July 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419300291.

Laub:2019:MCE

[1773] Patrick J. Laub, Robert Salomone, and Zdravko I. Botev. Monte Carlo estimation of the density of the sum of dependent random variables. Mathematics and Computers in Simulation, 161:23–31, July 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418303197.

Gagnon:2019:WCO

[1774] Philippe Gagnon, Mylène Bédard, and Alain Desgagné. Weak convergence and optimal tuning of the reversible jump algorithm. Mathematics and Computers in Simulation, 161:32–51, July 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301526.

Kritzer:2019:TAW

[1775] Peter Kritzer, Friedrich Pillichshammer, and G. W. Wasilkowski. Truncation in average and worst case settings for special classes of ∞- variate functions. Mathematics and Computers in Simulation, 161:52– 65, July 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418303136.

Haramoto:2019:CQA

[1776] Hiroshi Haramoto and Makoto Matsumoto. Checking the quality of approximation of p-values in statistical tests for random number generators by using a three-level test. Mathematics and Computers in Simulation, 161:66–75, July 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302039.

Harase:2019:CMT

[1777] Shin Harase. Conversion of Mersenne Twister to double-precision floatingpoint numbers. Mathematics and Computers in Simulation, 161:76– 83, July 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475418302040.

Tracheva:2019:NMC

[1778] Natalya Tracheva and Sergey Ukhinov. A new Monte Carlo method for estimation of time asymptotic parameters of polarized radiation. Mathematics and Computers in Simulation, 161:84–92, July 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418302702.

Huber:2019:FEM

[1779] Mark Huber and Bo Jones. Faster estimates of the mean of bounded random variables. Mathematics and Computers in Simulation, 161:93– 101, July 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847541930031X.

Ngo:2019:EMS

[1780] Hoang-Long Ngo and Dai Taguchi. On the Euler–Maruyama scheme for SDEs with bounded variation and Hölder continuous coefficients. Mathematics and Computers in Simulation, 161:102–112, July 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300321.

Lecot:2019:QMC

[1781] Christian Lécot, Pierre L’Ecuyer, Rami El Haddad, and Ali Tarhini. Quasi-Monte Carlo simulation of coagulation– fragmentation. Mathematics and Computers in Simulation, 161:113– 124, July 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419300515.

Anonymous:2019:NIe

[1782] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 161:125, July 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300916.

Anonymous:2019:ICEe

[1783] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 161:126–127, July 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475419300928.

Anonymous:2019:PA

[1784] Anonymous. Pages 1–270 (August 2019). Mathematics and Computers in Simulation, 162:1–270, August 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2019:EBj

[1785] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 162:ifc, August 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301144.

Hassani:2019:NCM

[1786] Hossein Hassani and Eskandar Naraghirad. A new computational method based on optimization scheme for solving variable-order time fractional Burgers’ equation. Mathematics and Computers in Simulation, 162: 1–17, August 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419300023.

Pal:2019:TSF

[1787] Shanoli Samui Pal and Samarjit Kar. Time series forecasting for stock market prediction through data discretization by fuzzistics and rule generation by rough set theory. Mathematics and Computers in Simulation, 162:18–30, August 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300011.

Youssef:2019:TPS

[1788] Maha Youssef and Gerd Baumann. Troesch’s problem solved by Sinc methods. Mathematics and Computers in Simulation, 162:31–44, August 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475419300059.

Zhao:2019:NAC

[1789] LiuWei Zhao, Jianguo Du, and Qi-Wei Wang. Nonlinear analysis and chaos control of the complex dynamics of multi-market Cournot game with bounded rationality. Mathematics and Computers in Simulation, 162:45–57, August 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300217.

Maiti:2019:GAD

[1790] Atasi Patra Maiti, B. Dubey, and A. Chakraborty. Global analysis of a delayed stage structure prey– predator model with Crowley–Martin type functional response. Mathematics and Computers in Simulation, 162:58–84, August 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541930028X.

Munoz-Vazquez:2019:ARC

[1791] Aldo Jonathan Muñoz-Vázquez, Manuel Benjamín Ortiz-Moctezuma, Anand Sánchez-Orta, and Vicente Parra-Vega. Adaptive robust control of fractionalorder systems with matched and mismatched disturbances. Mathematics and Computers in Simulation, 162:85–96, August 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300278.

Buonomo:2019:OHE

[1792] Bruno Buonomo and Rossella Della Marca. Oscillations and hysteresis in an epidemic model with information-dependent imperfect vaccination. Mathematics and Computers in Simulation, 162:97–114, August 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475419300242.

Li:2019:TNA

[1793] Shucai Li, Jing Wang, Liping Li, Shaoshuai Shi, and Zongqing Zhou. The theoretical and numerical analysis of water inrush through filling structures. Mathematics and Computers in Simulation, 162:115– 134, August 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475418303343.

Gonzalez:2019:MMC

[1794] Mario González, Víctor Cárdenas, Homero Miranda, and Ricardo Álvarez-Salas. Modular multilevel converter for large-scale photovoltaic generation with reactive power flow and unbalanced active power extraction capabilities. Mathematics and Computers in Simulation, 162:135– 154, August 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419300266.

Zhang:2019:COC

[1795] Limin Zhang, Chaofeng Zhang, and Zhirong He. Codimension-one and codimension-two bifurcations of a discrete predator–prey system with strong Allee effect. Mathematics and Computers in Simulation, 162:155– 178, August 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419300254.

Kong:2019:EAN

[1796] Linghua Kong, Pengfei Zhu, Yushun Wang, and Zhankuan Zeng. Efficient and accurate numerical methods for the multidimensional convection– diffusion equations. Mathematics and Computers in Simulation, 162:179– 194, August 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419300345.

Wang:2019:TCM

[1797] Kung-Jeng Wang, Kun-Shan Lee, and Jia-Hong Liao. Technology cooperation modeling of multiple profitcentered business units: a system dynamics framework. Mathematics and Computers in Simulation, 162: 195–220, August 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418303021.

Mazzucco:2019:ACW

[1798] G. Mazzucco, B. Pomaro, V. A.

Salomoni, and C. E. Majorana. Apex control within an elastoplastic constitutive model for confined concretes. Mathematics and Computers in Simulation, 162:221– 232, August 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419300485.

Kotb:2019:SIM

[1799] M. S. Kotb and M. Z. Raqab. Statistical inference for modified Weibull distribution based on progressively type-II censored data. Mathematics and Computers in Simulation, 162:233–248, August 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541930045X.

Drazic:2019:SFP

[1800] Ivan Dražić, Nelida Črnjarić-Žic, and Loredana Simčić. A shear flow problem for compressible viscous micropolar fluid: Derivation of the model and numerical solution. Mathematics and Computers in Simulation, 162:249–267, August 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419300333.

Anonymous:2019:NIf

[1801] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 162:268, August 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419301193.

Anonymous:2019:ICEf

[1802] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 162:269–270, August 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847541930120X.

Anonymous:2019:PS

[1803] Anonymous. Pages 1–206 (September 2019). Mathematics and Computers in Simulation, 163:1–206, September 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2019:EBk

[1804] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 163:ifc, September 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301284.

L:2019:MMN

[1805] Diego F. Aranda L., Gilberto González-Parra, and Tommaso Benincasa. Mathematical modeling and numerical simulations of Zika in Colombia considering mutation. Mathematics and Computers in Simulation, 163:1–18, September 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419300679.

Broto:2019:SII

[1806] Baptiste Broto, François Bachoc, Marine Depecker, and Jean-Marc Martinez. Sensitivity indices for independent groups of variables. Mathematics and Computers in Simulation, 163:19–31, September 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300564.

Liu:2019:PGF

[1807] Jia Liu. The parameterized Gallager’s first bounds based on conditional triplet-wise error probability. Mathematics and Computers in Simulation, 163:32–46, September 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300539.

DelaSen:2019:SEM

[1808] M. De la Sen, S. Alonso-Quesada, A. Ibeas, and R. Nistal. On an SEIADR epidemic model with vaccination, treatment and dead-infectious corpses removal controls. Mathematics and Computers in Simulation, 163: 47–79, September 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300709.

Xie:2019:HWM

[1809] Jiaquan Xie, Tao Wang, Zhongkai Ren, Jun Zhang, and Long Quan. Haar wavelet method for approximating the solution of a coupled system of fractional-order integral– differential equations. Mathematics and Computers in Simulation, 163: 80–89, September 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300680.

Mendy:2019:HBG

[1810] A. Mendy, J. J. Tewa, M. Lam, and P. Tchinda Mouofo. Hopf bifurcation in a grazing system with two delays. Mathematics and Computers in Simulation, 163:90–129, September 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419300540.

Chen:2019:NPE

[1811] Qihuai Chen, Tianliang Lin, Haoling Ren, and Shengjie Fu. Novel potential energy regeneration systems for hybrid hydraulic excavators. Mathematics and Computers in Simulation, 163: 130–145, September 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300758.

Jafari:2019:LDA

[1812] Nasrin Jafari, Mojtaba Azhari, and Bijan Boroomand. Large deformation analysis of moderately thick viscoelastic plates. Mathematics and Computers in Simulation, 163:146– 167, September 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419300746.

Blondin:2019:CTA

[1813] M. J. Blondin, P. Sicard, and P. M.

Pardalos. Controller Tuning Approach with robustness, stability and dynamic criteria for the original AVR System. Mathematics and Computers in Simulation, 163:168–182, September 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419300771.

Skjong:2019:NSD

[1814] Stian Skjong and Eilif Pedersen. On the numerical stability in dynamical distributed simulations. Mathematics and Computers in Simulation, 163: 183–203, September 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847541930076X.

Anonymous:2019:NIg

[1815] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 163:204, September 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419301338.

Anonymous:2019:ICEg

[1816] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 163:205–206, September 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847541930134X.

Anonymous:2019:EBl

[1817] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 164:ifc, October 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301600.

Amaziane:2019:MSI

[1818] Brahim Amaziane and Domingo Barrera. MATCOM Special Issue MAMERN VII-2017: 7th International Conference on Approximation Methods and Numerical Modeling in Environment and Natural Resources. Mathematics and Computers in Simulation, 164:1–2, October 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301235.

Ahusborde:2019:FIF

[1819] E. Ahusborde, M. El Ossmani, and M. Id Moulay. A fully implicit finite volume scheme for single phase flow with reactive transport in porous media. Mathematics and Computers in Simulation, 164:3–23, October 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302118.

Allouch:2019:DSD

[1820] C. Allouch, A. Boujraf, and M. Tahrichi.

Discrete superconvergent degenerate kernel method for Fredholm integral equations. Mathematics and Computers in Simulation, 164:24– 32, October 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419300205.

Bouhiri:2019:CQI

[1821] S. Bouhiri, A. Lamnii, and M. Lamnii. Cubic quasi-interpolation spline collocation method for solving convection– diffusion equations. Mathematics and Computers in Simulation, 164: 33–45, October 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475418302969.

Bouhiri:2019:SQI

[1822] S. Bouhiri, A. Lamnii, M. Lamnii, and A. Zidna. A C spline quasiinterpolant for fitting 3D data on the sphere and applications. Mathematics and Computers in Simulation, 164:46–62, October 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418301666.

Boret:2019:DSM

[1823] Saúl E. Buitrago Boret and Olivo Romero Marin. Development of Surrogate models for CSI probabilistic production forecast of a heavy oil field. Mathematics and Computers in Simulation, 164:63–77, October 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418303185.

Fortes:2019:FHU

[1824] M. A. Fortes, P. González, A. Palomares, and M. L. Rodríguez. Filling holes using a mesh of filled curves. Mathematics and Computers in Simulation, 164:78–93, October 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541830332X.

Gonzalez:2019:FDA

[1825] P. González, H. Idais, M. Pasadas, and M. Yasin. 3D fuzzy data approximation by fuzzy smoothing bicubic splines. Mathematics and Computers in Simulation, 164:94– 102, October 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418302751.

Guiraldello:2019:ISM

[1826] Rafael T. Guiraldello, Roberto F.

Ausas, Fabricio S. Sousa, Felipe Pereira, and Gustavo C. Buscaglia. Interface spaces for the Multiscale Robin Coupled Method in reservoir simulation. Mathematics and Computers in Simulation, 164:103– 119, October 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418302842.

Ibanez:2019:ERV

[1827] M. J. Ibáñez, F. Jiménez-Molinos, J. B. Roldán, and R. Yáñez. Estimation of the reset voltage in resistive RAMs using the charge–flux domain and a numerical method based on quasi-interpolation and discrete orthogonal polynomials. Mathematics and Computers in Simulation, 164: 120–130, October 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418303318.

Idais:2019:OKA

[1828] H. Idais, M. Yasin, M. Pasadas, and P. González. Optimal knots allocation in the cubic and bicubic spline interpolation problems. Mathematics and Computers in Simulation, 164: 131–145, October 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418302957.

Jimenez-Gamero:2019:TPT

[1829] M. D. Jiménez-Gamero and M. V. Alba-Fernández. Testing for the Poisson–Tweedie distribution. Mathematics and Computers in Simulation, 164:146–162, October 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418301988.

Mghazli:2019:GPE

[1830] Zoubida Mghazli and Ilyas Naji. Guaranteed a posteriori error estimates for a fractured porous medium. Mathematics and Computers in Simulation, 164:163–179, October 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419300503.

Silva:2019:SFH

[1831] Cristiana J. Silva and Delfim F. M. Torres. Stability of a fractional HIV/AIDS model. Mathematics and Computers in Simulation, 164:180– 190, October 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419301582.

Anonymous:2019:EBm

[1832] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 164:191, October 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301673.

Anonymous:2019:EBn

[1833] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 164:192–193, October 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301685.

Anonymous:2019:PN

[1834] Anonymous. Pages 1–320 (November 2019). Mathematics and Computers in Simulation, 165:1–320, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2019:EBo

[1835] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 165:ifc, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541930182X.

Abbasi:2019:HCT

[1836] Ghasem Abbasi and Alaeddin Malek. Hyperthermia cancer therapy by domain decomposition methods using strongly continuous semigroups. Mathematics and Computers in Simulation, 165:1–12, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300734.

Ahsan:2019:NHW

[1837] Muhammad Ahsan, Imtiaz Ahmad, Masood Ahmad, and Iltaf Hussian. A numerical Haar wavelet-finite difference hybrid method for linear and nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 165:13–25, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300692.

Muller:2019:NCM

[1838] Laura Müller, Axel Klar, and Florian Schneider. A numerical comparison of the method of moments for the population balance equation. Mathematics and Computers in Simulation, 165:26–55, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300783.

Singh:2019:VCL

[1839] Prem Kumar Singh. Vague concept lattice reduction using granular computing and vague entropy. Mathematics and Computers in Simulation, 165:56–73, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300552.

Arthi:2019:CFO

[1840] G. Arthi, Ju H. Park, and K. Suganya. Controllability of fractional order damped dynamical systems with distributed delays. Mathematics and Computers in Simulation, 165:74– 91, November 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419300813.

Wiesner:2019:SEI

[1841] A. Wiesner, R. Diez, D. Florez, and H. Piquet. System for experimental investigation of DBD excilamps in view of control and optimization of UV emission. Mathematics and Computers in Simulation, 165: 92–106, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419300795.

Wei:2019:MMI

[1842] Hsiu-Chuan Wei. A mathematical model of intraguild predation with prey switching. Mathematics and Computers in Simulation, 165:107– 118, November 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419301004.

Gie:2019:VCC

[1843] Gung-Min Gie, Chang-Yeol Jung, and Thien Binh Nguyen. Validation of a 2D cell-centered Finite Volume method for elliptic equations. Mathematics and Computers in Simulation, 165: 119–138, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419301041.

Ramos:2019:TDT

[1844] Higinio Ramos and M. A. Rufai. A third-derivative two-step block Falkner-type method for solving general second-order boundaryvalue systems. Mathematics and Computers in Simulation, 165:139– 155, November 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419300990.

Mbroh:2019:FOF

[1845] Nana A. Mbroh and Justin B. Munyakazi. A fitted operator finite difference method of lines for singularly perturbed parabolic convection– diffusion problems. Mathematics and Computers in Simulation, 165:156– 171, November 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847541930103X.

Capone:2019:TPR

[1846] F. Capone, M. F. Carfora, R. De Luca, and I. Torcicollo. Turing patterns in a reaction–diffusion system modeling hunting cooperation. Mathematics and Computers in Simulation, 165: 172–180, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301065.

Ullah:2019:MAT

[1847] Saif Ullah, Muhammad Altaf Khan, Muhammad Farooq, and Taza Gul. Modeling and analysis of Tuberculosis (TB) in Khyber Pakhtunkhwa, Pakistan. Mathematics and Computers in Simulation, 165:181–199, November 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419301089.

Clarke:2019:PCS

[1848] Lauren Elizabeth Clarke and Gautham Krishnamoorthy. Pre-conditioning strategies to accelerate the convergence of iterative methods in multiphase flow simulations. Mathematics and Computers in Simulation, 165: 200–222, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301053.

Erfanian:2019:SNI

[1849] Majid Erfanian and Amin Mansoori. Solving the nonlinear integrodifferential equation in complex plane with rationalized Haar wavelet. Mathematics and Computers in Simulation, 165:223–237, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419301028.

Zeghdane:2019:NSS

[1850] Rebiha Zeghdane. Numerical solution of stochastic integral equations by using Bernoulli operational matrix. Mathematics and Computers in Simulation, 165:238–254, November 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419301016.

Kumar:2019:CFS

[1851] Sunil Kumar, B. V. Rathish Kumar, and J. H. M. Ten Thije Boonkkamp. Complete flux scheme for elliptic singularly perturbed differential–difference equations. Mathematics and Computers in Simulation, 165:255–270, November 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419301119.

Wang:2019:CEP

[1852] Lichun Wang. Computing the estimator of a parameter vector via a competing Bayes method. Mathematics and Computers in Simulation, 165: 271–279, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419301077.

Khattri:2019:CFS

[1853] Khim B. Khattri and Shiva P. Pudasaini. Channel flow simulation of a mixture with a full-dimensional generalized quasi two-phase model. Mathematics and Computers in Simulation, 165:280–305, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301107.

Fu:2019:TDT

[1854] Zhuo jia Fu, Jin hong Shi, Wen Chen, and Li wen Yang. Three-dimensional transient heat conduction analysis by boundary knot method. Mathematics and Computers in Simulation, 165: 306–317, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301533.

Anonymous:2019:NIh

[1855] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 165:318, November 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301879.

Anonymous:2019:ICEh

[1856] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 165:319–320, November 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419301880.

Anonymous:2019:PD

[1857] Anonymous. Pages 1–530 (December 2019). Mathematics and Computers in Simulation, 166:1–530, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2019:EBp

[1858] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 166:ifc, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302411.

Sobieski:2019:NIT

[1859] Wojciech Sobieski. Numerical investigations of tortuosity in randomly generated pore structures. Mathematics and Computers in Simulation, 166:1–20, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419301399.

Yan:2019:SNF

[1860] Xiaoqiang Yan and Chengjian Zhang. Solving nonlinear functional–differential and functional equations with constant delay via block boundary value methods. Mathematics and Computers in Simulation, 166:21– 32, December 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419301387.

Hausken:2019:GCD

[1861] Kjell Hausken. Governmental combat of the dynamics of multiple competing terrorist organizations. Mathematics and Computers in Simulation, 166:33–55, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301375.

Hasan:2019:MMZ

[1862] Bushra Hasan, Manmohan Singh, David Richards, and Aaron Blicblau. Mathematical modelling of Zika virus as a mosquito-borne and sexually transmitted disease with diffusion effects. Mathematics and Computers in Simulation, 166:56– 75, December 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419301430.

Pulch:2019:MOR

[1863] Roland Pulch. Model order reduction for random nonlinear dynamical systems and low-dimensional representations for their quantities of interest. Mathematics and Computers in Simulation, 166:76–92, December 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419301442.

Dincer:2019:ISF

[1864] Hasan Dinçer and Serhat Yüksel. An integrated stochastic fuzzy MCDM approach to the balanced scorecardbased service evaluation. Mathematics and Computers in Simulation, 166: 93–112, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301454.

Wrobel:2019:MNA

[1865] Mateusz Wróbel. Mathematical and numerical analysis of initial boundary value problem for a linear nonlocal equation. Mathematics and Computers in Simulation, 166:113– 125, December 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847541930148X.

Tian:2019:BOP

[1866] Miao Tian, Xiangfeng Yang, and Yi Zhang. Barrier option pricing of mean-reverting stock model in uncertain environment. Mathematics and Computers in Simulation, 166: 126–143, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419301466.

Nazir:2019:BBI

[1867] Shah Nazir, Sara Shahzad, Rahmita Wirza, Rohul Amin, Muhammad Ahsan, Neelam Mukhtar, Iván García-Magariño, and Jaime Lloret. Birthmark based identification of software piracy using Haar wavelet. Mathematics and Computers in Simulation, 166:144–154, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419301478.

Babayan:2019:TAH

[1868] Narek Babayan and Mojtaba Tahani. Team Arrangement Heuristic Algorithm (TAHA): Theory and application. Mathematics and Computers in Simulation, 166:155–176, December 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419301569.

Mejia-Parra:2019:FDA

[1869] Daniel Mejia-Parra, Aitor Moreno, Jorge Posada, Oscar Ruiz-Salguero, Iñigo Barandiaran, Juan Carlos Poza, and Raúl Chopitea. Frequency-domain analytic method for efficient thermal simulation under curved trajectories laser heating. Mathematics and Computers in Simulation, 166: 177–192, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301557.

Fahmy:2019:DOS

[1870] Mohamed Abdelsabour Fahmy. Design optimization for a simulation of rotating anisotropic viscoelastic porous structures using time-domain OQBEM. Mathematics and Computers in Simulation, 166:193–205, December 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419301521.

Fu:2019:MNE

[1871] Yayun Fu, Yongzhong Song, and Yushun Wang. Maximum-norm error analysis of a conservative scheme for the damped nonlinear fractional Schrödinger equation. Mathematics and Computers in Simulation, 166:

206–223, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419301491.

Heinrich:2019:HSS

[1872] Claudio Heinrich, Mikko S. Pakkanen, and Almut E. D. Veraart. Hybrid simulation scheme for volatility modulated moving average fields. Mathematics and Computers in Simulation, 166:224–244, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419301405.

Qasim:2019:LDD

[1873] Tehreem Qasim and Naeem Bhatti. A low dimensional descriptor for detection of anomalies in crowd videos. Mathematics and Computers in Simulation, 166:245–252, December 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419301788.

Azhdari:2019:ANS

[1874] Ebrahim Azhdari and Aram Emami. Analytical and numerical study of drying of tomato in non-shrinkage and shrinkage model. Mathematics and Computers in Simulation, 166: 253–265, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419301776.

Williamson:2019:PEE

[1875] Kevin Williamson, Pavel Burda, and Bedřich Sousedík. A posteriori error estimates and adaptive mesh refinement for the Stokes–Brinkman problem. Mathematics and Computers in Simulation, 166:266–282, December 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419301910.

CastilloGonzalez:2019:STQ

[1876] Rodrigo Castillo González, Julio B.

Clempner, and Alexander S. Poznyak. Solving traffic queues at controlledsignalized intersections in continuoustime Markov games. Mathematics and Computers in Simulation, 166: 283–297, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302009.

Qiu:2019:IDS

[1877] Wenlin Qiu, Hongbin Chen, and Xuan Zheng. An implicit difference scheme and algorithm implementation for the one-dimensional timefractional Burgers equations. Mathematics and Computers in Simulation, 166:298–314, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301934.

Yang:2019:ESS

[1878] Xuxin Yang, Antti Rasila, and Tommi Sottinen. Efficient simulation of the Schrödinger equation with a piecewise constant positive potential. Mathematics and Computers in Simulation, 166:315–323, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301764.

Veeresha:2019:NTD

[1879] P. Veeresha and D. G. Prakasha. A novel technique for (2 + 1)- dimensional time-fractional coupled Burgers equations. Mathematics and Computers in Simulation, 166:324– 345, December 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419302034.

Horng:2019:BAA

[1880] Shih-Cheng Horng and Shieh-Shing Lin. Bat algorithm assisted by ordinal optimization for solving discrete probabilistic bicriteria optimization problems. Mathematics and Computers in Simulation, 166:346– 364, December 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419302010.

Paffuti:2019:GMD

[1881] Giampiero Paffuti. Galerkin method for discs capacitors. Mathematics and Computers in Simulation, 166: 365–381, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302071.

Syafiie:2019:NTV

[1882] S. Syafiie. Nonlinear time varying perturbation stability analysis of a double diabetes system. Mathematics and Computers in Simulation, 166: 382–394, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302058.

Moreno:2019:CSG

[1883] E. Moreno, P. Cruz-Hernández, Z. Hemmat, A. Mojab, and E. A. Michael. Comparative study of grids based on the cubic crystal system for the FDTD solution of the wave equation. Mathematics and Computers in Simulation, 166:395– 409, December 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419302125.

Madureira:2019:GEN

[1884] Rodrigo L. R. Madureira, Mauro A.

Rincon, and Moncef Aouadi. Global existence and numerical simulations for a thermoelastic diffusion problem in moving boundary. Mathematics and Computers in Simulation, 166: 410–431, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302162.

Shit:2019:PFH

[1885] G. C. Shit, S. Maiti, M. Roy, and J. C. Misra. Pulsatile flow and heat transfer of blood in an overlapping vibrating atherosclerotic artery: a numerical study. Mathematics and Computers in Simulation, 166: 432–450, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302137.

Bayati:2019:IVN

[1886] Mahsa Bayati, Miriam Leeser, Yijia Gu, and Thomas Wahl. Identifying volatile numeric expressions in numeric computing applications. Mathematics and Computers in Simulation, 166: 451–460, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302149.

Garcia-Nieto:2019:MAA

[1887] P. J. García-Nieto, E. García-Gonzalo, J. R. Alonso Fernández, and C. Díaz Muñiz. Modeling of the algal atypical increase in La Barca reservoir using the DE optimized least square support vector machine approach with feature selection. Mathematics and Computers in Simulation, 166:461– 480, December 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419302265.

Coskun:2019:IVI

[1888] Safa Bozkurt Coşkun, Mehmet Tarık Atay, and Erman Şentürk. Interpolated variational iteration method for solving the jamming transition problem. Mathematics and Computers in Simulation, 166:481–493, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302216.

Xu:2019:BDF

[1889] Yang Xu, Yanming Zhang, and Jingjun Zhao. Backward difference formulae and spectral Galerkin methods for the Riesz space fractional diffusion equation. Mathematics and Computers in Simulation, 166:494– 507, December 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419302228.

Zhang:2019:EEC

[1890] Yanfang Zhang. To enter or not to enter? A competitive analysis with minimum quality standards. Mathematics and Computers in Simulation, 166:508–527, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302186.

Anonymous:2019:NIi

[1891] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 166:528, December 2019. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302460.

Anonymous:2019:ICEi

[1892] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 166:529–530, December 2019. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419302472.